A Complete Bibliography of IEEE Transactions on Computers (2020–2029)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 September 2025 Version 1.58

Title word cross-reference

2.5 [NAP^{+}20]. 2^{m} [GMZ22]. \{2^{q}, 2^{q} ∓ 1, 2^{2q} + 1\} [JRL25]. 2^{q} ∓ \sqrt{−1} [JRL25]. 3 [FMM^{+}21, LSCX20, LQY^{+}20, WLQ^{+}21, WZD^{+}20]. 8k [AJ22]. ^{++} [BZW^{+}25b]. ^{2} [HYW^{+}21, YWQ^{+}25]. ^{RT} [JYF^{+}23]. O(n) [EDGR^{+}24]. D [LZC^{+}21]. GF (2^{m}) [Ima21]. FiLIPDSM [RKMR23]. K [ZGR^{+}25, ZWWY22, ZLC^{+}23b]. GF(2^{m}) [ERKP21]. n [ZLC^{+}23b]. O(N) [GR23]. p^{k} [Koç20]. π [LQY^{+}20]. PoisonedGNN [APH^{+}23].

-1 [JRL25]. -Ary [ZLC^{+}23b]. -BA [LQY^{+}20]. -Based [NKN^{+}25]. -bit [AJ22]. -Cubes [ZLC^{+}23b]. -D [WLQ^{+}21]. -Inspired [HXL^{+}23]. -Level [ZWWY22]. -MAB [ZWWY22]. -Tree [WBJC22].

1 [JRL25]. 1.0 [PCA^{+}24]. 10K [RvSP^{+}25]. 123.0 [BGHR^{+}25]. 123.0-B-2 [BGHR^{+}25]. 128 [ABP25]. 128a [DPS22]. 16-bit [AJ22]. 1687 [IIEKS24]. 192 [ABP25].

2 [BGHR^{+}25, HYW^{+}21, LCS^{+}25, PC24, SAJA21]. 2-D [PC24]. 2-Way [PLB22]. 2.0 [MDR^{+}24]. 22q [JRL25]. 256 [ABP25]. 29-Billion [WMG^{+}25]. 2D [SM24]. 2D/3D [SM24]. 2q [JRL25]. 2q∓[JRL25].

3 [FZM^{+}23, LZC^{+}21, SCL^{+}24]. 3- [LZC^{+}21]. 3-D [FZM^{+}23]. 3D [FXL^{+}24, HCC^{+}25, LL23, PLZ^{+}23, PZY^{+}23, SM24, TKN23, WTL^{+}21, WHL^{+}23, WHK24, YZG^{+}25]. 3D-Stacked [HCC^{+}25, WTL^{+}21]. 3DIC [SRB23]. 3RSeT [CFA22].

4-Way [NS22].

5G [NWL^{+}25, RPS^{+}21].

61131 [SCL^{+}24]. 61131-3 [SCL^{+}24]. 64-Bit [MDPM24].

7nm [RvSP^{+}25].

8 [LFP^{+}22].

A3C [ZHYJ21]. A3C-DO [ZHYJ21]. AAPP [LXW^{+}23]. AAV [LWZ^{+}25]. ABE [CLZG22, ZOH^{+}25]. Abnormal [BXW^{+}25]. Abort [XXZ^{+}25]. Abstraction [WDQ^{+}22]. Abstractions [KMH^{+}23]. Abuse [ZOH^{+}25]. Accelerate [LY21, MLW^{+}25]. Accelerated [GWG^{+}25, JCZ^{+}23, WYB^{+}24, SYL^{+}23]. Accelerating [CHC25, FWZ^{+}21, FNS^{+}22, GXL^{+}24, HLL^{+}20, HZM^{+}23, JCZ^{+}23, JLP^{+}25, KPS^{+}24, KCL^{+}20, KPD^{+}23, LLM^{+}25, LCM25, LMH^{+}25, LY20, LDW^{+}25, MHM^{+}23, NWH^{+}25, QJY^{+}25, SIR20, SZL^{+}22, SWR^{+}23, SZL^{+}24, TRG^{+}24, TBS^{+}25, XLW^{+}20, ZLL^{+}23, ZMS^{+}23, ZLZ^{+}23]. Acceleration [BCBS21, CWS^{+}24a, CLLdS25, DSCB25, DPQK^{+}23, FAFK21, FFG^{+}25, FAKM22, GWG^{+}24, GWX^{+}23, KSKK23, KZXR25, KBQ^{+}23, KXGS22, KYS^{+}22, LQC^{+}22, LDF^{+}24, LWYJ23, MC23, NTL^{+}24, SKH^{+}25, SYW^{+}22, TDZ^{+}22, WGL^{+}20, WDL^{+}25, WFH25, XXC^{+}25, YHH^{+}25, YQX^{+}25, ZCR22, ZGWY22, ZWX^{+}25, ZYZ^{+}25a, ZZW^{+}25, ZYZ^{+}25b]. Accelerations [WYZ^{+}22]. Accelerative [LXW^{+}23]. Accelerator [ABP25, ACG20, AGQ^{+}23, CNOS22, ESW^{+}23, EAMK22, FB20, GSY^{+}20, HPJK22, HLQ^{+}23, HLJ^{+}25, JLZ21, KKS^{+}22, KPL^{+}22, KAWR23, KJK24, KLR23, LWL^{+}21, LL23, LQY^{+}20, LHL^{+}23, LZW^{+}24a, LYW^{+}25, LGW^{+}22, LDT^{+}25, MS25, NWL^{+}25, NKeSK^{+}23, NKL^{+}23, RPB^{+}23, SM22, SM24, SSCK25, TRV20, WGM^{+}20, WGJZ21, XJY^{+}24, YYCR24, YLZ^{+}24, YLC^{+}21, YWC^{+}24a, ZCR23, ZLWJ23, ZLL^{+}24b, ZCS^{+}25]. Accelerators [AB20, AC22, BWL^{+}25, BFG^{+}21, CAC^{+}22, DMD^{+}23, DRP24, JLY^{+}21, KKH22, KBR^{+}23, LHXH24, LWL^{+}24, LWW^{+}24, LLY22, MHJ^{+}21, MCS^{+}22, NICY24, OKU^{+}24, PAR^{+}22, PN24, QWK20, SKK23, SMC^{+}25, TPWY23, WSLX24, ZAS^{+}22, ZML^{+}25, Lu21]. Access [CKJ^{+}22, DQ23, HPJK22, HJYL22, JWS^{+}23, LV23, RPB^{+}23, RMKO23, WLW^{+}22a, WLY^{+}23, XGMJ25, ZXZ^{+}24, ddAPdS21]. Access-Pattern [XGMJ25]. Accesses [HYS^{+}20, SAG22]. Account [BHW^{+}23]. Account-Model [BHW^{+}23]. Accumulate [SNT22, ZCK20]. Accumulative [NNH^{+}25]. Accuracy [AMM21, DA22, LLL25b, SNT22, WMG^{+}25, ZMH^{+}25]. Accurate [BAM^{+}24, BCMT23, DYJ20, NML25, SMZ^{+}20, TTG^{+}23, USS^{+}21]. aChain [WPL^{+}23]. Achieving [AG24, SZHB21, XQC^{+}22, ZXL^{+}24]. ACL [XJL^{+}25]. Across [LHA^{+}25, LZW25, WGD^{+}22, ZTY^{+}25]. Activations [WL24]. Active [EDGR^{+}24, MHS^{+}20]. Acyclic [MBP21]. Ada [GLW^{+}24]. Ada-WL [GLW^{+}24]. Adapt [HZR^{+}23]. Adaptation [BSM21, OAC^{+}21, SBP^{+}20, SZ22, ZZF^{+}24]. AdaptDQC [XLT^{+}25]. Adapting [ZLL^{+}23, ZHL^{+}24]. Adaptive [APV22, BZW^{+}25a, BCBS21, BJMKK23, CTY^{+}24, CXL^{+}23, DSK23, DPS^{+}20, FLS20, FWZ^{+}21, GWG^{+}25, GLW^{+}24, HLS^{+}23a, HWC^{+}22a, HLY^{+}25, HBB^{+}21, IIEKS24, KPS^{+}24, KKKC20, KLC20, LL21, LQL^{+}24, LHXH22, LXW^{+}23, LHZR25, MLW^{+}23, NK22, NHW^{+}24, NKN^{+}25, OAB^{+}23, PLZ^{+}23, PYDG22, QWT^{+}23, STZ^{+}24, SMY22, SKH^{+}25, TKN23, WCB23, WL24, XLT^{+}25, XXC^{+}25, XJL^{+}25, YHW^{+}25, YLT^{+}23, YLL^{+}24, YHV^{+}21, YYZ^{+}25, YDL^{+}25, ZZG20, ZGQ^{+}22, ZXMX25, ZCC^{+}23, ZLM^{+}24, ZDC^{+}25, ZTZ^{+}25, ZYZ^{+}25b, HXL^{+}25]. Adaptive-Length [FLS20]. Adaptively [BB22]. AdaptMD [YLL^{+}24]. ADC [KJK24]. ADC-Free [KJK24]. Adder [JYM20, KK25]. Adder/Multiplier [JYM20]. Adders [JRL25, Mik24, RRDB20]. Adding [OTTT22]. Addition [JRL25]. Additive [XGZ^{+}24]. Address [HKS20, HLK^{+}25, SZL^{+}22]. Addressable [CWWW20, KSL^{+}22, Rot24]. Addressing [LHW^{+}25, WLW^{+}22a, ZZZ^{+}20, ZGL^{+}21]. AdEle [TKN23]. Adjacency [SXZJ24]. Adjusted [LL23]. Adjustment [APV22, JMW^{+}24, LQY^{+}20, ZCY^{+}25]. ADLPT [PLZ^{+}23]. Advance [WL20]. Advanced [KD25, NCD^{+}25, PHL^{+}25, QZZ^{+}24, RvSP^{+}25, VAV^{+}20]. Advanced-Feature [NCD^{+}25]. Adversarial [CCY^{+}24, HHN^{+}23, IKAG^{+}22, JYH^{+}24, LLX^{+}24a, LGC^{+}23, QZZ^{+}24, RSR22, TZY^{+}24, WHQ^{+}24, XXJ^{+}24, ZTY^{+}23, ZML^{+}24, ZLT^{+}24, ZCY^{+}24, BGM^{+}23]. Adversary [RBSG23]. AEML [TDZ^{+}22]. AES [CLCL22, RMTA20, UMM^{+}20, ZSWS24]. Affinity [XZL^{+}21]. Affinity-Aware [XZL^{+}21]. Against [BGHR^{+}25, BCMT23, CLCC25, CMGD24, CCY^{+}24, EGP24, GLMZ25, LHFW24, MXY^{+}23, PPQBA21, SKK^{+}21, SXH^{+}24, TZY^{+}24, XXJ^{+}24, ZML^{+}24, ZXZ^{+}24, ZJJ25, ZLH^{+}21, ZCY^{+}24]. Age [XWL^{+}24]. Age-Aware [XWL^{+}24]. Agent [YBF^{+}25, ZDW^{+}23]. Aggregate [HC24]. Aggregation [AB20, LLD^{+}25, LWZZ25, YPL^{+}25]. Aggregator [XWL^{+}24]. Aggressive [PLH^{+}24]. Agile [FQYS23, TLLL25]. Aging [ESN20, HMMP23, MSZ22, RCS^{+}21]. Aging-Aware [RCS^{+}21]. Agnostic [CFW25]. Agreement [DGX^{+}25]. Ahead [HKS20, LPD^{+}21]. AI [GWD25, HHZ^{+}23, MS25, PM25]. Aided [GLGL23]. AILC [LY21]. AINNS [ZCR23]. Airborne [MNB20]. ALAMNI [DSK23]. Algebraic [BPM23]. Algorithm [BGHR^{+}25, GYH^{+}22, GQH21, GPQ22, GPQ23, HJX^{+}25, HWJ^{+}21, LHXH22, LL23, LGW^{+}22, LSXZ21, OLZ^{+}20, PAR^{+}22, SZ22, WYZ^{+}22, WDW^{+}23, WWL^{+}25, XLY^{+}25, ZCP23, ZLL^{+}22a, ZWM20]. Algorithm-Centric [HWJ^{+}21]. Algorithm-Oriented [ZLL^{+}22a]. Algorithm/Architecture [XLY^{+}25]. Algorithmics [DWYX20]. Algorithms [DVV23, EAMJ^{+}23, God20, GJ20, HWZ^{+}25, JLL^{+}20, Koç20, KB21, MÖS22, MTK25, PN24, SGS^{+}21, WGM^{+}20, WGJZ21, WLD^{+}22, YLL^{+}20, YLHL23, ZTZ^{+}25]. Alignment [BAM^{+}24, KXGS22, QHZ^{+}21, SXC^{+}24]. All-Digital [ZCF20]. All-Inclusive [ZCR23]. Alleviate [WDZ^{+}22]. Alleviating [CSH^{+}24]. Allocation [CTY^{+}24, GQZ21, HLZ^{+}25, HXGR20, LLL^{+}20, OKC^{+}20, SCC21, WTL^{+}24, WWS^{+}22, YHW^{+}25, YBF^{+}25, ZCF20]. Allocator [FCZ^{+}23, HZT^{+}23]. Allspark [GWC^{+}25]. ALPINE [KBQ^{+}23]. Alterable [CCYC22]. Alternative [CLCL22]. Amazon [TRV20]. AMBEA [PLH^{+}24]. Ameliorate [LFW21]. Amnesiac [SKK^{+}21]. Among [JZH^{+}24, LZX^{+}25]. Amplification [WHX^{+}25]. Amplitudes [ddAPdS21]. AMR [LLL^{+}23]. AMS [FV23]. Analog [FLF20, KBQ^{+}23, Rot24, TOF^{+}24]. Analog/Mixed [FLF20]. Analysis [AG22, BKS22, BY24, BFC20, CGS^{+}20, CHM25, CRJZ21, DA22, EGMW21, Fic22, GSB23, GSC^{+}23, HDL^{+}24, HZF^{+}24, JGD^{+}21, JCY^{+}23, JCKH22, JLP^{+}25, KLR^{+}20, LJY21, LWH20, LSXZ21, MYGA20, MDR^{+}24, BCLC25, RAD20, RSZ23, RACB24, SLS^{+}21, SYD^{+}24, SCY21, SKK23, SLLS25, SGL^{+}20, TTG^{+}23, UYZP22, WLR20, WWJ^{+}23, WNL^{+}23, WFH^{+}24, WLZ^{+}25, WZGT22, XLT^{+}25, XAP20, XNH^{+}25, YNJS21, YZG^{+}25, ZABHCG23, ZYD^{+}20, ZGR^{+}25]. Analytic [HMZ^{+}25]. Analytical [DGTVGG21, WPL^{+}23, WLF^{+}25b]. Analytics [CXY24, HLS^{+}23b, LWYJ23, MAM23, MCS^{+}22, QWT^{+}23, ZGB^{+}21, ZCJ^{+}20, ZZF^{+}24]. Analyze [TPWY23]. Analyzing [GVN25, ZCB23]. Ancilla [BYM22]. Ancilla-Free [BYM22]. Android [RWCC23, XXJ^{+}24]. Annealing [ATT22, HXL^{+}23, LSH25]. ANNs [LFW21]. Anole [HWZ^{+}25]. Anomaly [HLF^{+}23, TKM20, ZTY^{+}23]. Anti [CLCC25]. Anti-Tampering [CLCC25]. Any [YLHL23]. ApGAN [RSA^{+}20]. API [RWCC23]. Application [CTM^{+}25, CLY22, DGG^{+}22, JWG^{+}23, KCAL21, MRA^{+}21, PAR^{+}22, PE22, SKLR22, XZL^{+}23, Xu24]. Application-Level [PE22]. Application-Managed [KCAL21]. Applications [BN24, BGHR^{+}25, CBB^{+}21b, CQ22, CXW^{+}23, CWC^{+}24, DGTVGG21, DPQK^{+}23, DRY^{+}22, FWZ^{+}21, FTR23, FWR^{+}20, GKFF20, JWK^{+}23, JYM20, KH23, LD22, LLL^{+}23, LXL^{+}25, LWZ^{+}25, LSS25, LAPB21, MWJ^{+}24, NML25, PNK^{+}23, PB23b, SMFS21, SPDQ22, SSZ^{+}20, TWZ^{+}23, VKRK22, WS20, YNJS21, YHV^{+}21, YWC^{+}24a, ZGB^{+}21, ZLL^{+}22b]. Applying [ZJW^{+}24]. Approach [BKS22, BFG^{+}21, CDRS20, CTM^{+}25, CCG^{+}22, DKZ^{+}25, FWM^{+}23, GGZC22, GLW^{+}24, GZW^{+}21, HBB^{+}21, JDCL23, JLP^{+}24, JCZ^{+}23, KDE^{+}24, KB21, LYGC24, LLL^{+}25a, LYC^{+}23, MPYJ25, QZZ^{+}24, TPWY23, VCLN21, WYSL22, WZCM23, WCL^{+}23, WLD^{+}25, WWC21, WWL^{+}25, XSYL22, XLW^{+}20, YBF^{+}25, ZZZ^{+}20, ZTY^{+}23, ZNW^{+}24, ZG23, ZDW^{+}23]. Approaches [KKB^{+}22, VHL20]. Approximable [AMM21]. Approximate [AVK20, AZS^{+}23, BN24, BCV22, BSM21, CXW^{+}23, FLS20, FTR23, KK25, MWJ^{+}24, RSMMG^{+}23, RRDB20, RSA^{+}20, SCA^{+}25, SMFS21, TOM23, USS^{+}21, XWP^{+}21, YWC^{+}24b, ZAS^{+}22]. Approximate-Communication-Enabled [XWP^{+}21]. Approximated [CQI^{+}22]. Approximation [CZB^{+}22, HZK24, KvL22, NKA24]. Approximation- [NKA24]. Apps [TQL^{+}22]. AQA [WL24]. AR-Light [WSC^{+}25]. Ara2 [PCA^{+}24]. Arbitrarily [ZCH^{+}24]. Arbitrary [GPQ22, GPQ23, ZCWC23]. Arbitrary-Deadline [GPQ23]. Arbitration [CFC^{+}22]. Architecting [JKHL22, KHHK21, LWW^{+}24]. Architectural [CKK^{+}22, GPH20, TPWY23]. Architecture [AAB^{+}23, AHC^{+}20, BRS^{+}24, CCZ^{+}25, CDW^{+}25, CSK22, DGQ^{+}25, DLY21, FPX^{+}25b, FHL^{+}22, GWH^{+}23, GWG^{+}24, GWG^{+}25, GR23, GL24, HWX^{+}24, HCC^{+}25, HCKK25, HMJ24, IKTY22, JKK^{+}22, JKKP25, JLZ21, JPHY20, JLY^{+}21, JDB^{+}23, JDCL23, KJK^{+}22, KKL^{+}25, KLR23, KIY21, LCY^{+}25, LHW^{+}25, LQM^{+}24, LGZ^{+}25, LLK^{+}23, LZW25, MKY^{+}24, MZM^{+}25, MSSL21, MHJ^{+}21, MHS^{+}20, MC23, MSZ22, NICY24, RPS^{+}21, RGS22, RCAB23, RBC^{+}23, SQR^{+}20, VAV^{+}20, WLW^{+}22a, WDQ^{+}22, WLW^{+}22b, WYZ^{+}22, WWL^{+}23, XLS^{+}24, XJL^{+}25, XLY^{+}25, XNH^{+}25, YFC^{+}22, YGW^{+}23, YLC^{+}21, YWC^{+}24b, YNP^{+}24, ZDZ^{+}23, ZQY^{+}20, ZLS^{+}24, ZLZ24a, ZYZ^{+}25a, ZFD^{+}20, ZLL^{+}22a, ZHM20, ZYQ^{+}24, ZTLW23, ZCSJ23]. Architecture-Level [JDCL23]. Architecture-Mapping [MHJ^{+}21]. Architecture-Neutral [ZHM20]. Architectures [BBD^{+}20, DMG23, DTZ^{+}25, LL21, LAPB21, MRA^{+}21, MÁJG^{+}24, PMA^{+}24, PN24, QWK20, QHZ^{+}21, RPB^{+}23, TWZ^{+}23, TZZ^{+}21, UMM^{+}20, USS^{+}21, WS20, XQC^{+}22, YLL^{+}24, ZLL^{+}23, ZWX^{+}25, ZWC^{+}22, dSBS^{+}22]. Area [KZS^{+}25, RMTA20, USS^{+}21, WWL^{+}25, XLZ^{+}25, ZSHB21, ZQY^{+}20, ZCWC23, ZCP23]. Area-Aware [KZS^{+}25]. Area-Efficient [ZQY^{+}20, ZCWC23]. Area-Latency [XLZ^{+}25]. Area-Optimized [USS^{+}21]. ARETE [TTG^{+}23]. Argument [QCX^{+}23]. ARINC [DPCL22]. Arithmetic [BLM21, God20, LLFT23, MDPM24, VHL20]. ARM [JLD^{+}25, SAJA21, ZYD^{+}20, ZCB23, BPJ^{+}22, DMX^{+}22]. ARM-Based [ZYD^{+}20]. Arnold [LTFL22]. Array [CHL^{+}23, GLW^{+}24, HWX^{+}24, IKTY22, LWNC22, WFH25, ZCWC23, ZLL^{+}22a]. Arrayed [HS22]. Arrays [DPQK^{+}23, JPHY20, KAA20, LRB23, PHC24, WFT^{+}21]. Artifacts [WFH^{+}24]. Artificial [dSdCF22]. Ary [ZLC^{+}23b]. ASHL [STZ^{+}24]. ASIC [MÖS22]. Assessment [DGG^{+}22, TTG^{+}23]. Asset [SRB23]. Assets [DTL^{+}25]. Assignment [HZYY22, KKB^{+}22, LW22, ZYL^{+}22, ZCW^{+}21]. Assistance [KD25, LR22]. Assisted [AY24, BJM^{+}21, CTY^{+}24, FZM^{+}23, GLMZ25, LHW^{+}25, LLZ^{+}25, LYW^{+}23, LWYJ23, NK22, NTR21, NLC^{+}25, PSM22, WLHW25, YZY^{+}25, ZFQ^{+}23, ZXY^{+}24, WDZ^{+}23]. Associated [LLL^{+}23]. Associative [FDKK21, LWW^{+}24, SXH^{+}24]. Assurance [YHC^{+}20]. Asymmetric [IDFH22, LSCX20, LPC^{+}21, VJWZ^{+}21]. Asymmetrically [LZY^{+}25]. AsyncGBP [BZW^{+}25b]. Asynchronous [FZG^{+}22, LMW^{+}24, WHL^{+}21, YPL^{+}25]. AToM [SKH^{+}25]. Atomic [GXY^{+}23]. Atoms [WMG^{+}25]. Attack [AAM^{+}25, APH^{+}23, BMBM20, CLCC25, CPM^{+}23, GLMZ25, LLX^{+}24a, LGC^{+}23, LOM^{+}25, MXY^{+}23, NT23, PM25, RKMR23, RBM21, TZY^{+}24, TDZ^{+}25, WT25, XXZ^{+}25, XXJ^{+}24, ZZZ^{+}23, ZML^{+}24, ZLT^{+}24, ZLW^{+}24, ZWZZ24, ZXL^{+}23, ZG23, JYH^{+}24]. Attacking [RSMMG^{+}23]. Attacks [BY22, CMGD24, CCC23, CCY^{+}24, DMX^{+}22, FTV25, LCHL21, LG22, LHFW24, ODK20, OLZ^{+}20, PBBA25, QZZ^{+}24, RGD^{+}24, SKK^{+}21, SXH^{+}24, TDH^{+}23, WHC^{+}23, WWJ^{+}23, WHQ^{+}24, WZG^{+}24, XTWW25, ZXZ^{+}24, ZLH^{+}21]. Attention [LHXH24, LCM25, LQC^{+}22, WSQ^{+}25, YQX^{+}25, ZCS^{+}25]. Attestation [GJN^{+}25]. Attribute [GLGL23, KCS23]. Attribute-Based [GLGL23]. Audio [LWW^{+}25]. Audio-Visual [LWW^{+}25]. Auditing [SHZ^{+}25, TWaKo^{+}23, ZSC^{+}23, ZQC^{+}25]. Augmented [BLM21, NKN^{+}25, WSC^{+}25, ZLZ24b]. Aurora [LLR25]. Authenticated [LZS^{+}24, LHY^{+}21, LHR^{+}23, XXL^{+}23, XCG^{+}25]. Authentication [LYW^{+}23, PB23a, WHQ^{+}24, ZSS^{+}22]. Auto [BMM^{+}22, DTH^{+}24, HYQ^{+}25, MLW^{+}23, QHT^{+}24, ZWL^{+}25]. Auto-Scaling [HYQ^{+}25]. Auto-Snapshot [DTH^{+}24]. Auto-Tuner [MLW^{+}23, ZWL^{+}25]. Auto-Tuning [BMM^{+}22, QHT^{+}24]. AutoDiagn [DWN^{+}22]. Autoencoder [SZS^{+}22]. Automata [GWX^{+}23, MFRR20, RMR22]. Automated [BRPM22, BCRX23, CHM25, CPM^{+}23, DWN^{+}22, HSH^{+}25, LW22, SCL^{+}24, SSZ^{+}20, WZG^{+}23, WZJ^{+}24, WZG^{+}24, ZWC^{+}22]. Automatic [ABI^{+}25, BGB^{+}21, FLF20, GLZ^{+}24, LD22, WLW^{+}22a, XJY^{+}24, YAG20, ZWSF24]. Automatic-Addressing [WLW^{+}22a]. Automatically [CYKG23]. Automation [SMZ^{+}20]. Automotive [ABI^{+}25, MRA^{+}21]. Autonomous [CZC^{+}21, HXL^{+}25, LCHK22, MSW^{+}21]. AutoPipe [LLL^{+}25a]. AutoPipe-H [LLL^{+}25a]. AUV [ZML^{+}24]. Auxiliary [OTTT22]. Availability [LZW^{+}21]. Avalon [CYX^{+}23]. AVF [TPWY23]. AVL [ZXX^{+}25]. Avoidance [MKH^{+}21, PC24, WLD^{+}22]. Aware [AhRX^{+}20, ALC25, BY22, CWWW20, CSW^{+}21, CZJ21, CXL^{+}25, CSK22, CZW^{+}24, FPHW25, FFG^{+}25, FCZ^{+}23, GQZ21, GKFF20, GLW^{+}24, HHPB20, HYS^{+}20, HGK^{+}22, HIRB25, HF22, HYQ^{+}25, JKNK24, KD25, KLC20, KH23, KAWR23, KZS^{+}25, LDZ^{+}23, LCJ^{+}24, LGL^{+}24, LZZ^{+}22, LSU^{+}23, LLL^{+}25a, LMZ^{+}25, LLK^{+}23, LDT^{+}25, MTV^{+}21, NKA24, PE22, PYYG21, PSBB21, RSP^{+}20, RCS^{+}21, RGvS^{+}24, RvSP^{+}25, RRDB20, RJ24, SNN21, SSK22, SLS^{+}21, SLY^{+}22b, STYQ24, SKM^{+}23, TKN23, WTL^{+}21, WGT^{+}22, WRW^{+}23, WLD^{+}25, WLW^{+}25, WSC^{+}24, XLS^{+}24, XZL^{+}21, XWL^{+}24, XYM23, YWX^{+}23, YCS^{+}24, YDG^{+}24, ZAS^{+}22, ZYXD20, ZZH^{+}25, ZLL^{+}22b, ZYQ^{+}24, ZDW^{+}23, ZZW^{+}25, ZWC^{+}22, ZCS^{+}25, YAG20]. Awareness [WDL^{+}25, ZCZ^{+}22]. AWS [TRV20]. AXI [BOL^{+}25, JYF^{+}23]. AXI-IC [JYF^{+}23]. AXI-Interconnect [JYF^{+}23]. AXI-REALM [BOL^{+}25]. AxMAP [RRDB20].

B [BGHR^{+}25, WBJC22]. BA [LQY^{+}20]. Backdoor [APH^{+}23]. Background [WZW^{+}23]. Backward [SZ22, YWQ^{+}25]. BAFL [FZG^{+}22]. Baker [LT25]. Balance [GZW^{+}21, SZL^{+}24]. Balanced [JRL25, XZL^{+}25]. Balancing [CBB21a, CFWC23, TARK23, TDZ^{+}22, YLL^{+}24, ZMH^{+}25]. Bandits [Gha21]. Bandwidth [BZW^{+}25a, BB22, HWZ^{+}22, ZYL^{+}22]. Bandwidth-Efficient [BZW^{+}25a, ZYL^{+}22]. Bank [KKH22]. BaPa [GZW^{+}21]. Bare [TLC^{+}24, ZXL^{+}24]. Bare-Metal [TLC^{+}24, ZXL^{+}24]. Barrier [LLS^{+}23]. Barriers [PQG^{+}22]. Base [JCKH22]. Based [AT23, AAM^{+}25, AMJ^{+}23, APH^{+}23, ATT22, BLH^{+}21, BSRP21, BCCM22, BTEC20, BMM^{+}22, BZW^{+}25b, BBC^{+}20, BBD^{+}20, BL22, CCT^{+}20, CB22, CSH^{+}24, CJSY24, CYPC25, CDP21, CLY22, CMQ^{+}22, CYX^{+}23, CPL^{+}23, CTZ^{+}24, CXY24, CCZ^{+}25, CFW25, CLCC25, CMGD24, CPB21, CZZ^{+}25, CZW^{+}25, DVV23, DPCL22, DLW^{+}25, DDK22, DSK23, DRP24, DSP^{+}21, DGG^{+}22, DA22, DPS^{+}20, DQ23, DWLF25, DWW25, FPX^{+}25a, FZG^{+}22, FTV25, GGZC22, Gha21, GXZ^{+}23, GLGL23, GXL^{+}24, GL24, HKS20, HS22, HWZ^{+}25, HCKK25, HF23, HHZ^{+}23, HHN^{+}23, HYH^{+}25, HBS20a, HLZ^{+}25, HP23, HWC^{+}22b, HZM^{+}23, HWG^{+}23, HWL^{+}24, xHzLH^{+}24, HWJ^{+}21, HGC^{+}22, HLF^{+}23, HZW^{+}24, HSP^{+}25, Ima21, IWKB22, JWK^{+}23, JKK^{+}22, JKKP25, JYM20, JLZ21, JPHY20, JLZ^{+}23, JDCL23, JQK^{+}24, JKHL22, JJKP22, KKS^{+}22, KPS^{+}24, KCS23, KSB24, KPD^{+}23, KH23, KAWR23, KJK24, KDE^{+}24, KYS^{+}22, KASAG23, LKK^{+}21, LZW^{+}21, LZF21, LCZ22, LWL^{+}22, LWL^{+}23, LLL^{+}23, LDZ^{+}23, LZS^{+}24, LTL^{+}25]. Based [LJH^{+}25, LCM25, LQN^{+}21, LYC22, LMM^{+}23, LFW21, LCH22, LLCJ23, LL23, LQY^{+}20, LHR^{+}23, LGC^{+}23, LJY^{+}24, LZW^{+}24a, LDF^{+}24, LY20, LLY22, LYC^{+}23, LZC^{+}24, MSW^{+}21, MLL^{+}25, MSSL21, MPYJ25, MS25, MIY^{+}20, MC23, MÁJG^{+}24, MDM22, NTDH25, NT23, NICY24, NHW^{+}24, NKN^{+}25, OKC^{+}20, OJ23, PM25, PYS20, PB23a, PYS^{+}23, PNK^{+}23, PMA^{+}24, PKPR23, QCX^{+}23, QWT^{+}23, QZZ^{+}24, RPMH21, RGvS^{+}24, RGS22, RSZ23, ROPdlT22, SCA^{+}25, STW^{+}21, SSW^{+}24, SCL^{+}24, SKK23, SEM23, SXH^{+}24, SM24, SSCK25, SJYQ25, SKM^{+}23, TWZ^{+}23, TZY^{+}24, TDZ^{+}25, TGA23, TKM20, UMM^{+}20, UYZP22, VBA20, VAV^{+}20, WHC20, WLW^{+}21, WGJZ21, WCQW22, WHC^{+}23, WWM^{+}23, WZW^{+}23, WXL^{+}23, WMH^{+}24, WHQ^{+}24, WSLX24, WY25, WWJ^{+}25, WJL^{+}20, WDZ^{+}22, WSHJ23, WZG^{+}24, WWL^{+}25, WLHW25, XLWO23, XL25, XCZ^{+}22, XJL^{+}25, XPR^{+}22, XXL^{+}23, Xu24, XCG^{+}25, XNL^{+}23, XJY^{+}24, YZX^{+}24, YLZ^{+}24, YCS^{+}24, YHW^{+}25, YBG^{+}22, YPL^{+}25, YHH^{+}25, YHV^{+}21, YLHL23, YH24, YWC^{+}24b, YYZ^{+}25, YBF^{+}25]. Based [ZGLZ20, ZTY^{+}23, ZQY^{+}20, ZGWY22, ZCD^{+}22, ZSS^{+}22, ZCWC23, ZGD23, ZLWJ23, ZNW^{+}24, ZXZ^{+}24, ZQC^{+}25, ZOH^{+}25, ZXMX25, ZC24, ZML^{+}25, ZWB^{+}22, ZCY^{+}24, ZHM20, ZCX^{+}20, ZWSF24, ZLC^{+}23a, ZDW^{+}23, ZBT22, ZGG25, ZHYJ21, ddAPdS21, FBM21, HKC21, WHM^{+}22, GQJ^{+}22, ZYD^{+}20]. Bases [ERKP21]. Basic [NP20, XNH^{+}25]. Basis [VJWZ^{+}21]. Bayesian [CWC^{+}24, LHX^{+}25, RSR22]. BB [HSH^{+}25]. BCube [FXC^{+}23, LLCJ23]. BCube-Based [LLCJ23]. BE [BZW^{+}25a]. BE-NPU [BZW^{+}25a]. Beacon [HWL^{+}24]. Beacon-Based [HWL^{+}24]. BEAST [LDT^{+}25]. BEAST-GNN [LDT^{+}25]. Behavior [BFC20, HDL^{+}24, ZTY^{+}23]. Behaviors [ZZC^{+}23]. Being [PBBA25]. Benchmark [HHZ^{+}23, KLP^{+}21, WFH^{+}24]. Benchmarking [BFG^{+}21, DMG23, HWZ^{+}22, HXL^{+}23, RHF24, WZJ^{+}24, XSYL22]. Beneficial [LDW^{+}25]. Benefits [WHL^{+}23]. BERT [KSKK23]. Better [LHXH22, ZYXD20]. Between [NML25, YLL^{+}20]. Beyond [BCKS22, CCG^{+}22, RPS^{+}21, RCC^{+}25]. BFT [DGX^{+}25, DLW^{+}25, GXZ^{+}24, LZW^{+}24b, SLDZ25]. BFT-DSN [GXZ^{+}24]. BFV [SYD^{+}24]. BHerd [LDW^{+}25]. Bi [PHC24]. Bi-Directional [PHC24]. Bias [SKLR22]. Biased [MHM^{+}23]. Biclique [PLH^{+}24, PHL^{+}25]. Bidding [WTL^{+}24]. Big [BKHY22, DWN^{+}22, HWL^{+}21, JKKP25, LPYT22, LLT^{+}23, ZGB^{+}21, MDPM24]. Big-Computing [JKKP25]. Big-PERCIVAL [MDPM24]. big.LITTLE [LL21]. BIKE [RBMG22]. Billion [WMG^{+}25]. Binarized [FHW^{+}22, YBG^{+}22]. Binary [BCKS22, FB20, FRFM^{+}25, KSB24, KJK24, KGHRM23, LCL^{+}20, LLL25b, OLC^{+}22]. Binary-Compatible [OLC^{+}22]. Binary-Unary [KSB24]. Biochips [HGC^{+}22]. Bipartite [PLH^{+}24]. BiRD [PHC24]. Birds [PCA^{+}23]. Bisection [JXH^{+}22]. Bit [BSM21, BYM22, CCYC22, DVV23, FW23, Ima21, JJZW24, KKS^{+}22, KSL^{+}22, LCZ22, LHX^{+}25, LGC^{+}23, LZW^{+}24a, LDT^{+}25, MDPM24, OTTT22, SZL^{+}24, YNJS21, YZG^{+}25, YBG^{+}22, ZZW^{+}25, AJ22]. Bit-Alterable [CCYC22]. Bit-Balance [SZL^{+}24]. Bit-Cell [FW23]. Bit-Flip [LGC^{+}23]. Bit-Level [BSM21, KKS^{+}22, LZW^{+}24a, SZL^{+}24]. Bit-Parallel [LCZ22]. Bit-Serial [Ima21]. Bit-Sliced [DVV23]. Bit-Sparsity [ZZW^{+}25]. Bit-Trojan [JJZW24]. Bit-Width [OTTT22]. Bitcoin [ZLC^{+}23a]. Bitmaps [PHL^{+}25]. Black [SSM23, XXJ^{+}24, ZCY^{+}24]. Black-Box [SSM23, ZCY^{+}24]. BLADE [SQR^{+}20]. Blend [HWZ^{+}25]. Blinded [BMBM20]. Block [BLKK23, CKK^{+}22, Das23, DH20, GLZ^{+}24, HKC21, KvL22, LCH22, NP20, NKL^{+}23, TGS^{+}22, ZZG20, ZFH23, ZWSF24, WLY^{+}23]. Block-Based [LCH22, HKC21]. Block-Wise [Das23]. Blockchain [BHW^{+}23, CZZ^{+}25, FZG^{+}22, GXZ^{+}23, GLGL23, GZG^{+}23, HWL^{+}24, HZYY22, JQK^{+}24, JHMM23, LLT^{+}23, LZS^{+}24, LSW25, LHN^{+}22, LGX^{+}22, LHR^{+}23, LZW^{+}24b, MLL^{+}24, RHF24, SSW^{+}24, TDMP23, TWaKo^{+}23, WPL^{+}23, WLY^{+}23, XLY^{+}22, XZC^{+}23, XZL^{+}25, XNL^{+}23, ZCD^{+}22, ZSC^{+}23, ZFQ^{+}23, ZCH^{+}24, ZXZ^{+}24, ZQC^{+}25]. Blockchain-Aided [GLGL23].

Blockchain-Based

[FZG^{+}22, JQK^{+}24, LZS^{+}24, LHR^{+}23, SSW^{+}24, XNL^{+}23, ZXZ^{+}24, ZQC^{+}25]. Blockchain-Cloud [LHN^{+}22]. Blockchain-Empowered [JHMM23]. Blockchain-Enabled [XZC^{+}23]. Blockchains [LHA^{+}25, LR22, XXL^{+}23, YDW^{+}25]. BlockCompass [RHF24]. BlockExplorer [LLT^{+}23]. Blocking [FL21, JCY^{+}23, MIPQ22, YBW21]. Blocks [FXL^{+}24, LTL^{+}25, XSYL22]. Bloom [Alm23, BL22, KS24, LMM^{+}23, VKRK22, WHY^{+}22]. BlueField [LCS^{+}25]. BlueField-2 [LCS^{+}25]. BlueVisor [JWD^{+}22]. BM [XSYL22]. BM-RCGL [XSYL22]. BMT [ZCY^{+}25]. Bonsai [ZCY^{+}25]. Boolean [TWJ^{+}22, ZCWC23]. Boosting [CZR22, LZW^{+}24a, SZK^{+}22, WDZ^{+}24]. Boot [SKK^{+}21, VTM^{+}20]. Boot-Extended [VTM^{+}20]. Booth [HZK24, XL25]. Bootstrapping [KDE^{+}24]. Borders [ZLT^{+}24]. Borrowable [LW22]. Both [BXW^{+}25, WHL^{+}23]. Bound [CHZ^{+}25, JSTG20, SGS^{+}21]. Boundary [CWS^{+}24a]. Bounded [BSM21, LWC^{+}22]. Bounding [RPB^{+}23]. Box [CLCL22, JBK24, SSM23, TGS^{+}22, WCL^{+}23, XXJ^{+}24, ZCY^{+}24]. Boxes [RMTA20]. Brain [GA22, TGA23, ZCX^{+}23]. Brain-Inspired [GA22, TGA23]. Branch [BMBM20, CY22, MHA^{+}20, ZXD^{+}24]. Branches [SGS^{+}21]. Breaching [MDM22]. Breaking [DLG^{+}24, DGZ^{+}22, WCYK20, ZXL^{+}23]. Bridged [XST20]. Bridging [BZW^{+}25b, GLZ^{+}24]. Bruijn [MYGA20]. BS [LZW^{+}24a]. BSR [ZLL^{+}24a]. BSR-FL [ZLL^{+}24a]. BTI [MSZ22]. BTI-Induced [MSZ22]. Buddy [MIPQ22]. Budget [WTL^{+}24, ZCF20]. Budgeting [LLJ^{+}23, RSP^{+}20]. Buffer [HSH^{+}25, LDLK22, LLY22]. Buffered [SCY21]. Bufferless [XST20]. Bugs [HA25, LB22]. Building [HFT^{+}25, LWNC22]. Built [GXZ^{+}23, MHS^{+}20]. Built-in [MHS^{+}20]. Bundle [LQY^{+}20, WCB23]. Bundled [RACB24]. Burer [SEM23]. Burn [AAB^{+}23]. Burn-In [AAB^{+}23]. Burrows [GR23]. Burst [DT20, HSH^{+}25]. Burstable [HZR^{+}23]. Bus [XTWG23]. Butterfly [LLS^{+}22]. Bypass [PVB21]. Bypassing [LB22, MZZC22, VTM^{+}20]. Byte [CWWW20, ZCX^{+}23]. Byte-Addressable [CWWW20]. Byzantine [ZLL^{+}24a, CERMH23, DGX^{+}25, GXZ^{+}24, JZX^{+}25, LQL^{+}24, TCX^{+}23]. Byzantine-Resilient [TCX^{+}23]. Byzantine-Robust [ZLL^{+}24a].

Cache [BY22, CCC23, DMX^{+}22, FDKK21, Has23, HLLC21, HDL^{+}24, KHP21, KHHK21, LHFW24, LWL^{+}25, LMZ^{+}25, OKC^{+}20, PYS20, PE22, RSP^{+}20, RvSP^{+}25, SCFPM22, SQR^{+}20, SXXL24, SXH^{+}24, SST^{+}24, TSM^{+}21, TBS^{+}25, TRBM22, VLPS25, XKS21, ZZG20, ZWY^{+}23, ZWSF24]. Cache-Aware [RSP^{+}20]. Cache-Clustering [SCFPM22]. Cache-Coherent [TBS^{+}25]. Caches [CFA22, DMX^{+}22, HKC21, RGvS^{+}24, SXH^{+}24, VLPS25, XAP20, XKS21]. Caching [BB22, CHZ^{+}25, DKJP21, KCS23, STYQ24, WDW^{+}23, ZZH^{+}25]. Cacomp [LLZ^{+}25]. CACTI [RvSP^{+}25]. Calculating [God20]. Calculus [GSB23]. CAM [LDG^{+}22, MGFY24]. CAMDNN [HGK^{+}22]. Camouflaged [ZCY^{+}24].

Candidate [PCA^{+}23]. Canonical [PYW^{+}25, ZWM20]. CAP [ZWC^{+}22]. Capability [KKL^{+}25]. Capacitance [BLH^{+}21]. Capacitance-Based [BLH^{+}21]. Capacity [AG24]. CAPE [RGvS^{+}24]. Capsule [ZFZ^{+}21]. Carbon [HIRB25]. Carbon-Aware [HIRB25]. CARM [SYL^{+}23]. Case [Alm23, BY24, FL21, HWG^{+}23, KCAL21, RCC^{+}25, WHQ^{+}24, XPR^{+}22, ZMS^{+}23]. Cases [HXL^{+}23, SCL^{+}24]. CASTLE [LKMJ21]. Cat [LTFL22]. Cauchy [MCT22]. Causal [SKM^{+}23]. Causal-Aware [SKM^{+}23]. Causality [JMW^{+}24, LJY^{+}24]. Cause [TJG^{+}23]. Cause-Effect [TJG^{+}23]. Caused [GSY^{+}20]. CBANA [HDL^{+}24]. CBuild [HFT^{+}25]. CCSDS [BGHR^{+}25]. CCSL [HZM^{+}23]. CDS [WHK24]. Cell [FW23, WFT^{+}21]. Cells [AVK20, SLOM^{+}23]. Cellular [MFRR20]. CEnT [IKTY22]. Center [DWLF25, FXC^{+}23, FPX^{+}25a, LLCJ23, WCQW22]. Centers [CZD^{+}24, FPX^{+}25b, GSC^{+}23, HYQ^{+}25, LDZ^{+}23, SLLS25]. Centric [CNOS22, DGTVGG21, DGZ^{+}22, HWJ^{+}21, LD22, LCY^{+}25, ODK20]. Certification [BSRP21]. CFHider [WZSL22]. CFI [PL21]. Chain [CZZ^{+}25, DTL^{+}25, GQJ^{+}22, GXY^{+}23, HYL^{+}25, LGX^{+}22, SLY22a, XXZ^{+}25, ZCR22]. Chaining [KK25]. Chains [BPM23, TJG^{+}23]. Challenge [SNRB23]. Challenge-Response [SNRB23]. Challenged [LWZ^{+}25]. Chameleon [LSW25]. Change [Gha21, KIY21, LLS^{+}23, LKMJ21, NICY24, SLDZ25, WSC^{+}24]. Change-Detection-Based [Gha21]. Changeable [MLW^{+}25]. Channel [BWSG25, BKS22, BSM21, CLCC25, DYPZ22, GXY^{+}23, HPGM20, HGC^{+}22, JCKH22, KLKK23, KLR^{+}20, LHFW24, LOM^{+}25, MRB^{+}24, OD23, RAD20, SSP^{+}24, SXH^{+}24, UYZP22, WL20, WHC^{+}23, WWJ^{+}23, WTL^{+}24, XPR^{+}22, ZYD^{+}20, ZXL^{+}23, ZLC^{+}23a, ZGG25]. Channel-Level [KLKK23]. Channel-Wise [MRB^{+}24]. Channels [AJ22, FWM^{+}23, GXY^{+}23, SPMP20, WSG^{+}23]. Characteristics [LCS^{+}25, SLLS25, TWY^{+}25]. Characterization [AT23, EEA22, GA22, HDAS21, LPD^{+}21, WRT^{+}22, ZZC^{+}23]. Characterizing [ZTY^{+}25]. Charge [JKKP25]. Check [RSMMG^{+}23]. Checking [DDK22, GZC^{+}21, LWNC22]. Checkpointing [AB22, MPYJ25]. Chip [AT23, AKG^{+}20, BZW^{+}25a, BPJ^{+}22, FQYS23, GvSHA22, HPJK22, HYW^{+}21, HBB^{+}21, HZW^{+}24, JKK^{+}22, KKH22, KRB^{+}22, LY21, MKH^{+}21, SZAT22, TKN23, TSM^{+}21, WNP^{+}22, XNB21, XWP^{+}21, ZLS^{+}24, JPHY20]. Chip-Level [ZLS^{+}24]. Chiplet [DRP24, MS25, WWJ^{+}25, MS25]. Chiplet-Based [DRP24, MS25]. Chiplet-Gym [MS25]. Chiplets [NAP^{+}20]. Chips [ABdGG^{+}25, EGP24, FTR23, XST20]. Cholesky [YGW^{+}23]. Chosen [XPR^{+}22]. CID [KHHK21]. CiM [DSCB25]. CIMAT [JPHY20]. CIMUS [HCC^{+}25]. CINDA [TBS^{+}25]. Cipher [DLG^{+}24, TGS^{+}22]. Ciphers [BKS22, CKK^{+}22, DH20, FBH^{+}22, HPGM20, WT25]. Ciphertexts [XPR^{+}22]. Circ [WBJC22]. Circ-Tree [WBJC22]. Circuit [BBdTF25, GA22, JLY^{+}21, JLP^{+}25, PCA^{+}23, WAN^{+}25, YDG^{+}24, ZSWS24, ZDW^{+}23, ddAPdS21, vSDHA23]. Circuit-Based [ddAPdS21]. Circuits [AMR^{+}20, AZS^{+}23, BYM22, CGLS21, CFWC23, CWC^{+}24, FHW^{+}22, JBK24, LWH20, SXZJ24, WWL^{+}25]. Circular [WBJC22]. CISTs [LLCJ23]. Class [DT20, MYGA20, WMH^{+}24, ZTLW23]. Classic [HWZ^{+}25]. Classical [KGHRM23, KB21, dSdCF22, ddAPdS21]. Classification [AOM^{+}21, LQN^{+}21, RWCC23, WHC^{+}23, XLS^{+}24, ZCX^{+}23, ZWM20]. Classifiers [TZY^{+}24]. Classifying [HA25, LLS^{+}24]. Client [LR22, WLW^{+}21, ZSX^{+}24]. Client-Assistance [LR22]. Clinical [LQN^{+}21]. Clipped [QJY^{+}25]. Cloaking [WHC20]. Closing [CJSY24]. Cloud [AAM^{+}25, CLZG22, CZD^{+}24, CDW^{+}25, CDF^{+}21, DWYX20, FWZ^{+}21, GWZ^{+}21, GLGL23, HHZ^{+}23, IWKB22, JWG^{+}23, LWL^{+}22, LDZ^{+}23, LQN^{+}21, LMM^{+}23, LLZ^{+}25, LHN^{+}22, LYW^{+}23, MSLY24, NWH^{+}25, SHZ^{+}24, SKM^{+}23, TCJ23, WLR20, WLW^{+}21, WCB23, WDW^{+}23, XLY^{+}22, XCG^{+}25, YTD^{+}21, YZY^{+}25, YDL^{+}25, ZGB^{+}21, ZNW^{+}24, ZXL^{+}24, ZXW^{+}24, ZXZ^{+}24, ZQC^{+}25, ZZM^{+}22, ZWSF24, ZZC^{+}23]. Cloud-Assisted [LLZ^{+}25, LYW^{+}23, YZY^{+}25]. Cloud-Based [LQN^{+}21]. Cloud-Edge [CDW^{+}25, YDL^{+}25]. Cloud-Edge-Client [WLW^{+}21]. Cloud/Edge [HHZ^{+}23, SKM^{+}23]. CloudChain [XLY^{+}22]. Clouds [HIRB25, XZL^{+}21, ZDY^{+}23]. ClusPar [DKZ^{+}25]. Cluster [HFT^{+}25, SLS^{+}21, ZCC^{+}23]. Cluster-Aware [SLS^{+}21]. Cluster-Oriented [HFT^{+}25]. Clustered [LWL^{+}25]. Clustering [JXH^{+}22, SCFPM22, WGJZ21]. Clusters [GLW^{+}24, TARK23, WGL^{+}20, ZYZ^{+}23, ZFH23, ZFH^{+}25]. CMOS [SEM23]. CMP [ZWC^{+}22]. CMPs [HMMP23, SCA^{+}25]. CMR [LYC22]. CMR-Based [LYC22]. CNN [CNOS22, JWS^{+}21, KKH22, LL23, LWW^{+}24, SZS^{+}22, TPWY23, WGL^{+}20, WWM^{+}23, WWX^{+}24, WDL^{+}25, ZCR22]. CNNs [BFG^{+}21, BCMT23, DSP^{+}21, GSY^{+}20, LGW^{+}22, WLW^{+}22a, WLW^{+}22b, WL24]. Co [AZS^{+}23, DGQ^{+}25, DRP24, DLK25, GYH^{+}22, GWZ^{+}21, HBS20a, JLY^{+}21, JLZ^{+}23, KHHK21, LHXH24, LCY^{+}25, LMZ^{+}25, PYS^{+}23, TGA23, TLLL25, WZD^{+}20, XLY^{+}25, ZFZ^{+}21, ZLL^{+}22a, ZWM20]. Co-Architecting [KHHK21]. Co-Computing [DLK25]. Co-Design [AZS^{+}23, DGQ^{+}25, GYH^{+}22, GWZ^{+}21, HBS20a, JLZ^{+}23, LHXH24, PYS^{+}23, TGA23, TLLL25, WZD^{+}20, XLY^{+}25, ZFZ^{+}21, ZLL^{+}22a]. Co-Designed [LCY^{+}25]. Co-Designing [ZWM20]. Co-Exploration [JLY^{+}21]. Co-Optimisation [DRP24]. Co-Scheduling [LMZ^{+}25]. COALA [CTY^{+}24]. Coalescing [SAG22]. Coalition [WCB23]. Coarse [DPQK^{+}23, LB22]. Coarse-Grained [DPQK^{+}23, LB22]. CoDA [LHXH24]. Code [CYPC25, CFC^{+}22, CMGD24, DDK22, Das23, FMM^{+}21, HWG^{+}23, HWR^{+}24, LD22, NTR21, VBA20, ZZW^{+}25]. Code-Based [CMGD24, DDK22]. Coded [GSC^{+}23, LRRK^{+}22, SLS^{+}21, ZFH23, ZFH^{+}25]. Codes [DT20, FDKK21, HBS^{+}20b, KAA20, KGHRM23, TZ22, XLL^{+}22, YLL^{+}20, YH24, ZCX^{+}23]. Codesign [SZL^{+}24, TCJ23]. Coding [FLS20, IMQOP21, MCT22, PM20, WGT^{+}22, YLHL23]. COFAE [NK22]. COFFA [DGQ^{+}25]. Cognitive [LLC^{+}24]. Cognizance [JM21]. Coherence [Has23, KHP21, OD23]. Coherent [KRB^{+}22, TBS^{+}25]. Cold [CSH^{+}24, SKK^{+}21]. Collaboration [WLW^{+}21, WZF^{+}24, YDL^{+}25]. Collaborative [ACKA23, AB20, GYZC25, HFT^{+}25, JMW^{+}24, JZY^{+}23, LLZ^{+}25, LGZ^{+}25, PCMP21, SSW^{+}24, SBP^{+}20, WCD25, WSC^{+}25]. Collaborator [ZNW^{+}24]. Collection [FZM^{+}23, HZF^{+}24, LGX^{+}22, ZFH^{+}25]. Collectives [MYUK21]. Collision [LOM^{+}25, OLZ^{+}20]. Collision-Optimized [OLZ^{+}20]. COLM [BMLOM20]. Colocation [CPL^{+}23, ZCC^{+}23]. Colony [XHY^{+}22]. Comb [JCKH22]. Combinational [LLJ^{+}23, SXZJ24, XSYL22]. Combinatorial [ZTT22]. Combined [CZD^{+}24, HP23, RMTA20, WLW^{+}22b]. Combining [CDF^{+}21, Fic22, KKB^{+}22]. Commercial [CCC23, XKS21]. Commodity [FNS^{+}22, LHS^{+}25, LLL^{+}25a, SCFPM22]. Common [IDFH22, ZMS^{+}23]. Communication [ACKA23, ABdGG^{+}25, BSM21, CCZ^{+}25, CZH^{+}25, CZW^{+}24, DWLF25, KRB^{+}22, LHL^{+}21, MZZC22, TJG^{+}23, XWP^{+}21, YBF^{+}25, ZXMX25, ZWC^{+}22]. Communication-Aware [ZWC^{+}22]. Communication-Based [YBF^{+}25]. Communications [BSRP21, Lu21, PCBD23, WTL^{+}21, WWJ^{+}25, XNB21]. Compact [DYJ20, GL24, HLQ^{+}23, MSSL21, ZSS20, ZZW^{+}25]. Compaction [SAG22, SJYQ25]. Comparable [DGX^{+}25]. Comparative [ZGR^{+}25]. Comparing [TJG^{+}23, VBA20]. Comparison [CFA22, DVV23, SO23, WY25, WZJ^{+}24, ZBT22]. Compatibility [LLD^{+}25]. Compatible [OLC^{+}22]. Compensated [GJ20]. Compilation [DYC^{+}25, NM22, WGD^{+}22]. Compilation-Runtime [DYC^{+}25]. Compiler [CTY^{+}24, CDP21, LQM^{+}24, LLZ^{+}25, NTR21, ZWL^{+}25, ZLL^{+}22a]. Compiler-Architecture [ZLL^{+}22a]. Compiler-Assisted [CTY^{+}24, NTR21]. Compilers [MLW^{+}23]. Complete [EAMJ^{+}23, FHL^{+}22]. Completely [DWLF25, WCQW22]. Complex [ZCH^{+}24]. Complexity [BCCM22, CMQ^{+}22, JBK24, KNP^{+}20, NTPAB^{+}25, SSJ21]. Complexity-Effective [NTPAB^{+}25]. Compliant [PCA^{+}24]. Component [HZW^{+}24]. Components [BJMKK23, RSA^{+}20, WHC^{+}23]. Composable [LZS^{+}24]. Composited [MYGA20]. Composition [CGLS21]. Compositional [Has23]. Compositions [CHM25]. ComPreEND [KPL^{+}22]. Comprehension [QLH^{+}24]. Comprehensive [IBB^{+}25, LLK^{+}23, OLD^{+}23, SDR^{+}22, ZG23, ZSX^{+}24]. Compressed [MHT25, WNP^{+}22, BCF^{+}25]. Compressing [YCL^{+}24]. Compression [BGHR^{+}25, GCL^{+}21, HLL^{+}20, JKNK24, KLKK23, LLL^{+}23, LTL^{+}25, LLWZ23, LLX^{+}24b, NK22, UMM^{+}20, WCL^{+}23, ZJW^{+}24, ZXW^{+}24, ZSC^{+}25]. Computation [BLM20, CWNL22, Das23, DWYX20, FHH22, GFB^{+}24, KPL^{+}22, LHL^{+}21, LRL22, LZW23a, LZC^{+}24, RSZ23, TLLL25, WWX^{+}24, XGZ^{+}24, YHW^{+}25, ZWB^{+}22, ZWM20]. Computational [DSJ^{+}22, HTZ^{+}25, KPD^{+}23, LSS25, SIR20, ZLWG22]. Computations [HLJ^{+}25, LLL25b, TDMP23]. Compute [DSCB25, EDGR^{+}24, JPHY20, LY21, SZHB21, TARK23, TRG^{+}24, WSM^{+}24]. Compute-Enabled [TRG^{+}24]. Compute-In-Memory [JPHY20, DSCB25, LY21]. Compute-Programs [WSM^{+}24]. Computer [GPH20, LT25, MHS^{+}20, SMZ^{+}20]. Computers [EGMW21, EAMJ^{+}23, WGD^{+}22, ZFD^{+}20, Ano20a, Ano23]. Computing [ACH21, ACG20, BBJR21, BTEC20, BZW^{+}25b, CCT^{+}20, CSH^{+}24, CGS^{+}20, CVOJRH22, CCCW21, CQ22, CWC^{+}24, CCG^{+}22, DSP^{+}21, DGZ^{+}22, DLK25, FTR23, GQZ21, GA22, GHK^{+}25, GMT24, GZG^{+}23, GYZC25, HR22, HCC^{+}25, HHZ^{+}23, HLZ^{+}25, HJX^{+}25, HLS^{+}23b, HMK^{+}21, HYL^{+}25, IWKB22, JKKP25, JWG^{+}23, JLY^{+}21, JDB^{+}23, JZY^{+}23, JZX^{+}25, KKRK22, KZXR25, KSB24, KAWR23, LCHK22, LZS^{+}24, LGL^{+}24, LLW^{+}24, LHW^{+}25, LFW21, LYGC24, LWW^{+}25, LWW^{+}24, LWH20, LZC^{+}24, LCC^{+}24, MDPM24, MDM22, NLC^{+}25, PLH^{+}24, PCCK22, PYW^{+}25, PLZ20, PPQBA21, RPS^{+}21, ROPdlT22, SCA^{+}25, SMFS21, SIR20, SNT22, SQR^{+}20, SGS^{+}21, TC21, TGA23, TCJ23, WHM^{+}22, WYZ^{+}22, WMH^{+}24, XLT^{+}25, XWL^{+}24, YWC^{+}24a, YWC^{+}24b, ZGLZ20, ZLL^{+}24b, ZJJ25, ZXG^{+}24, ZLM^{+}24, dHBF^{+}21, ZZH^{+}25]. Computing-in-Memory [HCC^{+}25, JLY^{+}21, LWW^{+}24]. Computing-in-Memory-Based [WMH^{+}24]. Concatenation [BBdTF25]. Concentration [RSMMG^{+}23]. Concolic [LM21]. Concurrency [DQ23, GSK^{+}22, LWX^{+}25, YDW^{+}25]. Concurrent [JZH^{+}24, LZZ^{+}22, SKLR22, SPH^{+}23]. Condition [MDM22]. Conditional [CZZ^{+}25, DPS22, SGS^{+}21, ZLH^{+}21]. Confidence [BWSG25]. Confidentiality [LHA^{+}25, WZSL22]. Confidentiality-Preserving [LHA^{+}25]. Configurability [CQI^{+}22]. Configurable [GMT24]. Configuration [CXY24, CCZ^{+}25, LCJ^{+}25, LFP^{+}22]. Configurations [WDCC20]. Configuring [BKHY22]. Conflict [SXH^{+}24]. Conflict-Based [SXH^{+}24]. Conflicting [LLS^{+}25]. Conformance [GZC^{+}21].

Congestion [HWZ^{+}25, SLY22a, TKN23]. Congestion-and-Energy-Aware [TKN23]. Conjugate [YGW^{+}23]. Conjunctive [JQY^{+}25, LJZ^{+}25]. Connected [LLC^{+}24, MYUK21, TKN23]. Connected-Dominating-Set [LLC^{+}24]. Conquer [OLZ^{+}20]. Consensus [CERMH23, HZYY22, JZY^{+}23, JZX^{+}25, LQL^{+}24, LZX^{+}25, XLY^{+}22]. Considering [LCJ^{+}25, SCC21]. Consistency [WLW^{+}22c]. Consistent [LZX^{+}25, VKRK22]. Consolidation [CZD^{+}24]. Consortium [LHN^{+}22]. Constant [JYM20, KAA22, KLL21, ZSS20]. Constant-Time [KAA22, KLL21, ZSS20]. Constrained [AZS^{+}23, BTEC20, GHK^{+}25, GPQ22, KLW^{+}25, LFGD25, LZC^{+}24, OAB^{+}23, SWR^{+}23, WTL^{+}24, ZHLR22]. Constraint [ZGQ^{+}22, ZHY^{+}25]. Constraints [KH23, ST23b, ZZ25]. Constructing [DTL^{+}25, LLCJ23, LLC^{+}24, LSW^{+}23, WCQW22]. Construction [BCKS22, CLCL22, FXC^{+}23, LSH25, XLWO23, YYCR24, YH24, ZCY^{+}25, ZTZ^{+}25]. Constructions [BKS22]. Consumption [OAC^{+}21, TDH^{+}23]. Contact [LHXH22]. Container [CB22, HZF^{+}24, HWR^{+}24, LLW^{+}24, LWX^{+}25, WDZ^{+}24]. Containerized [HWL^{+}21]. Containers [HZR^{+}23, HFT^{+}25]. Containing [BLM20]. Content [HGK^{+}22, KSL^{+}22, LLM^{+}25, Rot24, WDZ^{+}22]. Content-Addressable [KSL^{+}22, Rot24]. Content-Aware [HGK^{+}22]. Contention [BY24, CCZ^{+}22, LMZ^{+}25]. Contention-Aware [LMZ^{+}25]. Context [WDL^{+}25]. Context-Awareness [WDL^{+}25]. Contiguous [HZT^{+}23].

Continual [RSR22, XWL^{+}24].

Continuous [CFWC23, ddAPdS21]. Continuum [MSLY24]. Contour [CSW^{+}21]. Contract [DYC^{+}25]. Contracts [HJYL22, ZCH^{+}24]. Contrastive [CDW^{+}25]. Control [CB22, CDF^{+}21, DDK22, DTZ^{+}25, DPQK^{+}23, DQ23, FWZ^{+}21, GKFF20, HWZ^{+}25, HLS^{+}23b, LHY^{+}21, MKH^{+}21, MRA^{+}21, PL21, RMKO23, SLY22a, WHC20, WZSL22, XWP^{+}21, YPL^{+}25, ZLWG22, ZML^{+}24, ZXZ^{+}24, ZXX^{+}25, ZCF20]. Control-Flow [PL21]. Control-Theoretic [ZCF20]. Controlled [CAC^{+}22]. Controller [LLS^{+}25]. Controlling [ZYXD20]. CONV [ZGWY22, ZCY^{+}24]. Conv-GANs [ZCY^{+}24]. Converters [SPMP20]. Convolution [DA22, JLL^{+}20, LGW^{+}22, PHC24]. Convolutional [ACG20, AGQ^{+}23, DA22, DLY21, DRA21, JKNK24, JCZ^{+}23, JLL22, KNP^{+}20, LCL^{+}20, OKU^{+}24, PPQBA21, RBC^{+}23, WWC21, YFC^{+}22, YWP^{+}23]. Convolutions [YWF^{+}25]. Cooling [SZK^{+}22]. Cooperation [ZDW^{+}23]. Cooperative [LWZ^{+}25, MPYJ25, ZZZ^{+}20, ZLT^{+}25]. Coordinated [GSC^{+}23, XXC^{+}25, ZDZ^{+}23]. Coordination [YCKW20]. Coordinative [LHL^{+}21]. COP [LLJ^{+}23]. COPA [BMLOM20]. Cope [BLP^{+}22]. Copy [HS22]. Copying [LWL^{+}25]. CORDIC [CQCL25, MC23]. CORDIC-Based [MC23]. Core [AhRX^{+}20, ABI^{+}25, CKK^{+}22, GLB21, GCR^{+}23, Has23, JWD^{+}22, JDCL23, JYM^{+}23, KHP21, KKKC20, LZZ^{+}22, LQM^{+}24, LAPB21, MIPQ22, MB21, MBP21, MDJ20, PYS20, PCBD23, PCA^{+}24, SNN21, SRP^{+}21, SKLR22, WWJ^{+}25, WWS^{+}22, WSG^{+}23, YGW^{+}23, YWP^{+}23, ZZL21, ZYL^{+}22, ZWX^{+}25, Lu21, WWJ^{+}23]. Cores [HIRB25, HLT^{+}23, LHK^{+}22, LZW^{+}23b, LHZ^{+}24, PFHD21, RSP^{+}20, RPMH21, SZHB21, WMG^{+}25, ZCF20]. Corner [HJX^{+}25]. Correct [YYCR24]. Correct-by-Construction [YYCR24]. Correcting [DT20]. Correction [FDKK21, FMM^{+}21, KGHRM23]. Corrections [WZG^{+}23]. Correctly [God20]. Corrector [HYW^{+}21]. Correlation [CMGD24, CWNL22]. Corruptions [PG23]. Cortex [LLS^{+}22, SAJA21]. Cortex-M4 [SAJA21]. Cosine [JLL22]. COSMO [BCF^{+}25]. CoSpMV [TLLL25]. Cost [AAB^{+}23, BGB^{+}21, DLG^{+}24, DWYX20, FLS20, IBB^{+}25, JKHL22, LLW^{+}24, LZW^{+}24b, PYW^{+}22, SZK^{+}22, SMZ^{+}20, SLY22a, WDW^{+}23, XXZ^{+}25, ZCR22]. Cost-Driven [DWYX20, WDW^{+}23]. Cost-Effective [LZW^{+}24b, XXZ^{+}25]. Cost-Effectiveness [SLY22a]. Cotransformation [APK20]. Count [RMO21]. COUNTDOWN [CBB^{+}21b]. Counter [BHK^{+}23]. Countermeasure [NT23, ZYD^{+}20]. Countermeasures [GLMZ25, RGD^{+}24, ZXL^{+}23]. Counters [GWCS23, LG22, SSM23]. Counting [WYZ^{+}22, ZWJ^{+}25]. Coupled [WHK24, YGW^{+}23]. Cover [CLCC25]. Coverage [WZCM23]. Covering [TOF^{+}24]. Covers [CLCC25]. Covert [WHC^{+}23, WWJ^{+}23, ZLC^{+}23a]. CP [CLZG22, ZOH^{+}25]. CP-ABE [CLZG22, ZOH^{+}25]. CPA [TDH^{+}23]. CPI [WCD25]. CPU [CKK^{+}22, HDAS21, HZR^{+}23, KMH^{+}23, LPW20, LL21, NML25, WCYK20, ZCC^{+}23]. CPUs [FRFM^{+}25, JDCL23, KASAG23, XNH^{+}25, YWF^{+}25, ZYL^{+}22]. Crane [GSY^{+}20]. CRC [NWL^{+}25]. Credit [EGMW21]. CRIME [PCMP21]. Critical [CSvdBU22, HLL^{+}20, JLZ^{+}23, SXZJ24, XSYL22, ZTY^{+}23]. Criticality [BOL^{+}25, JDB^{+}23, LL22, MNB20, MBP21, RGvS^{+}24, VSG^{+}23, ZZ25]. Criticality-Aware [RGvS^{+}24]. CROSC [DYC^{+}25]. Cross [DTL^{+}25, GXY^{+}23, VAV^{+}20, WFT^{+}21, WLY^{+}23, WTL^{+}24, XXZ^{+}25, YDW^{+}25, YZM^{+}25, ZCH^{+}24]. Cross-Architecture [VAV^{+}20]. Cross-Chain [DTL^{+}25, GXY^{+}23, XXZ^{+}25]. Cross-Channel [GXY^{+}23, WTL^{+}24]. Cross-Layer [WLY^{+}23]. Cross-Modal [YZM^{+}25]. Cross-Point [WFT^{+}21]. Cross-Shard [ZCH^{+}24]. Cross-Sharding [YDW^{+}25]. Crossbar [BTEC20, WWM^{+}23]. Crossbar-Constrained [BTEC20]. Crossbar-Level [WWM^{+}23]. Crossbars [HJX^{+}25]. Crossing [PAR^{+}22]. Crosstalk [IDFH22]. Crowdsensing [FHL^{+}23, FZM^{+}23, ZFQ^{+}23]. CRT [LCZ22]. CRT-Based [LCZ22]. Cryo [RvSP^{+}25]. Cryo-CACTI [RvSP^{+}25]. Cryogenic [RvSP^{+}25]. Cryogenic-Aware [RvSP^{+}25]. Cryptographic [HWG^{+}23, NCD^{+}25, PNK^{+}23, TDH^{+}23]. Cryptography [DVV23, HP23, JDH^{+}25, KAA22, KGHRM23, MDJ20, TWZ^{+}23]. Cryptoprocessor [AMJ^{+}23]. Cryptosystems [BRPM22, GMZ22, RSZ23, XPR^{+}22, ZBT22]. CryptSQLite [WSS^{+}20]. Crystals [CZW^{+}25, DMG23, GL24, JDH^{+}25, NKeSK^{+}23]. Crystals-Dilithium [CZW^{+}25]. CRYSTALS-Kyber [DMG23, GL24, NKeSK^{+}23]. CSD [ZZW^{+}25]. Cube [DPS22]. Cubes [ZLC^{+}23b]. CUDA [SYL^{+}23]. CUDA-Accelerated [SYL^{+}23]. CURE [KKL^{+}25]. CurIAs [PB23a]. Curiosity [HZM^{+}23]. Curiosity-Driven [HZM^{+}23]. Current [PB23a]. Current-Based [PB23a]. Curve [MDJ20]. CUSPX [WDC^{+}25]. Custom [CWS^{+}24b, PN25, SEM23, WGM^{+}20, WGJZ21, WWX^{+}24, YYCR24]. Customizable [WCZ^{+}24]. Customization [BCRX23, FQYS23]. Customized [CDW^{+}25, YHH^{+}25]. Cuts [Akr22]. Cyber [FV23, JLZ^{+}23, LYF^{+}22, PKPR23, TOF^{+}24]. Cyber-Physical [FV23, LYF^{+}22]. Cycle [FW23]. Cyclebite [WSM^{+}24]. Cyclic [Das23].

D [BCV22, LSCX20, FMM^{+}21, FZM^{+}23, LW22, LQY^{+}20, NAP^{+}20, PC24, SM22, WLQ^{+}21, WZD^{+}20]. D-Flip-Flop [LW22]. D-Point [LQY^{+}20]. D-Reducibility [BCV22]. DAG [DLW^{+}25, GQH21, GPQ22, GPQ23, GXZ^{+}23, HECC^{+}21, JSTG20, LSU^{+}23, SGL^{+}20, STQ^{+}24, UGvdBC23, WZGT22]. DAG-Based [DLW^{+}25, GXZ^{+}23]. DAG-Fluid [GQH21]. DAGs [GQH21]. DAHE [ZYZ^{+}25b]. Dandelion [CZR22]. Dark [LLJ^{+}23, WLQ^{+}21, WLW^{+}22c, WWS^{+}22]. Data [AB20, BJM^{+}21, BHE21, BKHY22, CCYC22, CLY22, CZD^{+}24, CXY24, CSK22, CKJ^{+}22, CPB21, DWN^{+}22, DWL^{+}22, DQ23, DWLF25, DLK25, FLS20, FXC^{+}23, FPX^{+}25b, FPX^{+}25a, FHL^{+}22, FZM^{+}23, GSK^{+}22, GSC^{+}23, GMT24, GLW^{+}24, GZG^{+}23, HKC21, HHPB20, HLQ^{+}23, HC24, HYQ^{+}25, HMK^{+}21, HWL^{+}21, HJYL22, HWR^{+}24, JM21, JDCL23, JZH^{+}24, JJKP22, KMVD22, KXGS22, LPYT22, LHL^{+}21, LLT^{+}23, LDZ^{+}23, LZS^{+}24, LWH^{+}24, LLS^{+}24, LHA^{+}25, LCY^{+}25, LTL^{+}25, LWL^{+}25, LWC^{+}22, LMM^{+}23, LCH22, LLCJ23, LHY^{+}21, LHN^{+}22, LGX^{+}22, LYW^{+}23, LV23, LJY^{+}24, LLL^{+}25a, LZX^{+}25, LWZZ25, MHDMEA22, MCT22, MLW^{+}25, NTR21, NNH^{+}25, NML25, PG23, PCCK22, PB23b, ROPdlT22, SSY^{+}21, SWR^{+}23, SLLS25, SXC^{+}24, SJYQ25, SKM^{+}23, TBS^{+}25, TWL^{+}22, WSS^{+}20, WHC20, WCQW22, WLZ^{+}21, WDW^{+}23, WHK24, WSC^{+}24, WSG^{+}25, XNB21, XLL^{+}22, XWL^{+}24, XNL^{+}23, YWC^{+}21, YZM^{+}25, YYQ^{+}24, YNP^{+}24, ZGB^{+}21, ZGQ^{+}22, ZZZ^{+}23, ZGG^{+}23, ZJW^{+}24, ZOH^{+}25, ZC24, ZCZW23, ZXZ^{+}21, ddAPdS21]. Data-Centric [LCY^{+}25]. Data-Driven [JDCL23]. Data-Flow [FHL^{+}22]. Data-Independent [ZGG^{+}23]. Data-Parallel [GSK^{+}22, LHL^{+}21, ROPdlT22]. Data-Paralleled [LLL^{+}25a]. Data-Reuse [HLQ^{+}23]. Data-Type [CSK22]. Data-Types [JM21]. Database [CKRP21, JWS^{+}23, MPYJ25, WPL^{+}23]. Databases [TBS^{+}25, XGMJ25]. Datacenters [LWL^{+}24, SCY^{+}23]. Dataflow [CHL^{+}23, GWG^{+}24, GWG^{+}25, HBS20a, KXGS22, PHC24, SMC^{+}25, WFH25, ZGWY22]. Dataflow-Based [HBS20a]. Dataflows [HWX^{+}24, OKU^{+}24]. DataFly [LHA^{+}25]. Datapath [HMJ24, UMM^{+}20]. Dataset [CZR22]. Datasets [ZJW^{+}24]. DBMS [AY24]. DBMS-Assisted [AY24]. DC [JLP^{+}25, LTL^{+}25]. DC-ORAM [LTL^{+}25]. DCAS [ZCY^{+}25]. DCAS-BMT [ZCY^{+}25]. DCFNoC [PFHD21]. DCGG [XXC^{+}25]. DCNNs [LMH^{+}25]. DDoS [FTV25]. DDR [BB20, LLS^{+}25]. Deadline [GPQ23, YWX^{+}23]. Deadline-Aware [YWX^{+}23]. Deadlines [GPQ22]. Deadlock [MKH^{+}21, PC24, WLD^{+}22].

Debugging [AB22]. Decentralized [CWY^{+}23, CZC^{+}21, CCY^{+}24, DLZ^{+}24, GXZ^{+}23, GXZ^{+}24, LZY^{+}25, LXL^{+}25, LSW25, LYGC24, LHR^{+}23, SHZ^{+}25, WLY^{+}23, WJLC24, XZC^{+}23, ZOH^{+}25, ZWB^{+}22]. Decision [HKC^{+}23, LQN^{+}21, NHW^{+}24, XJL^{+}25]. Decision-Tree-Based [XJL^{+}25]. Decoder [HZK24, HSP^{+}25]. Decoders [VBA20]. Decoding [DT20, LSCX20, TZ22]. Decomposition [GYH^{+}22, HKC^{+}23, JWK^{+}23, JLZ^{+}23, KKB^{+}22, PYW^{+}25, WDQ^{+}22]. Decomposition-Based [JLZ^{+}23]. Decompression [KHHK21]. Decoupled [LCX21]. Decoupling [CDW^{+}25, HKC21]. Decryption [CLZG22]. DECT [DLG^{+}24]. DedupHR [WDZ^{+}22]. Deduplication [CPB21, DZC^{+}24, LHR^{+}22, LJY^{+}24, SHZ^{+}24, WCD25, WDZ^{+}22, WDZ^{+}23, YLG^{+}23, YLL^{+}24, YTD^{+}21, ZSC^{+}23, ZQC^{+}25]. Deduplication-Based [WDZ^{+}22]. Deep [AAM^{+}25, AC22, BWL^{+}25, CWS^{+}24a, CLCC25, CXL^{+}23, CSK22, DGG^{+}22, DRA21, DGZ^{+}22, GLMZ25, GWCS23, GYS^{+}24, HLL^{+}20, HHPB20, HCC^{+}23, HLZ^{+}25, HGK^{+}22, HLC^{+}22, JM21, JJZW24, KKS^{+}22, KPL^{+}22, KH23, KBQ^{+}23, KYS^{+}22, KBR^{+}23, LJH^{+}25, LPC^{+}21, LLZ^{+}25, LZG^{+}24, LFX^{+}21, LY20, LLY22, LZC^{+}24, LFP^{+}22, LCC^{+}24, MSP^{+}21, MRB^{+}24, MHT25, MLW^{+}25, PBBA25, PCCK22, PN24, RRMS25, SPB^{+}21, STZ^{+}24, SM24, STQ^{+}24, SKM^{+}23, TZY^{+}24, WHM^{+}22, WZX^{+}22, WTL^{+}24, WSLX24, WWC21, WLHW25, XLW^{+}20, YWX^{+}23, YPL^{+}25, YBF^{+}25, ZGLZ20, ZCK20, ZNW^{+}24, ZCC^{+}23, ZWL^{+}25, ZWSF24, ZHYJ21, ZWC^{+}22]. Deep-Learning-as-a-Service [LCC^{+}24]. DeepFire2 [AGQ^{+}23]. Deeply [WGL^{+}20]. Deeply-Pipelined [WGL^{+}20]. DeepMD [DWW25]. DeepMD-Kit [DWW25]. DeepP [LFP^{+}22]. DeepWare [GWCS23]. Defects [WRT^{+}22]. Defender [LHFW24]. Defense [LCC^{+}24, MXY^{+}23, SKK^{+}21]. Defined [FTV25, GVN25, KNP^{+}20, YLC^{+}21]. Deflection [XST20]. Defying [MZM^{+}25]. Degenerate [BJMKK23]. Degree [LPC^{+}21]. Delay [GHK^{+}25, KK25, LW22, SKR^{+}20, SMY22, WLD^{+}25, WZG^{+}24]. Delay-Constrained [GHK^{+}25]. Delayed [TARK23]. Deletable [WHY^{+}22]. Deletion [LCH22]. Delivery [JZX^{+}25, LLM^{+}25, MHK^{+}22]. Delta [CPB21, NTPAB^{+}25, ZJW^{+}24]. Delta-Encoding [CPB21]. Demand [BZW^{+}25a, MHA^{+}20]. Demands [GYS^{+}24]. Denial [BY22, NT23]. Denial-of-Service [BY22, NT23]. Denoising [WHQ^{+}24]. Density [LWX^{+}25]. Dependability [ZGR^{+}25]. Dependence [ZLH^{+}21]. Dependencies [HZW^{+}24, XZL^{+}23]. Dependency [ZC24, ZZH^{+}25]. Dependency-Aware [ZZH^{+}25]. Dependent [BSM21, PCMP21, WHM^{+}22, YBG^{+}22, ZMS^{+}23]. Deploying [MHS^{+}20]. Deployment [BGB^{+}21, GLS^{+}25, JBS^{+}25, LCJ^{+}24, LWX^{+}25, WLD^{+}25, WFZ^{+}25, ZFL^{+}22, ZZM^{+}22]. Depth [LZF21]. Depth-Limited [LZF21]. Depthwise [PHC24]. Derivation [TWJ^{+}22]. DESA [WFH25]. Descent [LSL^{+}25]. DESCO [JLZ^{+}23]. Design [ABC^{+}24, ABP25, AZS^{+}23, BK23, DGQ^{+}25, DA22, DZC^{+}24, DTZ^{+}25, FHW^{+}22, GMZ22, GWG^{+}25, GYH^{+}22, GSC^{+}23, GWZ^{+}21, HBS20a, HGC^{+}22, HSP^{+}25, JLZ^{+}23, LLFT23, LHXH24, LWH20, MHJ^{+}21, MKÖ^{+}22, MS25, NCD^{+}25, PYS^{+}23, PSBB21, QHZ^{+}21, RPS^{+}21, RSZ23, RMKO23, SRP^{+}21, SMZ^{+}20, SLS^{+}21, SYD^{+}24, SMC^{+}25, TGA23, TLLL25, TOM23, VHL20, WBJC22, WHL^{+}23, WNL^{+}23, WCD25, WFL^{+}25, WZD^{+}20, XNB21, XLY^{+}25, YHC^{+}20, YYCR24, YLC^{+}21, ZAS^{+}22, ZYD^{+}20, ZFZ^{+}21, ZCP22, ZXD^{+}24, ZWL^{+}25, ZLL^{+}22a]. Designed [LCY^{+}25, ZLW^{+}24, ZWZZ24]. Designing [KHP21, ZWM20]. Designs [BJMKK23, DSTD22, FWM^{+}23, HA25, LW22, RMTA20, SRB23, SDR^{+}22]. Detailed [WFH^{+}24]. Detect [GWCS23]. Detecting [BLH^{+}21, FWM^{+}23, LG22, SZS^{+}22]. Detection [AAM^{+}25, Akr22, BJM^{+}21, BXW^{+}25, BBC^{+}22, DMD^{+}23, EEA22, FDKK21, FTV25, Gha21, HHN^{+}23, HYH^{+}25, HJX^{+}25, HLF^{+}23, KPL^{+}22, LRB23, LHXH22, MLL^{+}25, MSP^{+}21, MGFY24, OLZ^{+}20, PSM22, PM25, PK23, Rot24, SNA^{+}20, SCY^{+}23, SKA^{+}22, VAV^{+}20, WHC^{+}23, WMH^{+}24, XCZ^{+}22, YWP^{+}23, ZGB^{+}21, ZTY^{+}23]. Detection/Correction [FDKK21]. Detector [HYW^{+}21, TKM20]. Detectors [Fic22, IKAG^{+}22, ZCY^{+}24]. Deterministic [CTZ^{+}24]. Deviation [BSM21]. Device [BCRX23, CYPC25, JKK^{+}22, JLD^{+}25, JLY^{+}21, TKM20, ZHL^{+}24]. Device-Circuit-Architecture [JLY^{+}21]. Deviceless [HYL^{+}25]. Devices [CWS^{+}24b, DYPZ22, GZC^{+}21, HYL^{+}25, KLC20, KLW^{+}25, LFGD25, LPD^{+}21, PPQBA21, RBMG22, SQR^{+}20, SWR^{+}23, TKM20, YZJ23, YWC^{+}21, YH20, ZLC^{+}22, ZLS^{+}24, ZYQ^{+}24, ZTY^{+}25]. DHTS [LZW23a]. Diagnose [CDRS20]. Diagnosis [CFW25, DWN^{+}22]. Diagonal [DLY21]. Different [GYS^{+}24].

Differentiable [LLK^{+}23]. Differential [AHK^{+}21, FYR^{+}24, LXL^{+}25, RKMR23, RBM21, WT25, ZWB^{+}22, ZCS^{+}25]. Differentially [LWZZ25]. Differentiated [CZJ21]. Digit [ERKP21, HSE^{+}24, LMDC21, LLFT23, LSW^{+}23]. Digit-First [LLFT23]. Digit-Serial [ERKP21]. Digital [CMQ^{+}22, KJK^{+}22, LGL^{+}24, LCC^{+}24, ZXY^{+}24, ZCF20]. Dilithium [CZW^{+}25]. Dimension [MHDMEA22]. Dimensional [SZS^{+}22]. Diminishing [YHV^{+}21]. DIMM [LCM25, MIY^{+}20]. DIMM-Based [LCM25]. DIPER [MGFY24]. Direct [CCZ^{+}22, WLY^{+}23, YWF^{+}25, LWL^{+}24]. Directed [HYS^{+}20, MBP21, ZCX^{+}20]. Directional [PHC24]. Directly [CKRP21]. Disaggregated [CL20, LWL^{+}24]. Disaggregation [HMZ^{+}25]. Disassembly [KLR^{+}20]. Disaster [LHW^{+}25]. Disclosed [LBC24]. DISCO [Has23]. Discovery [SSY^{+}21]. Discrete [GPH20, JLL22, KLL21, LPW20, LTFL22, ZSS20]. Discriminator [RSR22]. Disjoint [FXC^{+}23, FPX^{+}25a]. Disk [CKRP21, HLS^{+}23a, LPC^{+}21]. Disks [HKS20]. Disparity [YCS^{+}24]. Disruptions [HXGR20]. Dissemination [LHY^{+}21, LYW^{+}23]. Distance [MGFY24]. Distance-Tolerant [MGFY24]. Distillation [YPD^{+}24]. Distributed [ACKA23, Akr22, CZJ21, CYX^{+}23, DNMS20, DWL^{+}22, DRA21, GGZC22, GKT^{+}22, GLS^{+}25, HHZ^{+}23, HGC^{+}22, HZMC24, JZX^{+}25, LCJ^{+}24, LLX^{+}24a, LV23, LWZZ25, NWH^{+}25, STZ^{+}24, SSW^{+}24, TDZ^{+}22, WZH^{+}23, WJLC24, XLT^{+}25, XGMJ25, XQC^{+}22, XLL^{+}22, YPL^{+}25, YCY^{+}24, ZLL^{+}23]. Distribution [CZW^{+}24, JMW^{+}24, LDZ^{+}23, LQY^{+}20, ZCZW23]. Disturbance [BGM^{+}23, CFA22, IKTY22, LLS^{+}23, WSC^{+}24]. Divergence [DLK25]. Diverse [GLZ^{+}24, HZF^{+}24, XLS^{+}24]. Divide [KPH^{+}25, OLZ^{+}20]. Divide-and-Conquer [OLZ^{+}20]. Divide/Remainder [KPH^{+}25]. Divide/Sqrt [KPH^{+}25]. Dividers [AVK20]. Division [ABC^{+}24, Bru20, Bru23, HSE^{+}24, LSW^{+}23]. DL [ZYL^{+}22]. DLaaS [CCY^{+}24, xHzLH^{+}24, LLX^{+}24a, XXJ^{+}24, ZML^{+}24]. DLPU [DGZ^{+}22]. DLPU-Centric [DGZ^{+}22]. DMA [BRS^{+}24, PCBD23]. DMACN [KCS23]. DML [DRY^{+}22]. DMRlib [IMQOP21]. DNA [HSP^{+}25, LYW^{+}25]. DNN [CWT^{+}22, CZR22, CHZ^{+}25, CXH^{+}25, DRP24, FRFM^{+}25, GLZ^{+}24, GCL^{+}21, LLZ^{+}25, LGC^{+}23, LTJS^{+}22, LWYJ23, MHJ^{+}21, NTL^{+}24, NKL^{+}23, SPH^{+}23, SKK23, XQC^{+}22, YLC^{+}21, ZDZ^{+}23, ZAS^{+}22, ZML^{+}25]. DNNs [BGB^{+}21, GHK^{+}25, GXL^{+}24, HLQ^{+}23, JKHL22, SMC^{+}25, WZG^{+}23, ZWX^{+}25]. DO [ZHYJ21]. Document [JWS^{+}23]. Domain [AKG^{+}20, JKKP25, MKY^{+}24, MHK^{+}22, OLD^{+}23, WS20]. Domain-Specific [AKG^{+}20, WS20]. Domain-Wall [OLD^{+}23]. Domains [ZHL^{+}24]. Dominance [LSS25]. Dominating [LLC^{+}24]. DORY [BGB^{+}21]. Double [God20, ZFH23]. Double-Precision [God20]. Down [RvSP^{+}25, WCYK20]. DownShift [KOH^{+}23]. DPA [HP23]. DPU [HTZ^{+}25, LCS^{+}25, LWL^{+}24]. DPU-Direct [LWL^{+}24]. DQN [ZLL^{+}22b]. DRAM [AG24, BB22, FPHW25, HKC21, KCL^{+}20, LCJ^{+}25, LYH^{+}24, MTV^{+}21, OAK^{+}23, SKK^{+}21]. DRAM-Like [AG24]. DRAM-PM [FPHW25]. DRF [SCC21]. Drift [NICY24]. Driven [CXY24, DWYX20, GPH20, GZC^{+}21, HIRB25, HZM^{+}23, JDCL23, JXH^{+}22, LWYJ23, QSC^{+}25, WDW^{+}23, YTD^{+}21, ZLWG22, ZLM^{+}24]. Drives [BMM^{+}22, HS22, LPC^{+}21, YCKW20, ZYXD20]. Driving [CZC^{+}21, KD25]. DRL [GQJ^{+}22]. DRL-Based [GQJ^{+}22]. DS3 [AKG^{+}20]. DSN [GXZ^{+}24]. DSP [KvL22]. DSTC [ZWX^{+}25]. Dual [ERKP21, HLS^{+}23b, LLY22, PCA^{+}23, VLPS25, WRW^{+}23, WWL^{+}25, ZSHB21, ZWX^{+}25]. Dual-Issue [ZSHB21]. Dual-Mode [LLY22, PCA^{+}23]. Dual-Privacy [WRW^{+}23]. Dual-Side [ZWX^{+}25]. Duo [CCYC22]. Duo-Phase [CCYC22]. Durable [PB23b]. DVFS [LJH^{+}25, WLD^{+}25, YL20, YHV^{+}21]. DVFS-Based [YHV^{+}21]. DVREI [LMM^{+}22]. DWC [dSBS^{+}22]. Dynamic [BBD^{+}20, CBB21a, CWY^{+}23, DRY^{+}22, GLMZ25, GPRV23, HJYL22, HTZ^{+}25, HXGR20, KCS23, KLW^{+}25, LMM^{+}22, LQL^{+}24, LXL^{+}25, LTL^{+}25, LQC^{+}22, LZW23a, LYW^{+}25, LYC^{+}23, MB21, MHA^{+}20, MLW^{+}23, ODK20, PYW^{+}22, RSP^{+}20, SCL^{+}24, TTG^{+}23, WLW^{+}21, WHX^{+}25, XZL^{+}25, YLT^{+}23, ZQG^{+}24, ZZL21, ZZF^{+}24, ZCY^{+}25, ZZH^{+}25, ZLT^{+}25, ZLZ^{+}23]. Dynamic-Timing [TTG^{+}23]. Dynamically [CFWC23, CAC^{+}22, HLQ^{+}23, XXC^{+}25]. Dynamics [AGSD25, HLK^{+}25, WMG^{+}25, WYB^{+}24]. DyNNamic [HLQ^{+}23].

E2CNNs [PPQBA21]. EaD [WDZ^{+}23]. Early [BPJ^{+}22, CR24, KPS^{+}24, KPL^{+}22, KPH^{+}25, NLC^{+}25, PBBA25]. Easy [IMQOP21]. Easy-Coding [IMQOP21]. EC2P [XXZ^{+}25]. ECC [FDKK21, LCJ^{+}25, LSXZ21, PD21, WDZ^{+}23]. ECC-Assisted [WDZ^{+}23]. ECC-United [FDKK21]. ECDR [HYW^{+}21]. ECDSA [JCKH22]. ECO [JDH^{+}25]. ECO-CRYSTALS [JDH^{+}25]. EcoFlow [OKU^{+}24]. Economy [YLT^{+}23]. Economy-Oriented [YLT^{+}23]. EDA [SSP^{+}24]. EDF [CBB21a, CKP^{+}22, JSTG20, WLZ^{+}25]. EDF-Like [WLZ^{+}25]. Edge [ACG20, BBJR21, CWY^{+}23, CDW^{+}25, CCY^{+}24, DKZ^{+}25, DWYX20, GHK^{+}25, GYZC25, HHZ^{+}23, HLZ^{+}25, HGK^{+}22, HLS^{+}23b, HZYY22, HYL^{+}25, JWG^{+}23, JHMM23, JZY^{+}23, KLW^{+}25, LD22, LZS^{+}24, LGL^{+}24, LLW^{+}24, LHW^{+}25, LFW21, LYGC24, LLZ^{+}25, LZX^{+}25, LZC^{+}24, LWYJ23, MSLY24, NHW^{+}24, NLC^{+}25, PPQBA21, QWT^{+}23, RPS^{+}21, RCC^{+}25, RBC^{+}23, ROPdlT22, SQR^{+}20, STQ^{+}24, SKM^{+}23, TCX^{+}23, TKM20, WLW^{+}21, WHM^{+}22, WWX^{+}24, WLD^{+}25, XZL^{+}21, XWL^{+}24, YYQ^{+}24, YWC^{+}24a, YYZ^{+}25, YCY^{+}24, YDL^{+}25, ZQG^{+}24, ZGLZ20, ZFL^{+}22, ZXY^{+}24, ZHL^{+}24, ZNW^{+}24, ZZF^{+}24, ZXG^{+}24, ZLM^{+}24, ZYQ^{+}24, ZTY^{+}25, ZHYJ21, ZSX^{+}24, ZXG^{+}24, ZZH^{+}25]. Edge-Assisted [LWYJ23]. Edge-Based [QWT^{+}23]. Edge-Centric [LD22]. Edge-Cloud [DWYX20, JWG^{+}23, MSLY24, ZNW^{+}24]. Edge-Computing [PPQBA21]. Edge-Enabled [YYQ^{+}24]. Edge-Intelligent [WLW^{+}21]. Edge-MPQ [ZXG^{+}24]. Edge-Side [HLS^{+}23b, YWC^{+}24a]. Edges [ZLC^{+}23b]. Edit [MGFY24]. Editorial [Ano23, BBJR21, CDP21, CQ22, FAKM22, Kar24, Lu21, WS20]. Editors [QWK20, AW20]. EDoS [ZXZ^{+}24]. Effect [LLS^{+}22, TJG^{+}23]. Effective [AAB^{+}23, EEA22, JZSD24, JDCL23, KJC^{+}21, LZW^{+}24b, LCGH25, NTPAB^{+}25, RRMS25, SPDQ22, WFT^{+}21, XXZ^{+}25]. Effectiveness [SLY22a]. Effects [BPJ^{+}22, DQ23, LSL^{+}25, MSZ22]. Efficiency [BSRP21, BB20, DWW25, FDKK21, FHH22, HXL^{+}25, KIY21, LPD^{+}21, LZW^{+}24a, PYW^{+}25, SZK^{+}22, SKLR22, WXL^{+}23, WDZ^{+}24, YWX^{+}23, YL20, ZCJ^{+}20]. Efficiency-Oriented [YWX^{+}23]. Efficient [AA20, AhRX^{+}20, AHC^{+}20, ACH21, ABP25, ACG20, BWSG25, BZW^{+}25a, BYZZ20, BHW^{+}23, BB22, BRS^{+}24, BWL^{+}25, BYM22, CZB^{+}22, CLZG22, CCZ^{+}25, CFC^{+}22, CXL^{+}23, CXL^{+}25, CKJ^{+}22, CZW^{+}24, DGQ^{+}25, DMD^{+}23, DSP^{+}21, DLY21, DSJ^{+}22, DKZ^{+}25, FXL^{+}24, FPX^{+}25a, FHL^{+}23, FTR23, FRFM^{+}25, FZM^{+}23, GMZ22, GWG^{+}24, GHK^{+}25, GLB21, GYH^{+}22, GWX^{+}23, HKC^{+}23, HPJK22, HCC^{+}25, HBS20a, HMJ24, HDL^{+}24, HWL^{+}24, HYL^{+}25, HLK^{+}25, IKTY22, IMQOP21, IKAG^{+}22, JYM20, JDH^{+}25, JLL^{+}20, JLD^{+}25, JYM^{+}23, JHMM23, JZY^{+}23, KJC^{+}21, KSKK23, KAA22, KK25, KLC20, KJK24, LCZ22, LHR^{+}22, LLFT23, LHXH24, LFGD25, LWL^{+}21, LRB23, LCH22, LAKS20, LHL^{+}23, LZG^{+}24, LSS25, LWW^{+}25, LWH20, LSXZ21, LLL25b, MYGA20, MSSL21, MPYJ25, MYUK21, MHA^{+}20, MFRR20, MRB^{+}24, MHK^{+}22, MKYP21, NS22, NKeSK^{+}23, NHW^{+}24, OKU^{+}24, PC24, PL21, PYS^{+}23, PCA^{+}24, PN25, QLH^{+}24, QZZ^{+}24, RGS22, SKR^{+}20, SNN21, SRP^{+}21, SAG22, SKH^{+}25, SBP^{+}20]. Efficient [SWR^{+}25, SXC^{+}24, SCY^{+}23, TWL^{+}22, TWaKo^{+}23, WFW^{+}20, WDCC20, WHC20, WLW^{+}22a, WDQ^{+}22, WLW^{+}22b, WYSL22, WWM^{+}23, WZCM23, WJLC24, WYX^{+}24, WDC^{+}25, WFH25, WLF^{+}25a, WFL^{+}25, WWL^{+}23, XL25, XXL^{+}23, XZL^{+}25, XYM23, YWQ^{+}25, YZM^{+}25, YLG^{+}23, YQX^{+}25, YLC^{+}21, YYQ^{+}24, YZY^{+}25, YNP^{+}24, ZSHB21, ZZG^{+}23, ZLL^{+}24a, ZQY^{+}20, ZFZ^{+}21, ZYL^{+}22, ZCWC23, ZSC^{+}23, ZLWJ23, ZJW^{+}24, ZCH^{+}24, ZXZ^{+}24, ZSC^{+}25, ZWL^{+}25, ZWSF24, ZYZ^{+}25b, dSBS^{+}22]. Efficiently [AHK^{+}21]. Efflux [SZ22]. EGCN [HPJK22]. EiC [Kar24]. Eidetic [ESW^{+}23]. Eigenmode [IDFH22]. EIHDP [WLW^{+}21]. Elaborate [WHL^{+}23]. Elastic [CB22, WHY^{+}22]. ElasticDNN [ZHL^{+}24]. Electrical [AAB^{+}23]. Electromagnetic [BGM^{+}23, DYPZ22, HDAS21, SNA^{+}20, UYZP22]. Electronic [SNRB23]. Elementary [CWS^{+}24b, MFRR20]. Elevator [TKN23]. Eliminate [DZC^{+}24, DLK25, IKTY22]. Eliminating [DSK23, LHK^{+}22, SZK^{+}22]. Elimination [HLLC21, IDFH22, LZG^{+}24, WWM^{+}23]. Elisabeth [WT25]. Elliptic [MDJ20]. ELmD [BMLOM20]. ELOFS [ZLC^{+}22]. Emanations [HDAS21]. Embedded [BHE21, BJMKK23, DPCL22, DPQK^{+}23, DYPZ22, FW23, Has23, HF22, HXL^{+}25, HWG^{+}23, JWD^{+}22, JYM^{+}23, KHHK21, LKK^{+}21, LJH^{+}25, LCL^{+}20, MHK^{+}22, SMP22, SKA^{+}22, TRBM22, ZLC^{+}22]. Embedding [FXL^{+}24, KOT^{+}23, ZLC^{+}23b]. Emergency [LHW^{+}25]. Emerging [BMM^{+}22, CCZ^{+}22, WS20]. Empirical [HZC^{+}25, RWCC23]. Employing [JCKH22]. Empowered [JHMM23, MSP^{+}21, SST^{+}24, SJYQ25, WPL^{+}23]. Empowering [XXL^{+}23]. Emulating [BLM21]. Emulation [BBdTF25, EAMJ^{+}23]. En-Route [HMK^{+}21]. En/Decoding [TZ22]. Enable [WZD^{+}20]. Enabled [CXL^{+}25, GSB23, LGL^{+}24, LV23, LJY^{+}24, MXY^{+}23, TRG^{+}24, XWP^{+}21, XZC^{+}23, YYQ^{+}24, YCL^{+}24, ZSC^{+}25]. Enabling [CCT^{+}20, CCYC22, CHL^{+}23, DTH^{+}24, FQYS23, GWX^{+}23, JLL^{+}20, KZS^{+}25, LDG^{+}22, LHR^{+}22, LWL^{+}25, LCH22, LZG^{+}24, LZX^{+}25, LTJS^{+}22, MÁJG^{+}24, NM22, SHZ^{+}25, TBS^{+}25, WFW^{+}20, WWM^{+}23, WGD^{+}22, WSC^{+}25, XZL^{+}25, ZZG^{+}23, ZFZ^{+}21]. Encapsulation [HBS^{+}20b, SAJA21]. Enclaves [GWZ^{+}21]. Enclavisor [GWZ^{+}21]. Encoding [CPB21, NK22, WSC^{+}24, YLL^{+}20]. Encrypted [GMT24, LMM^{+}22, LHR^{+}22, LMM^{+}23, LTJS^{+}22, SHZ^{+}25, XGMJ25, YZM^{+}25, ZXW^{+}24]. Encryption [CCT^{+}20, CDF^{+}21, GLGL23, KDE^{+}24, LJZ^{+}25, LAKS20, PYS^{+}23, SYL^{+}23, SYD^{+}24, TRG^{+}24, TRV20, XCG^{+}25, YWQ^{+}25, ZYZ^{+}25b]. End [BGB^{+}21, CAC^{+}22, GGZC22, LCHK22, WZX^{+}22, WSHJ23, ZCR22, ZLS^{+}24]. End-to-End [BGB^{+}21, CAC^{+}22, GGZC22, LCHK22, WZX^{+}22, WSHJ23, ZCR22]. Endogenous [GQJ^{+}22]. Endpoint [CWS^{+}24b]. Endurance [FCZ^{+}23, YH20]. Energy [AhRX^{+}20, AHC^{+}20, ACH21, BB20, BB22, BRS^{+}24, CBB^{+}21b, CXL^{+}25, DMD^{+}23, DLY21, DSJ^{+}22, DH20, DSTD22, FRFM^{+}25, FZM^{+}23, GQZ21, GHK^{+}25, GYH^{+}22, HMJ24, HF22, JYM^{+}23, KSKK23, KMAA25, KK25, KLC20, KH23, KJK24, KIY21, LWL^{+}21, LHL^{+}23, MSSL21, MHA^{+}20, MFRR20, MHK^{+}22, NML25, NHW^{+}24, QHZ^{+}21, RGvS^{+}24, RSA^{+}20, SMFS21, SNN21, SAG22, SKLR22, STK23, SBP^{+}20, TKN23, WFW^{+}20, WLW^{+}22a, WLW^{+}22b, WLD^{+}25, XST20, ZSHB21, ZZL21, ZLWJ23, ZXD^{+}24, ZCF20].

Energy-Aware [GQZ21, HF22, KH23]. Energy-Delay [KK25]. Energy-Delay-Aware [WLD^{+}25]. Energy-Efficiency [BB20]. Energy-Efficient [AhRX^{+}20, AHC^{+}20, ACH21, BB22, BRS^{+}24, CXL^{+}25, DMD^{+}23, DLY21, DSJ^{+}22, FRFM^{+}25, FZM^{+}23, GHK^{+}25, GYH^{+}22, HMJ24, JYM^{+}23, KSKK23, KJK24, LWL^{+}21, MFRR20, MHK^{+}22, NHW^{+}24, SAG22, SBP^{+}20, WFW^{+}20, WLW^{+}22a, WLW^{+}22b]. Energy-Resilient [STK23]. Enforcement [FHL^{+}22]. Enforcing [PFHD21]. Engine [BRS^{+}24, DSK23, FPHW25, SPH^{+}23, TDZ^{+}22, XJL^{+}25, YZJ23, GWX^{+}23, YFC^{+}22]. Engineering [HWG^{+}23, ST23a, XTWG23]. Engines [DA22, JJKP22, PCBD23]. EnGN [LWL^{+}21]. Enhance [WHK24]. Enhanced [AOM^{+}21, CWY^{+}23, JCKH22, JLP^{+}25, KIY21, LWL^{+}24, SMY22, WLZ^{+}23, WLHW25, YZX^{+}24]. Enhancement [BCRX23, LJY21, ZCY^{+}25]. Enhancing [DWW25, GWG^{+}24, HWL^{+}21, HWR^{+}24, PHC24, VCLN21, WZF^{+}24]. Enlarging [MHJ^{+}21]. Ensemble [VAV^{+}20]. Ensembles [PPQBA21]. Ensuring [ZDC^{+}25]. Entangling [RJ24]. Entropy [CLLdS25]. Enumeration [PLH^{+}24, PHL^{+}25]. Environment [DWYX20, KD25, LGX^{+}22, SKM^{+}23, TDZ^{+}22, WCZ^{+}24, XHY^{+}22, YHC^{+}20, ZFL^{+}22]. Environment-Aware [KD25]. Environments [GQZ21, HZYY22, HYL^{+}25, IWKB22, LWZ^{+}25, PLZ20, STZ^{+}24, TDMP23, TCJ23, ZQG^{+}24, ZLT^{+}25]. EPRICE [FHL^{+}23]. Equality [ST23b]. Equations [AHK^{+}21]. Equivalent [WZG^{+}23]. ERA [LZW^{+}24a]. ERA-BS [LZW^{+}24a].

Erasure [GSC^{+}23, KAA20, SLS^{+}21, XLL^{+}22, YH24, ZFH23, ZFH^{+}25]. Erasure-Coded [GSC^{+}23, SLS^{+}21, ZFH23, ZFH^{+}25]. Error [BN24, BPJ^{+}22, BCMT23, BFC20, CXW^{+}23, DT20, DSJ^{+}22, FDKK21, FTR23, FMM^{+}21, HYW^{+}21, JYM20, KGHRM23, LWC^{+}22, LRB23, LRL22, MWJ^{+}24, MTV^{+}21, RMO21, Rot24, TWY^{+}25, TTG^{+}23, YHC^{+}20, YZG^{+}25]. Error-Bounded [LWC^{+}22]. Error-Correction [KGHRM23]. Error-Detection [Rot24]. Error-Resilient [BN24, YHC^{+}20]. Error-Tolerant [CXW^{+}23, FTR23, LRL22]. Errors [BLP^{+}22, BCMT23, KB21, LSCX20, PAR^{+}22, PPQBA21, QHT^{+}24, WNL^{+}23, YBG^{+}22, ZJJ25]. Escrow [GLGL23]. Escrow-Free [GLGL23]. Estimating [KJC^{+}21]. Estimation [AVK20, BBC^{+}22, SMZ^{+}20, YCS^{+}24]. Estimations [BPJ^{+}22]. ETBench [ZTY^{+}25]. European [NCD^{+}25]. Evading [PM25]. Evaluating [FXJW25, TWY^{+}25]. Evaluation [ABC^{+}24, BWSG25, BFG^{+}21, CHM25, DMX^{+}22, GKT^{+}22, HHZ^{+}23, MLL^{+}24, MÖS22, RWCC23, SSP^{+}24, ZTLW23]. Evaluations [LFX^{+}21]. Evasion [IKAG^{+}22]. Evasive [JJZW24]. Even [BLM21]. Event [BGHR^{+}25, GPH20, KMVD22]. Event-Driven [GPH20]. Evolutionary [RSZ23]. Evolving [ZHL^{+}24]. Exact [BBL22, CXW^{+}23, ZWM20]. Example [ZCY^{+}24]. Exceptional [BFC20]. Exchange [AMJ^{+}23, PD21]. Executing [TQL^{+}22]. Execution [BLKK23, DYC^{+}25, DSP^{+}21, FWM^{+}23, FBM21, GCR^{+}23, HJYL22, KJC^{+}21, LPW20, LGX^{+}22, NM22, PS22, SKR^{+}20, SCL^{+}24, WCZ^{+}24, XHY^{+}22, YBW21, ZSHB21, ZCH^{+}24].

Executions [WLD^{+}22]. Exhaustive [FWM^{+}23]. Existing [TC21]. Exit [CR24, JQY^{+}25, KPS^{+}24, NLC^{+}25]. Expandable [WHY^{+}22]. Expediting [CXH^{+}25]. Experience [BCRX23, ZLWG22]. Experience-Driven [ZLWG22]. Experiment [CXY24]. Experiment-Driven [CXY24]. Expert [KLW^{+}25]. Explainable [GWD25, PSM22]. Explicit [BFC20, TZZ^{+}21]. Exploiting [BCV22, CXH^{+}25, CCC23, KKS^{+}22, KLKK23, LSCX20, PB23a, ROPdlT22, SIR20, SCY21, SZL^{+}24, WWM^{+}23, WLY^{+}23, WWS^{+}22, WDZ^{+}22, XNB21, XTWG23, YYW^{+}24, ZLH^{+}21, ZMS^{+}23, ZCCG23, ZG23]. Exploration [CPM^{+}23, HZM^{+}23, JLY^{+}21, LLFT23, LSXZ21, MHJ^{+}21, QHZ^{+}21, SMC^{+}25, YLC^{+}21, ZML^{+}25, ZWL^{+}25]. Exploring [DKJP21, KYS^{+}22, LLT^{+}23, LYC^{+}23, MDPM24, NKeSK^{+}23, PCA^{+}24, ZJJ25]. Exponential [God20]. Exponentially [ZLC^{+}23b]. Exponentially-Many [ZLC^{+}23b]. Exposed [HMJ24]. Extended [GFB^{+}24, KJK^{+}22, VTM^{+}20]. Extending [FRFM^{+}25, LGX^{+}22, ZZL21]. Extensibility [XHY^{+}22]. Extensible [ZLC^{+}22]. Extension [ABP22, BHK^{+}23, GFB^{+}24, KGHRM23, SZHB21, XZL^{+}25]. Extensive [MKÖ^{+}22]. External [MKYP21, SNA^{+}20]. Extracting [WSM^{+}24]. Extraction [ATT22]. Extrapolation [WCL^{+}23]. Extreme [JKHL22, YL20]. Extreme-Scale [JKHL22].

FaaS [CZH^{+}25]. FaaSBatch [WDZ^{+}24]. Fabric [DSCB25]. Fabrication [EGP24]. Fabrication-Time [EGP24]. FACCT [ZSS20]. Factored [Das23]. Factorization [GZW^{+}21, LWZZ25]. Factors [TPWY23, TDH^{+}23]. FadingBF [VKRK22]. Failure [HLS^{+}23a, LDZ^{+}23, PZY^{+}23, SLLS25, TZ22, XLL^{+}22, YZX^{+}24]. Failures [JZX^{+}25, LLCJ23]. Fair [GXY^{+}23, KMAA25, LHR^{+}23, SCFPM22, WZJ^{+}24, XNL^{+}23]. Fairness [BLH^{+}21, LHN^{+}22, SNN21]. Fairness-Aware [SNN21]. Falcon [ZGL^{+}21]. Falic [YLZ^{+}24]. False [RGS22]. Families [EEA22]. Family [WCQW22]. Farewell [Kar24]. FAS [ZLL^{+}22b]. FAS-DQN [ZLL^{+}22b]. FasDL [CCZ^{+}25]. Fast [AJ22, ACG20, BLKK23, BHE21, BWL^{+}25, BAM^{+}24, BCMT23, CCT^{+}20, CVOJRH22, CXY24, CLLdS25, CZW^{+}24, DYJ20, DYC^{+}25, DVA22, GPH20, GWX^{+}23, HF23, HJX^{+}25, JLL^{+}20, JYH^{+}24, LLW^{+}24, MZZC22, NTDH25, PLB22, PSBB21, PN24, SCY21, TWZ^{+}23, TZZ^{+}21, TZ22, VLPS25, WSC^{+}25, WHL^{+}21, XLL^{+}22, YZJ23, YLL^{+}20, ZCJ^{+}20, ZWM20, ZFH^{+}25, ZSS20]. Fast-Track [VLPS25]. Faster [LLWZ23]. Fault [BBC^{+}20, BBC^{+}22, CFW25, FXC^{+}23, FXL^{+}24, GXZ^{+}24, HP23, JLZ^{+}23, JZY^{+}23, JZX^{+}25, LWL^{+}23, LQL^{+}24, MCT22, MDR^{+}24, OLD^{+}23, RKMR23, RBSG23, RBM21, RRMS25, SKK23, SWR^{+}25, TOF^{+}24, VTM^{+}20, WT25, WFZ^{+}25, ZXL^{+}23]. Fault-Free [SKK23]. Fault-Tolerant [FXC^{+}23, FXL^{+}24, GXZ^{+}24, JZY^{+}23, MCT22, SWR^{+}25]. Faults [CDRS20, FMM^{+}21, RBSG23, YNJS21]. Faulty [FXL^{+}24, ZLC^{+}23b]. FAWA [JYH^{+}24]. FBF [BL22]. Feather [LHS^{+}25, PCA^{+}23]. Feather-Weight [LHS^{+}25]. Feature [NCD^{+}25, YYW^{+}24]. Features [JLL22, KMVD22, MWJ^{+}24, WHC^{+}23].

Featuring [GCR^{+}23]. Fed [LSL^{+}25]. Fed-OGD [LSL^{+}25]. Federated [CWY^{+}23, CDW^{+}25, CZW^{+}24, FZG^{+}22, FTV25, GPQ23, GZG^{+}23, JHMM23, LZY^{+}25, LSL^{+}25, LSU^{+}23, LGZ^{+}25, LHZR25, LDW^{+}25, MHM^{+}23, NNH^{+}25, QJY^{+}25, TCX^{+}23, WRW^{+}23, WHL^{+}21, XWL^{+}24, YZX^{+}24, YPD^{+}24, YYZ^{+}25, YDL^{+}25, ZLL^{+}24a, ZLWG22, ZGQ^{+}22, ZGG^{+}23, ZXY^{+}24, ZMH^{+}25, ZXMX25, ZSX^{+}24]. FedGKD [YPD^{+}24]. FedQClip [QJY^{+}25]. FedRFQ [YZX^{+}24]. Feedback [CB22, LWYJ23, TARK23]. Feedback-Driven [LWYJ23]. Feedforward [GSB23]. FeFET [GvSHA22, KZXR25, KSL^{+}22, YBG^{+}22]. FeFET-Based [YBG^{+}22]. Fenglin [FQYS23]. Fenglin-I [FQYS23]. Fermat [XLY^{+}25]. Ferroelectric [LHX^{+}25, LLY22]. Ferroelectric-Based [LLY22]. FET [LHX^{+}25]. Fetch [MAM23]. Feynman [BBdTF25]. Fidelity [RMR22]. FiDRL [LJH^{+}25]. Field [CLCL22, TRG^{+}24, YCS^{+}24, ZSC^{+}25, KGHRM23]. FIFA [MSZ22]. FIFO [GSB23]. File [CSW^{+}21, GGZC22, HWL^{+}21, LCH22, ZLC^{+}22]. FileDAG [GXZ^{+}23]. Filter [BL22, LMM^{+}23, NKN^{+}25, VKRK22, WHY^{+}22]. Filter-Based [LMM^{+}23]. Filtered [LZF21]. Filtering [ACKA23, BAM^{+}24, BBC^{+}22, LCHL21, SAG22, TWZ^{+}23, ZCJ^{+}20]. Filters [Alm23, KS24, RSMMG^{+}23, RGD^{+}24, VHL20]. Finance [WHQ^{+}24]. Finding [AMM21, JBK24]. Fine [CHL^{+}23, DDK22, DZC^{+}24, FCZ^{+}23, HWX^{+}24, HSH^{+}25, JWS^{+}21, LZW^{+}24a, MLL^{+}24, PL21, XNL^{+}23, ZZG20, ZXZ^{+}24, ZCCG23, ZXZ^{+}21]. Fine-Grained [CHL^{+}23, DDK22, DZC^{+}24, FCZ^{+}23, HWX^{+}24, HSH^{+}25, JWS^{+}21, LZW^{+}24a, MLL^{+}24, PL21, XNL^{+}23, ZZG20, ZXZ^{+}24, ZCCG23, ZXZ^{+}21]. FinFET [SKA^{+}22]. FinFETs [RvSP^{+}25]. Fingerprint [ZWY^{+}23]. Fingerprints [WHY^{+}22]. Finite [TRG^{+}24]. Finite-Field [TRG^{+}24]. Firm [GPRV23]. Firmware [TLC^{+}24]. First [LLFT23, MHS^{+}20, SMP22]. Fits [WGD^{+}22]. Fixed [BBL22, JCY^{+}23, JCKH22, VHL20, XNH^{+}25, ZABHCG23]. Fixed-Base [JCKH22]. Fixed-Point [VHL20]. Fixed-Priority [XNH^{+}25, ZABHCG23]. FL [ZLL^{+}24a]. FLALM [XLZ^{+}25]. Flash [BMM^{+}22, CCYC22, JKHL22, KKS^{+}22, KCAL21, KOT^{+}23, LSCX20, PLZ^{+}23, PZY^{+}23, SLLS25, WZW^{+}23, WHL^{+}23, WDZ^{+}22, WDZ^{+}23, WHK24, YWC^{+}21, YZG^{+}25, YH20, ZLC^{+}22, KKS^{+}22]. Flash-Based [BMM^{+}22, JKHL22, KKS^{+}22, WZW^{+}23]. Flash-Memory [YH20]. FlashDecoding [DHM^{+}25]. FLEX [LL22]. FlexBlock [NKL^{+}23]. Flexible [BLKK23, DTL^{+}25, GLW^{+}24, HDL^{+}24, JQK^{+}24, KAA22, LL22, LCX21, LJH^{+}25, LLD^{+}25, LFGD25, MKÖ^{+}22, NKL^{+}23, SLDZ25, WJLC24, XLZ^{+}25, ZCK20, ZLZ24a]. FlexiPair [BRPM22]. Flight [NKN^{+}25]. FLiMS [PLB22]. Flip [LW22, LGC^{+}23, ST23a, RBM21]. FLIXR [KOT^{+}23]. Floating [BLM21, Bru20, Bru23, CQI^{+}22, GNH20, GFB^{+}24, KPH^{+}25, LLL25b, Mik24, NKL^{+}23, TOM23, ZSHB21]. Floating-Point [BLM21, Bru20, Bru23, CQI^{+}22, GNH20, KPH^{+}25, LLL25b, Mik24, TOM23, ZSHB21]. Flock [PCA^{+}23, ZWC^{+}23]. Flop [LW22]. FLOPs [NHW^{+}24]. Flow [AB22, DDK22, DTZ^{+}25, FHL^{+}22, GSB23, HBS20a, xHzLH^{+}24, HWJ^{+}21, HGC^{+}22, MSLY24, PL21, SSP^{+}24, SPH^{+}23, WZSL22, YCKW20, ZYZ^{+}23, ZXX^{+}25]. Flow-Based [HGC^{+}22]. Fluid [GPQ22, GQH21]. Flux [FHW^{+}22]. Fly [BJMKK23]. Focused [AAM^{+}25]. Fog [GQZ21]. Folding [RBMG22]. Footprint [MIY^{+}20]. Footprint-Based [MIY^{+}20]. Forensics [ZHM20]. FORESEE [KJC^{+}21]. Forest [SWR^{+}23]. Forgetting [MZM^{+}25]. Fork [MPYJ25]. Fork-Based [MPYJ25]. Form [ZTT22, ZWM20]. Formalization [ZCR23]. Format [SLY^{+}22b]. Formats [GNH20]. Forward [RMTA20, XCG^{+}25, YWQ^{+}25]. Four [YLL^{+}20, YH24]. Fourier [LRRK^{+}22]. FP [RACB24]. FPDeep [WGL^{+}20]. FPGA [CDP21, CTZ^{+}24, CPM^{+}23, DMG23, DNMS20, ESN20, EAMK22, GMZ22, HWJ^{+}21, LDF^{+}24, LGW^{+}22, MYUK21, MSZ22, PYS^{+}23, RPB^{+}23, SPB^{+}21, SDR^{+}22, SM24, TRV20, WGL^{+}20, WDCC20, WGM^{+}20, WGJZ21, WCZ^{+}24, WAN^{+}25, WWL^{+}23, WYB^{+}24, XLZ^{+}25, XJL^{+}25, YLZ^{+}24, ZMS^{+}23, ZYZ^{+}25b, ZBT22, ZGG25, dHBF^{+}21]. FPGA-Accelerated [WYB^{+}24]. FPGA-Based [CDP21, LDF^{+}24, YLZ^{+}24, ZGG25]. FPGA-SoC [WCZ^{+}24]. FPGAs [AB22, AGQ^{+}23, BK23, CLLdS25, DRY^{+}22, HBS^{+}20b, HWZ^{+}22, KMAA25, LRB23, ROPdlT22, SIR20, ZDZ^{+}23]. Fractal [ZFD^{+}20]. Framework [AKG^{+}20, BRPM22, BKHY22, CTY^{+}24, CCCW21, CHM25, CPL^{+}23, CPM^{+}23, DGQ^{+}25, DWN^{+}22, DLK25, EGP24, FZG^{+}22, Gha21, GKFF20, GVN25, HLS^{+}23a, HCC^{+}23, HIRB25, HDL^{+}24, HYQ^{+}25, HMZ^{+}25, JMW^{+}24, JLP^{+}25, KJC^{+}21, KKRK22, KSKK23, KKKC20, KAA20, LL21, LCX21, LHXH24, LQL^{+}24, LLZ^{+}25, LZW25, MLL^{+}24, MKYP21, NTDH25, PAR^{+}22, PLZ20, QHT^{+}24, RCS^{+}21, RMKO23, SNA^{+}20, SSW^{+}24, SYD^{+}24, SKK23, SJYQ25, WZX^{+}22, WJLC24, WSLX24, WSQ^{+}25, WZG^{+}24, XXJ^{+}24, XJY^{+}24, YCS^{+}24, YTD^{+}21, YWC^{+}24a, YCL^{+}24, ZLL^{+}24a, ZML^{+}24, ZSC^{+}25, ZHM20, ZGG25, ZHYJ21, dHBF^{+}21]. Free [BYM22, BLM25, GR23, GLGL23, KJK24, LHY^{+}21, SKK23, ZWJ^{+}25]. Frequency [APV22, GLMZ25]. Frequent [NK22]. Freshness [ZLL^{+}22b]. Freshness-Aware [ZLL^{+}22b]. Friendly [WDL^{+}25, WT25, WLC^{+}24]. Frontend [XTWG23]. FSM [FHH22]. Full [CGLS21, KKL^{+}25, SZAT22, SZHB21, WSG^{+}23, ZLW^{+}24, ZWZZ24]. Full-Chip [SZAT22]. Full-Round [ZLW^{+}24, ZWZZ24]. Fully [DLZ^{+}24, LYC^{+}23, MYUK21, MSZ22, TRG^{+}24, WLW^{+}22a, WZG^{+}24, YWC^{+}24a]. Function [BYM22, FHH22, God20, GSS^{+}23, GQJ^{+}22, HYL^{+}25, LLW^{+}24, LBC24, MSLY24, NTDH25, SLY22a, SSM23, XZL^{+}23, ZZG^{+}23, ZFL^{+}22, ZCWC23, ZWC^{+}23, ZXX^{+}25]. Function-as-a-Service [ZWC^{+}23]. Functional [BL22, KBR^{+}23]. Functionalities [PCA^{+}23]. Functions [AMM21, AA20, BYZZ20, BCCM22, CWS^{+}24b, KSB24, KS24, LCHL21, LSW25, VJWZ^{+}21, XGMJ25]. Fused [DGQ^{+}25, SMC^{+}25, WYX^{+}24]. Fused-Grained [DGQ^{+}25]. Fusion [BLKK23, xHzLH^{+}24, ZDC^{+}25]. Future [LHK^{+}22]. FutureDID [DLZ^{+}24]. FVM [LHS^{+}25].

G [CHM25]. Galois [KGHRM23]. Game [BLH^{+}21, DKZ^{+}25, NHW^{+}24, WLF^{+}25a, ZGLZ20]. Game-Based [NHW^{+}24]. Game-Theoretic [DKZ^{+}25]. Gamepad [BLH^{+}21]. Games [LL21]. Gaming [LWL^{+}22]. GAN [KBR^{+}23, MSSL21, RSA^{+}20].

GANDAFL [KXGS22]. Gang [RACB24]. GANs [ZCY^{+}24]. Gap [CJSY24]. Garbage [ZFH^{+}25]. GAS [ZLL^{+}24b]. Gate [BBdTF25, FFG^{+}25, MLL^{+}25, UMM^{+}20, YQX^{+}25]. Gate-Circuit [BBdTF25]. Gate-Level [FFG^{+}25, MLL^{+}25]. GateKeeper [BAM^{+}24]. GateKeeper-GPU [BAM^{+}24]. Gates [XSYL22]. Gateway [PD21]. Gating [HBB^{+}21, HXL^{+}25]. Gaussian [ERKP21, KAA22, KLL21, SZS^{+}22, ZSS20]. GCN [HPJK22]. GCNs [SYW^{+}22]. GCONV [ZCR22]. Gem5 [QHT^{+}24]. Gem5Tune [QHT^{+}24]. GEMM [FRFM^{+}25, GXL^{+}24]. General [CNOS22, FHH22, HLS^{+}23a, KDE^{+}24, LYW^{+}25, MZZC22, Xu24, ZLL^{+}24b]. General-Purpose [ZLL^{+}24b]. Generalized [CQCL25, LTFL22, MLL^{+}24, MKY^{+}24, MBP21, ZWWY22]. Generated [DGTVGG21]. Generating [LGC^{+}23, XLS^{+}24]. Generation [ABI^{+}25, BWL^{+}25, BJMKK23, CDRS20, DVA22, FLF20, HCKK25, KPS^{+}24, MHT25, PMA^{+}24, QSC^{+}25, STQ^{+}24, XJY^{+}24, YGW^{+}23, ZLWJ23, ZLS^{+}24, ZG23, ZCY^{+}24]. Generations [BMM^{+}22]. Generative [RSR22, XLW^{+}20]. Generator [CMQ^{+}22, CTZ^{+}24, FHH22, HF23, SCL^{+}24]. Generic [CHM25, ZZW^{+}25]. GenoDedup [CPB21]. Genome [CPB21, QHZ^{+}21]. Geometry [LSS25]. GF [GMZ22]. GFBE [MLL^{+}24]. Global [BBL22, CKP^{+}22, JSTG20, JCY^{+}23, MHM^{+}23, SXZJ24, XAP20, YPD^{+}24]. Globally [BK23]. GNN [LDT^{+}25, WSHJ23, XXC^{+}25, ZLL^{+}23]. GNU [SPDQ22]. Good [JJKP22]. Google [SZK^{+}22]. Governing [LL21]. GP [JDCL23]. GPGPU [SAG22, SJYQ25, YNJS21].

GPGPU-Empowered [SJYQ25]. GPU [BZW^{+}25b, BAM^{+}24, BBD^{+}20, CKRP21, CKJ^{+}22, FFG^{+}25, FBM21, HCC^{+}23, HLT^{+}23, JWS^{+}21, KKRK22, LKK^{+}21, LPW20, LL21, LXW^{+}23, LZW^{+}23b, LHZ^{+}24, LLL^{+}25a, NML25, SYD^{+}24, SPH^{+}23, TWY^{+}25, TDZ^{+}22, WXL^{+}23, WDC^{+}25, YLT^{+}23, YCL^{+}24, ZZG^{+}23, ZWX^{+}25, ZDC^{+}25, ZSC^{+}25, ZWJ^{+}25, vSDHA23]. GPU-Based [BZW^{+}25b, WXL^{+}23, FBM21]. GPU-Enabled [YCL^{+}24, ZSC^{+}25]. GPU-Powered [KKRK22]. GPU-SPICE [vSDHA23]. GPUs [CXH^{+}25, JLL^{+}20, LZW23a, PHL^{+}25, WSHJ23, XXC^{+}25, XZL^{+}21, YWX^{+}23, YBW21, ZCZ^{+}22, ZCCG23]. Gradient [CCT^{+}20, LSL^{+}25, MZM^{+}25, WGT^{+}22, YGW^{+}23, ZMH^{+}25]. Gradients [LDW^{+}25]. Grain [DPS22]. Grain-128a [DPS22]. Grained [CHL^{+}23, DGQ^{+}25, DDK22, DZC^{+}24, DPQK^{+}23, FCZ^{+}23, HWX^{+}24, HSH^{+}25, HZF^{+}24, JWS^{+}21, LB22, LZW^{+}24a, MLL^{+}24, PL21, XNL^{+}23, ZZG20, ZXZ^{+}24, ZCCG23, ZXZ^{+}21]. GraNDe [YNP^{+}24]. Graph [APH^{+}23, DSP^{+}21, FNS^{+}22, FWR^{+}20, GWG^{+}24, HYH^{+}25, JCZ^{+}23, JJKP22, KPD^{+}23, KLR23, LZC^{+}21, LTFL22, LCY^{+}25, LXL^{+}25, LT25, LWL^{+}21, LZZ^{+}22, LDT^{+}25, MLL^{+}25, MAM23, MB21, MCS^{+}22, NKA24, PLH^{+}24, SCY21, WCQW22, WXL^{+}23, WWC21, XYM23, YQX^{+}25, YNP^{+}24, ZYQ^{+}24]. Graph-Based [KPD^{+}23]. Graphfire [MAM23]. Graphs [KZXR25, LYC^{+}23, MBP21, RDS23, SSK22, WSM^{+}24, ZLL^{+}23]. Gray [WCL^{+}23]. Gray-Box [WCL^{+}23]. Great [WCYK20]. GreedW [WJLC24]. Grid [ZWB^{+}22]. GRIP [KLR23]. GroPipe [LMH^{+}25]. Grouped [LMH^{+}25]. Groups [ZTLW23]. Grow [DYJ20]. Growing [YCKW20]. Growing-Scale [YCKW20]. Guarantee [ZDY^{+}23]. Guaranteeing [LCHK22]. Guarantees [GJN^{+}25, LXL^{+}25, VKRK22, ZQG^{+}24]. Guardauto [CZC^{+}21]. GuardBands [MTV^{+}21]. Guest [Ano23, BBJR21, CDP21, CQ22, FAKM22, Lu21, QWK20, WS20, AW20]. Guests [BXW^{+}25]. Guided [SLLS25, TTG^{+}23, YL20]. Guidelines [LCS^{+}25]. Gym [MS25].

H [LLL^{+}25a]. H3 [KS24]. HAM [WTL^{+}21]. Hamiltonian [ZLC^{+}23b]. Handheld [WZD^{+}20]. Handling [DGQ^{+}25, HMMP23, NNH^{+}25, ST23b]. HAOTuner [MLW^{+}23]. Hard [JYM^{+}23, LPC^{+}21, LW22, PS22, SM22]. Hardening [dSBS^{+}22]. Hardware [APH^{+}23, ALC25, Ano23, AHK^{+}21, AW20, BN24, BMLOM20, CGLS21, CQ22, CQCL25, CZW^{+}25, DMG23, DMD^{+}23, DRP24, DLY21, EGP24, FAFK21, FB20, FAKM22, FHW^{+}22, FWR^{+}20, GWG^{+}25, GR23, GYH^{+}22, GWZ^{+}21, GL24, HF23, HHN^{+}23, HYH^{+}25, HWG^{+}23, HXL^{+}23, HLJ^{+}25, HMZ^{+}25, IKAG^{+}22, JWD^{+}22, KAA22, KMH^{+}23, LG22, LCY^{+}25, LQY^{+}20, LRRK^{+}22, LLK^{+}23, LZW25, MLL^{+}25, MDJ20, MC23, MFRR20, MLW^{+}23, MLW^{+}25, NCD^{+}25, ODK20, OAC^{+}21, PAR^{+}22, PSM22, PM25, PYS^{+}23, PK23, PMA^{+}24, PN25, RBMG22, RBSG23, ROPdlT22, SNRB23, SZL^{+}22, SSM23, SXXL24, SZL^{+}24, TLLL25, UMM^{+}20, WFH^{+}24, WZF^{+}24, WDL^{+}25, WZD^{+}20, XXC^{+}25, XLS^{+}24, XCZ^{+}22, ZFZ^{+}21, ZJJ25, ZHM20, ZYQ^{+}24, ZTLW23, ZGG25, ZZZ^{+}20]. Hardware-Assisted [PSM22]. Hardware-Aware [ALC25, LLK^{+}23, XLS^{+}24, ZYQ^{+}24].

Hardware-Based [ROPdlT22, ZHM20]. Hardware-Friendly [WDL^{+}25]. Hardware-Mapping [DRP24]. Hardware-Software [GWZ^{+}21, XXC^{+}25]. Hardware-Trojan [HHN^{+}23]. Hardware/Software [PMA^{+}24, WZF^{+}24, XCZ^{+}22]. Harmonization [YWC^{+}21]. Harness [IDFH22]. Harnessing [AHC^{+}20, HDAS21]. Harsh [YHC^{+}20]. Harvesting [ZXD^{+}24]. Hash [FAFK21, KS24, LSW25, NT23, QCX^{+}23, RGS22]. Hash-Based [NT23, QCX^{+}23, RGS22]. Hashed [HLK^{+}25]. Hashing [WHX^{+}25]. HashScape [HLK^{+}25]. HASP [LMW^{+}24]. Hazard [BLM25]. Hazard-Free [BLM25]. HDC [CCCW21]. HDD [DGG^{+}22, LYC22]. HE-Friendly [WT25]. Head [NKN^{+}25]. Health [CXL^{+}25, DGG^{+}22, KCS23, LPC^{+}21]. Healthcare [LYW^{+}23]. HEAWS [TRV20]. HePREM [FBM21]. Herd [LDW^{+}25]. Heterogeneity [HYQ^{+}25, KMAA25, LLL^{+}25a, WGT^{+}22]. Heterogeneity- [HYQ^{+}25]. Heterogeneity-Aware [LLL^{+}25a, WGT^{+}22]. Heterogeneous [BY24, BOL^{+}25, BZW^{+}25b, BFG^{+}21, CTY^{+}24, CSY^{+}25, CYPC25, CDW^{+}25, CSK22, FBM21, GQZ21, GWG^{+}25, GSK^{+}22, HHPB20, HECC^{+}21, HZMC24, HTZ^{+}25, HYL^{+}25, HXGR20, JWG^{+}23, KKKC20, KH23, KIY21, LHA^{+}25, LSU^{+}23, LAPB21, MCD^{+}25, NML25, OAB^{+}23, PYW^{+}22, PYW^{+}25, RDS23, SCA^{+}25, SNN21, SRP^{+}21, SSK22, SKLR22, STZ^{+}24, SBP^{+}20, SMC^{+}25, TDZ^{+}22, VRR^{+}24, WLR20, WFL^{+}25, XCZ^{+}22, XJL^{+}25, XNH^{+}25, YWX^{+}23, YPD^{+}24, ZLL^{+}23, ZLWJ23, ZGG^{+}23, ZZM^{+}22, ZCX^{+}20, ZGL^{+}21, ZCZW23, ZFH23, ZWY^{+}23, ZDY^{+}23].

Heuristic [SEM23, YDG^{+}24]. HGNAS [ZYQ^{+}24]. HiBid [WTL^{+}24]. HiCoCS [YDW^{+}25]. Hidden [BYM22, DMD^{+}23]. Hiding [LJZ^{+}25, XGMJ25, ZLT^{+}24]. Hierarchical [CL20, GYH^{+}22, HLT^{+}23, HLF^{+}23, LMW^{+}24, LDF^{+}25, RCS^{+}21, WLW^{+}21, WTL^{+}24, WHX^{+}25]. Hierarchies [CSY^{+}25, MCD^{+}25]. Hierarchy [LHK^{+}22, LMZ^{+}25]. High [BHK^{+}23, BB22, BRS^{+}24, BWL^{+}25, CGS^{+}20, CQCL25, CZH^{+}25, CFC^{+}22, CZW^{+}25, DHM^{+}25, DMG23, DGTVGG21, DA22, DSTD22, EAMJ^{+}23, EAMK22, FHH22, FHL^{+}22, FTR23, GQZ21, HHPB20, HLQ^{+}23, HWZ^{+}22, HLT^{+}23, HLJ^{+}25, JDB^{+}23, JWS^{+}23, KLL21, KRB^{+}22, LDLK22, LWL^{+}25, LHX^{+}25, LWX^{+}25, LWL^{+}21, LSW^{+}23, LZW^{+}23b, LYH^{+}24, LHZ^{+}24, LMZ^{+}25, LGW^{+}22, LLR25, MHDMEA22, MCS^{+}22, NKeSK^{+}23, PYW^{+}25, PSBB21, RMR22, SNT22, STYQ24, SST^{+}24, TWZ^{+}23, TRBM22, UMM^{+}20, VCLN21, WSS^{+}20, WWM^{+}23, WZCM23, WHX^{+}25, WLW^{+}25, WDZ^{+}23, XCZ^{+}22, XLY^{+}25, YDW^{+}25, YZG^{+}25, YWF^{+}25, YLG^{+}23, YYW^{+}24, ZCP22, ZLWJ23, ZYZ^{+}25a, dHBF^{+}21]. High-Accuracy [SNT22]. High-Concurrency [LWX^{+}25]. High-Coverage [WZCM23]. High-Density [LWX^{+}25]. High-Efficiency [FHH22, PYW^{+}25]. High-Efficient [WWM^{+}23]. High-Intensity [HLJ^{+}25]. High-Level [CFC^{+}22, DSTD22, MCS^{+}22, PSBB21, VCLN21]. High-Performance [BRS^{+}24, BWL^{+}25, CZH^{+}25, CZW^{+}25, EAMJ^{+}23, EAMK22, FTR23, HHPB20, JWS^{+}23, KLL21, KRB^{+}22, LHX^{+}25, LZW^{+}23b, LHZ^{+}24, NKeSK^{+}23, STYQ24, SST^{+}24, TRBM22, WLW^{+}25, YWF^{+}25, ZYZ^{+}25a]. High-Performant [YYW^{+}24]. High-Precision [YZG^{+}25]. High-Radix [CQCL25, LDLK22, XLY^{+}25, ZCP22]. High-Radix/Mixed-Radix [XLY^{+}25]. High-Resilience [JDB^{+}23]. High-Resolution [MHDMEA22]. High-Speed [BHK^{+}23, DMG23, TWZ^{+}23]. High-Throughput [LWL^{+}21]. Higher [CMGD24, DVV23, RMR22]. Higher-Order [CMGD24, DVV23]. Highly [FPX^{+}25b, JYF^{+}23, JJZW24, LDF^{+}25, WDCC20, ZFZ^{+}21]. HIPEDAP [DMD^{+}23]. Historical [WFH^{+}24]. Hit [HA25]. Hit-Statements [HA25]. Hitchhiker [ZLZ^{+}23]. HLS [ZBT22]. Hogweed [MSP^{+}21]. Holes [SXC^{+}24]. Holistic [HCC^{+}23, LWX^{+}25, WZW^{+}23]. Holistic-Scalable [LWX^{+}25]. HOME [HCC^{+}23]. Homing [TWJ^{+}22]. Homomorphic [CDF^{+}21, KDE^{+}24, PYS^{+}23, RSZ23, SYL^{+}23, SYD^{+}24, TRG^{+}24, TRV20, ZYZ^{+}25b]. Homomorphically [LTJS^{+}22]. Homomorphism [DSP^{+}21]. Honeycomb [LDF^{+}24]. Hop [Akr22]. Host [ZYXD20]. Host-Aware [ZYXD20]. Hotness [FPHW25, HHPB20, LWL^{+}25]. Hotness- [HHPB20]. Hotness-Aware [FPHW25]. Hotplug [MIY^{+}20]. Hotspot [WTL^{+}21]. Hotspot-Aware [WTL^{+}21]. HPC [HECC^{+}21, SSZ^{+}20]. HPC-DAG [HECC^{+}21]. HPDK [LYH^{+}24]. HPKA [NKeSK^{+}23]. HQC [ABP25]. HQC-128 [ABP25]. HQC-192 [ABP25]. HQC-256 [ABP25]. HTDetector [HHN^{+}23]. Hubs [XXZ^{+}25]. Huge [HWC^{+}22a, JZSD24]. Human [LHXH22, LWW^{+}25, WZX^{+}22]. Human-Inspired [LWW^{+}25]. Humas [HYQ^{+}25]. HW [MÁJG^{+}24, TGA23]. HW-Based [MÁJG^{+}24]. HW/SW [TGA23]. Hybrid [ATT22, BN24, FPHW25, FB20, GWG^{+}24, GYZC25, HZYY22, IWKB22, JLP^{+}24, JXH^{+}22, KSB24, KB21, LDLK22, LQM^{+}24, LLL^{+}20, LZW23a, LYH^{+}24, LMH^{+}25, LY20, NICY24, NML25, NKN^{+}25, OKC^{+}20, PMA^{+}24, RPS^{+}21, RMR22, STW^{+}21, SSCK25, SKM^{+}23, TSM^{+}21, WLW^{+}25, ZGD23, ZNW^{+}24, ZTY^{+}25, CCG^{+}22, JLP^{+}24]. Hybrid-Memcached [JLP^{+}24]. Hybrid-SIMD [CCG^{+}22]. Hydra [YWX^{+}23, NICY24]. Hyperbolic [CQCL25]. Hypercall [BXW^{+}25]. Hypercall-Oriented [BXW^{+}25]. Hypercube [LDF^{+}25]. Hypercube-Structured [LDF^{+}25]. Hyperdimensional [CCCW21, KKRK22, KZXR25, SIR20, TGA23, WMH^{+}24, YWC^{+}24a, YWC^{+}24b, ZJJ25]. Hypergraph [LLX^{+}24b]. Hypervisor [BXW^{+}25, JWD^{+}22]. Hyte [FPHW25].

I/O [BMM^{+}22, CHZ^{+}25, DTH^{+}24, HYS^{+}20, HWL^{+}21, KJC^{+}21, LJY^{+}24, LYH^{+}24, NTDH25, PE22, WJL^{+}20, ZYXD20, ZXL^{+}24, ZZC^{+}23]. I/O-Bound [CHZ^{+}25]. I/Os [WZW^{+}23]. IBM [LFP^{+}22]. IC [JYF^{+}23, PB23a]. Idempotence [LKK^{+}21]. Idempotence-Based [LKK^{+}21]. Identification [HLL^{+}20, KLR^{+}20, LWL^{+}25, MGFY24]. Identifying [SXZJ24]. Identity [DLZ^{+}24]. Idler [ZYXD20]. IEC [SCL^{+}24]. IEEE [BCCM23, CQ22, FAKM22, IIEKS24, Ano20a, Ano23, BBJR21, CDP21, Lu21, WS20]. IID [ZGQ^{+}22]. IIoT [TDMP23]. Illuminating [ZCX^{+}23]. Illustrating [SXC^{+}24]. Image [BBC^{+}20, BBC^{+}22, CKJ^{+}22, DZC^{+}24, FLS20, HCC^{+}25, HFT^{+}25, KLC20, ZLT^{+}24]. Images [HZF^{+}24, LMM^{+}22, ZCX^{+}23]. Imaging [GWCS23, WZD^{+}20, YCS^{+}24]. Impact [BBC^{+}22, ESN20, SZK^{+}22, CHZ^{+}25]. Impeccable [AMR^{+}20]. Implementation [ABC^{+}24, BN24, BMLOM20, CQCL25, CLCL22, CZW^{+}25, JYM20, LAKS20, RBMG22, SMY22, TWL^{+}22, XLWO23, ZCWC23]. Implementations [CLCC25, DVV23, MYGA20, PN25, TGS^{+}22, WDC^{+}25]. Implemented [LT25]. Implementing [APK20]. Implications [KLP^{+}21]. Implicit [BSRP21, BCKS22, LWZZ25]. Importance [CHZ^{+}25]. Importance-Informed [CHZ^{+}25]. Imprecise [JYM20, JDB^{+}23, RSA^{+}20, ZZ25]. Improve [JLZ^{+}23, PPQBA21, ZCJ^{+}20]. Improved [ACH21, CZW^{+}25, MDR^{+}24, NP20, ZCZ^{+}22]. Improving [BSRP21, BB20, GZW^{+}21, HKC21, HXL^{+}25, HLF^{+}23, JWS^{+}21, KBR^{+}23, LWL^{+}22, LLL^{+}23, LR22, OAC^{+}21, PLZ^{+}23, PM20, WTL^{+}21, WFT^{+}21, WZGT22, WSHJ23, YH20, ZCC^{+}23, ZDC^{+}25]. IMR [LYC22]. IMR-Based [LYC22]. in-Cache [SQR^{+}20]. In-Container [WDZ^{+}24]. In-Line [LJY^{+}24]. In-Memory [BTEC20, CSY^{+}25, CXY24, DSP^{+}21, ESW^{+}23, GWC^{+}25, KSL^{+}22, KAWR23, KBQ^{+}23, KYS^{+}22, LHX^{+}25, LQM^{+}24, LSXZ21, MPYJ25, TGA23, WYZ^{+}22, YWC^{+}24b, ZCWC23, CCT^{+}20, WLW^{+}22b, WLF^{+}25b]. In-Memory-Computing [ZLL^{+}24b]. In-Module [LLS^{+}23]. In-Network [LLD^{+}25, NWL^{+}25]. In-Place [BCKS22]. In-ReRAM [KSKK23]. In-Situ [KJK24, LY20, LZW25]. Incendio [CSH^{+}24]. Incentive [FHL^{+}23, HZYY22, WRW^{+}23].

Incentivize [YDL^{+}25]. Incentivizing [WZH^{+}23, ZWWY22]. Inclusions [GJ20]. Inclusive [ZCR23]. Incomplete [YGW^{+}23]. Incremental [AB20, HXGR20, LY21, PYW^{+}22]. Independence [KS24]. Independent [DWLF25, WCQW22, ZGG^{+}23]. Index [Ano20a, KOT^{+}23]. Indexing [WCD25]. Indirect [CCZ^{+}22]. Indistinguishability [ZC24]. Induced [HDAS21, LWNC22, MSZ22, PAR^{+}22, VTM^{+}20]. Industrial [BY24, CDF^{+}21, YZY^{+}25, FHL^{+}23]. Industry [BGM^{+}23]. Inexact [AVK20]. Inference [ALC25, BCRX23, BBD^{+}20, CXH^{+}25, CXL^{+}23, CSK22, CWC^{+}24, CZW^{+}24, DHM^{+}25, DSK23, DA22, FRFM^{+}25, GLZ^{+}24, HLY^{+}25, JLD^{+}25, JKHL22, LMDC21, LHX^{+}25, LTJS^{+}22, LWYJ23, MDM22, PCMP21, WDCC20, WZX^{+}22, WDL^{+}25, WGD^{+}22, ZCK20, ZXG^{+}24, ZHY^{+}25, ZWC^{+}22]. Inference-Aware [CSK22]. Inference-Based [BBD^{+}20]. Influence [WFH^{+}24, YDG^{+}24]. Information [CCC23, xHzLH^{+}24, LHR^{+}23, PYW^{+}22, XKS21, Xu24, YHC^{+}20, YDL^{+}25, ZXL^{+}23, ZSX^{+}24, ZGK20]. Informed [CHZ^{+}25]. Infrastructure [SRB23]. Inheritance [FL21]. Initiative [NCD^{+}25]. Initio [WMG^{+}25]. Injection [RRMS25, VTM^{+}20]. Inner [PN24, SSCK25]. Inner-Outer-Hybrid [SSCK25]. Inner-Product [PN24]. Innovative [AAM^{+}25]. Inodes [CSW^{+}21]. Input [ESN20, LWW^{+}24, MLW^{+}25, PCMP21, PHC24, RRDB20, SLY^{+}22b, ZMS^{+}23]. Input-Aware [SLY^{+}22b]. Input-Dependent [PCMP21, ZMS^{+}23]. Input-Skippable [LWW^{+}24]. Inputs [CR24]. Insertion [LCH22, MAM23, ZXX^{+}25]. Insertion/Deletion [LCH22]. Insights [DTZ^{+}25, WZF^{+}24]. Insignificant [WWM^{+}23]. Inspection [TRBM22]. Inspired [GA22, HXL^{+}23, LWW^{+}25, TGA23, WZX^{+}22]. Instabilities [CYKG23]. Instance [KMVD22, LWL^{+}25, ZYL^{+}22]. Instances [ATT22, SZS^{+}22]. Instruction [AGB^{+}23, CKK^{+}22, HMJ24, KHHK21, KLR^{+}20, MHK^{+}22, OJ23, RJ24, TWY^{+}25, UYZP22, WGM^{+}20, WZCM23, WWX^{+}24]. Instruction-Level [TWY^{+}25, WZCM23]. Instructions [CWS^{+}24b, WGJZ21, WLW^{+}22c]. Insulation [GJN^{+}25]. Integer [ABC^{+}24, KPH^{+}25, WFH^{+}24]. Integers [DVA22, MÖS22, ZSS20]. Integrated [JYF^{+}23, RCS^{+}21, SM24, ZXG^{+}24]. Integration [KBQ^{+}23, MRA^{+}21, NAP^{+}20, SPMP20, WFZ^{+}25]. Integrity [FHL^{+}22, PL21, SHZ^{+}25, TWaKo^{+}23, ZSC^{+}23, ZQC^{+}25, ZXX^{+}25]. Intel [CDF^{+}21, JQY^{+}25, WZSL22, XTWG23]. Intelligence [NHW^{+}24, NLC^{+}25]. Intelligence-Assisted [NLC^{+}25]. Intelligent [GWH^{+}23, JMW^{+}24, LYF^{+}22, OAC^{+}21, WLW^{+}21, WLHW25]. Intensity [HLJ^{+}25, STYQ24]. Intensity-Aware [STYQ24]. Intensive [DPQK^{+}23, PB23b, ZSHB21]. Inter [HS22, LHK^{+}22, LPW20, MZZC22, MYUK21, PCBD23, WWJ^{+}25, ZML^{+}25]. Inter- [WWJ^{+}25]. Inter-Core [PCBD23]. Inter-FPGA [MYUK21]. Inter-Layer [ZML^{+}25]. Inter-Partition [HS22]. Inter-Process [MZZC22]. Inter-Processor [LPW20]. Inter-Warp [LHK^{+}22]. Interaction [DGZ^{+}22]. Interactive [QCX^{+}23, WDQ^{+}22]. Interconnect [JYF^{+}23]. Interconnects [TBS^{+}25]. Interface [BB20]. Interference [WDZ^{+}22, WZGT22]. Interlaced [WLZ^{+}21]. Interlacing [ZSWS24]. Interlacing-Uncompute [ZSWS24]. Intermediate [WRT^{+}22]. Internal [YWC^{+}21]. Internet [FHL^{+}23, DRA21, GKT^{+}22, HC24, LQL^{+}24, LHR^{+}23, QWT^{+}23, SYL^{+}23]. Internet-of-Things [DRA21, LHR^{+}23]. Interoperability [DTL^{+}25]. Interoperable [TDMP23]. Interpolation [APK20, HCKK25]. Interrupts [DYPZ22]. Intersection [ZYZ^{+}25a]. Interval [GJ20]. Intervals [BWSG25]. Intra [IKTY22, WWJ^{+}25, ZCCG23, ZML^{+}25]. Intra- [ZML^{+}25]. Intra-Array [IKTY22]. Intra-chiplet [WWJ^{+}25]. Intra-SM [ZCCG23]. Intricacies [GWD25]. Introduction [AW20, Kar24, QWK20]. Intrusive [BXW^{+}25]. Invariance [HJX^{+}25]. Invariant [BBD^{+}20]. Invariants [KMH^{+}23]. Inverse [CB22, RMTA20]. Inversion [GMZ22, Koç20]. Invertible [MSZ22]. Investigating [SSM23]. Invisible [TRBM22]. Invocation [LJH^{+}25]. Invocation-Based [LJH^{+}25]. IoMT [CXL^{+}25]. IOPS [SSCK25]. IoT [ABP22, BBJR21, BCBS21, BGB^{+}21, CWS^{+}24b, GQJ^{+}22, HZMC24, LD22, LFGD25, LAKS20, LYW^{+}23, LZX^{+}25, QSC^{+}25, VAV^{+}20, WLW^{+}21, YYQ^{+}24, YZY^{+}25, ZSS^{+}22, ZLS^{+}24, ZLW^{+}24, ZWZZ24, ZHM20]. IOV [ZXL^{+}24]. IP [RGS22]. Irreducible [Ima21]. Irregular [DGQ^{+}25]. Irregularities [GSY^{+}20]. Irregularity [ZZZ^{+}20]. ISA [ABP22, GFB^{+}24, JDH^{+}25, KGHRM23, SZHB21]. Ising [OTTT22, ST23a, ST23b, SEM23]. Islands [CFWC23]. Isogeny [SAJA21]. Isolation [HWR^{+}24, WHC20, YYW^{+}24].

Isomorphism [LZF21]. ISPA [ZCCG23]. ISSA [LWW^{+}24]. Issue [Ano23, AW20, BBJR21, BCCM23, CQ22, FAKM22, Lu21, QWK20, SZHB21, WS20, ZSHB21]. Iterated [JXH^{+}22]. Iterative [Bru23, LMDC21, RRMS25, TOM23, ZCP23]. Iteratively [WYSL22]. IVP [GWH^{+}23].

Jacobi [YGW^{+}23]. Java [WLC^{+}24]. JBNN [FHW^{+}22]. Jitter [BPM23]. Job [CPL^{+}23, IMQOP21, LLD^{+}25]. Jobs [BLP^{+}22, GYS^{+}24, IWKB22, WZH^{+}23, YWX^{+}23, ZCC^{+}23]. Join [WSG^{+}25]. Joint [DYC^{+}25, MSLY24, MHJ^{+}21, MRB^{+}24, WLD^{+}25, WCYK20, ZYZ^{+}23]. JointPS [ZYZ^{+}23]. Journal [Lou20]. Journaling [CWWW20]. JPEG [DZC^{+}24]. Juliet [GMT24].

KaratSaber [WWL^{+}23]. Karatsuba [PN25, WWL^{+}23]. Karnaugh [VJWZ^{+}21]. KDN [YHW^{+}25]. KDN-Based [YHW^{+}25]. Kernel [GSS^{+}23, LKK^{+}21, LCHL21, SXC^{+}24, WDL^{+}25, WLF^{+}25b, ZDC^{+}25, dOCC23]. Kernels [LPYT22, TLC^{+}24]. Key [AAM^{+}25, AMJ^{+}23, BL22, CJSY24, CSY^{+}25, CLY22, EDGR^{+}24, GJN^{+}25, HBS^{+}20b, LLWZ23, LYH^{+}24, LDF^{+}24, PD21, SAJA21, WLW^{+}25, YZJ23, ZGD23, ZXW^{+}24, ZLW^{+}24, ZWZZ24, ZOH^{+}25, ZWO^{+}25]. Key-Exchange [AMJ^{+}23]. Key-Value [CLY22, ZGD23]. Keys [DVA22]. Keyword [DWL^{+}22, JQY^{+}25, LJZ^{+}25, LMM^{+}23]. Kit [TSM^{+}21, DWW25]. KnightSim [GPH20]. Knob [SMFS21]. Knobs [OAC^{+}21]. Knowledge [BHW^{+}23, JMW^{+}24, QCX^{+}23, YLZ^{+}24, YPD^{+}24, ZGG^{+}23]. KPI [SCY^{+}23]. KPU [WLF^{+}25b].

Kreyvium [RBM21]. Kubernetes [CB22, MXY^{+}23]. Kutta [BFC20]. KV [SJYQ25]. KVSTL [CLY22]. Kyber [CLCC25, DMG23, GL24, NKeSK^{+}23, XPR^{+}22].

L1 [RCAB23, VLPS25]. L4L [ZLWG22]. LAC [STYQ24]. Ladder [NS22]. Lagrange [ZLM^{+}24]. LAMP [LLL^{+}23]. Land [WZW^{+}24]. Language [KPS^{+}24]. Large [ABI^{+}25, FTV25, HYQ^{+}25, IBB^{+}25, KPD^{+}23, LZC^{+}21, LDG^{+}22, LWH^{+}24, LCY^{+}25, LWL^{+}21, LMM^{+}23, LTJS^{+}22, MÖS22, MÁJG^{+}24, PLH^{+}24, SYW^{+}22, SCY^{+}23, WCD25, WAN^{+}25, WWL^{+}25, YCY^{+}24, ZCJ^{+}20, ZLL^{+}23, ZCSJ23]. Large-Scale [HYQ^{+}25, KPD^{+}23, LZC^{+}21, LDG^{+}22, LWH^{+}24, LCY^{+}25, SYW^{+}22, SCY^{+}23, WCD25, WAN^{+}25, WWL^{+}25, YCY^{+}24, ZCJ^{+}20, ZCSJ23]. Laser [VTM^{+}20]. Laser-Induced [VTM^{+}20]. LAShards [ZTZ^{+}25]. Last [HLLC21]. Last-Level [HLLC21]. Latch [YHC^{+}20]. Latching [LW22]. Latching-Delay [LW22]. Latencies [LCHK22]. Latency [AG24, ALC25, ABC^{+}24, Bru20, DHM^{+}25, DLW^{+}25, GLZ^{+}24, HMZ^{+}25, MÖS22, NHW^{+}24, BCLC25, PM20, PVB21, RCAB23, WZX^{+}22, WHL^{+}23, WDZ^{+}23, XLZ^{+}25, ZQG^{+}24, ZCP22, ZHY^{+}25, ZLL^{+}22b, ZTLW23]. Latency-Sensitive [ZLL^{+}22b]. Lattice [AMJ^{+}23, DVV23, HP23, KAA22, PYS^{+}23, TWZ^{+}23, XPR^{+}22, XCG^{+}25, ZBT22]. Lattice-Based [AMJ^{+}23, DVV23, HP23, PYS^{+}23, TWZ^{+}23, XPR^{+}22, XCG^{+}25, ZBT22]. Lattices [AA20]. Launch [LOM^{+}25]. LAWS [YAG20]. Layer [CLY22, KOT^{+}23, SMC^{+}25, WLY^{+}23, ZXG^{+}24, ZML^{+}25]. Layer-Fused [SMC^{+}25]. Layer-Wise [ZXG^{+}24].

Layered [SHZ^{+}24]. Layers [PAR^{+}22]. Layerwise [GLZ^{+}24]. Layout [HKC^{+}23, HYS^{+}20, JJKP22, TC21, XLL^{+}22]. Layout-Aware [HYS^{+}20]. Layouts [BCKS22]. LDPC [HBS^{+}20b, LSCX20]. Leader [LZW^{+}24b]. Leaderless [LLR25]. LeafHooks [ZXX^{+}25]. Leakage [CCC23, DH20, HLC^{+}22, LHY^{+}21, SSP^{+}24, XPR^{+}22]. Leakage-Free [LHY^{+}21]. Leaking [CY22, XKS21]. Leaks [OD23]. Learnability [CHM25]. Learnable [YWC^{+}24a]. Learned [BL22]. Learning [AAM^{+}25, APV22, BWL^{+}25, BJMKK23, BL22, CYPC25, CCCW21, CPL^{+}23, CWY^{+}23, CWS^{+}24a, CDW^{+}25, CLCC25, CZW^{+}24, CAC^{+}22, DGG^{+}22, DGZ^{+}22, DWW25, FZG^{+}22, FAKM22, FTV25, GLMZ25, GWCS23, GGZC22, GKT^{+}22, GYS^{+}24, GZG^{+}23, HLL^{+}20, HHPB20, HWZ^{+}25, HYH^{+}25, HCC^{+}23, HLZ^{+}25, HF22, HZM^{+}23, HLT^{+}23, JKK^{+}22, JHMM23, KZXR25, KH23, KAWR23, KBQ^{+}23, KASAG23, KBR^{+}23, LZW^{+}21, LWL^{+}22, LZY^{+}25, LJH^{+}25, LSL^{+}25, LYGC24, LLZ^{+}25, LZW^{+}23b, LZG^{+}24, LGZ^{+}25, LHZR25, LY21, LZC^{+}24, LDW^{+}25, LFP^{+}22, LCC^{+}24, MZM^{+}25, MSP^{+}21, MHM^{+}23, MS25, MWJ^{+}24, MLW^{+}25, NNH^{+}25, PBBA25, PSM22, PZY^{+}23, PM25, PYW^{+}22, PK23, QWK20, QJY^{+}25, RAD20, RWCC23, RSA^{+}20, RSR22, SZAT22, SMZ^{+}20, STZ^{+}24, SSW^{+}24, SZ22, SM24, SSZ^{+}20, STQ^{+}24, SST^{+}24, SKM^{+}23, TZY^{+}24, TCX^{+}23, TTG^{+}23, TKM20, VAV^{+}20, WDCC20, WHM^{+}22, WCB23, WLZ^{+}23, WZH^{+}23, WRW^{+}23, WJLC24, WTL^{+}24, WLD^{+}25, WGD^{+}22, WJL^{+}20, WHL^{+}21, WLHW25, XLW^{+}20, XZC^{+}23, XWL^{+}24, XJY^{+}24]. Learning [YZX^{+}24, YWX^{+}23, YPD^{+}24, YPL^{+}25, YYZ^{+}25, YCY^{+}24, YDL^{+}25, YBF^{+}25, ZLL^{+}24a, ZGLZ20, ZLWG22, ZTY^{+}23, ZGQ^{+}22, ZGG^{+}23, ZXY^{+}24, ZNW^{+}24, ZMH^{+}25, ZXMX25, ZFD^{+}20, ZYZ^{+}23, ZLZ24b, ZWL^{+}25, ZLL^{+}22b, ZWSF24, ZHYJ21, ZWC^{+}22, ZSX^{+}24]. Learning-Assisted [GLMZ25]. Learning-Augmented [ZLZ24b]. Learning-Based [AAM^{+}25, BL22, CYPC25, CLCC25, DWW25, GGZC22, HWZ^{+}25, HZM^{+}23, KASAG23, LZW^{+}21, LWL^{+}22, PM25, SM24, ZWSF24]. Learning-Empowered [SST^{+}24]. Learning-Guided [TTG^{+}23]. Ledger [CYX^{+}23]. LedgerMaze [BHW^{+}23]. Length [FLS20, MTK25, WYX^{+}24]. Less [LLWZ23]. LET [BPM23, PCBD23]. Level [ABI^{+}25, ABdGG^{+}25, AKG^{+}20, BSM21, CFC^{+}22, DGTVGG21, DSTD22, FFG^{+}25, HCKK25, HZC^{+}25, HLLC21, JDCL23, KKS^{+}22, KLKK23, KASAG23, LL22, LLL^{+}23, LLL^{+}20, LZW^{+}24a, MLL^{+}25, MCS^{+}22, NAP^{+}20, NM22, OLC^{+}22, PE22, PSBB21, SCFPM22, SZL^{+}24, TWY^{+}25, VCLN21, WFW^{+}20, WWM^{+}23, WZCM23, WFZ^{+}25, XYM23, YH20, YDL^{+}25, ZXL^{+}24, ZLS^{+}24, ZWWY22]. Leveling [CSW^{+}21, GLW^{+}24, NK22]. Leveraging [GSK^{+}22, HLK^{+}25, PVB21, RCS^{+}21, SZL^{+}22, SYD^{+}24, WFT^{+}21, WGT^{+}22, XZL^{+}21]. LFOC [SCFPM22]. LFSR [BHK^{+}23, Ima21]. LFSR-Based [Ima21]. Libraries [CGS^{+}20]. Library [BWL^{+}25, CTY^{+}24, CBB^{+}21b, TGS^{+}22, YWF^{+}25, ZTT22]. Lieu [Das23]. Life [ZCR22]. Lifetime [HHPB20, LLC^{+}24, PLZ^{+}23, RCS^{+}21, YWC^{+}21]. Lifetime-Aware [HHPB20]. Lifetime-Retention [YWC^{+}21]. Light [YCS^{+}24, ZSC^{+}25, WSC^{+}25].

Lightening [WSG^{+}25]. LightWarner [PZY^{+}23]. Lightweight [BOL^{+}25, BBC^{+}22, CZH^{+}25, CKJ^{+}22, EEA22, HWX^{+}24, HPGM20, HBS^{+}20b, HDL^{+}24, JHMM23, KLP^{+}21, LWX^{+}25, LYW^{+}23, LV23, LZX^{+}25, MYUK21, OLZ^{+}20, PLB22, PD21, PLZ20, RBC^{+}23, SZHB21, WSC^{+}25, ZSS^{+}22, ZLS^{+}24, GWX^{+}23]. Like [AG24, LYC22, QSC^{+}25, WLZ^{+}25]. Lime [PYW^{+}22]. Limited [LZF21, WZD^{+}20, WYB^{+}24]. Limon [YZJ23]. Line [LWL^{+}23, LJY^{+}24, WCQW22]. Line-Graph-Based [WCQW22]. Linear [HKC^{+}23, HCKK25, ST23b]. Linearity [SLDZ25]. Linearly [CQI^{+}22]. Links [STW^{+}21]. Linux [LCHL21, BCLC25, dOCC23]. List [Ano20b, Ano25]. Little [JKKP25]. Little-Storing [JKKP25]. Live [AY24, SXC^{+}24, TRBM22]. LLM [DHM^{+}25, HLY^{+}25, WFZ^{+}25]. LMChain [HWL^{+}24]. LNS [ZDV^{+}22]. LNS-Madam [ZDV^{+}22]. Load [CBB21a, GZW^{+}21, HWL^{+}24, TARK23, TDZ^{+}22]. Load-Balancing [TDZ^{+}22]. Load-Migratable [HWL^{+}24]. Loading [LZC^{+}24]. Local [KLC20, LDW^{+}25, LCGH25, NTPAB^{+}25, RSR22, SXZJ24, TQL^{+}22]. Locality [CWWW20, JJKP22, WDZ^{+}22, YAG20]. Locality-Aware [CWWW20, YAG20]. Localization [LHL^{+}23, PSM22, XSYL22]. Localizing [HA25]. Locally [FXL^{+}24]. Location [ZC24]. Lock [DQ23, TZZ^{+}21, ZWJ^{+}25]. Lock-Based [DQ23]. Lock-Free [ZWJ^{+}25]. Locking [ZG23]. Locks [JGD^{+}21, JCY^{+}23]. Log [HLF^{+}23, QLH^{+}24]. Log-Based [HLF^{+}23]. Logarithmic [ACH21, APK20, LLL25b, ZDV^{+}22].

Logging [BCF^{+}25, HKS20, LPYT22]. Logic [AA20, ABI^{+}25, BCV22, CWNL22, FFG^{+}25, LB22, VJWZ^{+}21, WWL^{+}25, XSYL22, ZG23]. Logical [ZGWY22]. LogSay [QLH^{+}24]. Long [CMQ^{+}22, WYX^{+}24]. Longest [RGS22]. Longevity [CCYC22, RCS^{+}21]. Look [Das23, SMP22, AB22]. Look-Up [Das23]. LookAside [DSK23]. Loop [DGQ^{+}25]. Loss [GCL^{+}21, LLM^{+}25]. Lossless [DZC^{+}24]. Lossy [WCL^{+}23]. Low [ABP22, ALC25, AAB^{+}23, BCBS21, Bru20, BGB^{+}21, CMQ^{+}22, DLW^{+}25, DDK22, FLS20, HMZ^{+}25, JKHL22, JBS^{+}25, KNP^{+}20, LLW^{+}24, MÖS22, OKU^{+}24, PYW^{+}22, PVB21, RMTA20, RCAB23, RSA^{+}20, TWY^{+}25, TDZ^{+}25, WHX^{+}25, WHL^{+}21, WDZ^{+}23, XLZ^{+}25, ZQY^{+}20, ZLC^{+}22, ZCP22, ZCP23, ZDV^{+}22, ZTZ^{+}25, ZTLW23]. Low-Area [RMTA20]. Low-Complexity [KNP^{+}20]. Low-Cost [AAB^{+}23, BGB^{+}21, FLS20, JKHL22, PYW^{+}22]. Low-Latency [DLW^{+}25, HMZ^{+}25, MÖS22, PVB21, RCAB23, ZTLW23]. Low-Overhead [DDK22, ZLC^{+}22, ZTZ^{+}25]. Low-Power [BCBS21, JBS^{+}25, OKU^{+}24, ZQY^{+}20]. Low-Precision [ZDV^{+}22]. Low-Rate [TDZ^{+}25]. Lower [DLG^{+}24]. LPAH [SXC^{+}24]. LPC [FMM^{+}21]. LrGAN [MSSL21]. LRU [XKS21]. Ls [WSG^{+}25]. Ls-Stream [WSG^{+}25]. LSDedup [SHZ^{+}24]. LSM [CJSY24, CLY22, SJYQ25, ZWO^{+}25]. LSM-Tree [CLY22]. LSM-Tree-Based [CJSY24, SJYQ25]. LSM-Trees [ZWO^{+}25]. LSTM [DGG^{+}22, GYH^{+}22, MC23]. LSTMs [DYJ20]. LTI [VHL20]. LUNA [DSCB25]. LUNA-CiM [DSCB25].

LUT [ESN20, HCKK25]. LUT-Based [HCKK25]. LUT6 [XL25]. LUT6-Based [XL25]. LUTNet [WDCC20]. LUTs [ESN20]. LWE [LAKS20, XLWO23].

M4 [SAJA21]. MAB [ZWWY22]. MAC [JKKP25, JYM20, SNT22]. MAC-Based [JYM20]. Machine [APV22, BJMKK23, CYPC25, CZD^{+}24, CWC^{+}24, CAC^{+}22, DNMS20, DWW25, FAKM22, GKT^{+}22, HCC^{+}25, HF22, JJKP22, KAWR23, KASAG23, LLS^{+}24, LYF^{+}22, LLR25, PSM22, PM25, PK23, QWK20, RWCC23, SZAT22, SMZ^{+}20, ST23a, SEM23, SZS^{+}22, SSZ^{+}20, SST^{+}24, TTG^{+}23, WLZ^{+}23, WZH^{+}23, WJLC24, WGD^{+}22, XNLX20, ZFD^{+}20, ZYZ^{+}23]. Machines [AB20, AY24, LSH25, MIPQ22, SSJ21, ST23b, ZWC^{+}23]. Madam [ZDV^{+}22]. MAGIC [LYC22]. Magnetic [WLZ^{+}21, ZYXD20]. Magnifying [XPR^{+}22]. Main [NK22]. Maintenance [LYC^{+}23]. Making [LYC22, RRDB20]. MalFox [ZCY^{+}24]. Malicious [LHXH22]. Malleability [IMQOP21]. Malware [EEA22, Fic22, IKAG^{+}22, PSM22, SNA^{+}20, TZY^{+}24, VAV^{+}20, YWP^{+}23, ZCX^{+}23, ZCY^{+}24]. Malware-on-the-Brain [ZCX^{+}23]. Man [MXY^{+}23]. Man-in-The-Middle [MXY^{+}23]. MANA [AGB^{+}23]. Managed [KCAL21, SXXL24]. Management [AAM^{+}25, AhRX^{+}20, BLKK23, BBC^{+}20, CCYC22, CLY22, GGZC22, HMMP23, HCC^{+}23, HIRB25, HZF^{+}24, HTZ^{+}25, IMQOP21, KKH22, KKKC20, LCX21, LHR^{+}22, LHN^{+}22, NTDH25, PYYG21, RMKO23, SRB23, SMFS21, SPDQ22, SST^{+}24, TCJ23, WCYK20, WLZ^{+}21, YH20, ZZ25, ZCC^{+}23, ZCCG23]. Manager [WTL^{+}21]. Mangrove [BBD^{+}20].

Manipulation [ZXZ^{+}21]. Many [AhRX^{+}20, JWD^{+}22, JYM^{+}23, LQM^{+}24, Lu21, MB21, PFHD21, RSP^{+}20, RPMH21, SRP^{+}21, WWJ^{+}23, WWJ^{+}25, WWS^{+}22, XNH^{+}25, YGW^{+}23, ZYL^{+}22, ZLC^{+}23b]. Many-Core [AhRX^{+}20, JWD^{+}22, JYM^{+}23, LQM^{+}24, LAPB21, MB21, SRP^{+}21, WWJ^{+}25, WWS^{+}22, YGW^{+}23, ZYL^{+}22, Lu21, WWJ^{+}23]. Many-Cores [PFHD21, RSP^{+}20, RPMH21]. Manycore [LLJ^{+}23, RCS^{+}21, RCAB23]. Map [LTFL22, LTL^{+}25, VJWZ^{+}21]. Mapped [DTH^{+}24]. Mapping [BTEC20, BAM^{+}24, DRP24, ESdP^{+}25, GWG^{+}25, HGK^{+}22, HSH^{+}25, KH23, LZF21, LLL^{+}23, LL23, LAPB21, LCGH25, MHJ^{+}21, RSP^{+}20, RCS^{+}21, WSLX24, WWJ^{+}25, YDG^{+}24, ZTT22, ZDZ^{+}23, ZLZ24a, ZDW^{+}23]. Mapping-Based [LLL^{+}23]. MapReduce [AB20]. Maps [LT25]. MarCNNet [YWP^{+}23]. Market [CYX^{+}23]. Marketing [LHN^{+}22]. Markovian [YWP^{+}23]. Masked [DVV23]. Masking [CMGD24]. Massive [HZMC24]. Massively [vSDHA23]. Masta [WT25]. Matching [GWX^{+}23, RGS22]. Matrices [JWK^{+}23, YH24]. Matricesin [Das23]. Matrix [ESW^{+}23, GLB21, GZW^{+}21, KCL^{+}20, KJK^{+}22, LRB23, LWZZ25, MCT22, PN25, SXZJ24, TGS^{+}22, TAP^{+}25, WDL^{+}25, ZLL^{+}24b, ZZW^{+}25]. Maximal [PLH^{+}24, PHL^{+}25]. Maximization [LGL^{+}24, YHV^{+}21, ZLT^{+}25]. Maximizing [FDKK21, HLY^{+}25, KLKK23, WJL^{+}20, YHW^{+}25]. Maximum [LLC^{+}24, LWH20]. Maximum/Minimum [LWH20]. MC [LL22]. MC-FLEX [LL22]. MCS [WLHW25]. MCU [BCF^{+}25, LZG^{+}24].

MCUs [BGB^{+}21, JBS^{+}25]. MDC [CZW^{+}25, MPYJ25]. MDC-NTT [CZW^{+}25]. MDev [PYDG22]. MDev-NVMe [PYDG22]. MDTUpdate [ZFH23]. Mean [JYM20]. Mean-Error [JYM20]. Measurement [KB21, LHW^{+}21, NML25]. Measurements [BPJ^{+}22, BCLC25, WLZ^{+}23]. Measures [Xu24]. MEC [CXL^{+}25, XZL^{+}23]. MEC-Enabled [CXL^{+}25]. Mechanism [CSW^{+}21, KCS23, NHW^{+}24, PYW^{+}25, SKK^{+}21, WRW^{+}23, ZCJ^{+}20, ZCD^{+}22, ZSC^{+}23, ZFQ^{+}23, ZWB^{+}22, ZWWY22]. Mechanisms [GWD25, HMMP23, LPYT22]. Media [SWR^{+}25]. Mediated [PYDG22]. Medical [LHW^{+}25]. Meets [BBdTF25]. MegaKernels [JLL^{+}20]. Membership [RSMMG^{+}23]. Memcached [CRJZ21, JLP^{+}24, JLP^{+}24]. Memories [FDKK21, FW23, FMM^{+}21, KSL^{+}22, KOH^{+}23, NK22, OLD^{+}23, Rot24]. Memory [BLKK23, BHE21, BY22, BB20, BB22, BTEC20, CCT^{+}20, CWT^{+}22, CSY^{+}25, CL20, CZB^{+}22, CWWW20, CSW^{+}21, CCYC22, CXY24, CHC25, CFC^{+}22, CJYC23, DHM^{+}25, DSK23, DGTVGG21, DSP^{+}21, DSCB25, DPS^{+}20, DTH^{+}24, DLK25, ESW^{+}23, EDGR^{+}24, FCZ^{+}23, GWG^{+}24, GWC^{+}25, GvSHA22, GR23, GWX^{+}23, HZT^{+}23, HKC^{+}23, HHPB20, HPJK22, HCC^{+}25, HCKK25, HCC^{+}23, HWC^{+}22a, HWZ^{+}22, JPHY20, JLY^{+}21, JLP^{+}24, JWS^{+}23, KKH22, KSL^{+}22, KAWR23, KMH^{+}23, KKL^{+}25, KBQ^{+}23, KYS^{+}22, KIY21, LB22, LHK^{+}22, LLS^{+}23, LLS^{+}25, LSCX20, LCX21, LDLK22, LLW^{+}24, LHX^{+}25, LCM25, LLWZ23, LQM^{+}24, LLL^{+}20, LJY^{+}24, LWW^{+}24, LKMJ21, LY20, LY21, LLY22, LSXZ21, MCD^{+}25, MPYJ25, MHK^{+}22, NICY24, NTL^{+}24, PLZ^{+}23, PZY^{+}23, PQG^{+}22, PPQBA21, PB23b, QCX^{+}23, RGS22, RPB^{+}23, RCAB23, SAG22, SMZ^{+}20, SMY22, SPH^{+}23, SWR^{+}23, SXXL24, TSM^{+}21, TGA23, WTL^{+}21, WLW^{+}22a, WBJC22, WLW^{+}22b, WYZ^{+}22, WNL^{+}23, WMH^{+}24, WYX^{+}24, WHX^{+}25, WLW^{+}25, WCYK20, WDZ^{+}23]. Memory [WHK24, WSC^{+}24, WLF^{+}25b, XLY^{+}22, YZG^{+}25, YLG^{+}23, YLL^{+}24, YQX^{+}25, YH20, YWC^{+}24b, ZHLR22, ZCWC23, ZLL^{+}24b, ZCY^{+}25, ZWY^{+}23, ZYZ^{+}25b, ZCSJ23, ddAPdS21]. Memory-Aware [BY22]. Memory-Based [NICY24]. Memory-Centric [DGTVGG21]. Memory-Constrained [SWR^{+}23, ZHLR22]. Memory-Efficient [GWX^{+}23, MPYJ25, WYX^{+}24]. Memory-Enhanced [SMY22]. Memory-Free [GR23]. Memory-Mapped [DTH^{+}24]. Memory-Optimality [CWT^{+}22]. MemPool [RCAB23]. Memristive [SMY22, VJWZ^{+}21]. Memristor [HJX^{+}25, JWK^{+}23, LFW21, YWC^{+}24b]. Memristor-Based [JWK^{+}23, LFW21, YWC^{+}24b]. MemUnison [WLW^{+}22b]. Merged [ZZG20]. Merger [PLB22]. Merging [LWL^{+}25, SKH^{+}25, WJL^{+}20]. Merkle [ZCY^{+}25]. Mesh [PC24]. Message [JZX^{+}25, TZZ^{+}21]. Messages [ZLC^{+}23a]. Meta [RSR22, VCLN21, WLY^{+}23]. Meta-Block [WLY^{+}23]. Meta-Learning [RSR22]. Meta-Programming [VCLN21]. Metadata [GGZC22, LHR^{+}22, LCJ^{+}25, WNP^{+}22]. Metal [TLC^{+}24, ZXL^{+}24]. Metastability [BLM20]. Metastability-Containing [BLM20]. Metaverse [ZOH^{+}25].

Method

[ATT22, JCKH22, LWNC22, LLJ^{+}23, LCJ^{+}25, LMH^{+}25, QCX^{+}23, SKR^{+}20, SXZJ24, ST23b, VJWZ^{+}21, WL24, WHX^{+}25, XL25, YDG^{+}24, ZCWC23]. Methodology [ABdGG^{+}25, DSTD22, GPH20, LLL25b, RRMS25, SDR^{+}22]. Methods [AA20, BFC20, LMDC21, LWC^{+}22, MLL^{+}25, MKÖ^{+}22, RWCC23]. Metric [ZSX^{+}24]. MGARD [LWC^{+}22]. MGen [KSKK23]. Microarchitecting [AGB^{+}23]. Microarchitectural [PG23, XTWW25]. Microarchitecture [GWD25, KASAG23, LDLK22, ODK20]. Microarchitecture-Level [KASAG23]. Microarchitectures [DSJ^{+}22, YYCR24]. Microcontrollers [LHS^{+}25]. Microfluidic [HGC^{+}22]. Microprocessors [BPJ^{+}22, FAFK21, WLQ^{+}21]. Microservice [HYQ^{+}25, LHW^{+}25, WLD^{+}25, ZQG^{+}24]. Middle [MXY^{+}23]. Middleboxes [LCS^{+}25]. Migratability [ZXL^{+}24]. Migratable [HWL^{+}24]. Migration [AY24, GLW^{+}24, HHPB20, HWC^{+}22a, LHA^{+}25, LLL^{+}20, OKC^{+}20, RPMH21, WWS^{+}22]. Migrations [YWC^{+}21]. Million [WMG^{+}25]. Min [RMO21]. Minded [KMAA25]. Minimal [GCL^{+}21, SSJ21, YZX^{+}24]. Minimization [CCZ^{+}22, JXH^{+}22, KZS^{+}25]. Minimizing [ACKA23, HPJK22, WJL^{+}20, YWC^{+}21, YBW21]. Minimum [Akr22, LWH20, ZGR^{+}25]. Mining [BBD^{+}20, HLS^{+}23a]. Miniservers [RCC^{+}25]. MINOTAuR [GCR^{+}23]. MIPSGPU [YBW21]. Mirroring [CHL^{+}23]. Miss [DKJP21, JZSD24]. Misuses [GSS^{+}23]. Mitigate [QZZ^{+}24]. Mitigating [FMM^{+}21, GSS^{+}23, GSY^{+}20, KB21, LLS^{+}23, LSL^{+}25, WHQ^{+}24]. Mitigation [AGSD25, BY24, SKK23, TDZ^{+}25]. Mitigations [XTWW25]. Mix [FRFM^{+}25]. Mix-GEMM [FRFM^{+}25]. Mixed [BOL^{+}25, FRFM^{+}25, FLF20, JDB^{+}23, LL22, LYC^{+}23, MNB20, MBP21, MRB^{+}24, VSG^{+}23, XLL^{+}22, YLG^{+}23, ZZ25, ZXG^{+}24, dSBS^{+}22]. Mixed-Criticality [BOL^{+}25, JDB^{+}23, LL22, MNB20, MBP21, VSG^{+}23, ZZ25]. Mixed-Precision [FRFM^{+}25, MRB^{+}24, ZXG^{+}24, dSBS^{+}22]. Mixed-Radix [XLY^{+}25]. Mixture [SZS^{+}22]. ML [CXY24, GSB23, JLP^{+}25]. ML-Based [CXY24]. ML-Enhanced [JLP^{+}25]. ML-PTA [JLP^{+}25]. MLC [HLLC21, SCA^{+}25]. MLC-MRAM [SCA^{+}25]. MLCD [CYPC25]. MM [FHH22]. MM-FSM [FHH22]. MMDataLoader [JZH^{+}24]. Mobile [BLH^{+}21, EEA22, FZM^{+}23, KCL^{+}20, KLC20, LL21, LZS^{+}24, LHW^{+}25, LPD^{+}21, SBP^{+}20, TQL^{+}22, XZL^{+}21, XWL^{+}24, ZFQ^{+}23, ZXY^{+}24, ZZF^{+}24, ZZH^{+}25]. Mobility [LGL^{+}24, SKM^{+}23]. Mobility-Aware [LGL^{+}24]. Mobility-Based [SKM^{+}23]. Mod [Koç20]. Modal [HXL^{+}25, SSY^{+}21, YZM^{+}25]. Mode [HMJ24, LL22, LLY22, NKL^{+}23, OAK^{+}23, PCA^{+}23, SZL^{+}22]. Model [ATT22, BHW^{+}23, CB22, CDW^{+}25, CRJZ21, DGTVGG21, DPS^{+}20, FBM21, GCL^{+}21, HLL^{+}20, HECC^{+}21, HLC^{+}22, HSP^{+}25, IBB^{+}25, JZH^{+}24, KYS^{+}22, MZM^{+}25, OTTT22, PMA^{+}24, PKPR23, SDR^{+}22, SZL^{+}24, WLW^{+}22c, WMH^{+}24, Xu24, ZZZ^{+}23, ZGG^{+}23]. Model-Based [CB22, DPS^{+}20, PMA^{+}24]. Model-Hardware [SZL^{+}24]. Model-Heterogeneous [ZGG^{+}23].

Modeled [WLD^{+}22]. Modeling [FV23, GVN25, IIEKS24, JM21, KAA20, KASAG23, LZW^{+}21, MTV^{+}21, RvSP^{+}25, SSZ^{+}20, WWJ^{+}23, WCL^{+}23, WRT^{+}22, YZG^{+}25]. Modelling [AT23, NML25]. Models [BCF^{+}25, HLJ^{+}25, KPS^{+}24, KLW^{+}25, LDZ^{+}23, LWH^{+}24, LPC^{+}21, LTJS^{+}22, LM21, MHA^{+}20, RBSG23, ROPdlT22, RSR22, TOF^{+}24, TDH^{+}23, YCY^{+}24]. Modern [CSY^{+}25, CY22, FAFK21, XTWW25, ZWX^{+}25]. Modular [AJ22, BRS^{+}24, BOL^{+}25, CCG^{+}22, JRL25, MKH^{+}21, MÖS22, MTK25, TWZ^{+}23, XL25, XLZ^{+}25, ZCP22, ZCP23]. Modularized [WWC21]. Modulation [NWH^{+}25]. Module [LLS^{+}23, MDJ20]. Moduli [JRL25]. Moduli- [JRL25]. Modulo [PNK^{+}23]. Modulus [XLY^{+}25]. MoE [KLW^{+}25]. Moldable [BLP^{+}22]. Molecular [DWW25, WMG^{+}25, WYB^{+}24]. Monarch [PB23b]. Monitoring [BOL^{+}25, CXL^{+}25, KCS23, PLZ20, YWP^{+}23]. Monitors [MDM22]. Monotonicity [Mik24]. Monteiro [SEM23]. Montgomery [AJ22, NS22, XL25, XLZ^{+}25, ZCP22, ZCP23]. Morphing [WWC21]. Most [LLFT23]. Most-Significant [LLFT23]. MPAM [ZCB23]. MPI [CBB^{+}21b]. MPQ [ZXG^{+}24]. MPSoC [KKB^{+}22]. MPSoCs [HGK^{+}22]. MRAM [CFA22, FTR23, HWC^{+}22b, JKKP25, KJK^{+}22, OAK^{+}23, SCA^{+}25, SMFS21, TSM^{+}21]. MRAMs [WRT^{+}22]. MRC [ZTZ^{+}25]. MTHAEL [VAV^{+}20]. MTTKRP [SLY^{+}22b]. Muller [YLHL23]. MulTa [CCCW21]. MulTa-HDC [CCCW21]. Multi [Akr22, BLKK23, BMM^{+}22, BCBS21, CCCW21, CFC^{+}22, CZZ^{+}25, DGX^{+}25, DRP24, DLZ^{+}24, DRY^{+}22, FPX^{+}25a, FNS^{+}22, FZM^{+}23, GXZ^{+}23, Has23, HXL^{+}25, xHzLH^{+}24, HZF^{+}24, JDCL23, JWS^{+}21, KKH22, KSKK23, KCS23, KMAA25, KHP21, KSL^{+}22, KKKC20, LMW^{+}24, LHX^{+}25, LHY^{+}21, LZW^{+}24b, LAPB21, LFP^{+}22, MZM^{+}25, MIPQ22, MBP21, Mik24, MCS^{+}22, NKL^{+}23, PYS20, PCA^{+}24, RPB^{+}23, SCC21, SNN21, SKLR22, SSY^{+}21, STZ^{+}24, ST23a, SYW^{+}22, TDZ^{+}22, WCB23, WZH^{+}23, WWJ^{+}25, WSC^{+}25, XCG^{+}25, YNJS21, YLZ^{+}24, YWP^{+}23, YTD^{+}21, YH20, YDL^{+}25, YBF^{+}25, ZLT^{+}25, ZZM^{+}22, ZFH23, ZDW^{+}23, ZCF20, LWZ^{+}25, ZDZ^{+}23, ZYL^{+}22]. Multi-AAV [LWZ^{+}25]. Multi-Accelerator [RPB^{+}23]. Multi-Agent [YBF^{+}25, ZDW^{+}23]. Multi-Applications [DRY^{+}22]. Multi-Attribute [KCS23]. Multi-Bank [KKH22]. Multi-Bit [KSL^{+}22, LHX^{+}25, YNJS21]. Multi-Block [ZFH23]. Multi-Chain [CZZ^{+}25]. Multi-chiplet [WWJ^{+}25]. Multi-Core [Has23, JDCL23, KHP21, KKKC20, MIPQ22, MBP21, PYS20, PCA^{+}24, SNN21, SKLR22, YWP^{+}23]. Multi-Cores [ZCF20]. Multi-DNN [DRP24, ZDZ^{+}23]. Multi-GPU [TDZ^{+}22]. Multi-Grained [HZF^{+}24]. Multi-Hop [Akr22]. Multi-Leader [LZW^{+}24b]. Multi-Level [YH20, YDL^{+}25]. Multi-Modal [HXL^{+}25, SSY^{+}21]. Multi-Mode [NKL^{+}23]. Multi-Model [MZM^{+}25]. Multi-NN [BLKK23, LMW^{+}24]. Multi-Node [SYW^{+}22]. Multi-NUMA [ZYL^{+}22]. Multi-Objective [DRP24]. Multi-Party [DLZ^{+}24, LHY^{+}21]. Multi-Path [FPX^{+}25a]. Multi-Program [LFP^{+}22]. Multi-Resource [SCC21, ZZM^{+}22]. Multi-Scalar [YLZ^{+}24]. Multi-Segment [CKP^{+}22]. Multi-Source [xHzLH^{+}24]. Multi-Spin-Flip [ST23a]. Multi-Stage [STZ^{+}24]. Multi-Stream [BMM^{+}22].

Multi-Streaming [JWS^{+}21].

Multi-Target [BCBS21]. Multi-Task [CCCW21, KSKK23]. Multi-Tenant [KMAA25, WZH^{+}23]. Multi-Term [Mik24]. Multi-Threaded [CFC^{+}22, MCS^{+}22]. Multi-Threading [FNS^{+}22]. Multi-Tierd [YTD^{+}21]. Multi-UAV [FZM^{+}23, YBF^{+}25, ZLT^{+}25]. Multi-User [WSC^{+}25, XCG^{+}25]. Multi-Valued [DGX^{+}25]. Multi-Version [GXZ^{+}23]. Multi-Workflow [WCB23]. Multi/Many [LAPB21]. Multi/Many-Core [LAPB21]. Multiagent [SSW^{+}24]. Multicasting [LHK^{+}22]. Multiclass [CWC^{+}24]. Multicore [BY22, BY24, HBS20a, HF22, LB22, LSU^{+}23, MÁJG^{+}24, SCFPM22, SXXL24, XAP20, XNLX20, YWF^{+}25]. Multidimensional [WFT^{+}21]. Multidomain [TOF^{+}24]. Multihop [HZMC24]. Multilayer [AZS^{+}23, KZS^{+}25]. Multilevel [LWC^{+}22]. Multimodal [YZM^{+}25]. Multiobject [SSW^{+}24]. Multiobjective [ZNW^{+}24]. Multiple [ATT22, CSvdBU22, CPL^{+}23, DNMS20, Fic22, HMMP23, HWX^{+}24, HA25, KvL22, LLCJ23, RDS23, XST20, XNH^{+}25, ZCK20, ZGL^{+}21]. Multiple-FPGA [DNMS20]. Multiple-Precision [ZCK20]. Multiplexing [LDF^{+}25, WDZ^{+}24]. Multiplication [AJ22, BMBM20, ESW^{+}23, KCL^{+}20, KJK^{+}22, LRB23, MDJ20, MTK25, PN25, TWZ^{+}23, TAP^{+}25, WDL^{+}25, WWL^{+}23, XL25, XLZ^{+}25, XLY^{+}25, YLZ^{+}24, ZQY^{+}20, ZCP23, ZLL^{+}24b, ZZW^{+}25, SYL^{+}23]. Multiplications [DSK23, KvL22, PCCK22]. Multiplicative [BCCM22, JBK24, ZDV^{+}22].

Multiplier [ACH21, CQI^{+}22, JYM20, LCZ22, PNK^{+}23, USS^{+}21, ZCP22]. Multipliers [BK23, ERKP21, HZK24, Ima21, MWJ^{+}24, TOM23, ZLM^{+}24]. Multipliers-Driven [ZLM^{+}24]. Multiply [SNT22, ZCK20]. Multiply-Accumulate [SNT22, ZCK20]. Multiprocessor [BBL22, CSvdBU22, CFW25, MNB20, ZCW^{+}21]. Multiprocessors [VSG^{+}23]. Multitasking [CHL^{+}23, WWL^{+}25, ZCZ^{+}22, ZWC^{+}23]. Multithreading [ROPdlT22]. Multivariate [BWSG25, PNK^{+}23]. MUSE [YTD^{+}21]. Mutation [GZC^{+}21, MXY^{+}23]. Mutation-Enabled [MXY^{+}23]. Mutual [YDG^{+}24]. Mutual-Influence-Aware [YDG^{+}24]. MViD [KCL^{+}20].

Namespace [WZW^{+}24]. NAND [KKS^{+}22, LSCX20, PLZ^{+}23, PZY^{+}23, WHL^{+}23, WHK24, YZG^{+}25]. Narrow [JKNK24]. Narrow-Width [JKNK24]. NAS [LZW25]. Native [FPHW25, MDPM24]. NBBS [MIPQ22]. NCFET [RGvS^{+}24, SRP^{+}21, SZK^{+}22]. NCFET-Based [RGvS^{+}24]. nDirect2 [YWF^{+}25]. NDN [KCS23]. NDN-Based [KCS23]. NDP [CKJ^{+}22]. NDRec [LWH^{+}24]. NDSTRNG [CTZ^{+}24]. Near [CJYC23, DSK23, HMK^{+}21, KvL22, LWH^{+}24, LCM25, LQM^{+}24, SMFS21, TBS^{+}25, YQX^{+}25, YNP^{+}24]. Near-Data [HMK^{+}21, LWH^{+}24, TBS^{+}25, YNP^{+}24]. Near-Memory [CJYC23, DSK23, LCM25, LQM^{+}24, YQX^{+}25]. Near-Optimum [SMFS21]. Near-Precise [KvL22]. Nearest [BLM21, KSL^{+}22, KPD^{+}23]. Nearly [RGS22]. Nebula [KLP^{+}21]. Negative [KPL^{+}22]. Neighbor [KSL^{+}22, KPD^{+}23]. Nested [JZSD24, SGL^{+}20]. NetCRC [NWL^{+}25]. NetCRC-NR [NWL^{+}25]. Netlists [CPM^{+}23, MLL^{+}25]. NetMod [NWH^{+}25]. Nets [CFWC23]. Network [AT23, AHC^{+}20, AGQ^{+}23, BYZZ20, BBC^{+}20, CCT^{+}20, CSK22, DA22, DSCB25, DLY21, FXL^{+}24, FPX^{+}25b, FYR^{+}24, FXJW25, FTV25, GSB23, GHK^{+}25, GXZ^{+}23, GXZ^{+}24, HGK^{+}22, HLC^{+}22, HYW^{+}21, HBB^{+}21, HZW^{+}24, JLZ21, KKS^{+}22, KPL^{+}22, KLR23, LLM^{+}25, LLD^{+}25, LHW^{+}21, LWZ^{+}25, LCL^{+}20, LYW^{+}25, LHZR25, LY20, LLY22, MSLY24, MYUK21, MWJ^{+}24, MC23, NWL^{+}25, NICY24, OKU^{+}24, PN24, RPMH21, RSR22, SLY22a, STW^{+}21, SKM^{+}23, TKM20, VBA20, VAV^{+}20, WFW^{+}20, WDCC20, WYSL22, WLZ^{+}23, WZF^{+}24, XST20, XJY^{+}24, YQX^{+}25, YWP^{+}23, ZZG^{+}23, ZCK20, ZCR23, ZZZ^{+}23, ZHL^{+}24, ZLZ24a, ZCC^{+}23, dSdCF22]. Network-Based [DA22, RPMH21, TKM20]. Network-Challenged [LWZ^{+}25]. Network-on-Chip [AT23, HYW^{+}21, HBB^{+}21, HZW^{+}24]. Network-on-Chips [XST20]. Networking [FQYS23, QSC^{+}25, XLS^{+}24]. Networks [Akr22, APH^{+}23, AC22, CDRS20, CR24, CXL^{+}23, DRA21, DWLF25, ESW^{+}23, ESdP^{+}25, FXC^{+}23, FPX^{+}25a, FHW^{+}22, GYH^{+}22, HGK^{+}22, HA25, HSP^{+}25, IIEKS24, JKNK24, JM21, JCZ^{+}23, JJZW24, JLL22, KCL^{+}20, KLP^{+}21, KJK24, KYS^{+}22, KNP^{+}20, LWL^{+}23, LWL^{+}21, LHXH22, LLCJ23, LLC^{+}24, LRL22, LDF^{+}25, LFX^{+}21, LDT^{+}25, MLL^{+}25, MRB^{+}24, MHT25, NKA24, PCMP21, PCCK22, PYW^{+}22, PPQBA21, PKPR23, RBC^{+}23, RSR22, RRMS25, SPB^{+}21, TKN23, WCQW22, WZX^{+}22, WNL^{+}23, WWC21, XWP^{+}21, XLS^{+}24, XLW^{+}20, YFC^{+}22, YBG^{+}22, YZY^{+}25, YBF^{+}25, YNP^{+}24, ZZZ^{+}20, ZFZ^{+}21, ZXY^{+}24, ZXZ^{+}21]. Networks-Based [APH^{+}23]. Networks-on-Chip [TKN23, XWP^{+}21]. Neumann [CCG^{+}22, ZFD^{+}20]. Neural [APH^{+}23, AHC^{+}20, ACH21, AGQ^{+}23, AC22, BZW^{+}25a, BBC^{+}20, CCT^{+}20, CZB^{+}22, CR24, CXL^{+}23, CSK22, DA22, DSCB25, DLY21, DRA21, ESW^{+}23, ESdP^{+}25, FYR^{+}24, FHW^{+}22, HGK^{+}22, HA25, HLC^{+}22, HSP^{+}25, JKNK24, JM21, JLZ21, JLY^{+}21, JJZW24, JLL22, KKS^{+}22, KCL^{+}20, KLP^{+}21, KPL^{+}22, KJK24, KLR23, KZS^{+}25, KYS^{+}22, KNP^{+}20, LWL^{+}23, LWL^{+}21, LPC^{+}21, LCL^{+}20, LRL22, LGZ^{+}25, LYW^{+}25, LHZR25, LFX^{+}21, LY20, LLY22, LLK^{+}23, LDT^{+}25, LZW25, MKY^{+}24, MZM^{+}25, MLL^{+}25, MWJ^{+}24, MRB^{+}24, MHT25, NICY24, NKA24, OKU^{+}24, PCMP21, PCCK22, PN24, PPQBA21, PKPR23, RPMH21, RBC^{+}23, RSR22, RRMS25, SPB^{+}21, SZ22, TKM20, VBA20, VAV^{+}20, WFW^{+}20, WDCC20, WZX^{+}22, WZF^{+}24, WWC21, XLS^{+}24, XLW^{+}20, XJY^{+}24, YFC^{+}22, YBG^{+}22, YWP^{+}23, YNP^{+}24, ZZZ^{+}20, ZCK20, ZCR23, ZZZ^{+}23, ZHL^{+}24, ZLZ24a, ZCC^{+}23, ZYQ^{+}24, dSdCF22, CZB^{+}22]. Neural-PIM [CZB^{+}22]. Neuroimaging [CTM^{+}25]. Neuromorphic [CQ22, LRRK^{+}22, ZGK20]. Neuron [LRL22, WFW^{+}20, WZF^{+}24]. Neuron-Level [WFW^{+}20]. Neutral [CBB^{+}21b, ZHM20]. Next [ZLS^{+}24, DHM^{+}25]. Next-Generation [ZLS^{+}24]. NFV [ZFL^{+}22]. NGS [KXGS22]. Nicaea [JZX^{+}25]. NN [BLKK23, LMW^{+}24]. NNs [SZL^{+}24]. No [HXGR20]. NoC [AGSD25, CCZ^{+}22, DKJP21, PVB21, XJY^{+}24, YL20]. NoC-Based [XJY^{+}24]. NoCs [PC24]. Node [HYH^{+}25, SYW^{+}22]. Node-Wise [HYH^{+}25]. Nodes [ZLW^{+}24, ZWZZ24]. Noise [HDAS21, YCS^{+}24, dOCC23]. Noise-Aware [YCS^{+}24]. Noisy [LLS^{+}24]. Non [BHE21, BXW^{+}25, CWWW20, CTZ^{+}24, FBH^{+}22, Gha21, GVN25, GNH20, KGHRM23, KRB^{+}22, LLS^{+}25, LJY^{+}24, LY20, MIPQ22, NK22, QCX^{+}23, WFT^{+}21, WHX^{+}25, YBW21, ZGQ^{+}22, ZCY^{+}25, ZTZ^{+}25]. Non-Binary [KGHRM23]. Non-Blocking [MIPQ22, YBW21]. Non-Coherent [KRB^{+}22]. Non-Conflicting [LLS^{+}25]. Non-Deterministic [CTZ^{+}24]. Non-IID [ZGQ^{+}22]. Non-Interactive [QCX^{+}23]. Non-Intrusive [BXW^{+}25]. Non-Normalized [GNH20]. Non-Preemptive [GVN25]. Non-Stack [ZTZ^{+}25]. Non-Stationary [Gha21]. Non-Triangular [FBH^{+}22]. Non-Uniformity [WFT^{+}21]. Non-Volatile [BHE21, CWWW20, LJY^{+}24, LY20, NK22, WHX^{+}25, ZCY^{+}25, LLY22]. Noninteractive [BHW^{+}23]. Nonlinear [DDK22, FHH22, JLP^{+}25, KSB24]. Nonuniform [HJYL22]. Nonvolatile [ZXD^{+}24]. Normal [ERKP21]. Normalized [GNH20, LLL25b]. NOSTalgy [SMFS21]. Novas [ZZF^{+}24]. Novel [AA20, AB22, BHK^{+}23, BFG^{+}21, CDRS20, ERKP21, GZW^{+}21, JLP^{+}24, LHW^{+}21, WY25, YHC^{+}20, YFC^{+}22, ZGWY22, ZLM^{+}24, ZLC^{+}23a]. NPC [LLS^{+}25]. NPN [ZWM20]. NPU [BZW^{+}25a]. NPUs [CHL^{+}23]. NR [NWL^{+}25, NWL^{+}25]. NS [LHZR25]. NStore [WLW^{+}25]. NTRU [DMG23, YHH^{+}25]. NTRU-Based [YHH^{+}25]. NTT [CZW^{+}25, XLY^{+}25, ZQY^{+}20]. NTT-Based [ZQY^{+}20]. NTT-Uncoupled [ZQY^{+}20]. NTTU [ZQY^{+}20]. NUCA [RPMH21]. NUMA [WLW^{+}25, YLL^{+}24, ZYL^{+}22]. NUMA-Aware [WLW^{+}25]. Number [AJ22, APK20, CMQ^{+}22, CTZ^{+}24, DSJ^{+}22, HF23, LMDC21, LFX^{+}21, MKÖ^{+}22, YLHL23, ZDV^{+}22]. Numbers [LLL25b]. Numerical [CYKG23, CTM^{+}25, QLH^{+}24]. NV [CWWW20]. NV-Journaling [CWWW20]. NVDIMM [WCYK20]. NVM [CCT^{+}20, ZGD23]. NVM-Based [CCT^{+}20]. NVM-Storage [ZGD23]. NVMe [MKYP21, NTDH25, PYYG21, PYDG22, YZJ23]. NVMe-Based [NTDH25]. NVMs [LJY21]. NVRAM [LV23].

O [BMM^{+}22, DTH^{+}24, HYS^{+}20, HWL^{+}21, KJC^{+}21, LJY^{+}24, LYH^{+}24, NTDH25, PE22, WJL^{+}20, ZYXD20, ZXL^{+}24, ZZC^{+}23]. O-Bound [CHZ^{+}25]. Obfuscated [ZXX^{+}25]. Object [LLL^{+}20, ZCX^{+}20, ZWY^{+}23]. Object-Based [ZCX^{+}20]. Object-Level [LLL^{+}20]. Objective [DRP24]. Objectives [CPL^{+}23]. Oblivious [JQY^{+}25]. Observation [Fic22, LPD^{+}21]. Observations [LQY^{+}20, YZG^{+}25]. Observing [TRBM22]. Occlusion [YCS^{+}24]. Occurrence [LLCJ23]. Octave [LGW^{+}22]. OctCNN [LGW^{+}22]. OFEI [XXJ^{+}24]. Off [AG24, BZW^{+}25a, BFC20, GXY^{+}23, HPJK22, LGX^{+}22, LZC^{+}24]. Off-Chain [GXY^{+}23, LGX^{+}22]. Off-Chip [BZW^{+}25a, HPJK22]. Off-Loading [LZC^{+}24]. Offline [JQK^{+}24, LYGC24, WTL^{+}24]. Offline-to-Online [LYGC24]. Offloading [CKRP21, DWYX20, GYZC25, HLY^{+}25, HZMC24, HTZ^{+}25, JWG^{+}23, KLC20, LCS^{+}25, LYGC24, WHM^{+}22, YHW^{+}25, ZGLZ20, ZLM^{+}24, ZZH^{+}25]. Offs [ZHLR22, ZZL21]. OGD [LSL^{+}25]. Olive [QSC^{+}25]. Olive-Like [QSC^{+}25]. Olympus [CWT^{+}22]. OmpSs [dHBF^{+}21]. On-Chain [LGX^{+}22]. On-Chip [GvSHA22, JKK^{+}22, KKH22, KRB^{+}22, LY21, TSM^{+}21, WNP^{+}22, JPHY20]. On-Core [WSG^{+}23]. On-Demand [MHA^{+}20]. On-Device [BCRX23, JLD^{+}25, TKM20, ZHL^{+}24]. On-Line [LWL^{+}23]. One [FFG^{+}25, LLL25b, MZM^{+}25, SHZ^{+}25, SZS^{+}22, WMH^{+}24, WGD^{+}22]. One-Class [WMH^{+}24]. One-Pass [FFG^{+}25]. One-Shot [MZM^{+}25]. One-Size-Fits-All [WGD^{+}22]. Online [BKHY22, CPL^{+}23, GVN25, HF22, LLW^{+}24, LYGC24, RCS^{+}21, SLY22a, VKRK22, WZH^{+}23, WDW^{+}23, ZZF^{+}24]. Onto [KH23, LCS^{+}25]. Opara [CXH^{+}25]. Open [FQYS23, KRB^{+}22, PCA^{+}24, VRR^{+}24, ZGG25]. Open-Source [FQYS23, KRB^{+}22, PCA^{+}24, ZGG25]. OpenHD [KKRK22]. OpenMP [DWW25, SPDQ22, SGL^{+}20, SGS^{+}21]. OpenVX [LAPB21]. Operand [CJYC23]. Operand-Oriented [CJYC23]. Operating [GSS^{+}23, MTV^{+}21, dOCC23]. Operation [JKKP25, OKC^{+}20]. Operations [BLM21, FAFK21, GXY^{+}23, NTL^{+}24, TLC^{+}24]. Operator [CXH^{+}25, DA22, MLW^{+}23, ZWL^{+}25]. Operators [MLW^{+}25, WSG^{+}25]. Opportunistic [DKJP21, WZW^{+}23, ZFL^{+}22]. Optical [FPX^{+}25b, STW^{+}21]. Optimal [BK23, BLM20, CDW^{+}25, CKP^{+}22, GSC^{+}23, IIEKS24, JBK24, JWG^{+}23, QHZ^{+}21, ZXW^{+}24, ZLZ24a]. Optimality [CWT^{+}22, TC21]. Optimisation [DRP24, PSBB21]. Optimisations [MCD^{+}25]. Optimiser [MHM^{+}23]. Optimistic [DLW^{+}25]. Optimization [BKHY22, BCF^{+}25, CWS^{+}24a, CCZ^{+}25, DYC^{+}25, DHM^{+}25, GKT^{+}22, HLZ^{+}25, HGC^{+}22, JLP^{+}24, LSCX20, LZW^{+}21, LLJ^{+}23, LPD^{+}21, LZZ^{+}22, MCT22, MRA^{+}21, QCX^{+}23, RCS^{+}21, RGvS^{+}24, SPB^{+}21, STW^{+}21, SWR^{+}23, WLQ^{+}21, WGD^{+}22, WWS^{+}22, WWL^{+}25, XWP^{+}21, XJY^{+}24, YHW^{+}25, YBF^{+}25, ZTT22, ZNW^{+}24, ZXMX25, ZML^{+}25]. Optimizations [CDP21]. Optimize [IBB^{+}25, SDR^{+}22, WWJ^{+}25, ZCF20]. Optimized [BFG^{+}21, CERMH23, CKRP21, HKC^{+}23, OLZ^{+}20, SJYQ25, TGS^{+}22, USS^{+}21, ZSWS24]. Optimizing [CRJZ21, HYS^{+}20, HSH^{+}25, JBS^{+}25, LWC^{+}22, LWZ^{+}25, LZW23a, MB21, MS25, PBBA25, PCBD23, SLY^{+}22b, TAP^{+}25, WZG^{+}23, WXL^{+}23, WWX^{+}24, WLF^{+}25a, ZCR22]. Optimum [SMFS21]. OPTIMUS [ODK20]. Optoelectronic [STW^{+}21]. OPTWEB [MYUK21]. OQ [PC24]. ORAM [LTL^{+}25, ZLZ^{+}23]. ORB [HJX^{+}25]. Orchestration [CL20, GWC^{+}25, GQJ^{+}22]. Orchid [CPL^{+}23]. Order [CMGD24, DVV23, RMR22]. Ordered [LDF^{+}24]. Orderless [WZW^{+}24]. Ordinal [LPC^{+}21]. Organizing [AGSD25, VLPS25]. Oriented [BXW^{+}25, CJYC23, DSTD22, HFT^{+}25, MXY^{+}23, PYYG21, YWX^{+}23, YLT^{+}23, ZLL^{+}22a]. Orthogonal [LSL^{+}25, MZM^{+}25]. OS-Level [SCFPM22]. OSC [BKHY22]. OsmoticGate [QWT^{+}23]. OurRocks [CKRP21]. Out-of-Core [GLB21, LZZ^{+}22]. Out-of-Place [GLB21]. Outer [SSCK25]. Outlier [SZS^{+}22, SCY^{+}23, WMH^{+}24, WFZ^{+}25]. Output [CZZ^{+}25, LLL25b]. Output-Based [CZZ^{+}25]. Over-Scaling-Based [ZSS^{+}22]. Overhead [ACKA23, DDK22, WHL^{+}21, WDZ^{+}23, ZLC^{+}22, ZTZ^{+}25]. Overheads [JQY^{+}25]. Overloading [CKK^{+}22]. Oversubscription [YLT^{+}23]. Oz [WZW^{+}24].

P [SLOM^{+}23]. Packaging [LZW^{+}24b]. PackCache [WDW^{+}23]. Packed [ZJW^{+}24]. Page [HWC^{+}22a, HLK^{+}25, JZSD24, OAK^{+}23, PYS20]. Pages [HWC^{+}22a, YLG^{+}23]. Paging [JQY^{+}25]. Paillier [RSZ23]. Pairing [BRPM22]. Pako [DGX^{+}25]. PAM [CQI^{+}22]. PANTHER [AHC^{+}20]. Paradigm [GZG^{+}23, PD21, PCBD23]. Paradigms [TJG^{+}23]. Parallel [BLP^{+}22, BCKS22, BLM20, CVOJRH22, DT20, FWZ^{+}21, GSK^{+}22, GZW^{+}21, HYS^{+}20, HJX^{+}25, JGD^{+}21, JCY^{+}23, JZX^{+}25, LPYT22, LHL^{+}21, LCZ22, LLT^{+}23, LMH^{+}25, MTK25, MC23, NML25, PYW^{+}25, PS22, ROPdlT22, SXXL24, SJYQ25, UGvdBC23, WLD^{+}22, vSDHA23]. Paralleled [LLL^{+}25a]. Parallelism [BYZZ20, CERMH23, CXH^{+}25, FFG^{+}25, KLKK23, KJK^{+}22, LMW^{+}24, SGL^{+}20, WDZ^{+}24, ZCCG23, ZGL^{+}21]. Parallelization [AhRX^{+}20, CKP^{+}22, DWW25, ZBT22, ZWC^{+}22]. Parallelization-Aware [AhRX^{+}20]. Parameter [KvL22, QHT^{+}24, ZYZ^{+}23, ZGL^{+}21, ZYZ^{+}25b]. Parameter-Adaptive [ZYZ^{+}25b]. Parameters [BMM^{+}22]. ParaX [ZYL^{+}22]. ParBFT [CERMH23]. Parities [YLHL23, YH24]. Parity [YLL^{+}20]. PARLE [CHM25]. PARLE-G [CHM25]. PARMA [AhRX^{+}20]. PaRTAA [MNB20]. Partial [AVK20, AHK^{+}21, DRY^{+}22, KLKK23, WCD25]. Partially [AHK^{+}21, DGX^{+}25, TKN23, WZG^{+}23]. Partially-Synchronous [DGX^{+}25]. Partition [HS22, NHW^{+}24]. Partitioned [Alm23, CBB21a, RACB24, WZGT22, ZCW^{+}21]. Partitioning [CPL^{+}23, DKZ^{+}25, LZC^{+}21, LD22, LWL^{+}22, MB21, ODK20, PYS20, WSG^{+}23, XYM23]. Partitions [SAG22]. Party [DLZ^{+}24, LHY^{+}21, WY25]. Pass [FFG^{+}25, PYDG22]. Pass-Through [PYDG22]. Passing [TZZ^{+}21]. Password [ZLWJ23]. Past [SCC21]. Pasta [WT25]. Patching [SXC^{+}24]. Path [CFW25, FXL^{+}24, FPX^{+}25a, LXW^{+}23, RJ24, SXZJ24, SM24, UGvdBC23]. Path-Based [CFW25]. Pathogens [MGFY24]. Paths [DLW^{+}25, FXC^{+}23, FPX^{+}25a, JWS^{+}23, ZLC^{+}23b]. Patient [PBBA25]. Pattern [BMM^{+}22, CKJ^{+}22, GWX^{+}23, HYS^{+}20, IIEKS24, MHDMEA22, MFRR20, MHT25, XGMJ25, ZG23, ZCX^{+}20]. Pattern-Directed [HYS^{+}20, ZCX^{+}20]. Patterns [DQ23, RRDB20]. Payment [JQK^{+}24]. Payments [XXZ^{+}25]. PCB [PK23]. PCM [AG24, IKTY22]. PE [WDQ^{+}22, ZGWY22]. PE-Interactive [WDQ^{+}22]. Penalty [DKJP21, SSK22]. Penalty-Aware [SSK22]. Pentanomials [LCZ22]. Per-Operation [OKC^{+}20]. Perception [LYF^{+}22]. Perceptron [AZS^{+}23]. Perceptrons [KZS^{+}25]. PERCIVAL [MDPM24]. Perform [LYC22]. Performance [AG22, BCF^{+}25, BRS^{+}24, BWL^{+}25, CJSY24, CGS^{+}20, CBB^{+}21b, CZH^{+}25, CZW^{+}25, DQ23, EAMJ^{+}23, EAMK22, FHL^{+}22, FTR23, GQZ21, GWCS23, GKFF20, HKC21, HHPB20, HLT^{+}23, JWS^{+}23, KD25, KLL21, KRB^{+}22, KIY21, LG22, LWL^{+}25, LHX^{+}25, LCS^{+}25, LFW21, LZW^{+}23b, LHZ^{+}24, LMZ^{+}25, MRA^{+}21, MLW^{+}25, NKeSK^{+}23, OAC^{+}21, OJ23, PHC24, QHZ^{+}21, RPMH21, RHF24, RGvS^{+}24, SSM23, SSZ^{+}20, STYQ24, SST^{+}24, TRBM22, WLR20, WFT^{+}21, WLQ^{+}21, WCL^{+}23, WFH^{+}24, WHX^{+}25, WLW^{+}25, WLF^{+}25a, WWS^{+}22, WJL^{+}20, WDZ^{+}23, WSHJ23, WHK24, XLT^{+}25, XWP^{+}21, YWF^{+}25, YLG^{+}23, YLL^{+}24, ZZL21, ZLWJ23, ZXL^{+}24, ZCY^{+}25, ZYZ^{+}25a, ZDY^{+}23, dHBF^{+}21]. Performance-Aware [GKFF20]. Performance-Energy [ZZL21]. Performance-Neutral [CBB^{+}21b]. Performance-Power [ZDY^{+}23]. Performant [YYW^{+}24]. Period [CMQ^{+}22]. Periodic [BPM23, GPRV23, MBP21, RDS23]. Periodicity [DMD^{+}23]. Peripherals [ABdGG^{+}25, CZB^{+}22]. PermCNN [DLY21]. Permissioned [YDW^{+}25]. Permutating [SL23]. Permutation [CMQ^{+}22]. Permutations [UYZP22]. Permutations-Based [UYZP22]. Permuted [DLY21]. Persistence [JLP^{+}24]. Persistent [CSY^{+}25, CSW^{+}21, DTH^{+}24, FCZ^{+}23, JWS^{+}23, PBBA25, WBJC22]. Personalized [LGZ^{+}25, LHZR25]. Perspective [SMP22, SZAT22, ZDZ^{+}23, ZCB23]. Perspectives [PG23]. Pervasive [HZYY22]. PEs [XNH^{+}25]. Petri [CFWC23]. PFed [LHZR25]. PFed-NS [LHZR25]. Phase [AGSD25, CCYC22, KIY21, LLS^{+}23, LKMJ21, NICY24, SLDZ25, WSC^{+}24]. Phase-Change [LLS^{+}23, LKMJ21]. Physical [FV23, HGC^{+}22, JLZ^{+}23, LYF^{+}22, PKPR23, TOF^{+}24]. Physics [JDCL23, WZD^{+}20]. Physics-Based [JDCL23]. Physics-Limited [WZD^{+}20]. Piecewise [CQI^{+}22]. Piecewise-Linearly-Approximated [CQI^{+}22]. PIETT [OLD^{+}23]. Pike [SLDZ25]. PIM [CZB^{+}22, HCKK25, JKKP25, JCZ^{+}23, KJK^{+}22, LZW^{+}24a, MSSL21]. PIM-Accelerated [JCZ^{+}23]. PIM-Based [MSSL21]. Pipeline [APV22, LPW20, LLL^{+}25a, LMH^{+}25, WLW^{+}22b, YBW21, ZHY^{+}25]. Pipelined [BMLOM20, Bru23, DSP^{+}21, WGL^{+}20]. Pipelining [NKeSK^{+}23]. PISO [ERKP21]. PIT [ZXZ^{+}21]. PITEM [UYZP22]. Place [BCKS22, GLB21]. Placement [BYZZ20, HHPB20, HYL^{+}25, MSLY24, SLY22a, SKM^{+}23, XZL^{+}21, XZL^{+}23, XWL^{+}24, ZYZ^{+}23]. Plain [ZLT^{+}24]. Plane [DTZ^{+}25, WZD^{+}20]. Plane-Wave [WZD^{+}20]. Planner [LXW^{+}23]. Planning [SM24]. Plant [BYZZ20]. Plasticity [XNB21]. Plasticity-on-Chip [XNB21]. Platform [CCY^{+}24, KPD^{+}23, KRB^{+}22, LPW20, MSP^{+}21, PYW^{+}25, RHF24, ZHLR22]. Platforms [FWZ^{+}21, FLF20, GLZ^{+}24, LSU^{+}23, LZW25, NML25, SSK22, SKLR22, WWX^{+}24]. Plug [LZW25]. Plug-and-Search [LZW25]. PM [FPHW25, LYH^{+}24]. PM-DRAM [LYH^{+}24]. PMDB [ZGD23]. PMLiteDB [JWS^{+}23]. PMP [TLC^{+}24]. PODTherm [JDCL23]. PODTherm-GP [JDCL23]. Point [BLM21, Bru20, Bru23, CQI^{+}22, GNH20, GFB^{+}24, KPH^{+}25, KBR^{+}23, LQY^{+}20, LLL25b, MDJ20, Mik24, NHW^{+}24, NKL^{+}23, TOM23, VHL20, WFT^{+}21, XGMJ25, ZSHB21]. Poisoning [ZZZ^{+}23]. Policies [GVN25, MAM23]. Policy [HWC^{+}22a, HLY^{+}25, OKC^{+}20, RGvS^{+}24, SCFPM22, ZWSF24]. Polling [PYDG22]. Polyadic [PYW^{+}25]. PolyLUT [ALC25]. Polymorphic [AC22, PB23b]. Polynomial [ALC25, CMQ^{+}22, CWNL22, PNK^{+}23, TWZ^{+}23, WWL^{+}23]. Popularity [CZJ21]. Popularity-Aware [CZJ21]. Port [TDZ^{+}25]. Portable [LZS^{+}24]. Posit [LFX^{+}21, MDPM24]. Position [LTL^{+}25]. Positive [RSMMG^{+}23]. Post [GMZ22, KGHRM23, LHW^{+}25, SZAT22, WL24, WDC^{+}25, YHH^{+}25]. Post-Disaster [LHW^{+}25]. Post-Quantum [GMZ22, KGHRM23, WDC^{+}25, YHH^{+}25]. Post-Silicon [SZAT22]. Post-Training [WL24]. Potential [KLKK23, RGD^{+}24]. POWER [LFP^{+}22, ABP22, BZW^{+}25b, BCBS21, FTR23, GKFF20, HBS20a, HBB^{+}21, JBS^{+}25, KLR^{+}20, KASAG23, LLJ^{+}23, OKU^{+}24, OAC^{+}21, OAB^{+}23, OJ23, PLZ20, RAD20, RSP^{+}20, SRP^{+}21, TDH^{+}23, YL20, YBF^{+}25, ZQY^{+}20, ZDY^{+}23]. Power- [GKFF20, RSP^{+}20]. Power-Constrained [OAB^{+}23]. Power-Efficient [FTR23, SRP^{+}21]. Power-Gating [HBB^{+}21]. Powered [KKRK22, LJY21, ZXD^{+}24]. Powerful [LOM^{+}25]. PPCC [YL20]. PQNTRU [YHH^{+}25]. PR [KLKK23]. PR-SSD [KLKK23]. Practical [HWG^{+}23, HXL^{+}23, LQL^{+}24, LHS^{+}25, LQN^{+}21, TZY^{+}24, WNP^{+}22, YNJS21]. Practice [RBSG23]. Pragmatic [HWZ^{+}25]. Pre [BAM^{+}24, HLF^{+}23, KNP^{+}20]. Pre-Alignment [BAM^{+}24]. Pre-Defined [KNP^{+}20]. Pre-Training [HLF^{+}23]. PRECIOUS [SCA^{+}25]. Precise [FL21, HJYL22, KvL22, VSG^{+}23]. Precision [FRFM^{+}25, God20, GFB^{+}24, LY20, MCD^{+}25, MRB^{+}24, YZG^{+}25, ZCK20, ZDV^{+}22, ZXG^{+}24, dSBS^{+}22]. Preconditioned [YGW^{+}23]. Preconditioning [LLL^{+}23]. Predictability [PFHD21, ZCB23]. Predictable [BHE21, FBM21, GCR^{+}23, JWD^{+}22, KHP21, VRR^{+}24, XQC^{+}22]. Predicting [AMM21, LPC^{+}21, PZY^{+}23]. Prediction [BMBM20, HLS^{+}23a, xHzLH^{+}24, LLX^{+}24a, LL23, MHA^{+}20, PLZ^{+}23, RPMH21, SLLS25, YZG^{+}25]. Predictive [KPL^{+}22, MTV^{+}21, OJ23, PS22]. Predictor [MKY^{+}24, ZXD^{+}24]. Predictors [CY22]. Preemptive [GVN25, LKK^{+}21]. Prefender [LHFW24]. Prefetch [LFP^{+}22]. Prefetcher [AGB^{+}23, NTPAB^{+}25, RJ24]. Prefetching [CHZ^{+}25, LHFW24, WNP^{+}22]. Prefix [BLM20, RGS22]. Preprocessed [JZH^{+}24]. Preprocessing [KLC20, QZZ^{+}24]. Preprocessing-Based [QZZ^{+}24]. Presence [YNJS21]. Preservation [XNL^{+}23]. Preserving [BHW^{+}23, CZZ^{+}25, FHL^{+}23, LHA^{+}25, LMM^{+}23, SO23, WRW^{+}23, XZC^{+}23, ZLL^{+}24a, ZFQ^{+}23, ZQC^{+}25, ZWB^{+}22]. Pressure [MB21]. PRESTO [SSK22]. Pretender [LHFW24]. Prevent [ODK20]. Preventing [OD23]. Prevention [WSG^{+}23]. Pricing [WLW^{+}21]. Primary [LLS^{+}22]. PRIMER [DYPZ22]. Primitives [LZW^{+}23b, LHZ^{+}24]. Printed [AZS^{+}23, KZS^{+}25]. Priorities [SPDQ22]. Priority [BBL22, CSH^{+}24, FL21, JCY^{+}23, XNH^{+}25, ZABHCG23, ZCW^{+}21]. Priority-Based [CSH^{+}24]. Privacy [AAM^{+}25, BHW^{+}23, CWY^{+}23, CXL^{+}25, CZZ^{+}25, FHL^{+}23, FYR^{+}24, GZG^{+}23, LXL^{+}25, LMM^{+}23, MDM22, RSMMG^{+}23, SO23, WRW^{+}23, XZC^{+}23, XNL^{+}23, YCY^{+}24, ZLL^{+}24a, ZQC^{+}25, ZMH^{+}25, ZC24, ZWB^{+}22]. Privacy-Aware [CXL^{+}25]. Privacy-Breaching [MDM22]. Privacy-Enhanced [CWY^{+}23]. Privacy-Preserving [BHW^{+}23, CZZ^{+}25, FHL^{+}23, LMM^{+}23, SO23, XZC^{+}23, ZLL^{+}24a, ZQC^{+}25, ZWB^{+}22]. PrivAim [WRW^{+}23]. Private [CGLS21, CXL^{+}23, HC24, HWR^{+}24, LWZZ25, WY25, ZYZ^{+}25a]. Privileged [MZZC22, XHY^{+}22]. Proactive [MXY^{+}23, SKK^{+}21]. Probabilistic [BSM21]. Probabilistically [LCHK22]. Probability [SXZJ24]. Problem [CWS^{+}24a, ZWWY22]. Problems [ZTT22]. Procedure [dSdCF22]. Process [CSW^{+}21, HSH^{+}25, KKB^{+}22, MZZC22]. Process-to-BB [HSH^{+}25]. Processing [ABP22, BZW^{+}25a, BBC^{+}20, BCBS21, CZB^{+}22, CZJ21, CHC25, CKJ^{+}22, CJYC23, DWL^{+}22, GWH^{+}23, GWC^{+}25, GWX^{+}23, HCKK25, HMK^{+}21, LLS^{+}25, LZC^{+}21, LLT^{+}23, LWH^{+}24, LCY^{+}25, LCM25, LZZ^{+}22, LLX^{+}24b, NTL^{+}24, PCA^{+}24, RGS22, SJYQ25, TBS^{+}25, WDQ^{+}22, WLF^{+}25b, WSG^{+}25, YQX^{+}25, YNP^{+}24, ZFZ^{+}21, ZCD^{+}22, ZLZ24a, ZMS^{+}23, ZCSJ23, ZXZ^{+}21, ZGK20]. Processing-In-Memory [CZB^{+}22, CHC25, HCKK25, LLS^{+}25, NTL^{+}24, ZCSJ23]. Processing-In-Transmission [ZXZ^{+}21]. Processor [DLK25, GWD25, GMT24, KBQ^{+}23, LPW20, MHS^{+}20, NCD^{+}25, PCA^{+}24, WLW^{+}22c, XTWG23, YHH^{+}25, ZSHB21, ZLL^{+}22a]. Processors [AG22, CWT^{+}22, CCC23, DMX^{+}22, FWM^{+}23, FNS^{+}22, KKKC20, KH23, LJY21, LAKS20, Lu21, MBP21, ODK20, RCS^{+}21, SNN21, SXXL24, TSM^{+}21, TAP^{+}25, WZCM23, XKS21, XLW^{+}20, XTWW25, ZXD^{+}24]. Product [PN24, SSCK25, ZCP23]. Production [FV23, IBB^{+}25]. Profiled [PBBA25]. Profiling [CYKG23, DYPZ22, OJ23]. Program [LCJ^{+}24, LFP^{+}22]. Programmable [AHC^{+}20, BRPM22, DSCB25, LCJ^{+}24, NWH^{+}25]. Programming [HWJ^{+}21, LD22, NM22, VCLN21]. Programs [WSM^{+}24]. Progression [UGvdBC23]. Prolongation [GSB23]. Proof [CYX^{+}23, DTL^{+}25, SHZ^{+}25, YLZ^{+}24]. Proof-of-Market [CYX^{+}23]. Proof-of-Work [DTL^{+}25]. Prophet [SLLS25]. Proportional [HWL^{+}21, PE22]. Protect [LCHL21]. Protecting [BLH^{+}21, BGHR^{+}25, LPYT22, SRB23, WZSL22]. Protection [CZC^{+}21, HP23, LWL^{+}23, LLX^{+}24a, SSP^{+}24, YCY^{+}24, ZMH^{+}25, ZC24, ZGR^{+}25]. Protocol [FL21, JQK^{+}24, LQL^{+}24, LV23, SL23, SO23, TWL^{+}22, WY25, WHL^{+}21, YPL^{+}25]. Protocols [KHP21]. Prototype [YZX^{+}24]. Prototype-Based [YZX^{+}24]. Provable [CHM25]. Provably [XLWO23]. Providers [BZW^{+}25b]. Provisioning [CB22, WCB23, ZQG^{+}24, ZXY^{+}24]. Proximal [WYSL22]. PRS [ZCX^{+}20]. Prune [BYZZ20, DYJ20]. PruneAug [GLZ^{+}24]. Pruning [ALC25, GLZ^{+}24, KPL^{+}22, MRB^{+}24, YYZ^{+}25]. Pruning-Based [YYZ^{+}25]. Pseudo [JLP^{+}25, ZSHB21]. Pseudo-Transient [JLP^{+}25]. Pseudorandom [HF23]. PSO [RSZ23]. PSO-Based [RSZ23]. PStream [CZJ21]. PTA [JLP^{+}25]. Public [LLWZ23]. Public-Key [LLWZ23]. Publication [LXL^{+}25]. Publish [DWL^{+}22]. Publish/Subscribe [DWL^{+}22]. PUF [CHM25, PCA^{+}23, XLWO23]. PUF-TRNG [PCA^{+}23]. PUF2 [HWC^{+}22b]. Pulse [NM22]. Pulse-Level [NM22]. Purpose [ZLL^{+}24b]. Pursuing [YL20]. PyLog [HWJ^{+}21]. PyQUBO [ZTT22]. PySchedCL [GSK^{+}22]. Python [CYKG23, HWJ^{+}21, ZTT22]. Python-Based [HWJ^{+}21]. PyTracer [CYKG23].

QLC [YZG^{+}25]. QoS [BKHY22, LLX^{+}24a, ZCZ^{+}22, ZDC^{+}25]. QoS-Awareness [ZCZ^{+}22]. Qu [WAN^{+}25]. Qu-Trefoil [WAN^{+}25]. Quality [BCRX23, SMFS21, WRW^{+}23, XWP^{+}21, YZX^{+}24, YHV^{+}21]. Quality-Aware [WRW^{+}23]. Quality-Energy [SMFS21]. Quality-of-Experience [BCRX23]. Quantification [CHC25]. Quantified [TWJ^{+}22]. Quantitative [CRJZ21, TPWY23, XLT^{+}25]. Quantization [GCL^{+}21, MRB^{+}24, NKA24, SSJ21, WL24, WZF^{+}24, WZJ^{+}24, ZXG^{+}24]. Quantization-Aware [NKA24]. Quantized [JKNK24, QJY^{+}25]. Quantum [BBdTF25, BYM22, CLCL22, EGMW21, EAMJ^{+}23, FHW^{+}22, GMZ22, HXL^{+}23, KGHRM23, KB21, LLFT23, LLS^{+}24, LSS25, LSH25, MHDMEA22, SL23, TC21, WDC^{+}25, WY25, WAN^{+}25, XLT^{+}25, YDG^{+}24, YHH^{+}25, ZYD^{+}20, ZSWS24, ZDW^{+}23, dSdCF22, ddAPdS21]. Quantum-Classical [KB21]. Quantum-Resistant [ZYD^{+}20]. Qubit [LZF21, LCGH25]. QUBO [LSH25, ZTT22]. Queries [HMZ^{+}25, XXL^{+}23]. Query [DWL^{+}22, TWaKo^{+}23, WLF^{+}25b, YWC^{+}24b].

Queueing [AGSD25, WLR20]. Queues [YH20]. Queuing [CB22]. Quotient [LSW^{+}23, RGD^{+}24].

R [HHN^{+}23]. R-HTDetector [HHN^{+}23]. Racetrack [HKC^{+}23, KOH^{+}23, WLW^{+}22a, WLW^{+}22b]. Racetrack-ReRAM-Combined [WLW^{+}22b]. Racetracks [VLPS25]. Rack [GSC^{+}23]. Rack-Coordinated [GSC^{+}23]. Rad [LW22]. Rad-Hard [LW22]. Radiation [YHC^{+}20]. Radio [LLC^{+}24]. Radix [Bru23, CQCL25, GL24, HSE^{+}24, LDLK22, LSW^{+}23, XL25, XLY^{+}25, ZCP22]. Radix-16 [XL25]. Radix-4 [HSE^{+}24]. Radix-64 [Bru23]. Radix/Mixed [XLY^{+}25]. RAM [HLLC21, TRG^{+}24]. RAN [NWH^{+}25]. Random [CMQ^{+}22, CTZ^{+}24, GWG^{+}24, LCH22, PL21, SWR^{+}23, WXL^{+}23, WZG^{+}24, ddAPdS21, PL21]. Randomize [LBC24]. Randomizing [SXH^{+}24]. Range [WYB^{+}24, ZGD23]. Range-Based [ZGD23]. Range-Limited [WYB^{+}24]. Rank [HWG^{+}23]. Rank-Code-Based [HWG^{+}23]. Ransomware [BJM^{+}21, GWCS23]. Raptor [WYX^{+}24]. Raptor-T [WYX^{+}24]. Rare [PCCK22]. Rasta [RKMR23]. Rate [CFA22, LL23, RGS22, TDZ^{+}25]. Rate-Adjusted [LL23]. Ratio [CWC^{+}24, LLL^{+}23]. RCFI [PL21]. RCGL [XSYL22]. RDMA [LWL^{+}24, LV23, XLY^{+}22]. Re [LRL22]. Re-Computation [LRL22]. ReAAP [ZLL^{+}22a]. Reaching [CWT^{+}22]. Read [BAM^{+}24, CFA22, KLKK23, KXGS22, LPD^{+}21, PM20, WDZ^{+}22, WHK24]. Read-Ahead [LPD^{+}21]. Read/Write [WDZ^{+}22]. Real [AOM^{+}21, BHE21, BY22, BCCM23, BGB^{+}21, CBB21a, CSvdBU22, CHC25, CKP^{+}22, DSK23, DWN^{+}22, FHL^{+}23, GPRV23, GQH21, Has23, HIRB25, HF22, HECC^{+}21, HLS^{+}23b, JSTG20, JGD^{+}21, JYF^{+}23, JYM^{+}23, KHP21, KH23, LJY21, LL22, LZW^{+}21, LWZ^{+}25, LYF^{+}22, MNB20, MSP^{+}21, BCLC25, PS22, QWT^{+}23, RDS23, SZAT22, SCA^{+}25, SSK22, STK23, SM22, SGL^{+}20, WZGT22, XNH^{+}25, YCL^{+}24, ZABHCG23, ZLW^{+}24, ZSC^{+}25, ZHY^{+}25, ZHM20]. Real-Time [BHE21, BY22, BCCM23, CBB21a, CSvdBU22, CKP^{+}22, DSK23, DWN^{+}22, FHL^{+}23, GPRV23, GQH21, Has23, HIRB25, HF22, HECC^{+}21, HLS^{+}23b, JSTG20, JGD^{+}21, JYF^{+}23, JYM^{+}23, KHP21, KH23, LJY21, LL22, LZW^{+}21, LWZ^{+}25, LYF^{+}22, MNB20, MSP^{+}21, BCLC25, PS22, QWT^{+}23, RDS23, SZAT22, SCA^{+}25, SSK22, STK23, SM22, SGL^{+}20, WZGT22, XNH^{+}25, YCL^{+}24, ZABHCG23, ZLW^{+}24, ZSC^{+}25, ZHY^{+}25, ZHM20]. Real-World [AOM^{+}21, BGB^{+}21]. Realistic [SKK23]. Reality [NKN^{+}25, WSC^{+}25]. Realization [AA20, CXL^{+}23, KZS^{+}25]. Realize [CCYC22]. REALM [BOL^{+}25]. Reaping [WHL^{+}23]. Reasoning [QLH^{+}24]. Recognition [GYH^{+}22, LWW^{+}25, MHDMEA22, MFRR20, SSM23, YAG20]. Recommendation [LWH^{+}24, WGM^{+}20, YYW^{+}24]. Recommender [GZW^{+}21]. Reconfigurable [AHK^{+}21, BCBS21, CDRS20, CCG^{+}22, DGQ^{+}25, DPQK^{+}23, EAMJ^{+}23, LB22, LHL^{+}23, RBMG22, ROPdlT22, SPB^{+}21, WLHW25, ZLS^{+}24, ZLL^{+}22a, ZCS^{+}25]. Reconfiguration [DRY^{+}22]. Reconstruction [WHQ^{+}24]. Recording [WLZ^{+}21, ZYXD20]. Records [WWL^{+}23]. Recovery [BJM^{+}21, LLM^{+}25, TZ22, XLL^{+}22, ZLWJ23]. Rectangular [GLB21]. Recurrence [HSE^{+}24, KPH^{+}25]. Recurrent [FYR^{+}24, KCL^{+}20, PCMP21]. Recursion [WLD^{+}22]. Recursion-Tree [WLD^{+}22]. Recursive [LSH25]. Redactable [LSW25, TWaKo^{+}23]. ReDas [HWX^{+}24]. Reduce [DKJP21, DH20, ESN20, MSZ22, OTTT22, QHT^{+}24, WZX^{+}22]. Reduced [BZW^{+}25a, BB22, DPS22, JLZ21, YZX^{+}24, dSBS^{+}22]. Reducibility [BCV22]. Reducing [JQY^{+}25]. Reduction [CFA22, HZK24, JZSD24, KOH^{+}23, LWC^{+}22, MHDMEA22, ST23b, XST20, ZJW^{+}24]. Redundancy [Das23, DZC^{+}24, GYZC25, LZG^{+}24, LSXZ21, YZX^{+}24]. Redundant [DSJ^{+}22, HLJ^{+}25, LHK^{+}22, LMDC21, PCCK22, YHC^{+}20]. Reed [MCT22, TZ22, YLL^{+}20, YLHL23, YH24]. Reference [ZZZ^{+}23]. Refreshable [FW23]. Refreshing [FW23]. Region [MSW^{+}21]. Region-Based [MSW^{+}21]. Regional [ZHYJ21]. Register [WHC20]. Register-Based [WHC20]. Registers [SZHB21]. Regular [BCCM22, NTR21]. Regularity [DPCL22]. Regularity-Based [DPCL22]. Regulation [BOL^{+}25]. Reinforcement [HLZ^{+}25, HZM^{+}23, LWL^{+}22, LJH^{+}25, LYGC24, LZC^{+}24, MS25, PZY^{+}23, SSW^{+}24, STQ^{+}24, WHM^{+}22, WCB23, WTL^{+}24, WLD^{+}25, WJL^{+}20, WLHW25, XJY^{+}24, YBF^{+}25, ZGLZ20, ZNW^{+}24, ZLL^{+}22b, ZWSF24, ZHYJ21]. Related [BKS22, ZLW^{+}24]. Related-Key [ZLW^{+}24]. Relaxed [SEM23]. Relevant [RWCC23]. RelHDx [KZXR25]. Reliability [DSTD22, GA22, GvSHA22, KOH^{+}23, KAA20, KIY21, LHW^{+}21, PLZ^{+}23, RCS^{+}21, SXZJ24, SLLS25, WHL^{+}23, WZF^{+}24, XSYL22, ZQG^{+}24, ZZ25]. Reliability-Critical [SXZJ24, XSYL22]. Reliability-Oriented [DSTD22]. Reliable [BBC^{+}20, DTH^{+}24, DWLF25, FPX^{+}25a, GYZC25, HWC^{+}22b, KYS^{+}22, LDF^{+}25, LKMJ21, NML25, TGA23, VHL20, WFW^{+}20, WFZ^{+}25]. Relieving [WSC^{+}24]. Relocation [HYW^{+}21]. ReLU [KPL^{+}22]. Remainder [KPH^{+}25]. Remainders [AVK20]. Remap [HS22]. Remap-Based [HS22]. Remapping [HKS20]. Remapping-Based [HKS20]. Remodeling [ZHL^{+}24]. Remora [DLW^{+}25]. Remote [GJN^{+}25, KCS23, LQN^{+}21, LWL^{+}24, MKH^{+}21, SNA^{+}20]. Remotely [TQL^{+}22]. Removing [PCCK22]. Rendering [YCL^{+}24, ZSC^{+}25]. Renewables [HIRB25]. Renewables-Driven [HIRB25]. Reordering [ESN20, LLX^{+}24b, NP20]. Repair [LSXZ21, SLS^{+}21]. Repetitive [PCCK22]. Replacement [CCC23, MAM23, ZWSF24]. Replication [HYS^{+}20, LLR25, ZCX^{+}20]. Representation [CHM25, LMDC21, TWL^{+}22]. Reputation [ZCD^{+}22]. Reputation-Based [ZCD^{+}22]. Request [NTDH25, WLD^{+}25, YCKW20]. Requests [PVB21, WLD^{+}22]. ReRAM [WLW^{+}22b, AHC^{+}20, BTEC20, JLZ21, KSKK23, KJK24, LWL^{+}23, LZW^{+}24a, SKK23, WWM^{+}23, ZML^{+}25]. ReRAM-Based [BTEC20, JLZ21, KJK24, LWL^{+}23, LZW^{+}24a, SKK23, WWM^{+}23, ZML^{+}25]. Rescue [LHW^{+}25]. Research [Xu24, ZGG25]. Reservation [WL20]. Reservoir [SMY22]. RESET [WSC^{+}24]. RESET-Aware [WSC^{+}24]. Reshaping [HWX^{+}24, HLQ^{+}23]. Residue [AJ22, APK20, CMQ^{+}22, DSJ^{+}22]. Resilience [HPGM20, JDB^{+}23, YNJS21]. Resilient [BN24, BLP^{+}22, BJMKK23, IKAG^{+}22, NICY24, STK23, TCX^{+}23, YHC^{+}20]. Resistance [LGC^{+}23, NICY24, ZOH^{+}25]. Resistant [ZYD^{+}20]. Resisting [XXZ^{+}25]. Resistive [MGFY24]. Resolution [MHDMEA22]. Resolving [WZW^{+}24]. Resonance [HDAS21]. Resonance-Induced [HDAS21]. Resource [AZS^{+}23, BY24, CPL^{+}23, CCY^{+}24, HMMP23, HLS^{+}23b, HTZ^{+}25, IMQOP21, KLW^{+}25, LWL^{+}22, LCJ^{+}24, LWL^{+}25, LFGD25, LZC^{+}24, PYS^{+}23, PSBB21, SCC21, TCJ23, WLF^{+}25a, WDZ^{+}24, XYM23, YHW^{+}25, ZLWG22, ZGQ^{+}22, ZLC^{+}22, ZNW^{+}24, ZCC^{+}23, ZCCG23, ZZM^{+}22, ZHYJ21, CCZ^{+}25]. Resource-Aware [LCJ^{+}24, XYM23]. Resource-Constrained [KLW^{+}25, LFGD25, LZC^{+}24]. Resource-Scarce [ZLC^{+}22]. Resources [HZR^{+}23, XTWW25, ZCW^{+}21]. Response [CRJZ21, LJZ^{+}25, RACB24, SNRB23, ZABHCG23]. Response-Hiding [LJZ^{+}25]. Response-Time [RACB24]. Responsiveness [LWL^{+}22, ZYXD20]. Restoring [AVK20]. Results [CTM^{+}25, GNH20]. Retargeting [IIEKS24]. Retention [YWC^{+}21, YH20]. Retention-Time [YH20]. Rethinking [DTZ^{+}25, OAK^{+}23]. Retire [XTWW25]. Retrieval [LMM^{+}22, YZM^{+}25]. Return [YHV^{+}21]. Reusability [OKC^{+}20, PYS20]. Reusability-Based [PYS20]. Reuse [CNOS22, HLQ^{+}23, JM21, PHC24, SIR20]. Reuse-Centric [CNOS22]. Reusing [JZH^{+}24]. Reveal [XGZ^{+}24]. Revealing [MTV^{+}21, WLW^{+}22c]. Reverse [SPMP20, XTWG23]. Reverse-Engineering [XTWG23]. Reversible [BYM22]. Reviewers [Ano20b, Ano25]. Revisit [WZJ^{+}24]. Revisited [LSW25]. Revisiting [DVV23, MCT22, RBSG23, ZWO^{+}25]. ReViT [ZCS^{+}25]. Revocable [GLGL23, ZXZ^{+}24]. Reweighted [WYSL22]. RGKV [SJYQ25]. Rigorous [VHL20]. Ring [LAKS20, VLPS25]. Ring-LWE [LAKS20]. Ring-Shaped [VLPS25]. Rings [BPM23]. RISC [ABP22, CWS^{+}24b, CKK^{+}22, FHL^{+}22, FRFM^{+}25, GCR^{+}23, GFB^{+}24, HMJ24, JDH^{+}25, KKL^{+}25, KGHRM23, SMP22, SZHB21, TAP^{+}25, TDH^{+}23, WLW^{+}22c, WWX^{+}24, ZHLR22]. RISC-V [ABP22, CWS^{+}24b, FHL^{+}22, FRFM^{+}25, GCR^{+}23, GFB^{+}24, HMJ24, JDH^{+}25, KKL^{+}25, KGHRM23, SMP22, SZHB21, TAP^{+}25, TDH^{+}23, WWX^{+}24, ZHLR22]. Risk [EGMW21]. Risks [HLC^{+}22]. RL [WSLX24]. RL-Based [WSLX24]. RLWE [KDE^{+}24]. RLWE-Based [KDE^{+}24]. RM-TB [WWL^{+}25]. RNA [CHC25]. RNA-Seq [CHC25]. RNS [SYL^{+}23, SPMP20]. RO [YWQ^{+}25]. Robotic [LHL^{+}23]. Robotics [HR22, PMA^{+}24]. Robots [LXW^{+}23]. Robust [EGP24, HHN^{+}23, JMW^{+}24, LGC^{+}23, LZW^{+}24b, LWW^{+}25, QSC^{+}25, RSA^{+}20, SNA^{+}20, SCY^{+}23, TSM^{+}21, YCS^{+}24, YWQ^{+}25, ZLL^{+}24a, ZTY^{+}23]. Robustness [FXJW25, PPQBA21, ZJJ25]. ROCKY [TSM^{+}21]. ROLLED [HKC^{+}23]. ROLLO [HWG^{+}23]. Rollup [WFL^{+}25]. ROLoad [TLC^{+}24]. ROLoad-PMP [TLC^{+}24]. Roofline [SDR^{+}22]. Root [Bru20, Bru23, HSE^{+}24, LSW^{+}23, NAP^{+}20]. Rooters [BN24]. Rootkit [LCHL21]. Rotation [GLB21, HJX^{+}25]. Round [BFC20, BLM21, SAJA21, WY25, ZLW^{+}24, ZWZZ24]. Round-Off [BFC20]. Round-to-Nearest [BLM21]. Rounded [God20]. Rounds [DPS22]. Route [HMK^{+}21]. Router [HYW^{+}21]. Routers [LDLK22, PC24]. Routes [CCZ^{+}22]. Routines [CTY^{+}24]. Routing [FXC^{+}23, MSLY24, WLD^{+}25]. Row [ZGWY22]. RowHammer [HZC^{+}25]. RRAM [KYS^{+}22, WFT^{+}21]. RRAM-Based [KYS^{+}22]. RRPN [XCZ^{+}22]. RRPN-Based [XCZ^{+}22]. RSA [DVA22, RSZ23]. RSQC [LSH25]. RT [JYF^{+}23]. RTL [ABP25, FWM^{+}23, HA25, LM21]. RTSA [WSQ^{+}25]. Rule [ZLWJ23]. Rule-Based [ZLWJ23]. Run [AhRX^{+}20, CBB^{+}21b, HMMP23, PK23, SMFS21, WSQ^{+}25]. Run-Through [WSQ^{+}25]. Run-Time [AhRX^{+}20, CBB^{+}21b, HMMP23, PK23, SMFS21]. Runge [BFC20]. Running [LBC24]. Runtime [CZC^{+}21, DYC^{+}25, JLD^{+}25, LWX^{+}25, LHL^{+}23, SPDQ22, WLQ^{+}21, WCZ^{+}24, WLC^{+}24, XXC^{+}25, YYW^{+}24]. RUPA [ZLWJ23]. RuYi [HSH^{+}25]. RV [KKL^{+}25]. RV-CURE [KKL^{+}25]. RvDfi [FHL^{+}22]. RVV [PCA^{+}24].

S [TRG^{+}24, CLCL22, JBK24, KKS^{+}22, PVB21, RPMH21, RMTA20]. S-Box [CLCL22, JBK24]. S-Boxes [RMTA20]. S-FLASH [KKS^{+}22]. S-NUCA [RPMH21]. S-SMART [PVB21]. S2 [YFC^{+}22]. SaaS [ZDY^{+}23]. Saber [DMG23, WWL^{+}23, ZHLR22]. Saca [TPWY23]. Saca-AVF [TPWY23]. SADIMM [LCM25]. SAFA [NNH^{+}25, WHL^{+}21]. Safe [BOL^{+}25]. SafeDRL [ZQG^{+}24]. Safety [KMH^{+}23, KKL^{+}25, KBR^{+}23]. SAFLA [RDS23]. SAL [HCKK25]. SAL-PIM [HCKK25]. SAMBA [KAWR23]. Same [PCA^{+}23]. Sample [CLLdS25, XXJ^{+}24]. Sampler [ZSS20]. Sampling [CTZ^{+}24, Gha21, KAA22, KLL21, WXL^{+}23, WSHJ23]. Sampling-Based [CTZ^{+}24, WSHJ23]. Sandbox [GZG^{+}23]. Sanitization [WHL^{+}23]. SAT [IIEKS24, ZG23]. SATA [WAN^{+}25]. Satisfaction [YHW^{+}25]. Satisfiability [TWJ^{+}22]. Saving [CBB^{+}21b, NLC^{+}25]. Savings [XST20]. Scalability [HR22, WXL^{+}23, ZXL^{+}24]. Scalable [CYX^{+}23, DSJ^{+}22, DRY^{+}22, DKZ^{+}25, EAMJ^{+}23, FPX^{+}25b, GMZ22, GXY^{+}23, LWX^{+}25, LQC^{+}22, LM21, MKYP21, RBMG22, RCAB23, WGL^{+}20, XNLX20, YZJ23, ZCP22]. Scalar [BMBM20, NTL^{+}24, YLZ^{+}24]. Scale [AB20, FTV25, HYQ^{+}25, JKHL22, KPD^{+}23, LZC^{+}21, LDG^{+}22, LWH^{+}24, LCY^{+}25, SYW^{+}22, SCY^{+}23, WCD25, WAN^{+}25, WWL^{+}25, YCKW20, YCY^{+}24, ZCJ^{+}20, ZCSJ23]. Scale-up [AB20]. Scaling [CSY^{+}25, GLMZ25, GLW^{+}24, HYQ^{+}25, LHK^{+}22, WFW^{+}20, ZZL21, ZSS^{+}22]. Scalpel [LMZ^{+}25]. Scan [CDRS20, CKRP21, IBB^{+}25]. Scarce [NNH^{+}25, ZLC^{+}22]. Scarcity [CZR22]. SCARF [EGP24]. Scattered [SLS^{+}21]. SCAUL [RAD20]. Scavenger [ZWO^{+}25]. Scenario [HHZ^{+}23, KH23, ZHYJ21]. Scenario-Based [HHZ^{+}23, KH23]. Scene [XCZ^{+}22]. Schedulability [LJY21, WLZ^{+}25, XAP20]. Scheduled [CFWC23]. Scheduler [GYS^{+}24, SSK22]. Schedulers [PMA^{+}24]. Schedules [SWR^{+}25].

Scheduling [BLP^{+}22, BBL22, CSH^{+}24, CR24, CSvdBU22, CCY^{+}24, DRY^{+}22, DPS^{+}20, GPRV23, GVN25, GQH21, GPQ22, GPQ23, IWKB22, JSTG20, JCY^{+}23, JWS^{+}21, KMAA25, KKB^{+}22, LKK^{+}21, LL22, LHL^{+}21, LLW^{+}24, LJH^{+}25, LLD^{+}25, LSU^{+}23, LYF^{+}22, LMZ^{+}25, LAPB21, MBP21, MÁJG^{+}24, BCLC25, RACB24, RDS23, SNN21, SKLR22, STW^{+}21, STK23, SGL^{+}20, STQ^{+}24, UGvdBC23, VSG^{+}23, WCB23, WZW^{+}23, WZH^{+}23, WLF^{+}25a, WLZ^{+}25, WZGT22, XAP20, XNH^{+}25, YWX^{+}23, ZDZ^{+}23, ZCR23, ZNW^{+}24, ZYZ^{+}23, ZLZ24b, ZML^{+}25, ZLT^{+}25, ZLL^{+}22b, ZLZ^{+}23, ZHYJ21]. Scheme [ABC^{+}24, BHW^{+}23, BBC^{+}20, CLZG22, CERMH23, DZC^{+}24, DWLF25, HZK24, HJX^{+}25, LTL^{+}25, LHR^{+}23, LYW^{+}23, LDF^{+}25, LHZR25, MSW^{+}21, MKH^{+}21, ODK20, PYS^{+}23, QSC^{+}25, SNRB23, STZ^{+}24, STYQ24, SST^{+}24, TDZ^{+}25, XLL^{+}22, YH20, YYQ^{+}24, ZCX^{+}20, ZCF20, YAG20]. Schemes [BZW^{+}25a, HWG^{+}23, NLC^{+}25, Rot24, SYL^{+}23, YHH^{+}25]. Schnorr [BSRP21]. Schnorr-Based [BSRP21]. Scientific [LWC^{+}22, MDPM24, WCL^{+}23]. SciNet [TCJ23]. Scrabble [ZZG20]. Screening [BCF^{+}25]. SDN [TDZ^{+}25]. SE [YWQ^{+}25]. Search [BCKS22, JQY^{+}25, JXH^{+}22, KSL^{+}22, KPD^{+}23, LZF21, LJZ^{+}25, LMM^{+}23, LGZ^{+}25, LLK^{+}23, LCGH25, LZW25, MKY^{+}24, MZM^{+}25, RBC^{+}23, SHZ^{+}25, XLS^{+}24, XGMJ25, YWQ^{+}25, ZYQ^{+}24, dSdCF22]. Search-Efficient [YWQ^{+}25]. Searchable [LJZ^{+}25, XCG^{+}25, YWQ^{+}25]. Secret [SL23, AOM^{+}21]. Secrets [CLCC25, CY22]. Section [CDP21]. Sections [CSvdBU22]. Secure [CCT^{+}20, CYX^{+}23, CZH^{+}25, CXL^{+}23, CCY^{+}24, DMX^{+}22, DTL^{+}25, GJN^{+}25, JMW^{+}24, JLD^{+}25, JQK^{+}24, JHMM23, LHR^{+}22, LZS^{+}24, LWX^{+}25, LQN^{+}21, LSS25, LKMJ21, LWZZ25, LCC^{+}24, NAP^{+}20, PD21, RCC^{+}25, SRB23, SKR^{+}20, SHZ^{+}24, VTM^{+}20, WHC20, WCZ^{+}24, XLWO23, XGZ^{+}24, XKS21, XZC^{+}23, XCG^{+}25, YZM^{+}25, YYQ^{+}24, YZY^{+}25, ZSC^{+}23, ZCY^{+}25]. Securing [EGP24, TLC^{+}24]. Security [AAM^{+}25, APH^{+}23, Ano23, AW20, BSRP21, FHL^{+}22, FWR^{+}20, JLZ^{+}23, MDJ20, MHS^{+}20, ODK20, RGD^{+}24, RWCC23, WSS^{+}20, YWQ^{+}25, ZSS^{+}22, ZLH^{+}21, ZGG25]. Security-Centric [ODK20]. Security-Critical [JLZ^{+}23]. Security-First [MHS^{+}20]. Segment [CKP^{+}22]. Segmentation [FRFM^{+}25, HLJ^{+}25, KMVD22, LHZR25, WSC^{+}25, ZGWY22]. Segmented [KK25]. Selecting [LDW^{+}25]. Selection [CYPC25, HSE^{+}24, LSW^{+}23, RCS^{+}21, SLY^{+}22b, TKN23, XWL^{+}24, ZSX^{+}24]. Selective [CFA22, LRL22, SKR^{+}20]. Self [AGSD25, BKHY22, ERKP21, FBH^{+}22, FW23, KSB24, WLZ^{+}25, XNB21, ZTZ^{+}25]. Self-Adaptive [ZTZ^{+}25]. Self-Configuring [BKHY22]. Self-Dual [ERKP21]. Self-Organizing [AGSD25]. Self-Refreshable [FW23]. Self-Similarity [XNB21]. Self-Similarity-Based [KSB24]. Self-Suspending [WLZ^{+}25]. Self-Synchronizing [FBH^{+}22]. Semantic [HLJ^{+}25, SZHB21, WSC^{+}25, ZCS^{+}25]. Semantic-Aware [ZCS^{+}25]. Semantically [AOM^{+}21]. Semi [CBB21a, GPRV23, WHL^{+}21, XXJ^{+}24]. Semi-Asynchronous [WHL^{+}21]. Semi-Black-Box [XXJ^{+}24]. Semi-Partitioned [CBB21a]. Semi-Periodic [GPRV23]. Sensing [BCF^{+}25, LZX^{+}25, SSY^{+}21]. Sensitive [FQYS23, TLC^{+}24, ZLL^{+}22b]. Sensor [HCC^{+}25]. Separable [LCL^{+}20]. Separated [ZWO^{+}25]. Seq [CHC25]. Sequence [CDRS20, CMQ^{+}22, KLR^{+}20, QHZ^{+}21, TZY^{+}24, TWJ^{+}22, ZTY^{+}23]. Sequence-Based [TZY^{+}24]. Sequences [MYGA20, WYX^{+}24]. Sequencing [CPB21]. Serial [ERKP21, Ima21]. Serialized [WLW^{+}22a]. Server [ZYZ^{+}23, ZGL^{+}21]. Serverless [CSH^{+}24, CCZ^{+}25, HMZ^{+}25, LGL^{+}24, LWX^{+}25, WLF^{+}25a, WDZ^{+}24, XQC^{+}22, XZL^{+}23, ZFL^{+}22]. Serverless-Based [CCZ^{+}25]. Servers [LLL^{+}25a, LZX^{+}25, WLR20, ZZF^{+}24]. Service [BY22, GQJ^{+}22, HYL^{+}25, LCC^{+}24, MXY^{+}23, NT23, SLY22a, ZXY^{+}24, ZZF^{+}24, ZWC^{+}23]. Service-Oriented [MXY^{+}23]. Services [JMW^{+}24, JZY^{+}23, LFGD25, LQN^{+}21, LLX^{+}24a]. Serving [KLW^{+}25, YYW^{+}24, ZDZ^{+}23, ZHY^{+}25]. Set [HLL^{+}20, HMJ24, JRL25, LLC^{+}24, LWW^{+}24, SXH^{+}24, ZGWY22, ZYZ^{+}25a]. Set-Associative [LWW^{+}24, SXH^{+}24]. Sets [WGM^{+}20, WWX^{+}24]. Setting [HXGR20]. Setup [vSDHA23]. Seven [YLL^{+}20]. SFPoW [DTL^{+}25]. SGD [ACKA23, QJY^{+}25]. SGX [CDF^{+}21, JQY^{+}25, WZSL22, WLC^{+}24]. SGX-Friendly [WLC^{+}24]. Shadow [ZLW^{+}24, ZWZZ24]. Shadows [TDMP23]. Shape [MLW^{+}23]. Shaped [VLPS25]. Shard [ZCH^{+}24]. Sharding [HWL^{+}24, NT23, XZL^{+}25, YDW^{+}25, ZCH^{+}24]. Share [XGZ^{+}24]. Share-Transform-Reveal [XGZ^{+}24]. Shared [BY22, BY24, CLZG22, KPH^{+}25, LHK^{+}22, LMZ^{+}25, RCAB23, SWR^{+}25, XAP20, XLY^{+}22, ZCW^{+}21]. Shares [XGZ^{+}24]. Sharing [GZG^{+}23, HWL^{+}21, LHR^{+}23, LZX^{+}25, PE22, WZH^{+}23, YYQ^{+}24, YDL^{+}25, ZZG^{+}23, ZOH^{+}25]. Shift [KOH^{+}23]. Shifting [OLD^{+}23, OAB^{+}23]. Shingled [CLY22, ZYXD20]. Shooting [KBR^{+}23]. Short [BAM^{+}24, KXGS22]. Shot [MZM^{+}25]. Shuhai [HWZ^{+}22]. Siamese [WNL^{+}23]. Side [BWSG25, BKS22, CLCC25, DYPZ22, FWM^{+}23, HPGM20, HLS^{+}23b, JCKH22, KLR^{+}20, LHFW24, LOM^{+}25, OD23, RAD20, SSP^{+}24, SXH^{+}24, TOF^{+}24, UYZP22, XPR^{+}22, YWC^{+}24a, ZYD^{+}20, ZWX^{+}25, ZXL^{+}23, ZGG25]. Side-Channel [BWSG25, CLCC25, DYPZ22, HPGM20, JCKH22, KLR^{+}20, LOM^{+}25, RAD20, SSP^{+}24, SXH^{+}24, UYZP22, XPR^{+}22, ZYD^{+}20, ZXL^{+}23, ZGG25]. Sidechains [DTL^{+}25]. SIDH [CVOJRH22, LLWZ23]. Sifter [ZWL^{+}25]. Sight [ZLT^{+}24]. Signal [ABP22, FLF20, KLC20, UYZP22, WHC^{+}23]. Signaling [IDFH22]. Signals [SNA^{+}20]. Signature [AMJ^{+}23, WDC^{+}25]. Signed [USS^{+}21, SNRB23]. Significant [LLFT23, ZMH^{+}25]. SIKE [SAJA21, EAMK22, TWL^{+}22, ZYD^{+}20]. Silent [BLP^{+}22, LV23, PG23]. Silicon [GZC^{+}21, LLJ^{+}23, SZAT22, WLQ^{+}21]. Sim [SM22]. Sim-D [SM22]. SimBU [KSB24]. SIMD [CCG^{+}22, SM22]. Similarity [CPB21, FWR^{+}20, KSB24, XNB21]. Similarity-Based [CPB21]. Simple [MKH^{+}21]. Simulation [AKG^{+}20, BCMT23, FFG^{+}25, GPH20, JDCL23, JLP^{+}25, LDG^{+}22, RPS^{+}21, RMR22, WMG^{+}25]. Simulations [DWW25, PS22]. Simulator [QHT^{+}24, WAN^{+}25]. Simulators [NKN^{+}25]. Simultaneously [LWNC22]. Single [BGHR^{+}25, CKP^{+}22, DT20, FW23, FHW^{+}22, JJKP22, KvL22, MKY^{+}24, PCA^{+}24, SZHB21, YNJS21, ZWZZ24]. Single- [PCA^{+}24, YNJS21]. Single-Cycle [FW23]. Single-Domain [MKY^{+}24]. Single-Event [BGHR^{+}25]. Single-Flux-Quantum [FHW^{+}22]. Single-Issue [SZHB21]. Single-Key [ZWZZ24]. Single-Machine [JJKP22]. Single/Multi [CKP^{+}22]. Single/Multi-Segment [CKP^{+}22]. Situ [KJK24, LY20, LZW25]. Size [WGD^{+}22]. Sized [MLW^{+}25]. Sketch [GLS^{+}25, ZXMX25, ZGR^{+}25]. Sketch-Based [ZXMX25]. Sketches [RMO21, WLZ^{+}23]. Skewed [WSG^{+}25, ZCY^{+}25]. Skippable [LWW^{+}24]. Sky [ZCSJ23]. Sky-Sorter [ZCSJ23]. Skyline [DWL^{+}22]. SLA [IWKB22, YTD^{+}21, ZGB^{+}21]. SLA-Based [IWKB22]. SLA-Driven [YTD^{+}21]. Slack [ZZ25]. Slice [FWZ^{+}21]. Sliced [DVV23]. Sliding [DWL^{+}22]. SLOpt [ZHY^{+}25]. SM [ZCCG23]. Small [BLM25, JYM20]. Smart [BBJR21, DYC^{+}25, FLF20, HJYL22, KK25, ROPdlT22, TOF^{+}24, ZWB^{+}22, CCZ^{+}22, PVB21]. SmartNIC [LDF^{+}24]. Smartphone [VTM^{+}20]. SmartZone [JLD^{+}25]. Smooth [DVA22]. SMR [MSW^{+}21]. Snapshot [DTH^{+}24]. SNARK [QCX^{+}23]. Snitch [ZSHB21]. SNN [JKK^{+}22]. SNRC [LRL22]. SNs [WNL^{+}23]. SoC [ABI^{+}25, CTZ^{+}24, RMKO23, TDH^{+}23, WCZ^{+}24]. Social [SSY^{+}21]. SoCs [FBM21, IBB^{+}25, JYF^{+}23, RPB^{+}23, SPB^{+}21, SBP^{+}20, VRR^{+}24]. Soft [BPJ^{+}22, BCMT23, RMO21]. Softcore [USS^{+}21]. Software [CQ22, DPS^{+}20, FTV25, GKFF20, GWZ^{+}21, GLS^{+}25, JLZ21, KMH^{+}23, LCY^{+}25, LAKS20, MZZC22, PYS^{+}23, PMA^{+}24, TWL^{+}22, TLLL25, WZF^{+}24, WZD^{+}20, XXC^{+}25, XCZ^{+}22, XLW^{+}20, YLC^{+}21, ZFZ^{+}21, ZGG25, ZZZ^{+}20]. Software-Defined [FTV25, YLC^{+}21]. Software-Hardware [LCY^{+}25, PYS^{+}23, ZFZ^{+}21]. Software/Hardware [WZD^{+}20, ZZZ^{+}20]. Solid [HKS20, HS22, YCKW20]. Solid-State [HKS20, HS22, YCKW20]. Solomon [MCT22, TZ22, YLL^{+}20, YLHL23, YH24]. Solution [ATT22, BXW^{+}25, HLJ^{+}25, LZS^{+}24, PYDG22]. Solutions [SMZ^{+}20]. Solvers [CWS^{+}24a]. Solving [AHK^{+}21]. Some [DPS22]. SongC [LQM^{+}24]. Sort [EDGR^{+}24, MKYP21]. Sorter [ZCSJ23]. Sorting [BLM20, LWNC22, PLB22, XNLX20, ZCSJ23]. SOT [JKK^{+}22, KJK^{+}22]. SOT-MRAM [KJK^{+}22]. Source [BSM21, FQYS23, xHzLH^{+}24, KRB^{+}22, PCA^{+}24, ZGG25]. Source-Dependent [BSM21]. Space [BGHR^{+}25, CPM^{+}23, LHR^{+}22, LLFT23, LL23, MHJ^{+}21, QHZ^{+}21, SMC^{+}25, YLL^{+}24, YLC^{+}21, ZWO^{+}25, ZWL^{+}25]. Space-Efficient [LHR^{+}22]. Space-Time [ZWO^{+}25]. Spanning [DWLF25, WCQW22]. Spark [IWKB22]. Sparse [CCT^{+}20, GLZ^{+}24, GXL^{+}24, HLJ^{+}25, KCL^{+}20, LCM25, LQC^{+}22, LSH25, NNH^{+}25, PCCK22, SLY^{+}22b, TAP^{+}25, WDQ^{+}22, WYX^{+}24, WDL^{+}25, WSQ^{+}25, YFC^{+}22, ZZZ^{+}20, ZLL^{+}24b, ZWX^{+}25]. Sparsification [WYSL22]. Sparsity [GSY^{+}20, KKS^{+}22, KAWR23, KNP^{+}20, LZW^{+}24a, LDT^{+}25, SZL^{+}24, ZZW^{+}25]. Sparsity-Aware [LDT^{+}25]. Spatial [CHL^{+}23, DWL^{+}22, DTZ^{+}25, ESdP^{+}25, LLD^{+}25, WSLX24, ZCZ^{+}22, ZGK20]. Spatial-Keyword [DWL^{+}22]. Spatial-Temporal [ESdP^{+}25, LLD^{+}25]. Spatio [BLKK23, MHDMEA22, SPB^{+}21, ZZG^{+}23]. Spatio-Spectral [MHDMEA22]. Spatio-Temporal [BLKK23, SPB^{+}21, ZZG^{+}23]. SPDL [XZC^{+}23]. SPEC [WFH^{+}24]. Special [Ano23, AW20, BBJR21, BCCM23, CDP21, CQ22, FAKM22, LCZ22, Lu21, QWK20, WS20]. Specialized [AHK^{+}21]. Specific [AKG^{+}20, WS20]. Specification [GZC^{+}21, HZM^{+}23]. Specification-Driven [GZC^{+}21]. Spectral [JWK^{+}23, MHDMEA22]. Spectre [LG22, ZLH^{+}21]. Speculation [ZLH^{+}21]. Speculative [AG22, CY22, GCR^{+}23, PQG^{+}22, PVB21, SKR^{+}20, ZWL^{+}25]. Speech [LWW^{+}25, YAG20]. Speed [BHK^{+}23, DMG23, DA22, TWZ^{+}23, VSG^{+}23, WWL^{+}23]. SPHINCS [WDC^{+}25]. SPICE [vSDHA23]. SpikeBASE [SZ22]. Spiking [AGQ^{+}23, ESdP^{+}25, KMVD22, LRRK^{+}22, PKPR23, RSR22, SZ22]. Spin [JGD^{+}21, JCY^{+}23, ST23a, ST23b]. Spin-Variable [ST23b]. Spins [OTTT22]. Spintronic [VJWZ^{+}21]. Split [FXJW25, GL24, QCX^{+}23]. Split-Radix [GL24]. Split-Star [FXJW25]. SplitDB [CJSY24]. Splitting [CBB21a]. SpMM [SSCK25]. SpMV [DLK25, TLLL25]. SQL [WPL^{+}23]. SQL-Empowered [WPL^{+}23]. SQLite [WSS^{+}20]. Sqrt [KPH^{+}25]. Square [BN24, Bru20, Bru23, HSE^{+}24, LSW^{+}23]. Squarers [CXW^{+}23]. Squaring [MÖS22]. SRAM [JPHY20, NTL^{+}24, ZCWC23]. SRT [LSW^{+}23]. SSD [BJM^{+}21, GLW^{+}24, HKS20, KJC^{+}21, KLKK23, KAA20, MKYP21, PM20, SLLS25]. SSD-Assisted [BJM^{+}21]. SSDs [KOT^{+}23, STYQ24, SST^{+}24, WZW^{+}23, WZW^{+}24, WJL^{+}20]. SSL [BZW^{+}25b]. SSL/TLS [BZW^{+}25b].

Stability

[CTM^{+}25, KYS^{+}22, LMDC21]. Stable [AGSD25, YZJ23, ZABHCG23]. Stack [JLZ21, ZTZ^{+}25]. Stacked [HCC^{+}25, WTL^{+}21]. Stage [APV22, JLP^{+}25, KPS^{+}24, STZ^{+}24]. Staleness [ACKA23]. Stalls [YBW21]. Standard [DPCL22, DLG^{+}24, JDH^{+}25]. Star [FXJW25]. Start [CSH^{+}24]. Startup [LLW^{+}24, LWX^{+}25]. State [BHK^{+}23, HKS20, HS22, Lou20, LLR25, NTDH25, OD23, ODK20, TDZ^{+}25, WRT^{+}22, XZL^{+}23, YCKW20]. State-Machine [LLR25]. Stateful [SCC21, XZL^{+}23]. Statements [HA25]. States [SLOM^{+}23, XKS21]. Static [CFWC23, MBP21, SWR^{+}25]. Stationary [Gha21, ZGWY22]. Statistical [CMGD24, RRMS25, WCL^{+}23, WFH^{+}24]. Statistics [HC24]. Status [BXW^{+}25]. STDP [JKK^{+}22, KMVD22]. STEMS [ESdP^{+}25]. Stencil [LZW23a]. Step [HLLC21]. Stereo [YCS^{+}24, KKB^{+}22]. STFL [BB20]. STFL-DDR [BB20]. STfusion [BLKK23]. STK [XXL^{+}23]. Stochastic [AT23, CWNL22, CWC^{+}24, FHH22, JKK^{+}22, KKB^{+}22, LWH20, SNT22, ZABHCG23]. Stop [AB22]. Stopping [PBBA25]. Storage [BMM^{+}22, BL22, CLZG22, DZC^{+}24, FLS20, FPHW25, GXZ^{+}23, GLGL23, GXZ^{+}24, HGC^{+}22, HXGR20, JKHL22, KPD^{+}23, LZZ^{+}22, LV23, LJY^{+}24, MSW^{+}21, MCT22, NTDH25, PYYG21, SLS^{+}21, SWR^{+}23, SHZ^{+}24, SHZ^{+}25, SLY^{+}22b, WLY^{+}23, WDZ^{+}22, WDZ^{+}23, WHK24, XLL^{+}22, XCG^{+}25, XZL^{+}25, YTD^{+}21, YH20, YZY^{+}25, ZSC^{+}23, ZGD23, ZXZ^{+}24, ZQC^{+}25, ZCX^{+}20, ZCZW23, ZWSF24, ZFH^{+}25, ZZC^{+}23]. Storages [WAN^{+}25]. Store [CLY22, LYH^{+}24, LDF^{+}24, WLW^{+}25, ZGD23, ZXW^{+}24]. Stores [CJSY24, CSY^{+}25, SJYQ25]. Storing [JKKP25]. STR [XGZ^{+}24]. Straggler [LSL^{+}25]. Stragglers [WGT^{+}22, WSG^{+}25, ZGL^{+}21]. Strategies [CVOJRH22, DWW25, GPRV23, JZSD24]. Strategy [CZD^{+}24, CFW25, GQZ21, LZW23a, PBBA25, SKM^{+}23, WLW^{+}22c, XGZ^{+}24, YHW^{+}25, ZGWY22, dSBS^{+}22]. Stratified [JMW^{+}24]. Stratified-Causality [JMW^{+}24]. Stream [BKS22, BMM^{+}22, CZJ21, FBH^{+}22, HLS^{+}23a, SZHB21, SAG22, WT25, WSG^{+}25, ZMS^{+}23, SMC^{+}25, WSG^{+}25]. StreamDFP [HLS^{+}23a]. Streaming [DWL^{+}22, DKZ^{+}25, GWH^{+}23, JWS^{+}21, NTR21, SCY21]. Streamlining [AB20, JWS^{+}23, LPW20]. Strength [KLC20]. Strength-Aware [KLC20]. Stress [AAB^{+}23, CWS^{+}24a]. Strict [ZHY^{+}25]. Strong [GJN^{+}25, PCA^{+}23, XLWO23]. Structure [DLY21, LTFL22, LT25, LZZ^{+}22, LYC^{+}23, ZSWS24]. Structure-Aware [LZZ^{+}22]. Structure-Based [LYC^{+}23]. Structured [ALC25, LHY^{+}21, LDF^{+}25, SCL^{+}24, TAP^{+}25, YYW^{+}24]. Structured-Sparse [TAP^{+}25]. Structures [NTR21]. STT [CFA22, HLLC21, HWC^{+}22b, JKKP25, OAK^{+}23, SMFS21, TSM^{+}21, WRT^{+}22]. STT-MRAM [CFA22, JKKP25, OAK^{+}23, SMFS21, TSM^{+}21]. STT-MRAM-Based [HWC^{+}22b]. STT-MRAMs [WRT^{+}22]. STT-RAM [HLLC21]. Stuck [SKK23]. Stuck-at-Fault [SKK23]. Study [BY24, HZC^{+}25, HWG^{+}23, MKÖ^{+}22, TC21, WHQ^{+}24, XPR^{+}22]. Style [ZLT^{+}24]. Subarray [HCKK25]. Subarray-Level [HCKK25]. Subgraph [LZF21, XYM23]. Subgraph-Level [XYM23]. Subnetworks [FXJW25, XST20]. subQUBO [ATT22]. Subscribe [DWL^{+}22]. Subterranean [MDR^{+}24]. Successive [ZC24]. Succinct [QCX^{+}23]. Suffix [LWNC22, XNLX20]. SUGAR [XYM23]. Suite [KLP^{+}21]. Sunway [LDG^{+}22, WMG^{+}25, YGW^{+}23]. Supercomputer [LDG^{+}22, WMG^{+}25]. Supersingular [SAJA21]. Supply [AAB^{+}23, KLR^{+}20, PB23a, TWY^{+}25]. Support [CLY22, CJYC23, DNMS20, JLD^{+}25, LLS^{+}24, NKL^{+}23]. Supporting [GXY^{+}23, HWX^{+}24, SGL^{+}20]. Supports [CKK^{+}22]. Surface [VBA20, WLHW25]. Surfacing [ZML^{+}24]. SurgeNAS [LLK^{+}23]. Surgery [LLK^{+}23]. Suspending [WLZ^{+}25]. SVDE [HMZ^{+}25]. Svelto [MCS^{+}22]. SVM [LQN^{+}21]. SW [TGA23]. Swapping [KLW^{+}25]. SWEL [NK22]. SWEL-COFAE [NK22]. Switch [LL22, LLCJ23]. Switches [GLS^{+}25, LCJ^{+}24, NWH^{+}25]. Switching [AA20, DH20]. Symbolic [HJYL22, SCL^{+}24]. Symbols [YLL^{+}20]. Symmetrization [BCV22]. Synapses [LY20]. Synaptic [SZ22]. Synchronization [LPYT22, WSC^{+}25]. Synchronizing [FBH^{+}22]. Synchronous [DGX^{+}25]. Synergies [GWD25]. Synergistic [MCD^{+}25]. Synergizing [MAM23]. Synthesis [BCV22, CFC^{+}22, CAC^{+}22, DGTVGG21, DSTD22, HZM^{+}23, HWJ^{+}21, HGC^{+}22, MCS^{+}22, PSBB21, TC21, VCLN21]. Synthesis-Generated [DGTVGG21]. Syscall [OLC^{+}22]. Syscall-Level [OLC^{+}22]. System [ABI^{+}25, ABdGG^{+}25, AKG^{+}20, APK20, BXW^{+}25, CZJ21, CMQ^{+}22, CHC25, CZC^{+}21, CXL^{+}25, CKRP21, CZZ^{+}25, DNMS20, DLZ^{+}24, DGZ^{+}22, FHL^{+}23, GKFF20, GSS^{+}23, HZC^{+}25, HLS^{+}23b, HWL^{+}24, JKHL22, KCS23, KHHK21, LZW^{+}21, LWH^{+}24, LCJ^{+}25, LFX^{+}21, LWYJ23, MKY^{+}24, MIPQ22, NAP^{+}20, QLH^{+}24, QSC^{+}25, TOF^{+}24, WZX^{+}22, WTL^{+}24, WHQ^{+}24, XXC^{+}25, XCZ^{+}22, XZC^{+}23, ZLC^{+}22, ZLWJ23, ZDV^{+}22, ZWY^{+}23, ZGK20, dOCC23]. System-Level [ABI^{+}25, ABdGG^{+}25, AKG^{+}20, HZC^{+}25, NAP^{+}20]. System-on-Chip [AKG^{+}20]. System-on-Chips [ABdGG^{+}25]. Systematic [HLC^{+}22, WSG^{+}23, YH24]. SystemC [FV23]. SystemC-AMS [FV23]. Systems [AJ22, AhRX^{+}20, APH^{+}23, BHE21, BY22, BOL^{+}25, BCCM23, BJMKK23, CB22, CTY^{+}24, CYPC25, CDP21, CSW^{+}21, CSvdBU22, CTZ^{+}24, CFW25, CRJZ21, CDF^{+}21, DWN^{+}22, DSJ^{+}22, DZC^{+}24, DPQK^{+}23, DTH^{+}24, FLF20, FV23, GGZC22, GWC^{+}25, GSK^{+}22, GZW^{+}21, HLL^{+}20, HHPB20, HHZ^{+}23, Has23, HLZ^{+}25, HWC^{+}22a, HF22, HXL^{+}25, HECC^{+}21, HLC^{+}22, HWL^{+}21, HZYY22, HZMC24, HTZ^{+}25, JWD^{+}22, JLZ^{+}23, JDB^{+}23, JYM^{+}23, JWS^{+}23, KD25, KHP21, KHHK21, LKK^{+}21, LCHK22, LLS^{+}25, LHL^{+}21, LLJ^{+}23, LJH^{+}25, LCH22, LL23, LLL^{+}20, LCL^{+}20, LYF^{+}22, LJY^{+}24, LZW^{+}24b, LZC^{+}24, LCC^{+}24, MNB20, MKH^{+}21, MB21, MPYJ25, MHK^{+}22, NLC^{+}25, OAB^{+}23, PYS20, PS22, PMA^{+}24, PKPR23, RMKO23, RMR22, RDS23, SMP22, SCFPM22, SMZ^{+}20, SLOM^{+}23, SM22, SGS^{+}21, SKA^{+}22, TRBM22, WLW^{+}21, WWJ^{+}23, WLY^{+}23, WZH^{+}23, WWJ^{+}25, WCD25, WWS^{+}22, XAP20, XLL^{+}22, XCG^{+}25, XZL^{+}25, YWP^{+}23, YTD^{+}21, ZABHCG23, ZCD^{+}22, ZGD23, ZXW^{+}24, ZZ25, ZCW^{+}21]. Systems [ZCZW23, ZWSF24, DPCL22]. Systems-on-Chip [MKH^{+}21]. Systolic [CHL^{+}23, HWX^{+}24, LRB23, PHC24, WFH25, YFC^{+}22]. Systolic-Array [CHL^{+}23].

T [WYX^{+}24]. Table

[Das23, FPHW25, ZXX^{+}25]. Tables [HLK^{+}25, LSW^{+}23]. Tackling [ZZF^{+}24]. Tag [CFA22, HKC21, SXXL24]. Tag-Data [HKC21]. Tagged [SXXL24]. TaihuLight [LDG^{+}22]. Tailored [SSP^{+}24]. Takes [HSP^{+}25]. Taking [JM21]. Tampering [CLCC25]. Tangram [XZL^{+}25]. Tapping [ZFL^{+}22]. Target [BCBS21]. Targeted [JJZW24]. Task [CBB21a, CCCW21, DRY^{+}22, GQZ21, GYZC25, HLZ^{+}25, HECC^{+}21, HZMC24, KSKK23, LL22, LYGC24, LYF^{+}22, LMZ^{+}25, LAPB21, MÁJG^{+}24, RSP^{+}20, RPMH21, RDS23, SSK22, SGS^{+}21, WHM^{+}22, WWJ^{+}25, WWS^{+}22, WSM^{+}24, ZZH^{+}25, ZLT^{+}25]. Task-Level [LL22]. Tasks [AOM^{+}21, BPM23, CSvdBU22, CKP^{+}22, FL21, GPRV23, GPQ22, GPQ23, JSTG20, JGD^{+}21, JLZ^{+}23, JCY^{+}23, JZH^{+}24, LMW^{+}24, LSU^{+}23, LLZ^{+}25, MDM22, NLC^{+}25, RACB24, SPDQ22, SGL^{+}20, STQ^{+}24, WLZ^{+}25, WZGT22, XLS^{+}24, XNH^{+}25]. TB [WWL^{+}25]. TC [BCCM23, CQ22, FAKM22, Lu21, BBJR21, CDP21, WS20]. TCAM [TGA23]. TCAM-Based [TGA23]. Technique [CWNL22, SNT22, ZFH23]. Techniques [BJM^{+}21, DH20, GKT^{+}22, KKH22, PLZ^{+}23]. Technology [BTEC20, LY21, SRP^{+}21, SZK^{+}22, SKA^{+}22, TSM^{+}21]. TEE [LGX^{+}22, WFL^{+}25]. TeeRollup [WFL^{+}25]. Telepathy [LV23]. Temperature [LL23, PAR^{+}22, YBG^{+}22]. Temperature-Dependent [YBG^{+}22]. Temperature-Induced [PAR^{+}22]. Temperature-Prediction [LL23]. Template [GLMZ25]. Temporal [AGB^{+}23, BLKK23, ESdP^{+}25, LLD^{+}25, RBC^{+}23, SPB^{+}21, WSG^{+}23, WNP^{+}22, ZZG^{+}23, ZGK20]. Tenant [KMAA25, WZH^{+}23]. Tensity [JSTG20]. Tensor [FYR^{+}24, GYH^{+}22, HLT^{+}23, LHK^{+}22, LZW^{+}23b, LHZ^{+}24, MLW^{+}23, MLW^{+}25, PNK^{+}23, SLY^{+}22b, WDQ^{+}22, WSLX24, ZWX^{+}25]. Tensor-Train [LHZ^{+}24]. Tensorized [ZLZ24a]. TensorMap [WSLX24]. Term [Mik24]. Termination [KPH^{+}25]. Test [ABI^{+}25, ABdGG^{+}25, BBL22, HZW^{+}24, IBB^{+}25, MHT25, SCL^{+}24, WRT^{+}22, ZG23]. Tester [AAB^{+}23]. Testers [DPS22]. Testing [GZC^{+}21, LM21, WLW^{+}22c, WZCM23]. Tests [CTM^{+}25]. Tetris [WZD^{+}20]. TetriX [ZLZ24a]. Text [HCKK25, SCL^{+}24, XCZ^{+}22]. Their [AMM21, HPGM20]. THEMIS [KMAA25]. Theoretic [DKZ^{+}25, MKÖ^{+}22, ZCF20]. Theory [RBSG23, WLR20, WLF^{+}25a]. Thermal [JDCL23, KKKC20, LDZ^{+}23, SZAT22, WHC^{+}23, WWJ^{+}23, ZAS^{+}22]. Thermal-Aware [LDZ^{+}23, ZAS^{+}22]. Things [DRA21, FHL^{+}23, GKT^{+}22, HC24, LQL^{+}24, LFGD25, LHR^{+}23, QWT^{+}23, SYL^{+}23]. Thompson [Gha21]. Threaded [CFC^{+}22, MCS^{+}22]. Threading [CGS^{+}20, FNS^{+}22]. Throttling [OJ23]. Throughput [DHM^{+}25, HLY^{+}25, HXGR20, LWL^{+}21, LYH^{+}24, LGW^{+}22, LLR25, PYYG21, UMM^{+}20, WJL^{+}20, XCZ^{+}22, ZLT^{+}25]. Throughput-Oriented [PYYG21].

Throughput/Gate [UMM^{+}20]. TIE [CXY24]. Tierd [YTD^{+}21]. Tiered [CSY^{+}25, HWC^{+}22a]. Ties [BLM21]. Ties-to-Even [BLM21]. Ties-to-Zero [BLM21]. Tight [GJ20, KBQ^{+}23]. TightLLM [HLY^{+}25]. Tightly [ZXG^{+}24]. Tile [NCD^{+}25]. Tiled [GXL^{+}24]. Tiler [MSW^{+}21]. Tiling [LZW23a]. Time [AhRX^{+}20, BHE21, BY22, BCCM23, CBB21a, CBB^{+}21b, CSvdBU22, CRJZ21, CKP^{+}22, DSK23, DWN^{+}22, DLG^{+}24, EGP24, FL21, FWZ^{+}21, FHL^{+}23, FQYS23, GPRV23, GQH21, HMMP23, Has23, HIRB25, HF22, HECC^{+}21, HLS^{+}23b, JSTG20, JGD^{+}21, JWD^{+}22, JYF^{+}23, JYM^{+}23, KAA22, KMAA25, KHP21, KH23, KLL21, LJY21, LL22, LZW^{+}21, LW22, LL23, LWZ^{+}25, LYF^{+}22, LRRK^{+}22, MNB20, MSP^{+}21, BCLC25, PS22, PK23, QWT^{+}23, RACB24, RDS23, SZAT22, SCA^{+}25, SMFS21, SSK22, SMY22, STK23, SM22, SGL^{+}20, WZGT22, XNH^{+}25, YH20, YCL^{+}24, ZABHCG23, ZHLR22, ZCP23, ZLW^{+}24, ZWO^{+}25, ZZ25, ZSC^{+}25, ZHY^{+}25, ZHM20, ZCB23, ZSS20]. Time-Borrowable [LW22]. Time-Coded [LRRK^{+}22]. Time-Compositional [Has23]. Time-Memory [ZHLR22]. Time-Predictable [JWD^{+}22]. Time-Sensitive [FQYS23]. Time-Slice [FWZ^{+}21]. TimeCache [OD23]. Timely [XWL^{+}24]. Timerlat [BCLC25]. Times [KJC^{+}21, RPB^{+}23]. Timing [AG22, FFG^{+}25, GCR^{+}23, PAR^{+}22, PSBB21, TTG^{+}23, WSG^{+}23, ZLC^{+}23a]. Timing-Aware [FFG^{+}25]. Timing-Speculative [AG22]. Tiny [JBS^{+}25, ZSHB21]. Tissue [SLOM^{+}23]. TLB [JZSD24]. TLC [WHL^{+}23, WHK24]. TLS [BZW^{+}25b]. TNU [YHC^{+}20]. Toeplitz [Das23]. ToEx [KPS^{+}24]. Together [PCA^{+}23].

Token [KPS^{+}24, SKH^{+}25]. Token-Adaptive [KPS^{+}24]. Tolerance [DSJ^{+}22, JLZ^{+}23, LQL^{+}24, OLD^{+}23, WNL^{+}23]. Tolerant [CXW^{+}23, FXC^{+}23, FXL^{+}24, FTR23, GXZ^{+}24, JZY^{+}23, JZX^{+}25, LRL22, MCT22, MGFY24, RMO21, SWR^{+}25]. Tolerating [WGT^{+}22]. Tool [HWZ^{+}22]. Tools [TC21]. TOP [VRR^{+}24]. Topology [CFW25, QSC^{+}25]. Topology-Agnostic [CFW25]. Torus [FXL^{+}24, TRG^{+}24]. Tower [CLCL22]. TOWERs [LKMJ21]. TPU [SZK^{+}22]. Trace [HZF^{+}24]. Traces [KJC^{+}21]. Tracing [BCLC25]. Track [VLPS25]. Tracking [LWZ^{+}25, NKN^{+}25, SZAT22, SSW^{+}24, UYZP22]. TrackLace [WLZ^{+}21]. Tractable [BBL22]. Trade [ZZL21, ZHLR22]. Trade-Offs [ZHLR22, ZZL21]. Tradeoff [NHW^{+}24, ZDY^{+}23]. Tradeoffs [ZWO^{+}25]. Trading [AG24, HC24, XNL^{+}23]. Traffic [AT23, BOL^{+}25, xHzLH^{+}24, LHK^{+}22, TDZ^{+}25]. Train [LHZ^{+}24]. Training [AHC^{+}20, CCZ^{+}25, CHZ^{+}25, CSK22, DNMS20, GYS^{+}24, HHN^{+}23, HLF^{+}23, JPHY20, JWS^{+}21, JZH^{+}24, LWH^{+}24, LMH^{+}25, LFX^{+}21, LY20, MSSL21, NKL^{+}23, NKA24, RSR22, STZ^{+}24, SPH^{+}23, WGL^{+}20, WWM^{+}23, WL24, WSHJ23, XQC^{+}22, XYM23, ZCK20, ZLL^{+}23, ZDV^{+}22, dSdCF22]. Trajectory [YBF^{+}25]. Transaction [CYX^{+}23, CZZ^{+}25, DQ23, NT23, XXL^{+}23, ZCD^{+}22]. Transaction-Based [XXL^{+}23]. Transactional [DPS^{+}20, PQG^{+}22]. Transactions [Ano20a, Ano23, TBS^{+}25]. Transducers [BLM25]. Transfer [JMW^{+}24, NML25, SJYQ25, ZGG^{+}23, ZLT^{+}24]. Transform [JLL22, LRRK^{+}22, MKÖ^{+}22, XGZ^{+}24, YLHL23, GR23]. Transformations [WZG^{+}23]. Transformer [GWG^{+}25, HCKK25, KPS^{+}24, LQC^{+}22, SKH^{+}25, WYX^{+}24, WSQ^{+}25, WLHW25, ZTY^{+}25, ZCS^{+}25]. Transformer-Based [HCKK25, KPS^{+}24]. Transformers [GWC^{+}25, HLF^{+}23, JBS^{+}25, WFH25]. Transient [FWM^{+}23, JLP^{+}25, LZG^{+}24]. Transiently [LJY21]. Transition [TDH^{+}23]. Transitions [AGSD25]. Translation [CLY22, KOT^{+}23, SZL^{+}22]. Transmission [FPX^{+}25a, LZS^{+}24, ZXZ^{+}21]. Transmissions [SWR^{+}25]. Transparency [ZXL^{+}24]. Transparent [LHN^{+}22]. Transparently [TQL^{+}22]. Transport [YPL^{+}25]. Transpose [JPHY20]. Transposition [GLB21]. Tree [BCKS22, CJSY24, CLY22, LHY^{+}21, SJYQ25, WBJC22, WLD^{+}22, XJL^{+}25, ZCY^{+}25, ZFH23, WBJC22]. Tree-Structured [LHY^{+}21]. TREEHOUSE [SRB23]. Trees [DWLF25, HKC^{+}23, WCQW22, ZWO^{+}25]. Trefoil [WAN^{+}25]. Tremors [MDM22]. Trend [CZD^{+}24]. Triangle [WYZ^{+}22, ZWJ^{+}25]. Triangular [FBH^{+}22]. Trident [ZYZ^{+}25a]. Trinomials [Ima21]. Trivial [CGLS21]. TRNG [PCA^{+}23]. Trojan [HHN^{+}23, HYH^{+}25, PM25, CPM^{+}23, JJZW24, MLL^{+}25, PK23, SKA^{+}22]. Trojans [EGP24, PM25]. Trouble [KBR^{+}23]. Trouble-Shooting [KBR^{+}23]. True [CTZ^{+}24]. Truss [LYC^{+}23]. Trust [LGX^{+}22, LHR^{+}23, NAP^{+}20, ZFQ^{+}23]. Trust-Preserving [ZFQ^{+}23]. Trusted [GQJ^{+}22, GZG^{+}23, LGX^{+}22, WCZ^{+}24, XHY^{+}22]. Trustworthiness [ZDY^{+}23]. Trustworthy [LGX^{+}22]. TrustZone [JLD^{+}25]. Truth [SSY^{+}21]. Truthful [ZWWY22]. TSE [HLLC21]. TTADF [HBS20a]. Tucker [GYH^{+}22, HLT^{+}23].

Tuner [MLW^{+}23, ZWL^{+}25]. Tuning [BMM^{+}22, CXY24, KOH^{+}23, MCD^{+}25, MLW^{+}25, QHT^{+}24]. TurboDL [JWS^{+}21]. TurboGNN [WSHJ23]. Turing [BBdTF25]. TVLA [BWSG25]. Twice [CVOJRH22]. Twin [GSS^{+}23, LGL^{+}24, ZXY^{+}24]. Twin-Assisted [ZXY^{+}24]. Twin-Enabled [LGL^{+}24]. Twins [LCC^{+}24]. Two [HLLC21, JLP^{+}25, SLDZ25, WY25]. Two-Party [WY25]. Two-Phase [SLDZ25]. Two-Round [WY25]. Two-Stage [JLP^{+}25]. Two-Step [HLLC21]. Type [CSK22, LSU^{+}23]. Type-Aware [LSU^{+}23]. Typed [LSU^{+}23]. Types [JM21].

UAV [FZM^{+}23, LHW^{+}25, MSP^{+}21, WLHW25, YBF^{+}25, ZLT^{+}25]. UAV-Assisted [LHW^{+}25]. UAV-MCS [WLHW25]. Ubiquitous [WGM^{+}20]. UKFaaS [CZH^{+}25]. ulp [LLL25b]. Ultra [ABP22, ALC25, AZS^{+}23, WDZ^{+}23]. Ultra-Low [ABP22, ALC25, WDZ^{+}23]. Ultra-Resource [AZS^{+}23]. Ultrasound [WZD^{+}20]. Ultratiny [YWC^{+}24a]. Un-Core [ABI^{+}25]. Un-IOV [ZXL^{+}24]. Unary [FB20, KSB24, SNT22]. Unbiasedness [CQI^{+}22]. Uncomputability [BBdTF25]. Uncompute [ZSWS24]. Uncoupled [ZQY^{+}20]. Uncover [CLCC25]. Uncovering [GWD25]. Under-utilization [GSY^{+}20]. Underlying [PAR^{+}22]. Understanding [GSS^{+}23, SKR^{+}20]. Unified [AMJ^{+}23, ABP25, GYS^{+}24, GVN25, HSE^{+}24, JCY^{+}23, MC23, SSCK25, WDQ^{+}22, WZG^{+}24, Xu24]. Uniformity [KS24, QSC^{+}25, WFT^{+}21]. Unikernel [CZH^{+}25, OLC^{+}22]. Unipolar [CWNL22]. UniSched [GYS^{+}24]. Unit [BZW^{+}25a, Bru20, NWH^{+}25, WLF^{+}25b, ZTY^{+}23, ZCK20]. United [LDT^{+}25, FDKK21]. Units [Bru23, ZXG^{+}24]. Universal [HSP^{+}25]. Universally [LZS^{+}24]. Unleashing [LWL^{+}24]. Unmodified [XLW^{+}20]. Unpredictable [JZX^{+}25]. Unrolled [DH20]. Unsigned [BN24]. Unspent [CZZ^{+}25]. Unstructured [WSM^{+}24]. Unsupervised [JKK^{+}22, KMVD22, RAD20]. Untrusted [GWZ^{+}21, NAP^{+}20, TQL^{+}22]. Update [TBS^{+}25, ZDV^{+}22, ZFH23]. Updates [GSC^{+}23, SCY21]. Upgrade [HYQ^{+}25]. Upgrade-Aware [HYQ^{+}25]. Upsets [BGHR^{+}25]. Usability [CZR22]. Usage [ZXL^{+}24]. Use [HXL^{+}23, KMAA25, MDPM24, PC24]. User [CZW^{+}24, GYS^{+}24, GVN25, PD21, SKM^{+}23, WSC^{+}25, XCG^{+}25, YHW^{+}25]. User-Defined [GVN25]. User-Distribution-Aware [CZW^{+}24]. User-Gateway [PD21]. Users [LHXH22]. Using [AVK20, AA20, APK20, BWSG25, BLKK23, BFG^{+}21, BLM21, BJMKK23, CWWW20, CFWC23, CHL^{+}23, CWNL22, Das23, DYPZ22, DVA22, DWW25, EGMW21, FXL^{+}24, FRFM^{+}25, FHW^{+}22, GWD25, GLZ^{+}24, GZC^{+}21, HA25, HIRB25, HBS^{+}20b, HZM^{+}23, HLT^{+}23, Ima21, KMVD22, KLR^{+}20, LHK^{+}22, LMDC21, LG22, LCM25, LGX^{+}22, LZW^{+}23b, LHZ^{+}24, LWW^{+}24, LGW^{+}22, LFX^{+}21, MZM^{+}25, MHS^{+}20, MGFY24, MTK25, MDM22, OD23, PSM22, PZY^{+}23, PHL^{+}25, PD21, RSZ23, SNA^{+}20, SSM23, SHZ^{+}25, SSZ^{+}20, STQ^{+}24, SKM^{+}23, TBS^{+}25, TWL^{+}22, VCLN21, WLR20, WWX^{+}24, WFL^{+}25, WZD^{+}20, WWL^{+}23, WHY^{+}22, XGZ^{+}24, XGMJ25, XL25, XLY^{+}22, ZMH^{+}25, ZDV^{+}22, ZBT22, dSdCF22, PCBD23, RMR22]. Utility [LGL^{+}24]. Utilization [JSTG20, LWL^{+}25, SZHB21, ZCZ^{+}22, ZCC^{+}23, ZDC^{+}25, GSY^{+}20]. Utilization-Tensity [JSTG20]. Utilizing [CKJ^{+}22, HJX^{+}25, SMZ^{+}20].

V [ABP22, CWS^{+}24b, FHL^{+}22, FCZ^{+}23, FRFM^{+}25, GCR^{+}23, GFB^{+}24, HMJ24, JDH^{+}25, KKL^{+}25, KGHRM23, SMP22, SZHB21, TAP^{+}25, TDH^{+}23, WWX^{+}24, ZHLR22]. V-WAFA [FCZ^{+}23]. VALIANT [SSP^{+}24]. Value [BSM21, BL22, CJSY24, CSY^{+}25, CLY22, CWS^{+}24a, EDGR^{+}24, JKNK24, LYH^{+}24, LDF^{+}24, WLW^{+}25, YZJ23, ZGD23, ZXW^{+}24, ZWO^{+}25, ZSX^{+}24]. Value-Aware [JKNK24]. Value-Deviation-Bounded [BSM21]. Valued [DGX^{+}25]. Values [ZGR^{+}25]. Vandermonde [YH24]. Variability [CTM^{+}25]. Variable [ABC^{+}24, GFB^{+}24, HZR^{+}23, MTK25, ST23b, WYX^{+}24]. Variable-Length [WYX^{+}24]. Variant [WBJC22]. Variants [SYD^{+}24]. Variation [CSW^{+}21, FCZ^{+}23, KKB^{+}22, WJL^{+}20, ZDW^{+}23]. Variation-Aware [ZDW^{+}23]. Variational [SZS^{+}22]. Variations [PB23a]. Varying [VSG^{+}23]. Varying-Speed [VSG^{+}23]. VCMalloc [HZT^{+}23]. VDF [MÖS22, ZTLW23]. ve [RGS22]. VecQ [GCL^{+}21]. Vector [DNMS20, KCL^{+}20, LLS^{+}24, LSS25, NTL^{+}24, PCA^{+}24, TAP^{+}25, ZC24]. Vector-Indistinguishability [ZC24]. Vectorizations [NS22]. Vectorized [GCL^{+}21]. Vehicle [HLZ^{+}25, LCHK22]. Vehicular [BSRP21, ZLM^{+}24]. Verifiable [LMM^{+}22, LJZ^{+}25, SHZ^{+}25, TWaKo^{+}23]. Verification [CGLS21, DSP^{+}21, DLZ^{+}24, PL21, PKPR23, RMKO23, YZY^{+}25]. Versatile [LHXH24, ZXG^{+}24]. Version [CYPC25, GXZ^{+}23, VTM^{+}20]. versus [BPJ^{+}22]. Vertex [JXH^{+}22, MB21]. Very [IBB^{+}25]. VF [RCS^{+}21]. VF-Selection [RCS^{+}21]. Via [DSJ^{+}22, GZG^{+}23, HKC21, LLT^{+}23, UYZP22, WWM^{+}23, ZZL21, ZCR23, ZGL^{+}21, BPM23, BLM20, BJMKK23, CCZ^{+}22, CLY22, CRJZ21, DLK25, FWZ^{+}21, GWG^{+}24, GWG^{+}25, HLL^{+}20, HLY^{+}25, HWR^{+}24, JRL25, JLL^{+}20, JMW^{+}24, KPS^{+}24, KZS^{+}25, KLW^{+}25, LPYT22, LLL^{+}23, LWL^{+}25, LSL^{+}25, LFGD25, LWL^{+}24, LZX^{+}25, LWZZ25, MZZC22, NTL^{+}24, PM20, QJY^{+}25, SDR^{+}22, TRG^{+}24, TWZ^{+}23, TTG^{+}23, WFW^{+}20, WDQ^{+}22, WCB23, WSC^{+}25, WJL^{+}20, XXZ^{+}25, XLS^{+}24, XZL^{+}21, XYM23, YWC^{+}21, YLG^{+}23, YPD^{+}24, YHH^{+}25, ZLT^{+}24, ZCCG23, ZLL^{+}22b]. Vibration [MDM22]. Vibration-Based [MDM22]. Video [GWH^{+}23, GYH^{+}22, HLS^{+}23b, HMZ^{+}25, LWYJ23, QWT^{+}23, WSQ^{+}25, ZCJ^{+}20, ZZF^{+}24]. Videos [YCL^{+}24]. View [HLC^{+}22, SLDZ25, WSC^{+}25]. Violation [ZGB^{+}21]. Virtual [AY24, BYZZ20, CZD^{+}24, CJYC23, FLF20, GZC^{+}21, HLK^{+}25, MSLY24, NTDH25, WL20, ZWC^{+}23, LCHL21]. Virtual/Silicon [GZC^{+}21]. Virtualization [DPCL22, JZSD24, JYM^{+}23, LHS^{+}25, PYYG21, PYDG22, SMP22, SZL^{+}22, TDMP23, YLT^{+}23, ZZG^{+}23, ZXL^{+}24]. Virtualized [FNS^{+}22, PLZ20]. Virtually [HZT^{+}23]. VISE [CDF^{+}21]. Vision [GWG^{+}25, HCC^{+}25, SMZ^{+}20, SKH^{+}25, ZHL^{+}24, ZTY^{+}25, ZCS^{+}25]. Visual [GWC^{+}25, LLS^{+}22, LWW^{+}25, WZX^{+}22]. VisualNet [WZX^{+}22]. vKernel [HWR^{+}24]. VLSI [JYM20, TWZ^{+}23]. VM [BXW^{+}25]. VMT [FNS^{+}22]. VNF [XZL^{+}21, ZZM^{+}22]. Vol [Ano20a]. Volatile [BHE21, CWWW20, LJY^{+}24, LY20, NK22, WHX^{+}25, ZCY^{+}25, LLY22]. Volatile/Non [LLY22]. Volatile/Non-Volatile [LLY22]. Voltage [HDAS21, WFW^{+}20, WFT^{+}21, ZSS^{+}22]. Voltage-Noise [HDAS21]. Voltages [TWY^{+}25]. Volume [LJZ^{+}25]. Volume-Hiding [LJZ^{+}25]. Volumetric [YCL^{+}24]. VOQ [PC24]. Voting [HBB^{+}21]. VRBC [TWaKo^{+}23]. VSPIM [NTL^{+}24]. vTrust [TQL^{+}22]. Vulnerabilities [WLW^{+}22c]. Vulnerability [NT23, TPWY23, WZF^{+}24].

WAFA [FCZ^{+}23]. WAL [HKS20]. WAL-SSD [HKS20]. Walk [GWG^{+}24, WXL^{+}23]. Walks [WY25, dSdCF22]. Wall [DGZ^{+}22, MHK^{+}22, OLD^{+}23, LCHL21, WCYK20]. Warp [LHK^{+}22]. Wash [HGC^{+}22]. Watermark [JYH^{+}24]. Watermarking [SNRB23]. Waterwave [SPH^{+}23]. Wave [WZD^{+}20]. Wavelet [WZG^{+}24]. Wavelet-Based [WZG^{+}24]. WaWoT [LFGD25]. Way [NS22, PLB22]. Ways [DKJP21]. WBMatrix [TGS^{+}22]. WCRT [SGS^{+}21]. Weak [HWC^{+}22b]. Wear [CSW^{+}21, GLW^{+}24, NK22]. Wear-Leveling [CSW^{+}21, GLW^{+}24]. Web [CB22, LFGD25]. WebAssembly [LFGD25]. Weight [GCL^{+}21, LHS^{+}25, LGC^{+}23, PE22, SSJ21, WFZ^{+}25, ZDV^{+}22]. Weight-Aware [PE22]. Weight-Level [WFZ^{+}25]. Weighted [BYM22, CWC^{+}24]. Wheeler [GR23]. While [ZDC^{+}25]. WHISTLE [KMH^{+}23]. WhistleBlower [HZC^{+}25].

White [TGS^{+}22]. White-Box [TGS^{+}22]. Whole [ZCR22]. Whole-Life [ZCR22]. Width [JKNK24, OTTT22]. Window [DWL^{+}22]. Windows [Fic22]. Winner [HSP^{+}25]. Winner-Takes-All [HSP^{+}25]. Wireless [Akr22, ABP22, YBF^{+}25]. Wise [Das23, HYH^{+}25, MRB^{+}24, SYL^{+}23, ZXG^{+}24]. Within [LLL25b, NWH^{+}25, NCD^{+}25]. Without [ZZZ^{+}23]. WL [GLW^{+}24]. WOLF [WFZ^{+}25]. WooKong [WGM^{+}20]. Word [SYL^{+}23]. Word-Wise [SYL^{+}23]. Words [NK22]. Work [DTL^{+}25, VLPS25]. Workflow [WCB23]. Working [WAN^{+}25]. Workload [GWC^{+}25, HIRB25, LDZ^{+}23, MTV^{+}21, PYYG21, STYQ24, ZYXD20, ZHM20]. Workload-Aware [MTV^{+}21, PYYG21]. Workloads [CBB21a, DRP24, FNS^{+}22, GWX^{+}23, ZSHB21, ZDZ^{+}23, ZTY^{+}25, ZZC^{+}23]. World [AOM^{+}21, BGB^{+}21, CY22]. Worst [FL21]. Worst-Case [FL21]. Wrapped [DTL^{+}25]. Write [CKRP21, HKS20, IKTY22, LLS^{+}23, WFT^{+}21, WHX^{+}25, WDZ^{+}22, WSC^{+}24]. Write-Ahead-Logging [HKS20]. Write-Optimized [CKRP21]. Writes [WZW^{+}24]. Writing [WWM^{+}23]. Wrong [RJ24]. Wrong-Path-Aware [RJ24].

x86 [WZCM23]. XACC [NM22].

XeFlow [LPW20]. XMeter [AMM21]. XOR [BCCM22]. xoroshiro128 [HF23]. Xvpfloat [GFB^{+}24].

Zero [BHW^{+}23, BPM23, BLM21, JKNK24, LHR^{+}23, QCX^{+}23, RGS22, YLZ^{+}24]. Zero-Jitter [BPM23]. Zero-Knowledge [BHW^{+}23, QCX^{+}23, YLZ^{+}24]. Zhang [SEM23]. ZigZag [MHJ^{+}21]. zk [QCX^{+}23]. zk-SNARK [QCX^{+}23]. Zoned [WZW^{+}24]. zPerf [WCL^{+}23]. Zweilous [LCX21].

References

Aksoy:2020:NME

[AA20] L. Aksoy and M. Altun. Novel methods for efficient realization of logic functions using switching lattices. IEEE Transactions on Computers, 69(3):427–440, March 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Angione:2023:LCB

[AAB^{+}23] Francesco Angione, Davide Appello, Paolo Bernardi, Claudia Bertani, Giovambattista Gallo, Stefano Littardi, Giorgio Pollaccia, Walter Ruggeri, Matteo Sonza Reorda, Vincenzo Tancorre, and Roberto Ugioli. A lowcost burn-in tester architecture to supply effective electrical stress. IEEE Transactions on Computers, 72(5): 1447–1459, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ahmad:2025:DLB

[AAM^{+}25] Shahnawaz Ahmad, Mohd Arif, Shabana Mehfuz, Javed Ahmad, and Mohd Nazim. Deep learning-based cloud security: Innovative attack detection and privacy focused key management. IEEE Transactions on Computers, 74(6):1978–1989, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Addisie:2020:CAS

[AB20] A. Addisie and V. Bertacco. Collaborative accelerators for streamlining MapReduce on scale-up machines with incremental data aggregation. IEEE Transactions on Computers, 69(8):1233–1247, August 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Attia:2022:SLN

[AB22] Sameh Attia and Vaughn Betz. Stop and Look: a novel checkpointing and debugging flow for FPGAs. IEEE Transactions on Computers, 71(10): 2513–2526, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Angioli:2024:DIE

[ABC^{+}24] Marco Angioli, Marcello Barbirotta, Abdallah Cheikh, Antonio Mastrandrea, Francesco Menichelli, Saeid Jamili, and Mauro Olivieri. Design, implementation and evaluation of a new variable latency integer division scheme. IEEE Transactions on Computers, 73(7): 1767–1779, July 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Angione:2025:SLT

[ABdGG^{+}25] Francesco Angione, Paolo Bernardi, Nicola di Gruttola Giardino, Gabriele Filipponi, Claudia Bertani, and Vincenzo Tancorre. A systemlevel test methodology for communication peripherals in system-on-chips. IEEE Transactions on Computers, 74(2): 731–739, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Angione:2025:AGS

[ABI^{+}25] Francesco Angione, Paolo Bernardi, Giusy Iaria, Claudia Bertani, and Vincenzo Tancorre. Automatic generation of system-level test for un-core logic of large automotive SoC. IEEE Transactions on Computers, 74(9): 3195–3209, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Amor:2022:RVI

[ABP22] Hela Belhadj Amor, Carolynn Bernier, and Zdeněk Přikryl. A RISC-V ISA extension for ultra-low power IoT wireless signal processing. IEEE Transactions on Computers, 71(4):766–778, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Antognazza:2025:EUR

[ABP25] Francesco Antognazza, Alessandro Barenghi, and Gerardo Pelosi. An efficient and unified RTL accelerator design for HQC-128, HQC-192, and HQC-256. IEEE Transactions on Computers, 74(7): 2306–2320, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Azizimazreah:2022:PAD

[AC22] Arash Azizimazreah and Lizhong Chen. Polymorphic accelerators for deep neural networks. IEEE Transactions on Computers, 71(3): 534–546, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ardakani:2020:FEC

[ACG20] A. Ardakani, C. Condo, and W. J. Gross. Fast and efficient convolutional accelerator for edge computing. IEEE Transactions on Computers, 69(1): 138–152, January 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ansari:2021:ILM

[ACH21] M. S. Ansari, B. F. Cockburn, and J. Han. An improved logarithmic multiplier for energy-efficient neural computing. IEEE Transactions on Computers, 70(4):

614–625, April 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Abubaker:2023:MSC

[ACKA23] Nabil Abubaker, Orhun Caglayan, M. Ozan Karsavuran, and Cevdet Aykanat. Minimizing staleness and communication overhead in distributed SGD for collaborative filtering. IEEE Transactions on Computers, 72(10):2925– 2937, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Assare:2022:PAT

[AG22] Omid Assare and Rajesh K.

Gupta. Performance analysis of timing-speculative processors. IEEE Transactions on Computers, 71(2):407– 420, February 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Alam:2024:ADL

[AG24] Irina Alam and Puneet Gupta.

Achieving DRAM-like PCM by trading off capacity for latency. IEEE Transactions on Computers, 73(4):1180–1189, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ansari:2023:MMT

[AGB^{+}23] Ali Ansari, Fatemeh Golshan, Rahil Barati, Pejman Lotfi-Kamran, and Hamid Sarbazi-Azad. MANA: Microarchitecting a temporal instruction prefetcher. IEEE Transactions on Computers, 72(3): 732–743, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aung:2023:DCS

[AGQ^{+}23] Myat Thu Linn Aung, Daniel Gerlinghoff, Chuping Qu, Liwei Yang, Tian Huang, Rick Siow Mong Goh, Tao Luo, and Weng-Fai Wong. Deep-Fire2: a convolutional spiking neural network accelerator on FPGAs. IEEE Transactions on Computers, 72(10): 2847–2857, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Agarwal:2025:MPT

[AGSD25] Sneha Agarwal, Keshav Goel, Mitali Sinha, and Sujay Deb. Mitigation of phase transitions in self-organizing NoC for stable queueing dynamics. IEEE Transactions on Computers, 74(2):623–636, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ankit:2020:PPA

[AHC^{+}20] A. Ankit, I. E. Hajj, S. R. Chalamalasetti, S. Agarwal, M. Marinella, M. Foltin, J. P. Strachan, D. Milojicic, W. Hwu, and K. Roy. PAN-THER: A programmable architecture for neural network training harnessing energyefficient ReRAM. IEEE Transactions on Computers, 69(8):1128–1142, August 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Asgari:2021:ESP

[AHK^{+}21] B. Asgari, R. Hadidi, T. Krishna, H. Kim, and S. Yalamanchili. Efficiently solving partial differential equations in a partially reconfigurable specialized hardware. IEEE Transactions on Computers, 70(4):524–538, April 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Al-hayanni:2020:PPA

[AhRX^{+}20] M. A. N. Al-hayanni, A. Rafiev, F. Xia, R. Shafik, A. Romanovsky, and A. Yakovlev. PARMA: Parallelization-aware run-time management for energy-efficient many-core systems. IEEE Transactions on Computers, 69(10):1507– 1518, October 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ahmadpour:2022:BMM

[AJ22] Zabihollah Ahmadpour and Ghassem Jaberipur. Up to 8k-bit modular Montgomery multiplication in residue number systems with fast 16-bit residue channels. IEEE Transactions on Computers, 71(6): 1399–1410, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Arda:2020:DSL

[AKG^{+}20] S. E. Arda, A. Krishnakumar, A. A. Goksoy, N. Kumbhare, J. Mack, A. L. Sartor, A. Akoglu, R. Marculescu, and U. Y. Ogras. DS3: A systemlevel domain-specific systemon-chip simulation framework. IEEE Transactions on Computers, 69(8):1248–1262, August 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Akram:2022:DDM

[Akr22] Vahid Khalilpour Akram. Distributed detection of minimum cuts in wireless multihop networks. IEEE Transactions on Computers, 71(4): 919–932, April 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Andronic:2025:PUL

[ALC25] Marta Andronic, Jiawen Li, and George A. Constantinides. PolyLUT: Ultra-low latency polynomial inference with hardware-aware structured pruning. IEEE Transactions on Computers, 74(9): 3181–3194, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Almeida:2023:CPB

[Alm23] Paulo Sérgio Almeida. A case for partitioned Bloom filters. IEEE Transactions on Computers, 72(6):1681–1691, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aikata:2023:UCL

[AMJ^{+}23] Aikata Aikata, Ahmet Can Mert, David Jacquemin, Amitabh Das, Donald Matthews, Santosh Ghosh, and Sujoy Sinha Roy. A unified cryptoprocessor for lattice-based signature and key-exchange. IEEE Transactions on Computers, 72(6):1568–1580, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Akram:2021:XFA

[AMM21] Riad Akram, Shantanu Mandal, and Abdullah Muzahid. XMeter: Finding approximable functions and predicting their accuracy. IEEE Transactions on Computers, 70(7):1081–1093, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Aghaie:2020:IC

[AMR^{+}20] A. Aghaie, A. Moradi, S. Rasoolzadeh, A. R. Shahmirzadi, F. Schellenberg, and T. Schneider. Impeccable circuits. IEEE Transactions on Computers, 69(3): 361–376, March 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2020:IIT

[Ano20a] Anonymous. 2019 index IEEE Transactions on Computers vol. 68. IEEE Transactions on Computers, 69(1):1–22, January 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2020:RL

[Ano20b] Anonymous. 2019 reviewers list. IEEE Transactions on Computers, 69(1): 153–157, January 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2023:GEI

[Ano23] Anonymous. Guest editorial: IEEE Transactions on Computers, special issue on hardware security. IEEE Transactions on Computers, 72(2): 305, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Anonymous:2025:RL

[Ano25] Anonymous. 2024 reviewers list. IEEE Transactions on Computers, 74(1): 334–340, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Akmandor:2021:SSE

[AOM^{+}21] A. O. Akmandor, J. Ortiz, I. Manotas, B. Ko, and N. K. Jha. SECRET: Semantically enhanced classification of realworld tasks. IEEE Transactions on Computers, 70(3): 440–456, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Alrahis:2023:BAG

[APH^{+}23] Lilas Alrahis, Satwik Patnaik, Muhammad Abdullah Hanif, Muhammad Shafique, and Ozgur Sinanoglu. PoisonedGNN: Backdoor attack on graph neural networks-based hardware security systems. IEEE Transactions on Computers, 72(10):2822–2834, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Arnold:2020:IRL

[APK20] M. G. Arnold, V. Paliouras, and I. Kouretas. Implementing the residue logarithmic number system using interpolation and cotransformation. IEEE Transactions on Computers, 69(12):1719– 1732, December 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ajirlou:2022:MLP

[APV22] Arash Fouman Ajirlou and Inna Partin-Vaisband. A machine learning pipeline stage for adaptive frequency adjustment. IEEE Transactions on Computers, 71(3): 587–598, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Adusumilli:2023:TCB

[AT23] Vijaya Bhaskar Adusumilli and Venkatesh TG. Traffic characterization based stochastic modelling of networkon-chip. IEEE Transactions on Computers, 72(4): 1215–1222, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Atobe:2022:HAM

[ATT22] Yuta Atobe, Masashi Tawada, and Nozomu Togawa. Hybrid annealing method based on subQUBO model extraction with multiple solution instances. IEEE Transactions on Computers, 71(10):2606– 2619, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Adams:2020:ARD

[AVK20] E. Adams, S. Venkatachalam, and S. Ko. Approximate restoring dividers using inexact cells and estimation from partial remainders.

IEEE Transactions on Computers, 69(4):468–474, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Awad:2020:GEI

[AW20] A. Awad and R. Wang. Guest Editors’ introduction to the special issue on hardware security. IEEE Transactions on Computers, 69(11):1556– 1557, November 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Asanuma:2024:DAL

[AY24] Kota Asanuma and Hiroshi Yamada. DBMS-assisted live migration of virtual machines. IEEE Transactions on Computers, 73(2):380–393, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Armeniakos:2023:CDA

[AZS^{+}23] Giorgos Armeniakos, Georgios Zervakis, Dimitrios Soudris, Mehdi B. Tahoori, and Jörg Henkel. Co-design of approximate multilayer perceptron for ultra-resource constrained printed circuits. IEEE Transactions on Computers, 72(9): 2717–2725, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bingol:2024:GGF

[BAM^{+}24] Zülal Bingöl, Mohammed Alser, Onur Mutlu, Ozcan Ozturk, and Can Alkan. GateKeeper-GPU: Fast and accurate pre-alignment filtering in short read mapping. IEEE Transactions on Computers, 73(5):1206–1218, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Behnam:2020:SDI

[BB20] P. Behnam and M. N. Bojnordi. STFL-DDR: Improving the energy-efficiency of memory interface. IEEE Transactions on Computers, 69(12):1823–1834, December 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Behnam:2022:ARD

[BB22] Payman Behnam and Mahdi Nazm Bojnordi. Adaptively reduced DRAM caching for energyefficient high bandwidth memory. IEEE Transactions on Computers, 71(10):2675– 2686, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Biasielli:2020:NNB

[BBC^{+}20] M. Biasielli, C. Bolchini, L. Cassano, E. Koyuncu, and A. Miele. A neural network based fault management scheme for reliable image processing. IEEE Transactions on Computers, 69(5):764–776, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.
org/document/8955912.

Bolchini:2022:FIE

[BBC^{+}22] Cristiana Bolchini, Giacomo Boracchi, Luca Cassano, Antonio Miele, and Diego Stucchi. Fault impact estimation for lightweight fault detection in image filtering. IEEE Transactions on Computers, 71(2):282–295, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bombieri:2020:MIB

[BBD^{+}20] N. Bombieri, F. Busato, A. Danese, L. Piccolboni, and G. Pravadelli. Mangrove: An inference-based dynamic invariant mining for GPU architectures. IEEE Transactions on Computers, 69(4):606–620, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Boche:2025:FMT

[BBdTF25] Holger Boche, Yannik N.

Böck, Zoe Garcia del Toro, and Frank H. P. Fitzek. Feynman meets Turing: The uncomputability of quantum gate-circuit emulation and concatenation. IEEE Transactions on Computers, 74(3): 1053–1065, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benini:2021:GEI

[BBJR21] Luca Benini, Simone Benatti, Taekwang Jang, and Abbas Rahimi. Guest editorial: IEEE TC special issue on smart edge computing and IoT. IEEE Transactions on Computers, 70(8):1146– 1147, August 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Burmyakov:2022:TTE

[BBL22] Artem Burmyakov, Enrico Bini, and Chang-Gun Lee. Towards a tractable exact test for global multiprocessor fixed priority scheduling. IEEE Transactions on Computers, 71(11):2955–2967, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Brandalero:2021:MTA

[BCBS21] M. Brandalero, L. Carro, A. C. S. Beck, and M. Shafique. Multi-target adaptive reconfigurable acceleration for lowpower IoT processing. IEEE Transactions on Computers, 70(1):83–98, January 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bernasconi:2022:MCX

[BCCM22] Anna Bernasconi, Stelvio Cimato, Valentina Ciriani, and Maria Chiara Molteni. Multiplicative complexity of XOR based regular functions. IEEE Transactions on Computers, 71(11):2927– 2939, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bini:2023:ITS

[BCCM23] Enrico Bini, Tam Chantem, Bruce Childers, and Daniel Mosse. IEEE TC special issue on real-time systems. IEEE Transactions on Computers, 72(1):1–2, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bellarmino:2025:CCS

[BCF^{+}25] Nicolò Bellarmino, Riccardo Cantoro, Sophie M. Fosson, Martin Huch, Tobias Kilian, Ulf Schlichtmann, and Giovanni Squillero. COSMO: COmpressed sensing for models and logging optimization in MCU performance screening. IEEE Transactions on Computers, 74(2):652–664, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Berney:2022:BBS

[BCKS22] Kyle Berney, Henri Casanova, Ben Karsin, and Nodari Sitchinava. Beyond binary search: Parallel in-place construction of implicit search tree layouts. IEEE Transactions on Computers, 71(5): 1104–1116, May 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Oliveira:2025:TRT

[BCLC25] Daniel Bristot De Oliveira, Daniel Casini, Juri Lelli, and Tommaso Cucinotta. Timerlat: Real-time Linux scheduling latency measurements, tracing, and analysis. IEEE Transactions on Computers, 74(8):2608–2620, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bolchini:2023:FAE

[BCMT23] Cristiana Bolchini, Luca Cassano, Antonio Miele, and Alessandro Toschi. Fast and accurate error simulation for CNNs against soft errors. IEEE Transactions on Computers, 72(4):984–997, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bai:2023:ACD

[BCRX23] Yang Bai, Lixing Chen, Shaolei Ren, and Jie Xu. Automated customization of ondevice inference for qualityof-experience enhancement. IEEE Transactions on Computers, 72(5):1329–1342, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bernasconi:2022:ESD

[BCV22] Anna Bernasconi, Valentina Ciriani, and Tiziano Villa. Exploiting symmetrization and d-reducibility for approximate logic synthesis. IEEE Transactions on Computers, 71(1): 121–133, January 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Boldo:2020:REE

[BFC20] S. Boldo, F. Faissole, and A. Chapoutot. Round-off error and exceptional behavior analysis of explicit Runge–Kutta methods. IEEE Transactions on Computers, 69(12):1745– 1756, December 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Blott:2021:EOC

[BFG^{+}21] Michaela Blott, Nicholas J.

Fraser, Giulio Gambardella, Lisa Halder, Johannes Kath, Zachary Neveu, Yaman Umuroglu, Alina Vasilciuc, Miriam Leeser, and Linda Doyle. Evaluation of optimized CNNs on heterogeneous accelerators using a novel benchmarking approach. IEEE Transactions on Computers, 70(10):1654–1669, October 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Burrello:2021:DAE

[BGB^{+}21] Alessio Burrello, Angelo Garofalo, Nazareno Bruschi, Giuseppe Tagliavini, Davide Rossi, and Francesco Conti. DORY: Automatic end-to-end deployment of real-world DNNs on low-cost IoT MCUs. IEEE Transactions on Computers, 70(8):1253–1268, August 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bascones:2025:PCB

[BGHR^{+}25] Daniel Báscones, Francisco García-Herrero, Óscar Ruano, Carlos González, Daniel Mozos, and Juan Antonio Maestro. Protecting the CCSDS 123.0-B-2 compression algorithm against single-event upsets for space applications. IEEE Transactions on Computers, 74(3):944–954, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Beckers:2023:AED

[BGM^{+}23] Arthur Beckers, Sylvain Guilley, Philippe Maurine, Colin O’Flynn, and Stjepan Picek. (Adversarial) electromagnetic disturbance in the industry. IEEE Transactions on Computers, 72(2):414–422, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bazzaz:2021:FPN

[BHE21] M. Bazzaz, A. Hoseinghorban, and A. Ejlali. Fast and predictable non-volatile data memory for real-time embedded systems. IEEE Transactions on Computers, 70(3): 359–371, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bae:2023:HSC

[BHK^{+}23] Hyungjoon Bae, Yujin Hyun, Suchang Kim, Sangsoo Park, Jaeyoung Lee, Boseon Jang, Suyoung Choi, and In-Cheol Park. High-speed counter with novel LFSR state extension. IEEE Transactions on Computers, 72(3):893–899, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bao:2023:LEP

[BHW^{+}23] Zijian Bao, Debiao He, Wei Wei, Cong Peng, and Xinyi Huang. LedgerMaze: an efficient privacy-preserving noninteractive zero-knowledge scheme over account-model blockchain. IEEE Transactions on Computers, 72(12): 3489–3502, December 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Baek:2021:SAR

[BJM^{+}21] Sungha Baek, Youngdon Jung, David Mohaisen, Sungjin Lee, and DaeHun Nyang. SSDassisted ransomware detection and data recovery techniques. IEEE Transactions on Computers, 70(10):1762–1776, October 2021. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Butts:2023:RES

[BJMKK23] Corey Butts, Rashmi Jha, Temesguen Messay-Kebede, and David Kapp. Resilient embedded systems designs via on the fly generation of adaptive degenerate components using machine learning. IEEE Transactions on Computers, 72(5):1236–1246, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bottcher:2023:TGO

[BK23] Andreas Böttcher and Martin Kumm. Towards globally optimal design of multipliers for FPGAs. IEEE Transactions on Computers, 72(5): 1261–1273, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bei:2022:OOS

[BKHY22] Zhendong Bei, Nam Sung Kim, Kai Hwang, and Zhibin Yu. OSC: An online selfconfiguring big data framework for optimization of QoS. IEEE Transactions on Computers, 71(4):809–823, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Baksi:2022:NAS

[BKS22] Anubhab Baksi, Satyam Kumar, and Santanu Sarkar. A new approach for side channel analysis on stream ciphers and related constructions. IEEE Transactions on Computers, 71(10):2527– 2537, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Byun:2022:LFL

[BL22] Hayoung Byun and Hyesook Lim. Learned FBF: Learningbased functional Bloom filter for key value storage. IEEE Transactions on Computers, 71(8):1928–1938, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bai:2021:DCB

[BLH^{+}21] Shilei Bai, Bin Liang, Jianjun Huang, Wei You, Jiachun Li, Yaping Li, and Wenchang Shi. Detecting the capacitancebased gamepad for protecting mobile game fairness. IEEE Transactions on Computers, 70(9):1374–1387, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Baek:2023:SFF

[BLKK23] Eunjin Baek, Eunbok Lee, Taehun Kang, and Jangwoo Kim. STfusion: Fast and flexible multi-NN execution using spatio-temporal block fusion and memory management. IEEE Transactions on Computers, 72(4):1194–1207, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bund:2020:OMC

[BLM20] J. Bund, C. Lenzen, and M. Medina. Optimal metastabilitycontaining sorting via parallel prefix computation. IEEE Transactions on Computers, 69(2):198–211, February 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Boldo:2021:ERN

[BLM21] Sylvie Boldo, Christoph Lauter, and Jean-Michel Muller. Emulating round-to-nearest tiesto-zero augmented floatingpoint operations using roundto-nearest ties-to-even arithmetic. IEEE Transactions on Computers, 70(7):1046–1058, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bund:2025:SHF

[BLM25] Johannes Bund, Christoph Lenzen, and Moti Medina. Small hazard-free transducers. IEEE Transactions on Computers, 74(5):1549–1564, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benoit:2022:RSM

[BLP^{+}22] Anne Benoit, Valentin Le Fèvre, Lucas Perotin, Padma Raghavan, Yves Robert, and Hongyang Sun. Resilient scheduling of moldable parallel jobs to cope with silent errors. IEEE Transactions on Computers, 71(7):1696–1710, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bhattacharya:2020:BPA

[BMBM20] S. Bhattacharya, C. Maurice, S. Bhasin, and D. Mukhopadhyay. Branch prediction attack on blinded scalar multiplication. IEEE Transactions on Computers, 69(5):633–648, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bossuet:2020:PHI

[BMLOM20] L. Bossuet, C. Mancillas-López, and B. Ovilla-Martínez. Pipelined hardware implementation of COPA, ELmD, and COLM. IEEE Transactions on Computers, 69(10):1533– 1543, October 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bhimani:2022:ATP

[BMM^{+}22] Jaki Bhimani, Adnan Maruf, Ningfang Mi, Rajinikanth Pandurangan, and Vijay Balakrishnan. Auto-tuning parameters for emerging multistream flash-based storage drives through new I/O pattern generations. IEEE Transactions on Computers, 71(2): 309–322, February 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bandil:2024:HIU

[BN24] Lalit Bandil and Bal Chand Nagar. Hardware implementation of unsigned approximate hybrid square rooters for error-resilient applications. IEEE Transactions on Computers, 73(12):2734–2746, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benz:2025:ARS

[BOL^{+}25] Thomas Benz, Alessandro Ottaviano, Chaoqun Liang, Robert Balas, Angelo Garofalo, Francesco Restuccia, Alessandro Biondi, Davide Rossi, and Luca Benini. AXI-REALM: Safe, modular and lightweight traffic monitoring and regulation for heterogeneous mixed-criticality systems. IEEE Transactions on Computers, 74(9):3072– 3086, September 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bodmann:2022:SEE

[BPJ^{+}22] Pablo R. Bodmann, George Papadimitriou, Rubens L. Rech Junior, Dimitris Gizopoulos, and Paolo Rech. Soft error effects on Arm microprocessors: Early estimations versus chip measurements. IEEE Transactions on Computers, 71(10): 2358–2369, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bini:2023:ZJC

[BPM23] Enrico Bini, Paolo Pazzaglia, and Martina Maggio. Zerojitter chains of periodic LET tasks via algebraic rings. IEEE Transactions on Computers, 72(11):3057–3071, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bag:2022:PFA

[BRPM22] Arnab Bag, Debapriya Basu Roy, Sikhar Patranabis, and Debdeep Mukhopadhyay. FlexiPair: An automated programmable framework for pairing cryptosystems. IEEE Transactions on Computers, 71(3): 506–519, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Benz:2024:HPE

[BRS^{+}24] Thomas Benz, Michael Rogenmoser, Paul Scheffler, Samuel Riedel, Alessandro Ottaviano, Andreas Kurth, Torsten Hoefler, and Luca Benini. A highperformance, energy-efficient modular DMA engine architecture. IEEE Transactions on Computers, 73(1): 263–277, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bruguera:2020:LLF

[Bru20] J. D. Bruguera. Low latency floating-point division and square root unit. IEEE Transactions on Computers, 69(2):274–287, February 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bruguera:2023:RFP

[Bru23] Javier D. Bruguera. Radix-64 floating-point division and square root: Iterative and pipelined units. IEEE Transactions on Computers, 72(10): 2990–3001, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bilgin:2021:PVD

[BSM21] Bilgesu Arif Bilgin and Phillip Stanley-Marbell. Probabilistic value-deviation-bounded source-dependent bit-level channel adaptation for approximate communication. IEEE Transactions on Computers, 70(11):1949–1961, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Barreto:2021:SBI

[BSRP21] P. S. L. M. Barreto, M. A. Simplicio, J. E. Ricardini, and H. K. Patil. Schnorr-based implicit certification: Improving the security and efficiency of vehicular communications. IEEE Transactions on Computers, 70(3):393–399, March 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bhattacharjee:2020:CCT

[BTEC20] D. Bhattacharjee, Y. Tavva, A. Easwaran, and A. Chattopadhyay. Crossbar-constrained technology mapping for ReRAMbased in-memory computing. IEEE Transactions on Computers, 69(5):734–748, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.
org/document/8951116.

Bi:2025:EFH

[BWL^{+}25] Jun Bi, Yuanbo Wen, Xiaqing Li, Yongwei Zhao, Yuxuan Guo, Enshuai Zhou, Xing Hu, Zidong Du, Ling Li, Huaping Chen, Tianshi Chen, and Qi Guo. Efficient and fast high-performance library generation for deep learning accelerators. IEEE Transactions on Computers, 74(1): 155–169, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bache:2025:MTE

[BWSG25] Florian Bache, Jonas Wloka, Pascal Sasdrich, and Tim Güneysu. Multivariate TVLA — efficient side-channel evaluation using confidence intervals. IEEE Transactions on Computers, 74(3): 790–804, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bi:2025:HOA

[BXW^{+}25] Fangqi Bi, Guoqi Xie, Yuan Wang, Hao Wen, Zhenli He, Shaowen Yao, Sirong Zhao, Chenglai Xiong, Xingyu Hu, Bo Wan, and Yiwen Jiang. Hypercall-oriented abnormal VM status detection system: a non-intrusive solution for both hypervisor and guests. IEEE Transactions on Computers, 74(9):3032–3045, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bechtel:2022:MAD

[BY22] Michael Bechtel and Heechul Yun. Memory-aware denialof-service attacks on shared cache in multicore real-time systems. IEEE Transactions on Computers, 71(9):2351– 2357, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bechtel:2024:AMS

[BY24] Michael Bechtel and Heechul Yun. Analysis and mitigation of shared resource contention on heterogeneous multicore: an industrial case study. IEEE Transactions on Computers, 73(7):1753–1766, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bravyi:2022:EAF

[BYM22] Sergey Bravyi, Theodore J.

Yoder, and Dmitri Maslov. Efficient ancilla-free reversible and quantum circuits for the hidden weighted bit function. IEEE Transactions on Computers, 71(5):1170–1180, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bao:2020:PPE

[BYZZ20] W. Bao, D. Yuan, B. B. Zhou, and A. Y. Zomaya. Prune and plant: Efficient placement and parallelism of virtual network functions. IEEE Transactions on Computers, 69(6):800–811, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bai:2025:NBE

[BZW^{+}25a] Yichuan Bai, Xiaopeng Zhang, Qian Wang, Yaqing Li, Yuan Du, and Li Du. BE-NPU: a bandwidth-efficient neural processing unit with adaptive processing schemes for reduced off-chip bandwidth demand. IEEE Transactions on Computers, 74(7):2376–2388, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bian:2025:ABS

[BZW^{+}25b] Yi Bian, Fangyu Zheng, Yuewu Wang, Lingguang Lei, Yuan Ma, Tian Zhou, Jiankuo Dong, Guang Fan, and Jiwu Jing. AsyncGBP^{++}: Bridging SSL/TLS and heterogeneous computing power with GPUbased providers. IEEE Transactions on Computers, 74(2): 356–370, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Curzel:2022:EES

[CAC^{+}22] Serena Curzel, Nicolas Bohm Agostini, Vito Giovanni Castellana, Marco Minutoli, Ankur Limaye, Joseph Manzano, Jeff Zhang, David Brooks, Gu-Yeon Wei, Fabrizio Ferrandi, and Antonino Tumeo. Endto-end synthesis of dynamically controlled machine learning accelerators. IEEE Transactions on Computers, 71(12): 3074–3087, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cai:2022:IQM

[CB22] Zhicheng Cai and Rajkumar Buyya. Inverse queuing model-based feedback control for elastic container provisioning of web systems in Kubernetes. IEEE Transactions on Computers, 71(2): 337–348, February 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Casini:2021:TSL

[CBB21a] Daniel Casini, Alessandro Biondi, and Giorgio Buttazzo.

Task splitting and load balancing of dynamic real-time workloads for semi-partitioned EDF. IEEE Transactions on Computers, 70(12):2168– 2181, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cesarini:2021:CRT

[CBB^{+}21b] D. Cesarini, A. Bartolini, P. Bonfà, C. Cavazzoni, and L. Benini. COUNT-DOWN: a run-time library for performance-neutral energy saving in MPI applications. IEEE Transactions on Computers, 70(5):682–695, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2023:ILA

[CCC23] Yujie Cui, Hongwei Cui, and Xu Cheng. Information leakage attacks exploiting cache replacement in commercial processors. IEEE Transactions on Computers, 72(9): 2536–2547, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chang:2021:MHM

[CCCW21] Cheng-Yang Chang, Yu-Chuan Chuang, En-Jui Chang, and An-Yeu Andy Wu. MulTa-HDC: A multi-task learning framework for hyperdimensional computing. IEEE Transactions on Computers, 70(8):1269–1284, August 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Coluccio:2022:HSM

[CCG^{+}22] Andrea Coluccio, Umberto Casale, Angela Guastamacchia, Giovanna Turvani, Marco Vacca, Massimo Ruo Roch, Maurizio Zamboni, and Mariagrazia Graziano. Hybrid-SIMD: a modular and reconfigurable approach to beyond von Neumann computing. IEEE Transactions on Computers, 71(9):2287– 2299, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cai:2020:ESN

[CCT^{+}20] Y. Cai, X. Chen, L. Tian, Y. Wang, and H. Yang. Enabling secure NVM-based inmemory neural network computing by sparse fast gradient encryption. IEEE Transactions on Computers, 69(11): 1596–1610, November 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2024:SDD

[CCY^{+}24] Laizhong Cui, Ziteng Chen, Shu Yang, Ruiyu Chen, and Zhong Ming. A secure and decentralized DLaaS platform for edge resource scheduling against adversarial attacks. IEEE Transactions on Computers, 73(3):631–644, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:EDP

[CCYC22] Tseng-Yi Chen, Shao-Hung Chi, Ming-Chang Yang, and Ting-Ying Chien. Enabling the duo-phase data management to realize longevity bit-alterable flash memory. IEEE Transactions on Computers, 71(8):1982–1997, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:CME

[CCZ^{+}22] Peng Chen, Hui Chen, Jun Zhou, Mengquan Li, Weichen Liu, Chunhua Xiao, Yiyuan Xie, and Nan Guan. Contention minimization in emerging SMART NoC via direct and indirect routes. IEEE Transactions on Computers, 71(8):1874–1888, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:FES

[CCZ^{+}25] Xinglei Chen, Zinuo Cai, Hanwen Zhang, Ruhui Ma, and Rajkumar Buyya. FasDL: an efficient serverless-based training architecture with communication optimization and Resource configuration. IEEE Transactions on Computers, 74(2):468–482, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Coppolino:2021:VCI

[CDF^{+}21] L. Coppolino, S. D’Antonio, V. Formicola, G. Mazzeo, and L. Romano. VISE: Combining Intel SGX and homomorphic encryption for cloud industrial control systems. IEEE Transactions on Computers, 70(5): 711–724, May 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cardoso:2021:GEI

[CDP21] João M. P. Cardoso, André DeHon, and Laura Pozzi. Guest editorial: IEEE TC special section on compiler optimizations for FPGA-based systems. IEEE Transactions on Computers, 70(12):2013– 2014, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cantoro:2020:NSG

[CDRS20] R. Cantoro, A. Damljanovic, M. S. Reorda, and G. Squillero. A novel sequence generation approach to diagnose faults in reconfigurable scan networks. IEEE Transactions on Computers, 69(1):87–98, January 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:TOC

[CDW^{+}25] Xingyan Chen, Tian Du, Mu Wang, Tiancheng Gu, Yu Zhao, Gang Kou, Changqiao Xu, and Dapeng Oliver Wu. Towards optimal customized architecture for heterogeneous federated learning with contrastive cloud-edge model decoupling. IEEE Transactions on Computers, 74(4): 1123–1137, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2023:POB

[CERMH23] Xiao Chen, Btissam Er-Rahmadi, Tiejun Ma, and Jane Hillston. ParBFT: an optimized Byzantine consensus parallelism scheme. IEEE Transactions on Computers, 72(12):3354–3369, December 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheshmikhani:2022:RDR

[CFA22] Elham Cheshmikhani, Hamed Farbeh, and Hossein Asadi. 3RSeT: Read disturbance rate reduction in STT-MRAM caches by selective tag comparison. IEEE Transactions on Computers, 71(6): 1305–1319, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2022:EMA

[CFC^{+}22] Jianyi Cheng, Shane T. Fleming, Yu Ting Chen, Jason Anderson, John Wickerson, and George A. Constantinides. Efficient memory arbitration in high-level synthesis from multi-threaded code. IEEE Transactions on Computers, 71(4):933–946, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:PBT

[CFW25] Lin Chen, Hao Feng, and Jiong Wu. A path-based topology-agnostic fault diagnosis strategy for multiprocessor systems. IEEE Transactions on Computers, 74(6): 1886–1896, June 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2023:BSI

[CFWC23] Jianyi Cheng, Estibaliz Fraca, John Wickerson, and George A. Constantinides. Balancing static islands in dynamically scheduled circuits using continuous Petri nets. IEEE Transactions on Computers, 72(11):3300–3313, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cassiers:2021:HPC

[CGLS21] Gaëtan Cassiers, Benjamin Grégoire, Itamar Levi, and François-Xavier Standaert. Hardware private circuits: From trivial composition to full verification. IEEE Transactions on Computers, 70(10): 1677–1690, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Castello:2020:ATL

[CGS^{+}20] A. Castelló, R. M. Gual, S. Seo, P. Balaji, E. S. Quintana-Ortí, and A. J. Peña. Analysis of threading libraries for high performance computing. IEEE Transactions on Computers, 69(9): 1279–1292, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:ARS

[CHC25] Liang-Chi Chen, Chien-Chung Ho, and Yuan-Hao Chang. Accelerating RNA-Seq quantification on a real processing-in-memory system. IEEE Transactions on Computers, 74(7): 2334–2347, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2023:EFG

[CHL^{+}23] Jinwoo Choi, Yeonan Ha, Jounghoo Lee, Sangsu Lee, Jinho Lee, Hanhwi Jang, and Youngsok Kim. Enabling finegrained spatial multitasking on systolic-array NPUs using dataflow mirroring. IEEE Transactions on Computers, 72(12):3383–3398, December 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chatterjee:2025:PGP

[CHM25] Durba Chatterjee, Aritra Hazra, and Debdeep Mukhopadhyay. PARLE-G: Provable automated representation and analysis framework for learnability evaluation of generic PUF compositions. IEEE Transactions on Computers, 74(3):820–834, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:III

[CHZ^{+}25] Weijian Chen, Shuibing He, Ruidong Zhang, Xuechen Zhang, Ping Chen, Siling Yang, Haoyang Qu, and Xuan Zhan. ImPACT: Importanceinformed prefetching and caching for I/O-bound DNN training. IEEE Transactions on Computers, 74(8):2649– 2662, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cai:2024:SCP

[CJSY24] Miao Cai, Xuzhen Jiang, Junru Shen, and Baoliu Ye. SplitDB: Closing the performance gap for LSM-tree-based key–value stores. IEEE Transactions on Computers, 73(1): 206–220, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2023:OOV

[CJYC23] Duheon Choi, Taeyang Jeong, Joonhyeok Yeom, and Eui-Young Chung. Operandoriented virtual memory support for near-memory processing. IEEE Transactions on Computers, 72(8):2250– 2263, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2022:LEG

[CKJ^{+}22] Jungwoo Choi, Boyeal Kim, Ji-Ye Jeon, Hyuk-Jae Lee, Euicheol Lim, and Chae Eun Rhee. A lightweight and efficient GPU for NDP utilizing data access pattern of image processing. IEEE Transactions on Computers, 71(1): 13–26, January 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2022:ASB

[CKK^{+}22] P. Choi, W. Kong, J.-H. Kim, M.-K. Lee, and Dong Kyue Kim. Architectural supports for block ciphers in a RISC CPU core by instruction overloading. IEEE Transactions on Computers, 71(11):2844– 2857, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cho:2022:OPS

[CKP^{+}22] Youngeun Cho, Do Hyung Kim, Daechul Park, Seung Su Lee, and Chang-Gun Lee. Optimal parallelization of single/ multi-segment real-time tasks for global EDF. IEEE Transactions on Computers, 71(5): 1077–1091, May 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2021:OOD

[CKRP21] Won Gi Choi, Doyoung Kim, Hongchan Roh, and Sanghyun Park. OurRocks: Offloading disk scan directly to GPU in write-optimized database system. IEEE Transactions on Computers, 70(11):1831– 1844, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cao:2020:HOD

[CL20] W. Cao and L. Liu. Hierarchical orchestration of disaggregated memory. IEEE Transactions on Computers, 69(6): 844–855, June 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:UST

[CLCC25] Peng Chen, Jinnuo Li, Wei Cheng, and Chi Cheng. Uncover secrets through the cover: a deep learning-based side-channel attack against kyber implementations with anti-tampering covers. IEEE Transactions on Computers, 74(6):2159–2167, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chung:2022:ATF

[CLCL22] Doyoung Chung, Seungkwang Lee, Dooho Choi, and Jooyoung Lee. Alternative tower field construction for quantum implementation of the AES S-box. IEEE Transactions on Computers, 71(10): 2553–2564, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:AFS

[CLLdS25] Chao Chen, Chengyu Liu, Jianqing Li, and Bruno da Silva. Acceleration of fast sample entropy for FPGAs. IEEE Transactions on Computers, 74(1):1–14, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:KAS

[CLY22] Shuo-Han Chen, Yuhong Liang, and Ming-Chang Yang. KVSTL: An application support to LSM-tree based keyvalue store via shingled translation layer data management. IEEE Transactions on Computers, 71(7):1598–1611, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:ECA

[CLZG22] Ningyu Chen, Jiguo Li, Yichen Zhang, and Yuyan Guo. Efficient CP-ABE scheme with shared decryption in cloud storage. IEEE Transactions on Computers, 71(1):175–184, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2024:SHO

[CMGD24] Wei Cheng, Jingdian Ming, Sylvain Guilley, and Jean-Luc Danger. Statistical higherorder correlation attacks against code-based masking. IEEE Transactions on Computers, 73(10):2364–2377, October 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:LCL

[CMQ^{+}22] Shilin Chen, Shang Ma, Zhuo Qin, Bixin Zhu, Ziqian Xiao, and Meiqing Liu. A low complexity and long period digital random sequence generator based on residue number system and permutation polynomial. IEEE Transactions on Computers, 71(11):3008– 3017, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cicek:2022:GRC

[CNOS22] Nihat Mert Cicek, Lin Ning, Ozcan Ozturk, and Xipeng Shen. General reuse-centric CNN accelerator. IEEE Transactions on Computers, 71(4):880–891, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cogo:2021:GSB

[CPB21] V. Cogo, J. Paulo, and A. Bessani. GenoDedup: Similarity-based deduplication and delta-encoding for genome sequencing data. IEEE Transactions on Computers, 70(5):669–681, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2023:OOL

[CPL^{+}23] Ruobing Chen, Wangqi Peng, Yusen Li, Xiaoguang Liu, and Gang Wang. Orchid: an online learning based resource partitioning framework for job colocation with multiple objectives. IEEE Transactions on Computers, 72(12):3443– 3457, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cruz:2023:FAE

[CPM^{+}23] Jonathan Cruz, Christopher Posada, Naren Vikram Raj Masna, Prabuddha Chakraborty, Pravin Gaikwad, and Swarup Bhunia. A framework for automated exploration of Trojan attack space in FPGA netlists. IEEE Transactions on Computers, 72(10):2740–2751, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:GEI

[CQ22] Yiran Chen and Qinru Qiu. Guest editorial: IEEE TC special issue on software, hardware and applications for neuromorphic computing. IEEE Transactions on Computers, 71(11):2705–2706, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:HRG

[CQCL25] Hui Chen, Lianghua Quan, Ke Chen, and Weiqiang Liu. High-radix generalized hyperbolic CORDIC and its hardware implementation. IEEE Transactions on Computers, 74(3):983–995, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:PPL

[CQI^{+}22] Chuangtao Chen, Weikang Qian, Mohsen Imani, Xunzhao Yin, and Cheng Zhuo. PAM: a piecewise-linearly-approximated floating-point multiplier with unbiasedness and configurability. IEEE Transactions on Computers, 71(10):2473–2486, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Casale:2024:SIE

[CR24] Giuliano Casale and Manuel Roveri. Scheduling inputs in early exit neural networks. IEEE Transactions on Computers, 73(2):451–465, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2021:ORT

[CRJZ21] Wenxue Cheng, Fengyuan Ren, Wanchun Jiang, and Tong Zhang. Optimizing the response time of memcached systems via model and quantitative analysis. IEEE Transactions on Computers, 70(9): 1458–1471, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cai:2024:IPB

[CSH^{+}24] Xinquan Cai, Qianlong Sang, Chuang Hu, Yili Gong, Kun Suo, Xiaobo Zhou, and Dazhao Cheng. Incendio: Priority-based scheduling for alleviating cold start in serverless computing. IEEE Transactions on Computers, 73(7): 1780–1794, July 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Choi:2022:DNN

[CSK22] Seungkyu Choi, Jaekang Shin, and Lee-Sup Kim. A deep neural network training architecture with inference-aware heterogeneous data-type. IEEE Transactions on Computers, 71(5):1216–1229, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:SRT

[CSvdBU22] Jian-Jia Chen, Junjie Shi, Georg von der Brüggen, and Niklas Ueter. Scheduling of real-time tasks with multiple critical sections in multiprocessor systems. IEEE Transactions on Computers, 71(1): 146–160, January 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2021:CPV

[CSW^{+}21] Xianzhang Chen, Edwin H.-M. Sha, Xinxin Wang, Chaoshu Yang, Weiwen Jiang, and Qingfeng Zhuge. Contour: a process variation aware wear-leveling mechanism for inodes of persistent memory file systems. IEEE Transactions on Computers, 70(7):1034–1045, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cai:2025:SPM

[CSY^{+}25] Miao Cai, Junru Shen, Yifan Yuan, Zhihao Qu, and Baoliu Ye. Scaling persistent inmemory key–value stores over modern tiered, heterogeneous memory hierarchies. IEEE Transactions on Computers, 74(2):495–509, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chatelain:2025:NVA

[CTM^{+}25] Yohan Chatelain, Loïc Tetrel, Christopher J. Markiewicz, Mathias Goncalves, Gregory Kiar, Oscar Esteban, Pierre Bellec, and Tristan Glatard. A numerical variability approach to results stability tests and its application to neuroimaging. IEEE Transactions on Computers, 74(1):200–209, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cai:2024:CCA

[CTY^{+}24] Qinyun Cai, Guanghua Tan, Wangdong Yang, Xianhao He, Yuwei Yan, Keqin Li, and Kenli Li. COALA: a compilerassisted adaptive library routines allocation framework for heterogeneous systems. IEEE Transactions on Computers, 73(7):1724–1737, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2024:NND

[CTZ^{+}24] Yucong Chen, Yanshan Tian, Rui Zhou, Diego Martínez Castro, Deke Guo, and Qingguo Zhou. NDSTRNG: Nondeterministic sampling-based true random number generator on SoC FPGA systems. IEEE Transactions on Computers, 73(5):1313–1326, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cervantes-Vazquez:2022:PSS

[CVOJRH22] Daniel Cervantes-Vázquez, Eduardo Ochoa-Jiménez, and Francisco Rodríguez-Henríquez. Parallel strategies for SIDH: Toward computing SIDH twice as fast. IEEE Transactions on Computers, 71(6): 1249–1260, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chu:2024:SCC

[CWC^{+}24] Shao-I Chu, Chi-Long Wu, Tzu-Heng Chien, Bing-Hong Liu, and Tu N. Nguyen. Stochastic circuits for computing weighted ratio with applications to multiclass Bayesian inference machine. IEEE Transactions on Computers, 73(2):621–630, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chu:2022:PCU

[CWNL22] Shao-I Chu, Chi-Long Wu, Tu N. Nguyen, and Bing-Hong Liu. Polynomial computation using unipolar stochastic logic and correlation technique. IEEE Transactions on Computers, 71(6):1358–1373, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2024:DLA

[CWS^{+}24a] Yongsheng Chen, Zhuowei Wang, Xiaoyu Song, Zhe Yan, and Lianglun Cheng. Deep learning acceleration optimization of stress boundary value problem solvers. IEEE Transactions on Computers, 73(12):2844–2854, 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2024:RVC

[CWS^{+}24b] Yuxing Chen, Xinrui Wang, Suwen Song, Lang Feng, and Zhongfeng Wang. RISC-V custom instructions of elementary functions for IoT endpoint devices. IEEE Transactions on Computers, 73(2): 523–535, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cai:2022:ORM

[CWT^{+}22] Xuyi Cai, Ying Wang, Kaijie Tu, Chengsi Gao, and Lei Zhang. Olympus: Reaching memory-optimality on DNN processors. IEEE Transactions on Computers, 71(8): 1939–1951, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2020:NJL

[CWWW20] C. Chen, Q. Wei, W. Wong, and C. Wang. NV-journaling: Locality-aware journaling using byte-addressable nonvolatile memory. IEEE Transactions on Computers, 69(2): 288–299, February 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2023:PED

[CWY^{+}23] Shuzhen Chen, Yangyang Wang, Dongxiao Yu, Ju Ren, Congan Xu, and Yanwei Zheng. Privacy-enhanced decentralized federated learning at dynamic edge. IEEE Transactions on Computers, 72(8): 2165–2180, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:OEO

[CXH^{+}25] Aodong Chen, Fei Xu, Li Han, Yuan Dong, Li Chen, Zhi Zhou, and Fangming Liu. Opara: Exploiting operator parallelism for expediting DNN inference on GPUs. IEEE Transactions on Computers, 74(1):325–333, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2023:PID

[CXL^{+}23] Ke Cheng, Ning Xi, Ximeng Liu, Xinghui Zhu, Haichang Gao, Zhiwei Zhang, and Yulong Shen. Private inference for deep neural networks: a secure, adaptive, and efficient realization. IEEE Transactions on Computers, 72(12): 3519–3531, December 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2025:EEP

[CXL^{+}25] Xiaolu Cheng, Xiaoshuang Xing, Wei Li, Hong Xue, and Tong Can. An energyefficient and privacy-aware MEC-enabled IoMT health monitoring system. IEEE Transactions on Computers, 74(9):2936–2949, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2023:EAS

[CXW^{+}23] Ke Chen, Chenyu Xu, Haroon Waris, Weiqiang Liu, Paolo Montuschi, and Fabrizio Lombardi. Exact and approximate squarers for error-tolerant applications. IEEE Transactions on Computers, 72(7): 2120–2126, July 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2024:TFE

[CXY24] Chao Chen, Jinhan Xin, and Zhibin Yu. TIE: Fast experiment-driven ML-based configuration tuning for inmemory data analytics. IEEE Transactions on Computers, 73(5):1233–1247, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chowdhuryy:2022:LST

[CY22] Md Hafizul Islam Chowdhuryy and Fan Yao. Leaking secrets through modern branch predictors in the speculative world. IEEE Transactions on Computers, 71(9):2059– 2072, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chatelain:2023:PAP

[CYKG23] Yohan Chatelain, Nigel Yong Sao Young, Gregory Kiar, and Tristan Glatard. PyTracer: Automatically profiling numerical instabilities in Python. IEEE Transactions on Computers, 72(6):1792–1803, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cao:2025:MML

[CYPC25] Kaiwen Cao, Hanchen Ye, Yihan Pang, and Deming Chen. MLCD: Machine learningbased code version and device selection for heterogeneous systems. IEEE Transactions on Computers, 74(7): 2417–2430, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2023:ASS

[CYX^{+}23] Weilin Chen, Wei Yang, Lide Xue, Bingren Chen, Youwen Zhu, and Liusheng Huang. Avalon: a scalable and secure distributed transaction ledger based on proof-of-market. IEEE Transactions on Computers, 72(12):3576– 3589, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cao:2022:NPE

[CZB^{+}22] Weidong Cao, Yilong Zhao, Adith Boloor, Yinhe Han, Xuan Zhang, and Li Jiang. Neural-PIM: Efficient processingin-memory with neural approximation of peripherals. IEEE Transactions on Computers, 71(9):2142– 2155, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cheng:2021:GDR

[CZC^{+}21] Kun Cheng, Yuan Zhou, Bihuan Chen, Rui Wang, Yuebin Bai, and Yang Liu. Guardauto: A decentralized runtime protection system for autonomous driving. IEEE Transactions on Computers, 70(10):1569–1581, October 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2024:CTV

[CZD^{+}24] Yuxuan Chen, Zhen Zhang, Yuhui Deng, Geyong Min, and Lin Cui. A combined trend virtual machine consolidation strategy for cloud data centers. IEEE Transactions on Computers, 73(9): 2150–2164, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2025:ULH

[CZH^{+}25] Zhenqian Chen, Yuchun Zhan, Peng Hu, Xinkui Zhao, Muyu Yang, Siwei Tan, Lufei Zhang, Liqiang Lu, Jianwei Yin, and Zuoning Chen. UKFaaS: Lightweight, highperformance and secure FaaS communication with unikernel. IEEE Transactions on Computers, 74(10):3305– 3318, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2021:PPA

[CZJ21] Hanhua Chen, Fan Zhang, and Hai Jin. PStream: A popularity-aware differentiated distributed stream processing system. IEEE Transactions on Computers, 70(10): 1582–1597, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:DBD

[CZR22] Xiangru Chen, Jiaqi Zhang, and Sandip Ray. Dandelion: Boosting DNN usability under dataset scarcity. IEEE Transactions on Computers, 71(10): 2487–2498, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2024:UDA

[CZW^{+}24] Yangguang Cui, Zhixing Zhang, Nuo Wang, Liying Li, Chunwei Chang, and Tongquan Wei. Userdistribution-aware federated learning for efficient communication and fast inference. IEEE Transactions on Computers, 73(4):1004–1018, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2025:HPH

[CZW^{+}25] Yijun Cui, Junjie Zhong, Bei Wang, Tianyu Xu, Chenghua Wang, and Weiqiang Liu. High-performance hardware implementation of crystalsdilithium based on improved MDC-NTT. IEEE Transactions on Computers, 74(9): 2896–2908, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Cui:2025:CPP

[CZZ^{+}25] Jie Cui, Wenting Zhuang, Hong Zhong, Qingyang Zhang, Fengqun Wang, and Debiao He. Conditional privacypreserving transaction for the unspent transaction outputbased multi-chain blockchain system. IEEE Transactions on Computers, 74(7):2473–2486, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deepika:2022:ADC

[DA22] S. Deepika and V. Arunachalam. Analysis & design of convolution operator for high speed and high accuracy convolutional neural networkbased inference engines. IEEE Transactions on Computers, 71(2):390–396, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Das:2023:BWC

[Das23] Arindam Das. Block-wise computation of cyclic redundancy code using factored Toeplitz matricesin lieu of look-up table. IEEE Transactions on Computers, 72(4): 1110–1121, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deVeras:2021:CBQ

[ddAPdS21] Tiago M. L. de Veras, Ismael C. S. de Araujo, Daniel K. Park, and Adenilton J. da Silva. Circuit-based quantum random access memory for classical data with continuous amplitudes. IEEE Transactions on Computers, 70(12): 2125–2135, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dar:2022:NCB

[DDK22] Gilad Dar, Giorgio Di Natale, and Osnat Keren. Nonlinear code-based low-overhead finegrained control flow checking. IEEE Transactions on Computers, 71(3):658–669, March 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DeSanto:2022:DLH

[DGG^{+}22] Aniello De Santo, Antonio Galli, Michela Gravina, Vincenzo Moscato, and Giancarlo Sperlì. Deep learning for HDD health assessment: an application based on LSTM. IEEE Transactions on Computers, 71(1):69–80, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2025:CCD

[DGQ^{+}25] Yuan Dai, Xuchen Gao, Yunhui Qiu, Jingyuan Li, Yuhang Cao, Yiqing Mao, Sichao Chen, Wenbo Yin, Wai-Shing Luk, and Lingli Wang. COFFA: a co-design framework for fused-grained reconfigurable architecture towards efficient irregular loop handling. IEEE Transactions on Computers, 74(9):3099– 3113, September 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Davila-Guzman:2021:AMM

[DGTVGG21] Maria Angélica Dávila-Guzmán, Rubén Gran Tejero, María Villarroya-Gaudó, and Darío Suárez Gracia. Analytical model for memory-centric high level synthesis-generated applications. IEEE Transactions on Computers, 70(12):2056– 2069, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2025:PMV

[DGX^{+}25] Xiaohai Dai, Zhengxuan Guo, Jiang Xiao, Guanxiong Wang, Yifei Liang, Chen Yu, and Hai Jin. Pako: Multi-valued Byzantine agreement comparable to partially-synchronous BFT. IEEE Transactions on Computers, 74(3):887– 900, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Du:2022:BIW

[DGZ^{+}22] Zidong Du, Qi Guo, Yongwei Zhao, Xi Zeng, Ling Li, Limin Cheng, Zhiwei Xu, Ninghui Sun, and Yunji Chen. Breaking the interaction wall: a DLPU-centric deep learning computing system. IEEE Transactions on Computers, 71(1):209–222, January 2022.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dhanuskodi:2020:TRS

[DH20] S. N. Dhanuskodi and D. Holcomb. Techniques to reduce switching and leakage energy in unrolled block ciphers. IEEE Transactions on Computers, 69(10):1414–1423, October 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deHaro:2021:OFH

[dHBF^{+}21] Juan Miguel de Haro, Jaume Bosch, Antonio Filgueras, Miquel Vidal, Daniel Jiménez-González, Carlos Álvarez, Xavier Martorell, Eduard Ayguadé, and Jesús Labarta. OmpSs@FPGA framework for high performance FPGA computing. IEEE Transactions on Computers, 70(12):2029– 2042, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2025:FHT

[DHM^{+}25] Guohao Dai, Ke Hong, Qiuli Mao, Xiuhong Li, Jiaming Xu, Haofeng Huang, Hongtu Xia, Xuefei Ning, Shengen Yan, Yun Liang, and Yu Wang. FlashDecoding++Next: High throughput LLM inference with latency and memory optimization. IEEE Transactions on Computers, 74(10):3263– 3276, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Das:2021:OCN

[DKJP21] Abhijit Das, Abhishek Kumar, John Jose, and Maurizio Palesi. Opportunistic caching in NoC: Exploring ways to reduce miss penalty. IEEE Transactions on Computers, 70(6):892–905, June 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ding:2025:CGT

[DKZ^{+}25] Zezhong Ding, Deyu Kong, Zhuoxu Zhang, Xike Xie, and Jianliang Xu. ClusPar: a game-theoretic approach for efficient and scalable streaming edge partitioning. IEEE Transactions on Computers, 74(1):116–130, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ding:2024:BDS

[DLG^{+}24] Lin Ding, Zhengting Li, Ziyu Guan, Xinhai Wang, and Zheng Wu. Breaking the DECT standard cipher with lower time cost. IEEE Transactions on Computers, 73(5): 1290–1299, May 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dunbo:2025:EDD

[DLK25] Zhang Dunbo, Shen Li, and Lu Kai. Eliminate data divergence in SpMV via processor and memory co-computing framework. IEEE Transactions on Computers, 74(6): 2017–2030, June 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2025:RLL

[DLW^{+}25] Xiaohai Dai, Wei Li, Guanxiong Wang, Jiang Xiao, Haoyang Chen, Shufei Li, Albert Y. Zomaya, and Hai Jin. Remora: a low-latency DAGbased BFT through optimistic paths. IEEE Transactions on Computers, 74(1):57–70, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deng:2021:PEE

[DLY21] C. Deng, S. Liao, and B. Yuan.

PermCNN: Energy-efficient convolutional neural network hardware architecture with permuted diagonal structure. IEEE Transactions on Computers, 70(2):163–173, February 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deng:2024:FFD

[DLZ^{+}24] Haotian Deng, Jinwen Liang, Chuan Zhang, Ximeng Liu, Liehuang Zhu, and Song Guo. FutureDID: a fully decentralized identity system with multi-party verification. IEEE Transactions on Computers, 73(8):2051–2065, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Das:2023:HEE

[DMD^{+}23] Arghadip Das, Chandrachur Majumder, Debaprasad De, Arnab Raha, and Mrinal Kanti Naskar. HIPEDAP: Energyefficient hardware accelerators for hidden periodicity detection. IEEE Transactions on Computers, 72(10):2781– 2794, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dang:2023:HSH

[DMG23] Viet Ba Dang, Kamyar Mohajerani, and Kris Gaj. Highspeed hardware architectures and FPGA benchmarking of

CRYSTALS-Kyber, NTRU, and

Saber. IEEE Transactions on Computers, 72(2):306– 320, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deng:2022:ECA

[DMX^{+}22] Shuwen Deng, Nikolay Matyunin, Wenjie Xiong, Stefan Katzenbeisser, and Jakub Szefer. Evaluation of cache attacks on Arm processors and secure caches. IEEE Transactions on Computers, 71(9):

2248–2262, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dass:2020:DTS

[DNMS20] J. Dass, Y. Narawane, R. N.

Mahapatra, and V. Sarin. Distributed training of support vector machine on a multiple-FPGA system. IEEE Transactions on Computers, 69(7): 1015–1026, July 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deOliveira:2023:OSN

[dOCC23] Daniel Bristot de Oliveira, Daniel Casini, and Tommaso Cucinotta. Operating system noise in the Linux kernel. IEEE Transactions on Computers, 72(1):196–207, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2022:RBV

[DPCL22] Guangli Dai, Pavan Kumar Paluri, Albert Mo Kim Cheng, and Bozheng Liu. Regularitybased virtualization under the ARINC 653 Standard for Embedded Systems. IEEE Transactions on Computers, 71(10): 2592–2605, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Denkinger:2023:ACI

[DPQK^{+}23] Benoît Walter Denkinger, Miguel Peón-Quirós, Mario Konijnenburg, David Atienza, and Francky Catthoor. Acceleration of control intensive applications on coarse-grained reconfigurable arrays for embedded systems. IEEE Transactions on Computers, 72(9): 2548–2560, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DiSanzo:2020:AMB

[DPS^{+}20] P. Di Sanzo, A. Pellegrini, M. Sannicandro, B. Ciciani, and F. Quaglia. Adaptive model-based scheduling in software transactional memory. IEEE Transactions on Computers, 69(5):621–632, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dalai:2022:SCC

[DPS22] Deepak Kumar Dalai, Santu Pal, and Santanu Sarkar. Some conditional cube testers for Grain-128a of reduced rounds. IEEE Transactions on Computers, 71(6):1374–1385, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DiSanzo:2023:ETD

[DQ23] Pierangelo Di Sanzo and Francesco Quaglia. On the effects of transaction data access patterns on performance in lock-based concurrency control. IEEE Transactions on Computers, 72(6):1718–1732, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Disabato:2021:DDC

[DRA21] Simone Disabato, Manuel Roveri, and Cesare Alippi. Distributed deep convolutional neural networks for the Internet-of-Things. IEEE Transactions on Computers, 70(8):1239–1252, August 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Das:2024:MOH

[DRP24] Abhijit Das, Enrico Russo, and Maurizio Palesi. Multiobjective hardware-mapping co-optimisation for multi-DNN workloads on chipletbased accelerators. IEEE Transactions on Computers, 73(8):1883–1898, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dhar:2022:DDP

[DRY^{+}22] Ashutosh Dhar, Edward Richter, Mang Yu, Wei Zuo, Xiaohao Wang, Nam Sung Kim, and Deming Chen. DML: Dynamic partial reconfiguration with scalable task scheduling for multiapplications on FPGAs. IEEE Transactions on Computers, 71(10):2577–2591, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

dosSantos:2022:RPD

[dSBS^{+}22] Fernando F. dos Santos, Marcelo Brandalero, Michael B. Sullivan, Pedro M. Basso, Michael Hübner, Luigi Carro, and Paolo Rech. Reduced precision DWC: An efficient hardening strategy for mixedprecision architectures. IEEE Transactions on Computers, 71(3):573–586, March 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dehghanzadeh:2025:LCP

[DSCB25] Peyman Dehghanzadeh, Ovishake Sen, Baibhab Chatterjee, and Swarup Bhunia. LUNA-CiM: a programmable compute-in-memory fabric for neural network acceleration. IEEE Transactions on Computers, 74(4):1348–1361, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deSouza:2022:CAN

[dSdCF22] Luciano S. de Souza, Jonathan H. A. de Carvalho, and Tiago A. E. Ferreira. Classical artificial neural network training using quantum walks as a search procedure. IEEE Transactions on Computers, 71(2):378–389, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deng:2022:SEE

[DSJ^{+}22] Bobin Deng, Sriseshan Srikanth, Anirudh Jain, Thomas M. Conte, Erik DeBenedictis, and Jeanine Cook. Scalable energy-efficient microarchitectures with computational error tolerance via redundant residue number systems. IEEE Transactions on Computers, 71(3):613–627, March 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Das:2023:AAL

[DSK23] Palash Das, Shashank Sharma, and Hemangee K. Kapoor. ALAMNI: Adaptive LookAside memory based nearmemory inference engine for eliminating multiplications in real-time. IEEE Transactions on Computers, 72(3): 693–706, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dazzi:2021:EPE

[DSP^{+}21] Martino Dazzi, Abu Sebastian, Thomas Parnell, Pier Andrea Francese, Luca Benini, and Evangelos Eleftheriou. Efficient pipelined execution of CNNs based on inmemory computing and graph homomorphism verification. IEEE Transactions on Computers, 70(6):922–935, June 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dilek:2022:HLS

[DSTD22] Selma Dilek, Rawan Smri, Suleyman Tosun, and Deniz Dal. A high-level synthesis methodology for energy and reliability-oriented designs. IEEE Transactions on Computers, 71(1):161–174, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Das:2020:NCS

[DT20] A. Das and N. A. Touba. A new class of single burst error correcting codes with parallel decoding. IEEE Transactions on Computers, 69(2): 253–259, February 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ding:2024:ERM

[DTH^{+}24] Bo Ding, Wei Tong, Yu Hua, Zhangyu Chen, Xueliang Wei, and Dan Feng. Enabling reliable memory-mapped I/O with auto-snapshot for persistent memory systems. IEEE Transactions on Computers, 73(9):2290–2304, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deng:2025:SCS

[DTL^{+}25] Zhihong Deng, Chunming Tang, Taotao Li, Zhikang Zeng, Parhat Abla, and Debiao He. SFPoW: Constructing secure and flexible proof-of-work sidechains for cross-chain interoperability with wrapped assets. IEEE Transactions on Computers, 74(7):2278–2292, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deng:2025:RCF

[DTZ^{+}25] Jinyi Deng, Xinru Tang, Jiahao Zhang, Yuxuan Li, Linyun Zhang, Fengbin Tu, Shaojun Wei, Yang Hu, and Shouyi Yin. Rethinking control flow in spatial architectures: Insights into control flow plane design. IEEE Transactions on Computers, 74(1):185– 199, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dimitrov:2022:FGR

[DVA22] Vassil Dimitrov, Luigi Vigneri, and Vidal Attias. Fast generation of RSA keys using smooth integers. IEEE Transactions on Computers, 71(7): 1575–1585, July 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

DAnvers:2023:RHO

[DVV23] Jan-Pieter D’Anvers, Michiel Van Beirendonck, and Ingrid Verbauwhede. Revisiting higher-order masked comparison for lattice-based cryptography: Algorithms and bitsliced implementations. IEEE Transactions on Computers, 72(2):321–332, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deng:2022:SKS

[DWL^{+}22] Ze Deng, Yue Wang, Tao Liu, Schahram Dustdar, Rajiv Ranjan, Albert Zomaya, Yizhi Liu, and Lizhe Wang. Spatial-keyword skyline publish/subscribe query processing over distributed sliding window streaming data. IEEE Transactions on Computers, 71(10):2659–2674, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dong:2025:RCS

[DWLF25] Hui Dong, Huaqun Wang, Mengjie Lv, and Weibei Fan. Reliable communication scheme based on completely independent spanning trees in data center networks. IEEE Transactions on Computers, 74(6):2003–2016, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Demirbaga:2022:AAR

[DWN^{+}22] Umit Demirbaga, Zhenyu Wen, Ayman Noor, Karan Mitra, Khaled Alwasel, Saurabh Garg, Albert Y. Zomaya, and Rajiv Ranjan. AutoDiagn: An automated real-time diagnosis framework for big data systems. IEEE Transactions on Computers, 71(5):1035–1048, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Du:2025:PSD

[DWW25] Qi Du, Feng Wang, and Chengkun Wu. Parallelization strategies for DeepMD-Kit using OpenMP: Enhancing efficiency in machine learningbased molecular simulations. IEEE Transactions on Computers, 74(10):3534–3545, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Du:2020:ACD

[DWYX20] M. Du, Y. Wang, K. Ye, and C. Xu. Algorithmics of cost-driven computation offloading in the edge-cloud environment. IEEE Transactions on Computers, 69(10): 1519–1532, October 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2025:CCR

[DYC^{+}25] Surong Dai, Jinni Yang, Wenyang Cui, Yaozheng Fang, and Ye Lu. CROSC: Compilation-runtime joint optimization for fast smart contract execution. IEEE Transactions on Computers, 74(9): 2962–2976, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dai:2020:GPC

[DYJ20] X. Dai, H. Yin, and N. K. Jha. Grow and prune compact, fast, and accurate LSTMs. IEEE Transactions on Computers, 69(3):441–452, March 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dey:2022:PPI

[DYPZ22] Moumita Dey, Baki Berkay Yilmaz, Milos Prvulovic, and Alenka Zaji . PRIMER: Profiling interrupts using electromagnetic side-channel for embedded devices. IEEE Transactions on Computers, 71(8): 1824–1838, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Deng:2024:DLD

[DZC^{+}24] Cai Deng, Xiangyu Zou, Qi Chen, Bo Tang, and Wen Xia. The design of a lossless deduplication scheme to eliminate fine-grained redundancy for JPEG image storage systems. IEEE Transactions on Computers, 73(5):1385–1399, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

El-Araby:2023:TCS

[EAMJ^{+}23] Esam El-Araby, Naveed Mahmud, Mingyoung Joshua Jeng, Andrew MacGillivray, Manu Chaudhary, Md. Alvir Islam Nobel, SM Ishraq Ul Islam, David Levy, Dylan Kneidel, Madeline R. Watson, Jack G. Bauer, and Andrew E. Riachi. Towards complete and scalable emulation of quantum algorithms on highperformance reconfigurable computers. IEEE Transactions on Computers, 72(8): 2350–2364, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

ElKhatib:2022:HPF

[EAMK22] Rami El Khatib, Reza Azarderakhsh, and Mehran Mozaffari-Kermani. High-performance FPGA accelerator for SIKE. IEEE Transactions on Computers, 71(6):1237–1248, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Esmaili-Dokht:2024:KVS

[EDGR^{+}24] Pouya Esmaili-Dokht, Miquel Guiot, Petar Radojkovi , Xavier Martorell, Eduard Ayguadé, Jesus Labarta, Jason Adlard, Paolo Amato, and Marco Sforzin. O(n) key value sort with active compute memory. IEEE Transactions on Computers, 73(5): 1341–1356, May 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Elish:2022:LED

[EEA22] Karim O. Elish, Mahmoud O. Elish, and Hussain M. J. Almohri. Lightweight, effective detection and characterization of mobile malware families. IEEE Transactions on Computers, 71(11):2982– 2995, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Egger:2021:CRA

[EGMW21] Daniel J. Egger, Ricardo García Gutiérrez, Jordi Cahué Mestre, and Stefan Woerner. Credit risk analysis using quantum computers. IEEE Transactions on Computers, 70(12): 2136–2145, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Eslami:2024:SSC

[EGP24] Mohammad Eslami, Tara Ghasempouri, and Samuel Pagliarini. SCARF: Securing chips with a robust framework against fabrication-time hardware trojans. IEEE Transactions on Computers, 73 (12):2761–2775, 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

El-Razouk:2021:NDS

[ERKP21] Hayssam El-Razouk, Kirthi Kotha, and Mahidhar Puligunta. Novel GF(2^{m}) digitserial PISO multipliers for the self-dual Gaussian normal bases. IEEE Transactions on Computers, 70(10): 1732–1746, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Eissa:2025:SST

[ESdP^{+}25] Sherif Eissa, Sander Stuijk, Floran de Putter, Andrea Nardi-Dei, Federico Corradi, and Henk Corporaal. STEMS: Spatial-temporal mapping for spiking neural networks. IEEE Transactions on Computers, 74(9):2991–3002, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ebrahimi:2020:LIR

[ESN20] M. Ebrahimi, R. Sadeghi, and Z. Navabi. LUT input reordering to reduce aging impact on FPGA LUTs. IEEE Transactions on Computers, 69(10): 1500–1506, October 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Eckert:2023:EMM

[ESW^{+}23] Charles Eckert, Arun Subramaniyan, Xiaowei Wang, Charles Augustine, Ravishankar Iyer, and Reetuparna Das. Eidetic: an in-memory matrix multiplication accelerator for neural networks. IEEE Transactions on Computers, 72(6):1539–1553, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fairouz:2021:HAH

[FAFK21] Abbas A. Fairouz, Monther Abusultan, Viacheslav V. Fedorov, and Sunil P. Khatri. Hardware acceleration of hash operations in modern microprocessors. IEEE Transactions on Computers, 70(9): 1412–1426, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ferdman:2022:GEI

[FAKM22] Michael Ferdman, Jorge Albericio, Tushar Krishna, and Peter Milder. Guest editorial: IEEE TC special issue: Hardware acceleration of machine learning. IEEE Transactions on Computers, 71(12):3072– 3073, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Faraji:2020:HBU

[FB20] S. R. Faraji and K. Bazargan. Hybrid binary–unary hardware accelerator. IEEE Transactions on Computers, 69(9): 1308–1319, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Francq:2022:NTS

[FBH^{+}22] Julien Francq, Loïc Besson, Paul Huynh, Philippe Guillot, Gilles Millerioux, and Marine Minier. Non-triangular selfsynchronizing stream ciphers. IEEE Transactions on Computers, 71(1):134–145, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Forsberg:2021:HPE

[FBM21] B. Forsberg, L. Benini, and A. Marongiu. HePREM: A predictable execution model for GPU-based heterogeneous SoCs. IEEE Transactions on Computers, 70(1):17–29, January 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2023:VWE

[FCZ^{+}23] Xiaoliu Feng, Xianzhang Chen, Qingfeng Zhuge, Duo Liu, Edwin H.-M. Sha, and Chun Jason Xue. V-WAFA: an endurance variation aware fine-grained allocator for persistent memory. IEEE Transactions on Computers, 72(4): 998–1010, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Farbeh:2021:EUC

[FDKK21] H. Farbeh, L. Delshadtehrani, H. Kim, and S. Kim. ECC-United cache: Maximizing efficiency of error detection/ correction codes in associative cache memories. IEEE Transactions on Computers, 70(4): 640–654, April 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fang:2025:ATA

[FFG^{+}25] Weijie Fang, Yanggeng Fu, Jiaquan Gao, Longkun Guo, Gregory Gutin, and Xiaoyan Zhang. Acceleration of timingaware gate-level logic simulation through one-pass GPU parallelism. IEEE Transactions on Computers, 74(8): 2675–2686, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2022:MFH

[FHH22] Xincheng Feng, Ke Hu, and Kaining Han. MM-FSM: a high-efficiency general nonlinear function generator for stochastic computation. IEEE Transactions on Computers, 71(9):1998–2009, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2022:RRV

[FHL^{+}22] Lang Feng, Jiayi Huang, Luyi Li, Haochen Zhang, and Zhongfeng Wang. RvDfi: a RISC-V architecture with security enforcement by high performance complete dataflow integrity. IEEE Transactions on Computers, 71(10): 2499–2512, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2023:EEP

[FHL^{+}23] Qi Feng, Debiao He, Min Luo, Xinyi Huang, and Kim-Kwang Raymond Choo. EPRICE: an efficient and privacy-preserving real-time incentive system for crowdsensing in Industrial Internet of Things. IEEE Transactions on Computers, 72(9): 2482–2495, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fu:2022:JHD

[FHW^{+}22] Rongliang Fu, Junying Huang, Haibin Wu, Xiaochun Ye, Dongrui Fan, and Tsung-Yi Ho. JBNN: a hardware design for binarized neural networks using single-flux-quantum circuits. IEEE Transactions on Computers, 71(12):3203– 3214, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ficco:2022:MAC

[Fic22] Massimo Ficco. Malware analysis by combining multiple detectors and observation windows. IEEE Transactions on Computers, 71(6):1276–1290, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Faldella:2021:PWC

[FL21] Eugenio Faldella and Daniela Loreti. Precise worst-case blocking time of tasks under priority inheritance protocol. IEEE Transactions on Computers, 70(11):1901– 1913, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fraccaroli:2020:AGA

[FLF20] E. Fraccaroli, M. Lora, and F. Fummi. Automatic generation of analog/mixed signal virtual platforms for smart systems. IEEE Transactions on Computers, 69(9):1263– 1278, September 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2020:ALC

[FLS20] Q. Fan, D. J. Lilja, and S. S. Sapatnekar. Adaptive-length coding of image data for low-cost approximate storage. IEEE Transactions on Computers, 69(2):239–252, February 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Freitas:2021:LEC

[FMM^{+}21] David C. C. Freitas, David F. M. Mota, César Marcon, Jarbas A. N. Silveira, and João C. M. Mota. LPC: An error correction code for mitigating faults in 3D memories. IEEE Transactions on Computers, 70(11):2001– 2012, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feliu:2022:VVM

[FNS^{+}22] Josué Feliu, Ajeya Naithani, Julio Sahuquillo, Salvador Petit, Moinuddin Qureshi, and Lieven Eeckhout. VMT: Virtualized multi-threading for accelerating graph workloads on commodity processors. IEEE Transactions on Computers, 71(6):1386–1398, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2025:HHA

[FPHW25] Xiaopeng Fan, Xiaoshuang Peng, Kaixin Huang, and Chuliang Weng. Hyte: a hotness-aware hybrid DRAM-PM native table storage engine. IEEE Transactions on Computers, 74(8):2593– 2607, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2025:REM

[FPX^{+}25a] Weibei Fan, Yao Pan, Fu Xiao, Mengjie Lv, Lei Han, and Shui Yu. Reliable and efficient multi-path transmission based on disjoint paths in data center networks. IEEE Transactions on Computers, 74(10): 3362–3376, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2025:HSN

[FPX^{+}25b] Weibei Fan, Yao Pan, Fu Xiao, Pinchang Zhang, Lei Han, and Sun-Yuan Hsieh. A highly scalable network architecture for optical data centers. IEEE Transactions on Computers, 74(10):3433– 3447, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fu:2023:FOS

[FQYS23] Wenwen Fu, Wei Quan, Jinli Yan, and Zhigang Sun. Fenglin-I: An open-source time-sensitive networking chip enabling agile customization. IEEE Transactions on Computers, 72(1):140–153, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fornt:2025:MGE

[FRFM^{+}25] Jordi Fornt, Enrico Reggiani, Pau Fontova-Musté, Narcís Rodas, Alessandro Pappalardo, Osman Sabri Unsal, Adrián Cristal Kestelman, Josep Altet, Francesc Moll, and Jaume Abella. Mix-GEMM: Extending RISC-V CPUs for energy-efficient mixed-precision DNN inference using binary segmentation. IEEE Transactions on Computers, 74(2):582– 596, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ferdaus:2023:AMH

[FTR23] Farah Ferdaus, B. M. S. Bahar Talukder, and Md Tauhidur Rahman. Approximate MRAM: High-performance and power-efficient computing with MRAM chips for errortolerant applications. IEEE Transactions on Computers, 72(3):668–681, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fotse:2025:FLB

[FTV25] Yannis Steve Nsuloun Fotse, Vianney Kengne Tchendji, and Mthulisi Velempini. Federated learning based DDoS attacks detection in large scale software-defined network. IEEE Transactions on Computers, 74(1):101– 115, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fraccaroli:2023:MCP

[FV23] Enrico Fraccaroli and Sara Vinco. Modeling cyberphysical production systems with SystemC-AMS. IEEE Transactions on Computers, 72(7):2039–2051, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Frankel:2023:SRB

[FW23] Binyamin Frankel and Shmuel Wimer. A self-refreshable bitcell for single-cycle refreshing of embedded memories. IEEE Transactions on Computers, 72(2):513–519, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fadiheh:2023:EAD

[FWM^{+}23] Mohammad Rahmani Fadiheh, Alex Wezel, Johannes Müller, Jörg Bormann, Sayak Ray, Jason M. Fung, Subhasish Mitra, Dominik Stoffel, and Wolfgang Kunz. An exhaustive approach to detecting transient execution side channels in RTL designs of processors. IEEE Transactions on Computers, 72(1): 222–235, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fyrbiak:2020:GSA

[FWR^{+}20] M. Fyrbiak, S. Wallat, S. Reinhard, N. Bissantz, and C. Paar. Graph similarity and its applications to hardware security. IEEE Transactions on Computers, 69(4):505–519, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2021:APA

[FWZ^{+}21] Hao Fan, Song Wu, Xinyu Zhao, Zhenjiang Xie, Sheng Di, Jiang Xiao, Chen Yu, and Hai Jin. Accelerating parallel applications in cloud platforms via adaptive timeslice control. IEEE Transactions on Computers, 70(7): 992–1005, July 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2023:DPC

[FXC^{+}23] Weibei Fan, Fu Xiao, Hui Cai, Xiaobai Chen, and Shui Yu. Disjoint paths construction and fault-tolerant routing in BCube of data center networks. IEEE Transactions on Computers, 72(9): 2467–2481, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2025:ERS

[FXJW25] Kai Feng, Guodong Xie, Zhangjian Ji, and Dajin Wang. Evaluating robustness of subnetworks for the splitstar network. IEEE Transactions on Computers, 74(9): 3087–3098, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fan:2024:EFT

[FXL^{+}24] Weibei Fan, Fu Xiao, Mengjie Lv, Lei Han, and Shui Yu. Efficient fault-tolerant path embedding for 3D torus network using locally faulty blocks. IEEE Transactions on Computers, 73(9):2305– 2319, September 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2024:TRN

[FYR^{+}24] Jun Feng, Laurence T. Yang, Bocheng Ren, Deqing Zou, Mianxiong Dong, and Shunli Zhang. Tensor recurrent neural network with differential privacy. IEEE Transactions on Computers, 73(3): 683–693, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Feng:2022:BBB

[FZG^{+}22] Lei Feng, Yiqi Zhao, Shaoyong Guo, Xuesong Qiu, Wenjing Li, and Peng Yu. BAFL: A blockchain-based asynchronous federated learning framework. IEEE Transactions on Computers, 71(5): 1092–1103, May 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Fu:2023:EED

[FZM^{+}23] Luwei Fu, Zhiwei Zhao, Geyong Min, Wang Miao, Liang Zhao, and Wenjie Huang. Energy-efficient 3-D data collection for multi-UAV assisted mobile crowdsensing. IEEE Transactions on Computers, 72(7):2025–2038, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Genssler:2022:BIC

[GA22] Paul R. Genssler and Hussam Amrouch. Brain-inspired computing for circuit reliability characterization. IEEE Transactions on Computers, 71(12): 3336–3348, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gong:2021:VML

[GCL^{+}21] C. Gong, Y. Chen, Y. Lu, T. Li, C. Hao, and D. Chen. VecQ: Minimal loss DNN model compression with vectorized weight quantization. IEEE Transactions on Computers, 70(5):696–710, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gruin:2023:MTP

[GCR^{+}23] Alban Gruin, Thomas Carle, Christine Rochange, Hugues Cassé, and Pascal Sainrat. MINOTAuR: A timing predictable RISC-V core featuring speculative execution. IEEE Transactions on Computers, 72(1):183–195, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guthmuller:2024:XRV

[GFB^{+}24] Eric Guthmuller, César Fuguet, Andrea Bocco, Jérôme Fereyre, Riccardo Alidori, Ihsane Tahir, and Yves Durand. Xvpfloat: RISC-V ISA extension for variable extended precision floating point computation. IEEE Transactions on Computers, 73(7):1683–1697, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2022:EEL

[GGZC22] Yuanning Gao, Xiaofeng Gao, Ruisi Zhang, and Guihai Chen. An end-to-end learningbased metadata management approach for distributed file systems. IEEE Transactions on Computers, 71(5): 1021–1034, May 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ghatak:2021:CDB

[Gha21] Gourab Ghatak. A changedetection-based Thompson sampling framework for nonstationary bandits. IEEE Transactions on Computers, 70(10):1670–1676, October 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ghasemi:2025:EED

[GHK^{+}25] Mehdi Ghasemi, Soroush Heidari, Young Geun Kim, Carole-Jean Wu, and Sarma Vrudhula. Energy-efficient, delay-constrained edge computing of a network of DNNs. IEEE Transactions on Computers, 74(2):569–581, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Graillat:2020:TII

[GJ20] S. Graillat and F. Jézéquel.

Tight interval inclusions with compensated algorithms. IEEE Transactions on Computers, 69(12):1774–1783, December 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gurevin:2025:SRA

[GJN^{+}25] Deniz Gurevin, Chenglu Jin, Phuong Ha Nguyen, Omer Khan, and Marten van Dijk. Secure remote attestation with strong key insulation guarantees. IEEE Transactions on Computers, 74(3): 848–859, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Giardino:2020:PPA

[GKFF20] M. Giardino, E. Klawitter, B. Ferri, and A. Ferri. A power- and performanceaware software framework for control system applications. IEEE Transactions on Computers, 69(10):1544–1555, October 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2022:EOD

[GKT^{+}22] Yansong Gao, Minki Kim, Chandra Thapa, Alsharif Abuadbba, Zhi Zhang, Seyit Camtepe, Hyoungshick Kim, and Surya Nepal. Evaluation and optimization of distributed machine learning techniques for Internet of Things. IEEE Transactions on Computers, 71(10):2538– 2552, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2024:SRB

[GL24] Wenbo Guo and Shuguo Li. Split-radix based compact hardware architecture for CRYSTALS-Kyber. IEEE Transactions on Computers, 73(1):97–108, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Godard:2021:ECP

[GLB21] Paul Godard, Vincent Loechner, and Cédric Bastoul. Efficient out-of-core and outof-place rectangular matrix transposition and rotation. IEEE Transactions on Computers, 70(11):1942–1948, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2023:RBA

[GLGL23] Yuyan Guo, Zhenhua Lu, Hui Ge, and Jiguo Li. Revocable blockchain-aided attributebased encryption with escrowfree in cloud storage. IEEE Transactions on Computers, 72(7):1901–1912, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Galli:2025:DLA

[GLMZ25] Davide Galli, Francesco Lattari, Matteo Matteucci, and Davide Zoni. A deep learning-assisted template attack against dynamic frequency scaling countermeasures. IEEE Transactions on Computers, 74(1):293– 306, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2025:DSD

[GLS^{+}25] Kejun Guo, Fuliang Li, Jiaxing Shen, Xingwei Wang, and Jiannong Cao. Distributed sketch deployment for software switches. IEEE Transactions on Computers, 74(4): 1210–1223, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2024:AWA

[GLW^{+}24] Yunfei Gu, Linhui Liu, Chentao Wu, Jie Li, and Minyi Guo. Ada-WL: an adaptive wear-leveling aware data migration approach for flexible SSD array scaling in clusters. IEEE Transactions on Computers, 73(8):1967–1982, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Geng:2024:PBD

[GLZ^{+}24] Hanfei Geng, Yifei Liu, Yujie Zheng, Li Lyna Zhang, Jingwei Sun, Yujing Wang, Yang Wang, Guangzhong Sun, Mao Yang, Ting Cao, and Yunxin Liu. PruneAug: Bridging DNN pruning and inference latency on diverse sparse platforms using automatic layerwise block pruning. IEEE Transactions on Computers, 73(11):2576–2589, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gouert:2024:JCP

[GMT24] Charles Gouert, Dimitris Mouris, and Nektarios Georgios Tsoutsos. Juliet: a configurable processor for computing on encrypted data. IEEE Transactions on Computers, 73(9):2335–2349, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Galimberti:2022:ESF

[GMZ22] Andrea Galimberti, Gabriele Montanaro, and Davide Zoni. Efficient and scalable FPGA design of GF(2^{m}) inversion for post-quantum cryptosystems. IEEE Transactions on Computers, 71(12):3295– 3307, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gonzalez-Navarro:2020:NRN

[GNH20] S. González-Navarro and J. Hormigo. New results on non-normalized floating-point formats. IEEE Transactions on Computers, 69(12):1733– 1744, December 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Godunov:2020:ACC

[God20] A. Godunov. Algorithms for calculating correctly rounded exponential function in doubleprecision arithmetic. IEEE Transactions on Computers, 69(9):1388–1400, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Giles:2020:KFD

[GPH20] C. E. Giles, C. L. Peterson, and M. A. Heinrich. Knight-Sim: A fast discrete eventdriven simulation methodology for computer architectural simulation. IEEE Transactions on Computers, 69(1): 65–71, January 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guan:2022:FSA

[GPQ22] Fei Guan, Long Peng, and Jiaqing Qiao. A fluid scheduling algorithm for DAG tasks with constrained or arbitrary deadlines. IEEE Transactions on Computers, 71(8): 1860–1873, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guan:2023:NFS

[GPQ23] Fei Guan, Long Peng, and Jiaqing Qiao. A new federated scheduling algorithm for arbitrary-deadline DAG tasks. IEEE Transactions on Computers, 72(8):2264–2277, August 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2023:DSS

[GPRV23] Yiqin Gao, Guillaume Pallez, Yves Robert, and Frédéric Vivien. Dynamic scheduling strategies for firm semiperiodic real-time tasks. IEEE Transactions on Computers, 72(1):55–68, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guan:2021:DFR

[GQH21] F. Guan, J. Qiao, and Y. Han. DAG-Fluid: A real-time scheduling algorithm for DAGs. IEEE Transactions on Computers, 70(3): 471–482, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2022:ETD

[GQJ^{+}22] Shaoyong Guo, Yuanyuan Qi, Yi Jin, Wenjing Li, Xuesong Qiu, and Luoming Meng. Endogenous trusted DRL-Based service function chain orchestration for IoT. IEEE Transactions on Computers, 71(2):

397–406, February 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gai:2021:EAH

[GQZ21] K. Gai, X. Qin, and L. Zhu.

An energy-aware high performance task allocation strategy in heterogeneous fog computing environments. IEEE Transactions on Computers, 70(4):626–639, April 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ghosh:2023:MFH

[GR23] Surajeet Ghosh and Sanchita Saha Ray. O(N) memory-free hardware architecture for Burrows–Wheeler Transform. IEEE Transactions on Computers, 72(7): 2080–2093, July 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Geyer:2023:NCF

[GSB23] Fabien Geyer, Alexander Scheffler, and Steffen Bondorf. Network calculus with flow prolongation: a feedforward FIFO analysis enabled by ML. IEEE Transactions on Computers, 72(1):97–110, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gong:2023:ORC

[GSC^{+}23] Guowen Gong, Zhirong Shen, Liang Chen, Suzhen Wu, Xiaolu Li, Patrick P. C. Lee, Zhiguo Wan, and Jiwu Shu. Optimal rack-coordinated updates in erasure-coded data centers: Design and analysis. IEEE Transactions on Computers, 72(7):1871–1885, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ghose:2022:PLC

[GSK^{+}22] Anirban Ghose, Siddharth Singh, Vivek Kulaharia, Lokesh Dokara, Srijeeta Maity, and Soumyajit Dey. PySchedCL: Leveraging concurrency in heterogeneous data-parallel systems. IEEE Transactions on Computers, 71(9):2234– 2247, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2023:UMT

[GSS^{+}23] Jinyu Gu, Jiacheng Shi, Haroran Su, Wentai Li, Binyu Zang, Haibing Guan, and Haibo Chen. Understanding and mitigating twin function misuses in operating system kernel. IEEE Transactions on Computers, 72(8): 2181–2193, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guan:2020:CMA

[GSY^{+}20] Y. Guan, G. Sun, Z. Yuan, X. Li, N. Xu, S. Chen, J. Cong, and Y. Xie. Crane: Mitigating accelerator underutilization caused by sparsity irregularities in CNNs. IEEE Transactions on Computers, 69(7):931–943, July 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gohari:2025:TUF

[GVN25] Pourya Gohari, Jeroen Voeten, and Mitra Nasri. Towards a unified framework for modeling and analyzing userdefined online non-preemptive scheduling policies. IEEE Transactions on Computers, 74(10):3347–3361, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Genssler:2022:RFC

[GvSHA22] Paul R. Genssler, Victor M.

van Santen, Jörg Henkel, and Hussam Amrouch. On the reliability of FeFET on-chip memory. IEEE Transactions on Computers, 71(4):947–958, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ge:2025:AWO

[GWC^{+}25] Mengke Ge, Junpeng Wang, Binhan Chen, Yingjian Zhong, Haitao Du, Song Chen, and Yi Kang. Allspark: Workload orchestration for visual transformers on processing in-memory systems. IEEE Transactions on Computers, 74(2):427–441, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ganfure:2023:DIP

[GWCS23] Gaddisa Olani Ganfure, Chun-Feng Wu, Yuan-Hao Chang, and Wei-Kuan Shih. Deep-Ware: Imaging performance counters with deep learning to detect ransomware. IEEE Transactions on Computers, 72(3):600–613, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gamatie:2025:UIS

[GWD25] Abdoulaye Gamatié, Yuyang Wang, and Diego Valdez Duran. Uncovering the intricacies and synergies of processor microarchitecture mechanisms using explainable AI. IEEE Transactions on Computers, 74(2):637–651, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2024:EGR

[GWG^{+}24] Yingxue Gao, Teng Wang, Lei Gong, Chao Wang, Yiqing Hu, Yi Yang, Zhongming Liu, Xi Li, and Xuehai Zhou. Enhancing graph random walk acceleration via efficient dataflow and hybrid memory architecture. IEEE Transactions on Computers, 73(3):887–901, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2025:HAV

[GWG^{+}25] Yingxue Gao, Teng Wang, Lei Gong, Chao Wang, Dong Dai, Yang Yang, Xianglan Chen, Xi Li, and Xuehai Zhou. Hardware accelerated vision transformer via heterogeneous architecture design and adaptive dataflow mapping. IEEE Transactions on Computers, 74(4):1224–1238, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2023:IIV

[GWH^{+}23] Chengsi Gao, Ying Wang, Yinhe Han, Weiwei Chen, and Lei Zhang. IVP: An intelligent video processing architecture for video streaming. IEEE Transactions on Computers, 72(1):264–277, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gong:2023:EFM

[GWX^{+}23] Lei Gong, Chao Wang, Haojun Xia, Xianglan Chen, Xi Li, and Xuehai Zhou. Enabling fast and memoryefficient acceleration for pattern matching workloads: The Lightweight Automata Processing Engine. IEEE Transactions on Computers, 72(4):

1011–1025, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2021:EHS

[GWZ^{+}21] Jinyu Gu, Xinyue Wu, Bojun Zhu, Yubin Xia, Binyu Zang, Haibing Guan, and Haibo Chen. Enclavisor: A hardware-software co-design for enclaves on untrusted cloud. IEEE Transactions on Computers, 70(10):1598– 1611, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2024:ASD

[GXL^{+}24] Cong Guo, Fengchen Xue, Jingwen Leng, Yuxian Qiu, Yue Guan, Weihao Cui, Quan Chen, and Minyi Guo. Accelerating sparse DNNs based on tiled GEMM. IEEE Transactions on Computers, 73(5): 1275–1289, May 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2023:CCS

[GXY^{+}23] Yihao Guo, Minghui Xu, Dongxiao Yu, Yong Yu, Rajiv Ranjan, and Xiuzhen Cheng. Cross-channel: Scalable offchain channels supporting fair and atomic cross-chain operations. IEEE Transactions on Computers, 72(11):3231– 3244, November 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2023:FMV

[GXZ^{+}23] Hechuan Guo, Minghui Xu, Jiahao Zhang, Chunchi Liu, Dongxiao Yu, Schahram Dustdar, and Xiuzhen Cheng. FileDAG: a multi-version decentralized storage network built on DAG-based blockchain. IEEE Transactions on Computers, 72(11): 3191–3202, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2024:BDB

[GXZ^{+}24] Hechuan Guo, Minghui Xu, Jiahao Zhang, Chunchi Liu, Rajiv Ranjan, Dongxiao Yu, and Xiuzhen Cheng. BFT-DSN: a Byzantine fault-tolerant decentralized storage network. IEEE Transactions on Computers, 73(5):1300–1312, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gong:2022:AHC

[GYH^{+}22] Yu Gong, Miao Yin, Lingyi Huang, Chunhua Deng, and Bo Yuan. Algorithm and hardware co-design of energyefficient LSTM networks for video recognition with hierarchical Tucker tensor decomposition. IEEE Transactions on Computers, 71(12):3101– 3114, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gao:2024:UUS

[GYS^{+}24] Wei Gao, Zhisheng Ye, Peng Sun, Tianwei Zhang, and Yonggang Wen. UniSched: a unified scheduler for deep learning training jobs with different user demands. IEEE Transactions on Computers, 73(6):1500–1515, June 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2025:HRR

[GYZC25] Hao Guo, Lei Yang, Qingfeng Zhang, and Jiannong Cao. Hybrid redundancy for reliable task offloading in collaborative edge computing. IEEE Transactions on Computers, 74(9): 3238–3250, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Gu:2021:SDC

[GZC^{+}21] H. Gu, J. Zhang, M. Chen, T. Wei, L. Lei, and F. Xie. Specification-driven conformance checking for virtual/ silicon devices using mutation testing. IEEE Transactions on Computers, 70(3): 400–413, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2023:SCD

[GZG^{+}23] Shaoyong Guo, Keqin Zhang, Bei Gong, Liandong Chen, Yinlin Ren, Feng Qi, and Xuesong Qiu. Sandbox computing: a data privacy trusted sharing paradigm via blockchain and federated learning. IEEE Transactions on Computers, 72(3): 800–810, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Guo:2021:BNA

[GZW^{+}21] Ruixin Guo, Feng Zhang, Lizhe Wang, Wusheng Zhang, Xinya Lei, Rajiv Ranjan, and Albert Y. Zomaya. BaPa: A novel approach of improving load balance in parallel matrix factorization for recommender systems. IEEE Transactions on Computers, 70(5):789–802, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Heidari:2025:LMB

[HA25] Mahsa Heidari and Bijan Alizadeh. Localizing multiple bugs in RTL designs by classifying hit-statements using neural networks. IEEE Transactions on Computers, 74(5): 1786–1799, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hassan:2023:DTC

[Has23] Mohamed Hassan. DISCO: Time-compositional cache coherence for multi-core realtime embedded systems. IEEE Transactions on Computers, 72(4):1163–1177, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2021:VAA

[HBB^{+}21] Jiayi Huang, Shilpa Bhosekar, Rahul Boyapati, Ningyuan Wang, Byul Hur, Ki Hwan Yum, and Eun Jung Kim. A voting approach for adaptive network-on-chip powergating. IEEE Transactions on Computers, 70(11):1962– 1975, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hautala:2020:TPE

[HBS20a] I. Hautala, J. Boutellier, and O. Silvén. TTADF: Power efficient dataflow-based multicore co-design flow. IEEE Transactions on Computers, 69(1): 51–64, January 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2020:LKE

[HBS^{+}20b] J. Hu, M. Baldi, P. Santini, N. Zeng, S. Ling, and H. Wang. Lightweight key encapsulation using LDPC codes on FPGAs. IEEE Transactions on Computers, 69(3):

327–341, March 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2024:TAS

[HC24] Zaobo He and Zhipeng Cai.

Trading aggregate statistics over private Internet of Things data. IEEE Transactions on Computers, 73(2):394– 407, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2023:HHG

[HCC^{+}23] Shuibing He, Ping Chen, Shuaiben Chen, Zheng Li, Siling Yang, Weijian Chen, and Lidan Shou. HOME: a holistic GPU memory management framework for deep learning. IEEE Transactions on Computers, 72(3):826–838, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2025:CSC

[HCC^{+}25] Lixia Han, Yiyang Chen, Siyuan Chen, Haozhang Yang, Ao Shi, Guihai Yu, Jiaqi Li, Zheng Zhou, Yijiao Wang, Yanzhi Wang, Xiaoyan Liu, Jinfeng Kang, and Peng Huang. CIMUS: 3D-stacked computing-in-memory under image sensor architecture for efficient machine vision. IEEE Transactions on Computers, 74(7):2321–2333, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2025:SPS

[HCKK25] Wontak Han, Hyunjun Cho, Donghyuk Kim, and Joo-Young Kim. SAL-PIM: a subarray-level processing-in-memory architecture with LUT-based linear interpolation for transformer-based text generation. IEEE Transactions on Computers, 74(9): 2909–2922, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hadjilambrou:2021:HCE

[HDAS21] Zacharias Hadjilambrou, Shidhartha Das, Marco A. Antoniades, and Yiannakis Sazeides. Harnessing CPU electromagnetic emanations for resonance-induced voltagenoise characterization. IEEE Transactions on Computers, 70(9):1338–1349, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2024:CLE

[HDL^{+}24] Qilin Hu, Yan Ding, Chubo Liu, Keqin Li, Kenli Li, and Albert Y. Zomaya. CBANA: a lightweight, efficient, and flexible cache behavior analysis framework. IEEE Transactions on Computers, 73(9): 2262–2274, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Houssam-Eddine:2021:HDT

[HECC^{+}21] Zahaf Houssam-Eddine, Nicola Capodieci, Roberto Cavicchioli, Giuseppe Lipari, and Marko Bertogna. The HPC-DAG task model for heterogeneous real-time systems. IEEE Transactions on Computers, 70(10):1747–1761, October 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hoffmann:2022:OML

[HF22] José Luis Conradi Hoffmann and Antônio Augusto Fröhlich. Online machine learning for energy-aware multicore real-time embedded systems. IEEE Transactions on Computers, 71(2):493– 505, February 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hanlon:2023:FHP

[HF23] James Hanlon and Stephen Felix. A fast hardware pseudorandom number genera-

tor based on xoroshiro128.

IEEE Transactions on Computers, 72(5):1518–1524, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2025:CCO

[HFT^{+}25] Zhuo Huang, Hao Fan, Bin Tang, Song Wu, Chen Yu, and Hai Jin. CBuild: Clusteroriented collaborative image building for containers. IEEE Transactions on Computers, 74(9):2870–2881, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2022:FBM

[HGC^{+}22] Xing Huang, Wenzhong Guo, Zhisheng Chen, Bing Li, Tsung-Yi Ho, and Ulf Schlichtmann. Flow-based microfluidic biochips with distributed channel storage: Synthesis, physical design, and wash optimization. IEEE Transactions on Computers, 71(2): 464–478, February 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Heidari:2022:CCA

[HGK^{+}22] Soroush Heidari, Mehdi Ghasemi, Young Geun Kim, Carole-Jean Wu, and Sarma Vrudhula. CAMDNN: Contentaware mapping of a network of deep neural networks on edge MPSoCs. IEEE Transactions on Computers, 71(12):3191– 3202, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hasegawa:2023:PRH

[HHN^{+}23] Kento Hasegawa, Seira Hidano, Kohei Nozawa, Shinsaku Kiyomoto, and No-

zomu Togawa. R-HTDetector:

Robust hardware-trojan detection based on adversarial training. IEEE Transactions on Computers, 72(2): 333–345, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2020:HLA

[HHPB20] M. Han, J. Hyun, S. Park, and W. Baek. Hotnessand lifetime-aware data placement and migration for highperformance deep learning on heterogeneous memory systems. IEEE Transactions on Computers, 69(3): 377–391, March 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hao:2023:SBA

[HHZ^{+}23] Tianshu Hao, Kai Hwang, Jianfeng Zhan, Yuejin Li, and Yong Cao. Scenario-based AI benchmark evaluation of distributed cloud/edge computing systems. IEEE Transactions on Computers, 72(3): 719–731, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hewage:2025:FCA

[HIRB25] Tharindu B. Hewage, Shashikant Ilager, Maria A. Rodriguez, and Rajkumar Buyya. A framework for carbon-aware real-time workload management in clouds using renewables-driven cores. IEEE Transactions on Computers, 74(8):2757–2771, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hong:2025:PCS

[HJX^{+}25] Qinghui Hong, Haoyou Jiang, Pingdan Xiao, Sichun Du, and Tao Li. A parallel computing scheme utilizing memristor crossbars for fast corner detection and rotation invariance in the ORB algorithm. IEEE Transactions on Computers, 74(3):996–1010, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2022:PDS

[HJYL22] Jianjun Huang, Jiasheng Jiang, Wei You, and Bin Liang. Precise dynamic symbolic execution for nonuniform data access in smart contracts. IEEE Transactions on Computers, 71(7):1551–1563, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hameed:2021:IPB

[HKC21] Fazal Hameed, Asif Ali Khan, and Jeronimo Castrillon. Improving the performance of block-based DRAM caches via tag-data decoupling. IEEE Transactions on Computers, 70(11):1914–1927, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hakert:2023:RRM

[HKC^{+}23] Christian Hakert, Asif Ali Khan, Kuan-Hsun Chen, Fazal Hameed, Jeronimo Castrillon, and Jian-Jia Chen. ROLLED: Racetrack memory optimized linear layout and efficient decomposition of decision trees. IEEE Transactions on Computers, 72(5): 1488–1502, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2020:WSA

[HKS20] K. Han, H. Kim, and D. Shin. WAL-SSD: Address remapping-based write-ahead-logging solid-state disks. IEEE Transactions on Computers, 69(2):260–273, February 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2022:SVM

[HLC^{+}22] Xing Hu, Ling Liang, Xiaobing Chen, Lei Deng, Yu Ji, Yufei Ding, Zidong Du, Qi Guo, Timothy Sherwood, and Yuan Xie. A systematic view of model leakage risks in deep neural network systems. IEEE Transactions on Computers, 71(12):3254– 3267, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2023:ILB

[HLF^{+}23] Shaohan Huang, Yi Liu, Carol Fung, He Wang, Hailong Yang, and Zhongzhi Luan. Improving log-based anomaly detection by pre-training hierarchical transformers. IEEE Transactions on Computers, 72(9):2656–2667, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2025:HIS

[HLJ^{+}25] Jiahui Huang, Zhan Li, Yuxian Jiang, Zhihan Zhang, Hao Wang, and Sheng Chang. A high-intensity solution of hardware accelerator for sparse and redundant computations in semantic segmentation models. IEEE Transactions on Computers, 74(9): 3129–3142, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hur:2025:HLV

[HLK^{+}25] Won Hur, Jiwon Lee, Jaewon Kwon, Minjae Kim, and Won Woo Ro. HashScape: Leveraging virtual address dynamics for efficient hashed page tables. IEEE Transactions on Computers, 74(6): 1872–1885, June 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2020:ADL

[HLL^{+}20] R. Han, C. H. Liu, S. Li, S. Wen, and X. Liu. Accelerating deep learning systems via critical set identification and model compression. IEEE Transactions on Computers, 69(7):1059–1070, July 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hsieh:2021:TTS

[HLLC21] Jen-Wei Hsieh, Yi-Yu Liu, Hung-Tse Lee, and Tai Chang. TSE: Two-step elimination for MLC STT-RAM last-level cache. IEEE Transactions on Computers, 70(9):1498– 1510, September 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hanson:2023:DDR

[HLQ^{+}23] Edward Hanson, Shiyu Li, Xuehai Qian, Hai Helen Li, and Yiran Chen. DyN-Namic: Dynamically reshaping, high data-reuse accelerator for compact DNNs. IEEE Transactions on Computers, 72(3):880–892, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2023:SGS

[HLS^{+}23a] Shujie Han, Patrick P. C.

Lee, Zhirong Shen, Cheng He, Yi Liu, and Tao Huang. StreamDFP: a general stream mining framework for adaptive disk failure prediction. IEEE Transactions on Computers, 72(2):520–534, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2023:ESR

[HLS^{+}23b] Chuang Hu, Rui Lu, Qianlong Sang, Huanghuang Liang, Dan Wang, Dazhao Cheng, Jin Zhang, Qing Li, and JunKun Peng. An edge-side realtime video analytics system with dual computing resource control. IEEE Transactions on Computers, 72(12):3399– 3415, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2023:HPH

[HLT^{+}23] Hao Huang, Xiao-Yang Liu, Weiqin Tong, Tao Zhang, Anwar Walid, and Xiaodong Wang. High performance hierarchical Tucker tensor learning using GPU tensor cores. IEEE Transactions on Computers, 72(2):452–465, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2025:TMT

[HLY^{+}25] Yitao Hu, Xiulong Liu, Guotao Yang, Linxuan Li, Kai Zeng, Zhixin Zhao, Sheng Chen, Laiping Zhao, Wenxin Li, and Keqiu Li. TightLLM: Maximizing throughput for LLM inference via adaptive offloading policy. IEEE Transactions on Computers, 74(7): 2195–2209, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2025:TOA

[HLZ^{+}25] Qiang He, Quanwei Li, Chuangchuang Zhang, Xingwei Wang, Yuanguo Bi, Liang Zhao, Ammar Hawbani, and Keping Yu. Task optimization allocation in vehicle based edge computing systems with deep reinforcement learning. IEEE Transactions on Computers, 74(9):3156– 3167, September 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hepola:2024:EEE

[HMJ24] Kari Hepola, Joonas Multanen, and Pekka Jääskeläinen. Energy-efficient exposed datapath architecture with a RISC-V instruction set mode. IEEE Transactions on Computers, 73(2):560–573, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2021:CRN

[HMK^{+}21] Jiayi Huang, Pritam Majumder, Sungkeun Kim, Troy Fulton, Ramprakash Reddy Puli, Ki Hwan Yum, and Eun Jung Kim. Computing en-route for near-data processing. IEEE Transactions on Computers, 70(6):906–921, June 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Haghbayan:2023:RTR

[HMMP23] Hashem Haghbayan, Antonio Miele, Onur Mutlu, and Juha Plosila. Run-time resource management in CMPs handling multiple aging mechanisms. IEEE Transactions on Computers, 72(10):2872– 2887, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huo:2025:SSF

[HMZ^{+}25] Pingyi Huo, Theodore Michailidis, Yi Zheng, Prapti Panigrahi, Kiwan Maeng, Jishen Zhao, and Vijaykrishnan Narayanan. SVDE: Serverless framework for lowlatency video analytic queries with hardware disaggregation. IEEE Transactions on Computers, 74(8):2717–2730, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Heinz:2023:CFD

[HP23] Daniel Heinz and Thomas Pöppelmann. Combined fault and DPA protection for lattice-based cryptography. IEEE Transactions on Computers, 72(4):1055–1066, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Heuser:2020:LCT

[HPGM20] A. Heuser, S. Picek, S. Guilley, and N. Mentens. Lightweight ciphers and their side-channel resilience. IEEE Transactions on Computers, 69(10):1434– 1448, October 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2022:EEG

[HPJK22] Yunki Han, Kangkyu Park, Youngbeom Jung, and Lee-Sup Kim. EGCN: an efficient GCN accelerator for minimizing off-chip memory access. IEEE Transactions on Computers, 71(12):3127– 3139, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hamann:2022:SCR

[HR22] Heiko Hamann and Andreagiovanni Reina. Scalability in computing and robotics. IEEE Transactions on Computers, 71(6):1453–1465, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2022:RBI

[HS22] Kyuhwa Han and Dongkun Shin. Remap-based interpartition copy for arrayed solid-state drives. IEEE Transactions on Computers, 71(7):1640–1654, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Harris:2024:UDS

[HSE^{+}24] David Harris, James Stine, Milo Ercegovac, Alberto Nannarelli, Katherine Parry, and Cedar Turek. Unified digit selection for radix-4 recurrence division and square root. IEEE Transactions on Computers, 73(1):292–300, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hua:2025:ROB

[HSH^{+}25] Yusheng Hua, Xuanhua Shi, Ligang He, Kang He, Teng Zhang, Hai Jin, and Yong Chen. RuYi: Optimizing burst buffer through automated, fine-grained processto-BB mapping. IEEE Transactions on Computers, 74(3): 955–967, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2025:DUD

[HSP^{+}25] Chun Huang, Jiaying Shao, Baolei Peng, Qingshuang Guo, Panlong Li, Junwei Sun, and Yanfeng Wang. Design of a universal decoder model based on DNA winner-takes-all neural networks. IEEE Transactions on Computers, 74(4): 1267–1277, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2025:DDO

[HTZ^{+}25] Zhaoyang Huang, Yanjie Tan, Yifu Zhu, Huailiang Tan, and Keqin Li. Dynamic DPU offloading and computational resource management in heterogeneous systems. IEEE Transactions on Computers, 74(9):3046–3058, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Heo:2022:APM

[HWC^{+}22a] Taekyung Heo, Yang Wang, Wei Cui, Jaehyuk Huh, and Lintao Zhang. Adaptive page migration policy with huge pages in tiered memory systems. IEEE Transactions on Computers, 71(1):53–68, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2022:SMB

[HWC^{+}22b] Yupeng Hu, Linjun Wu, Zhuojun Chen, Yun Huang, Xiaolin Xu, Keqin Li, and Jiliang Zhang. STT-MRAM-based reliable weak PUF2. IEEE Transactions on Computers, 71(7):1564–1574, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2023:EPR

[HWG^{+}23] Jingwei Hu, Wen Wang, Kris Gaj, Liping Wang, and Huaxiong Wang. Engineering practical rank-code-based cryptographic schemes on embedded hardware. A case study on ROLLO. IEEE Transactions on Computers, 72(7): 2094–2110, July 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2021:PAC

[HWJ^{+}21] Sitao Huang, Kun Wu, Hyunmin Jeong, Chengyue Wang, Deming Chen, and Wen-Mei Hwu. PyLog: An algorithmcentric Python-based FPGA programming and synthesis flow. IEEE Transactions on Computers, 70(12):2015– 2028, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2021:EPS

[HWL^{+}21] Dan Huang, Jun Wang, Qing Liu, Nong Xiao, Huafeng Wu, and Jiangling Yin. Enhancing proportional I/O sharing on containerized big data file systems. IEEE Transactions on Computers, 70(12):2083– 2097, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2024:LEL

[HWL^{+}24] Dengcheng Hu, Jianrong Wang, Xiulong Liu, Qi Li, and Keqiu Li. LMChain: an efficient load-migratable beaconbased sharding blockchain system. IEEE Transactions on Computers, 73(9):2178– 2191, September 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2024:VEC

[HWR^{+}24] Hang Huang, Honglei Wang, Jia Rao, Song Wu, Hao Fan, Chen Yu, Hai Jin, Kun Suo, and Lisong Pan. vKernel: Enhancing container isolation via private code and data. IEEE Transactions on Computers, 73(7):1711–1723, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2024:RLA

[HWX^{+}24] Meng Han, Liang Wang, Limin Xiao, Tianhao Cai, Zeyu Wang, Xiangrong Xu, and Chenhao Zhang. ReDas: a lightweight architecture for supporting fine-grained reshaping and multiple dataflows on systolic array. IEEE Transactions on Computers, 73(8): 1997–2011, August 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2022:STB

[HWZ^{+}22] Hongjing Huang, Zeke Wang, Jie Zhang, Zhenhao He, Chao Wu, Jun Xiao, and Gustavo Alonso. Shuhai: A tool for benchmarking high bandwidth memory on FPGAs. IEEE Transactions on Computers, 71(5):1133–1144, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Han:2025:APB

[HWZ^{+}25] Feixue Han, Yike Wang, Yunbo Zhang, Qing Li, Dayi Zhao, and Yong Jiang. Anole: a pragmatic blend of classic and learning-based algorithms in congestion control. IEEE Transactions on Computers, 74(7):2501–2514, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huo:2020:ITA

[HXGR20] Z. Huo, L. Xiao, M. Guo, and X. Rong. Incremental throughput allocation of heterogeneous storage with no disruptions in dynamic setting. IEEE Transactions on Computers, 69(5):679–698, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8946738.

Huang:2023:BQI

[HXL^{+}23] Tian Huang, Jun Xu, Tao Luo, Xiaozhe Gu, Rick Goh, and Weng-Fai Wong. Benchmarking quantum(-inspired) annealing hardware on practical use cases. IEEE Transactions on Computers, 72(6): 1692–1705, June 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hou:2025:IEM

[HXL^{+}25] Xiaofeng Hou, Cheng Xu, Chao Li, Jiacheng Liu, Xuehan Tang, Kwang-Ting Cheng, and Minyi Guo. Improving efficiency in multimodal autonomous embedded systems through Adaptive Gating. IEEE Transactions on Computers, 74(2): 691–704, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hasegawa:2025:NWH

[HYH^{+}25] Kento Hasegawa, Kazuki Yamashita, Seira Hidano, Kazuhide Fukushima, Kazuo Hashimoto, and Nozomu Togawa. Node-wise hardware trojan detection based on graph learning. IEEE Transactions on Computers, 74(3): 749–761, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2025:ESF

[HYL^{+}25] Yaodong Huang, Tingting Yao, Zelin Lin, Xiaojun Shang, Yukun Yuan, Laizhong Cui, and Yuanyuan Yang. Efficient service function chain placement over heterogeneous devices in deviceless edge computing environments. IEEE Transactions on Computers, 74(1):222–236, January 2025.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hua:2025:HHU

[HYQ^{+}25] Qin Hua, Dingyu Yang, Shiyou Qian, Jian Cao, Guangtao Xue, and Minglu Li. Humas: a heterogeneityand upgrade-aware microservice auto-scaling framework in large-scale data centers. IEEE Transactions on Computers, 74(3):968–982, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2020:OPA

[HYS^{+}20] S. He, Y. Yin, X. Sun, X. Zhang, and Z. Li. Optimizing parallel I/O accesses through pattern-directed and layout-aware replication. IEEE Transactions on Computers, 69(2):212–225, February 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2021:EEC

[HYW^{+}21] L. Huang, C. Yuan, J. Wang, M. Ebrahimi, X. Xie, and Q. Li. ECDR^{2}2: Error corrector and detector relocation router for network-on-chip. IEEE Transactions on Computers, 70(4):606–613, April 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

He:2025:WSL

[HZC^{+}25] Wei He, Zhi Zhang, Yueqiang Cheng, Wenhao Wang, Wei Song, Yansong Gao, Qifei Zhang, Kang Li, Dongxi Liu, and Surya Nepal. Whistle-Blower: a system-level empirical study on RowHammer. IEEE Transactions on Computers, 74(3):805–819, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2024:MGT

[HZF^{+}24] Zhuo Huang, Qi Zhang, Hao Fan, Song Wu, Chen Yu, Hai Jin, Jun Deng, Jing Gu, and Zhimin Tang. Multi-grained trace collection, analysis, and management of diverse container images. IEEE Transactions on Computers, 73(7): 1698–1710, July 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Haider:2024:DRA

[HZK24] Muhammad Hamis Haider, Hao Zhang, and Seok-Bum Ko. Decoder reduction approximation scheme for Booth multipliers. IEEE Transactions on Computers, 73(3): 735–746, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2023:ARL

[HZM^{+}23] Ming Hu, Min Zhang, Frédéric Mallet, Xin Fu, and Mingsong Chen. Accelerating reinforcement learning-based CCSL specification synthesis using curiosity-driven exploration. IEEE Transactions on Computers, 72(5):1431–1446, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2024:DMT

[HZMC24] Wenjie Huang, Zhiwei Zhao, Geyong Min, and Jiajun Chen. Distributed multihop task offloading in massive heterogeneous IoT systems. IEEE Transactions on Computers, 73(4):1126–1137, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2023:ABC

[HZR^{+}23] Hang Huang, Yuqing Zhao, Jia Rao, Song Wu, Hai Jin, Duoqiang Wang, Suo Kun, and Lisong Pan. Adapt burstable containers to variable CPU resources. IEEE Transactions on Computers, 72(3):614–626, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hadjadj:2023:VVC

[HZT^{+}23] Yacine Hadjadj, Chakib Mustapha Anouar Zouaoui, Nasreddine Taleb, Sarah Mazari, Mohamed El Bahri, and Miloud Chikr El Mezouar. VCMalloc: a virtually contiguous memory allocator. IEEE Transactions on Computers, 72(12):3431– 3442, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2024:CDB

[HZW^{+}24] Letian Huang, Tianjin Zhao, Ziren Wang, Junkai Zhan, Junshi Wang, and Xiaohang Wang. Component dependencies based network-on-chip test. IEEE Transactions on Computers, 73(12):2805–2816, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Huang:2022:IAH

[HZYY22] Yaodong Huang, Yiming Zeng, Fan Ye, and Yuanyuan Yang. Incentive assignment in hybrid consensus blockchain systems in pervasive edge environments. IEEE Transactions on Computers, 71(9): 2102–2115, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Iaria:2025:CST

[IBB^{+}25] Giusy Iaria, Paolo Bernardi, Claudia Bertani, Lorenzo Cardone, Giuseppe Garozzo, and Vincenzo Tancorre. A comprehensive scan test cost model to optimize the production of very large SoCs. IEEE Transactions on Computers, 74(4): 1278–1292, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Iparraguirre:2022:ACH

[IDFH22] Daniel Iparraguirre, José G. Delgado-Frias, and Howard Heck. Asymmetric crosstalk harness signaling for common eigenmode elimination. IEEE Transactions on Computers, 71(9):2048–2058, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ibrahim:2024:ASM

[IIEKS24] Abrar A. Ibrahim, Ahmed M. Y. Ibrahim, M. Watheq El-Kharashi, and Mona Safar. Adaptive SAT modeling for optimal pattern retargeting in IEEE 1687 networks. IEEE Transactions on Computers, 73(2):536–547, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Islam:2022:EHM

[IKAG^{+}22] Md Shohidul Islam, Khaled N. Khasawneh, Nael Abu-Ghazaleh, Dmitry Ponomarev, and Lei Yu. Efficient hardware malware detectors that are resilient to adversarial evasion. IEEE Transactions on Computers, 71(11):2872– 2887, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Imran:2022:CEA

[IKTY22] Muhammad Imran, Taehyun Kwon, Nur A. Touba, and Joon-Sung Yang. CEnT: An efficient architecture to eliminate intra-array write disturbance in PCM. IEEE Transactions on Computers, 71(5): 992–1007, May 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Imana:2021:LBB

[Ima21] J. L. Imaña. LFSR-based bit-serial GF (2^{m}) multipliers using irreducible trinomials. IEEE Transactions on Computers, 70(1):156–162, January 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Iserte:2021:DEC

[IMQOP21] Sergio Iserte, Rafael Mayo, Enrique S. Quintana-Ortí, and Antonio J. Peña. DMRlib: Easy-coding and efficient resource management for job malleability. IEEE Transactions on Computers, 70(9): 1443–1457, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Islam:2022:SBS

[IWKB22] Muhammed Tawfiqul Islam, Huaming Wu, Shanika Karunasekera, and Rajkumar Buyya. SLA-based scheduling of Spark jobs in hybrid cloud computing environments. IEEE Transactions on Computers, 71(5):1117–1132, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jeon:2024:TFB

[JBK24] Yongjin Jeon, Seungjun Baek, and Jongsung Kim. Toward finding S-box circuits with optimal multiplicative complexity. IEEE Transactions on Computers, 73(8): 2036–2050, August 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jung:2025:ODT

[JBS^{+}25] Victor Jean-Baptiste Jung, Alessio Burrello, Moritz Scherer, Francesco Conti, and Luca Benini. Optimizing the deployment of tiny transformers on low-power MCUs. IEEE Transactions on Computers, 74(2):526–541, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2022:ESC

[JCKH22] Sunghyun Jin, Sung Min Cho, HeeSeok Kim, and Seokhie Hong. Enhanced sidechannel analysis on ECDSA employing fixed-base comb method. IEEE Transactions on Computers, 71(9):2341– 2350, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2023:UBA

[JCY^{+}23] Xu Jiang, Zewei Chen, Maolin Yang, Nan Guan, Yue Tang, and Yi Wang. A unified blocking analysis for parallel tasks with spin locks under global fixed priority scheduling. IEEE Transactions on Computers, 72(1):15–28, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2023:AGC

[JCZ^{+}23] Hai Jin, Dan Chen, Long Zheng, Yu Huang, Pengcheng Yao, Jin Zhao, Xiaofei Liao, and Wenbin Jiang. Accelerating graph convolutional networks through a PIMaccelerated approach. IEEE Transactions on Computers, 72(9):2628–2640, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2023:HRI

[JDB^{+}23] Zhe Jiang, Xiaotian Dai, Alan Burns, Neil Audsley, Zonghua Gu, and Ian Gray. A highresilience imprecise computing architecture for mixedcriticality systems. IEEE Transactions on Computers, 72(1):29–42, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2023:PGP

[JDCL23] Lin Jiang, Anthony Dowling, Ming-C. Cheng, and Yu Liu. PODTherm-GP: a physicsbased data-driven approach for effective architecture-level thermal simulation of multicore CPUs. IEEE Transactions on Computers, 72(10): 2951–2962, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ji:2025:ECE

[JDH^{+}25] Xinyi Ji, Jiankuo Dong, Junhao Huang, Zhijian Yuan, Wangchen Dai, Fu Xiao, and Jingqiang Lin. ECO-CRYSTALS: Efficient cryptography CRYSTALS on standard RISC-V ISA. IEEE Transactions on Computers, 74(2):401–413, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2021:APR

[JGD^{+}21] X. Jiang, N. Guan, H. Du, W. Liu, and W. Yi. On the analysis of parallel realtime tasks with spin locks. IEEE Transactions on Computers, 70(2):199–211, February 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2023:LBE

[JHMM23] Rui Jin, Jia Hu, Geyong Min, and Jed Mills. Lightweight blockchain-empowered secure and efficient federated edge learning. IEEE Transactions on Computers, 72(11):3314– 3325, November 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jo:2022:DLG

[JJKP22] Yong-Yeon Jo, Myung-Hwan Jang, Sang-Wook Kim, and Sunju Park. A data layout with good data locality for single-machine based graph engines. IEEE Transactions on Computers, 71(8): 1784–1793, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2024:HET

[JJZW24] Lingxin Jin, Wei Jiang, Jinyu Zhan, and Xiangyu Wen. Highly evasive targeted bit-Trojan on deep neural networks. IEEE Transactions on Computers, 73(9):2350– 2363, September 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2022:AFB

[JKHL22] Yunho Jin, Shine Kim, Tae Jun Ham, and Jae W. Lee. Architecting a flashbased storage system for lowcost inference of extreme-scale DNNs. IEEE Transactions on Computers, 71(12):3153– 3164, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jang:2022:SSD

[JKK^{+}22] Yunho Jang, Gyuseong Kang, Taehwan Kim, Yeongkyo Seo, Kyung-Jin Lee, Byong-Guk Park, and Jongsun Park. Stochastic SOT device based SNN architecture for on-chip unsupervised STDP learning. IEEE Transactions on Computers, 71(9):2022– 2035, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jang:2025:BCL

[JKKP25] Yunho Jang, Dongsu Kim, Yeseul Kim, and Jongsun Park. Big-computing and little-storing STT-MRAM PIM architecture with charge domain based MAC operation. IEEE Transactions on Computers, 74(4):1239–1252, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jang:2024:ZNW

[JKNK24] Myeongjae Jang, Jinkwon Kim, Haejin Nam, and Soontae Kim. Zero and narrowwidth value-aware compression for quantized convolutional neural networks. IEEE Transactions on Computers, 73(1):249–262, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jian:2025:SRS

[JLD^{+}25] Zhaolong Jian, Xu Liu, Qiankun Dong, Longkai Cheng, Xueshuo Xie, and Tao Li. SmartZone: Runtime support for secure and efficient on-device inference on ARM TrustZone. IEEE Transactions on Computers, 74(6): 2144–2158, June 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jia:2020:EEF

[JLL^{+}20] L. Jia, Y. Liang, X. Li, L. Lu, and S. Yan. Enabling efficient fast convolution algorithms on GPUs via MegaKernels. IEEE Transactions on Computers, 69(7):986–997, July 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ju:2022:CNN

[JLL22] Sanghyeon Ju, Youngjoo Lee, and Sunggu Lee. Convolutional neural networks with discrete cosine transform features. IEEE Transactions on Computers, 71(12):3389– 3395, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2024:HMN

[JLP^{+}24] Zhang Jiang, Xianduo Li, Tianxiang Peng, Haoran Li, Jingxuan Hong, Jin Zhang, and Xiaoli Gong. Hybrid-Memcached: a novel approach for memcached persistence optimization with hybrid memory. IEEE Transactions on Computers, 73(7):1866–1874, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2025:MPT

[JLP^{+}25] Zhou Jin, Wenhao Li, Haojie Pei, Xiaru Zha, Yichao Dong, Xiang Jin, Xiao Wu, Dan Niu, and Wei W. Xing. ML-PTA: a two-stage MLenhanced framework for accelerating nonlinear DC circuit simulation with pseudotransient analysis. IEEE Transactions on Computers, 74(10):3319–3331, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2021:DCA

[JLY^{+}21] W. Jiang, Q. Lou, Z. Yan, L. Yang, J. Hu, X. S. Hu, and Y. Shi. Device-circuit-architecture co-exploration for computing-in-memory neural accelerators. IEEE Transactions on Computers, 70(4): 595–605, April 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ji:2021:RAR

[JLZ21] Y. Ji, Z. Liu, and Y. Zhang. A reduced architecture for ReRAM-based neural network accelerator and its software stack. IEEE Transactions on Computers, 70(3): 316–331, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2023:DDB

[JLZ^{+}23] Wei Jiang, Xinke Liao, Jinyu Zhan, Deepak Adhikari, and Ke Jiang. DE-SCO: Decomposition-based co-design to improve fault tolerance of security-critical tasks in cyber physical systems. IEEE Transactions on Computers, 72(6):1652–1665, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jha:2021:MDR

[JM21] Nandan Kumar Jha and Sparsh Mittal. Modeling data reuse in deep neural networks by taking data-types into cognizance. IEEE Transactions on Computers, 70(9):1526– 1538, September 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jia:2024:SRK

[JMW^{+}24] Ju Jia, Siqi Ma, Lina Wang, Yang Liu, and Robert H. Deng. A secure and robust knowledge transfer framework via stratified-causality distribution adjustment in intelligent collaborative services.

IEEE Transactions on Computers, 73(1):58–72, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2020:CCM

[JPHY20] H. Jiang, X. Peng, S. Huang, and S. Yu. CIMAT: A compute-in-memory architecture for on-chip training based on transpose SRAM arrays. IEEE Transactions on Computers, 69(7):944–954, July 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jie:2024:SFB

[JQK^{+}24] Wanqing Jie, Wangjie Qiu, Arthur Sandor Voundi Koe, Jianhong Li, Yin Wang, Yaqi Wu, Jin Li, and Zhiming Zheng. A secure and flexible blockchain-based offline payment protocol. IEEE Transactions on Computers, 73(2): 408–421, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2025:RPE

[JQY^{+}25] Qin Jiang, Saiyu Qi, Xu Yang, Yong Qi, Jianfeng Wang, Youshui Lu, Bochao An, and Ee-Chien Chang. Reducing paging and exit overheads in Intel SGX for oblivious conjunctive keyword search. IEEE Transactions on Computers, 74(3):776–789, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jaberipur:2025:BMA

[JRL25] Ghassem Jaberipur, Elham Rahman, and Jeong-A Lee. Balanced modular addition for the moduli set {2^{q}, 2^{q} ∓ 1, 2^{2q} + 1}2q,2q∓1, 22q + 1viamoduli−( 2^{q} ∓\sqrt{−1}2q∓− 1)adders.IEEETransactionsonComputers, 74

(1) : 316−−324, January2025.CODENIT COB4.ISSN0018− 9340(print), 1557−9956(electronic).

Jiang:2020:UTB

[JSTG20] X. Jiang, J. Sun, Y. Tang, and N. Guan. Utilizationtensity bound for real-time DAG tasks under global EDF scheduling. IEEE Transactions on Computers, 69(1): 39–50, January 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2022:BTP

[JWD^{+}22] Zhe Jiang, Ran Wei, Pan Dong, Yan Zhuang, Neil C. Audsley, and Ian Gray. Blue-Visor: Time-predictable hardware hypervisor for many-core embedded systems. IEEE Transactions on Computers, 71(9):2205–2218, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ji:2023:TOA

[JWG^{+}23] Tingxiang Ji, Xili Wan, Xinjie Guan, Aichun Zhu, and Feng Ye. Towards optimal application offloading in heterogeneous edge-cloud computing. IEEE Transactions on Computers, 72(11):3259– 3272, November 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jalali:2023:MBS

[JWK^{+}23] Zeinab S. Jalali, Chenghong Wang, Griffin Kearney, Geng Yuan, Caiwen Ding, Yinan Zhou, Yanzhi Wang, and Sucheta Soundarajan. Memristor-based spectral decomposition of matrices and its applications. IEEE Transactions on Computers, 72(5): 1460–1472, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2021:TIC

[JWS^{+}21] H. Jin, W. Wu, X. Shi, L. He, and B. B. Zhou. TurboDL: Improving the CNN training on GPU with fine-grained multistreaming scheduling. IEEE Transactions on Computers, 70(4):552–565, April 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2023:PSA

[JWS^{+}23] Hai Jin, Shuo Wei, Yan Sha, Chencheng Ye, Haikun Liu, and Xiaofei Liao. PM-LiteDB: Streamlining access paths for high-performance persistent memory document database systems. IEEE Transactions on Computers, 72(6):1778–1791, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2022:CDI

[JXH^{+}22] Yan Jin, Bowen Xiong, Kun He, Jin-Kao Hao, Chu-Min Li, and Zhang-Hua Fu. Clustering driven iterated hybrid search for vertex bisection minimization. IEEE Transactions on Computers, 71(10): 2370–2380, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2023:AIR

[JYF^{+}23] Zhe Jiang, Kecheng Yang, Nathan Fisher, Ian Gray, Neil C. Audsley, and Zheng Dong. AXI-IC^{RT} RT: Towards a real-time AXI-interconnect for highly integrated SoCs. IEEE Transactions on Computers, 72(3):786–799, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2024:FFA

[JYH^{+}24] Hao Jiang, Jintao Yang, Guang Hua, Lixia Li, Ying Wang, Shenghui Tu, and Song Xia. FAWA: Fast Adversarial Watermark Attack. IEEE Transactions on Computers, 73(2):301–313, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Javadi:2020:SCM

[JYM20] M. H. S. Javadi, M. H. Yalame, and H. R. Mahdiani. Small constant mean-error imprecise adder/multiplier for efficient VLSI implementation of MAC-based applications. IEEE Transactions on Computers, 69(9):1376– 1387, September 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jiang:2023:THR

[JYM^{+}23] Zhe Jiang, Kecheng Yang, Yunfeng Ma, Nathan Fisher, Neil Audsley, and Zheng Dong. Towards hard realtime and energy-efficient virtualization for many-core embedded systems. IEEE Transactions on Computers, 72(1): 111–126, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jin:2024:MRP

[JZH^{+}24] Hai Jin, Zhanyang Zhu, Ligang He, Yuhao Li, Yusheng Hua, and Xuanhua Shi. MM-DataLoader: Reusing preprocessed data among concurrent model training tasks. IEEE Transactions on Computers, 73(2):510–522, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jia:2024:EHP

[JZSD24] Weiwei Jia, Jiyuan Zhang, Jianchen Shan, and Xiaoning Ding. Effective huge page strategies for TLB miss reduction in nested virtualization. IEEE Transactions on Computers, 73(8):1983–1996, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jing:2025:NBF

[JZX^{+}25] Guanlin Jing, Yifei Zou, Minghui Xu, Yanqiang Zhang, Dongxiao Yu, Zhiguang Shan, Xiuzhen Cheng, and Rajiv Ranjan. Nicaea: a Byzantine fault tolerant consensus under unpredictable message delivery failures for parallel and distributed computing. IEEE Transactions on Computers, 74(3):915–928, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jing:2023:EFT

[JZY^{+}23] Guanlin Jing, Yifei Zou, Dongxiao Yu, Chuanwen Luo, and Xiuzhen Cheng. Efficient fault-tolerant consensus for collaborative services in edge computing. IEEE Transactions on Computers, 72(8): 2139–2150, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kishani:2020:MFR

[KAA20] M. Kishani, S. Ahmadian, and H. Asadi. A modeling framework for reliability of erasure codes in SSD arrays. IEEE Transactions on Computers, 69(5):649–665, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8944129.

Karabulut:2022:EFC

[KAA22] Emre Karabulut, Erdem Alkim, and Aydin Aysu. Efficient, flexible, and constanttime Gaussian sampling hardware for lattice cryptography. IEEE Transactions on Computers, 71(8):1810–1823, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Karanth:2024:EEF

[Kar24] Avinash Karanth. Editorial: EiC farewell and introduction of new EiC. IEEE Transactions on Computers, 73(1): 1–2, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kumar:2023:MLB

[KASAG23] Ajay Krishna Ananda Kumar, Sami Al-Salamin, Hussam Amrouch, and Andreas Gerstlauer. Machine learningbased microarchitecture-level power modeling of CPUs.

IEEE Transactions on Computers, 72(4):941–956, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2023:SSA

[KAWR23] Dong Eun Kim, Aayush Ankit, Cheng Wang, and Kaushik Roy. SAMBA: Sparsity aware in-memory computing based machine learning accelerator. IEEE Transactions on Computers, 72(9): 2615–2627, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kwon:2021:HQC

[KB21] Hyeokjea Kwon and Joonwoo Bae. A hybrid quantumclassical approach to mitigating measurement errors in quantum algorithms. IEEE Transactions on Computers, 70(9):1401–1411, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Klein:2023:AAM

[KBQ^{+}23] Joshua Klein, Irem Boybat, Yasir Mahmood Qureshi, Martino Dazzi, Alexandre Levisse, Giovanni Ansaloni, Marina Zapater, Abu Sebastian, and David Atienza. ALPINE: Analog in-memory acceleration with tight processor integration for deep learning. IEEE Transactions on Computers, 72(7):1985–1998, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kundu:2023:TSG

[KBR^{+}23] Shamik Kundu, Suvadeep Banerjee, Arnab Raha, Fei Su, Suriyaprakash Natarajan, and Kanad Basu. Troubleshooting at GAN point: Improving functional safety in deep learning accelerators. IEEE Transactions on Computers, 72(8):2194–2208, August 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Koo:2021:CAM

[KCAL21] J. Koo, C. Chung, Arvind, and S. Lee. A case for applicationmanaged flash. IEEE Transactions on Computers, 70(2): 240–254, February 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2020:MSM

[KCL^{+}20] B. Kim, J. Chung, E. Lee, W. Jung, S. Lee, J. Choi, J. Park, M. Wi, S. Lee, and J. H. Ahn. MViD: Sparse matrix–vector multiplication in mobile DRAM for accelerating recurrent neural networks. IEEE Transactions on Computers, 69(7):955–967, July 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kar:2023:DDM

[KCS23] Pushpendu Kar, Kewei Chen, and Jiayi Shi. DMACN: a dynamic multi-attribute caching mechanism for NDN-based remote health monitoring system. IEEE Transactions on Computers, 72(5):1301–1313, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kasarapu:2025:PEA

[KD25] Sreenitha Kasarapu and Sai Manoj Pudukotai Dinakarrao. Performance and environment-aware advanced driving assistance systems. IEEE Transactions on Computers, 74(1):131–142, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2024:GBA

[KDE^{+}24] Andrey Kim, Maxim Deryabin, Jieun Eom, Rakyong Choi, Yongwoo Lee, Whan Ghang, and Donghoon Yoo. General bootstrapping approach for RLWE-based homomorphic encryption. IEEE Transactions on Computers, 73(1): 86–96, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kuo:2023:RVG

[KGHRM23] Yao-Ming Kuo, Francisco García-Herrero, Oscar Ruano, and Juan Antonio Maestro.

RISC-V Galois Field ISA extension for non-binary errorcorrection codes and classical and post-quantum cryptography. IEEE Transactions on Computers, 72(3): 682–692, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2023:EAS

[KH23] Jangryul Kim and Soonhoi Ha. Energy-aware scenariobased mapping of deep learning applications onto heterogeneous processors under realtime constraints. IEEE Transactions on Computers, 72(6): 1666–1680, June 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2021:CCA

[KHHK21] Jinkwon Kim, Seokin Hong, Jeongkyu Hong, and Soontae Kim. CID: Co-architecting instruction cache and decompression system for embedded systems. IEEE Transactions on Computers, 70(7): 1132–1145, July 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kaushik:2021:DPC

[KHP21] Anirudh Mohan Kaushik, Mohamed Hassan, and Hiren Patel. Designing predictable cache coherence protocols for multi-core real-time systems. IEEE Transactions on Computers, 70(12):2098– 2111, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kwon:2021:REH

[KIY21] Taehyun Kwon, Muhammad Imran, and Joon-Sung Yang. Reliability enhanced heterogeneous phase change memory architecture for performance and energy efficiency. IEEE Transactions on Computers, 70(9):1388–1400, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2021:FEE

[KJC^{+}21] Yoonsuk Kang, Yong-Yeon Jo, Jaehyuk Cha, Wan D. Bae, Wonjun Lee, and Sang-Wook Kim. FORESEE: An effective and efficient framework for estimating the execution times of I/O traces on the SSD. IEEE Transactions on Computers, 70(12):2146– 2160, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2022:SMD

[KJK^{+}22] Taehwan Kim, Yunho Jang, Min-Gu Kang, Byong-Guk Park, Kyung-Jin Lee, and Jongsun Park. SOT-MRAM digital PIM architecture with extended parallelism in matrix multiplication. IEEE Transactions on Computers, 71(11): 2816–2828, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2024:AFR

[KJK24] Hyeonuk Kim, Youngbeom Jung, and Lee-Sup Kim. ADC-free ReRAM-based insitu accelerator for energyefficient binary neural networks. IEEE Transactions on Computers, 73(2):353– 365, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Karimi:2025:EDE

[KK25] Tayebeh Karimi and Arezoo Kamran. Energy-delay efficient segmented approximate adder with smart chaining. IEEE Transactions on Computers, 74(2):597–608, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Khodabandeloo:2022:PSA

[KKB^{+}22] Behnam Khodabandeloo, Ahmad Khonsari, Payman Behnam, Alireza Majidi, and Mohammad Hassan Hajiesmaili. Stereo: Assignment and scheduling in MPSoC under process variation by combining stochastic and decomposition approaches. IEEE Transactions on Computers, 71(11): 2940–2954, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2022:MBC

[KKH22] Duseok Kang, Donghyun Kang, and Soonhoi Ha. Multibank on-chip memory management techniques for CNN accelerators. IEEE Transactions on Computers, 71(5): 1181–1193, May 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2020:ATM

[KKKC20] Y. G. Kim, M. Kim, J. Kong, and S. W. Chung. An adaptive thermal management framework for heterogeneous multicore processors. IEEE Transactions on Computers, 69(6): 894–906, June 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2025:RCR

[KKL^{+}25] Yonghae Kim, Anurag Kar, Jaewon Lee, Jaekyu Lee, and Hyesoon Kim. RV-CURE: a RISC-V capability architecture for full memory safety. IEEE Transactions on Computers, 74(10):3291–3304, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2022:OGP

[KKRK22] Jaeyoung Kang, Behnam Khaleghi, Tajana Rosing, and Yeseong Kim. OpenHD:

A GPU-powered framework for hyperdimensional computing. IEEE Transactions on Computers, 71(11):2753– 2765, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2022:FNF

[KKS^{+}22] Myeonggu Kang, Hyeonuk Kim, Hyein Shin, Jaehyeong Sim, Kyeonghan Kim, and Lee-Sup Kim. S-FLASH: a NAND flash-based deep neural network accelerator exploiting bit-level sparsity. IEEE Transactions on Computers, 71(6):1291–1304, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2020:SSA

[KLC20] Y. G. Kim, Y. S. Lee, and S. W. Chung. Signal strengthaware adaptive offloading with local image preprocessing for energy efficient mobile devices. IEEE Transactions on Computers, 69(1):99–111, January 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2023:PSM

[KLKK23] Mincheol Kang, Wonyoung Lee, Jinkwon Kim, and Soontae Kim. PR-SSD: Maximizing partial read potential by exploiting compression and channel-level parallelism. IEEE Transactions on Computers, 72(3):772–785, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kong:2021:HPC

[KLL21] Liang Kong, Shuguo Li, and Ruirui Liu. High-performance constant-time discrete Gaussian sampling. IEEE Transactions on Computers, 70(7): 1019–1033, July 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2021:NBS

[KLP^{+}21] Bogil Kim, Sungjae Lee, Chanho Park, Hyeonjin Kim, and William J. Song. The Nebula Benchmark Suite: Implications of lightweight neural networks. IEEE Transactions on Computers, 70(11):1887– 1900, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Krishnankutty:2020:ISI

[KLR^{+}20] D. Krishnankutty, Z. Li, R. Robucci, N. Banerjee, and C. Patel. Instruction sequence identification and disassembly using power supply sidechannel analysis. IEEE Transactions on Computers, 69(11): 1639–1653, November 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kiningham:2023:GGN

[KLR23] Kevin Kiningham, Philip Levis, and Christopher Ré. GRIP: a graph neural network accelerator architecture. IEEE Transactions on Computers, 72(4):914–925, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kong:2025:SMM

[KLW^{+}25] Rui Kong, Yuanchun Li, Weijun Wang, Linghe Kong, and Yunxin Liu. Serving MoE models on resourceconstrained edge devices via dynamic expert swapping. IEEE Transactions on Computers, 74(8):2799–2811, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Karabulut:2025:TTH

[KMAA25] Emre Karabulut, Arsalan Ali Malik, Amro Awad, and Aydin Aysu. THEMIS: Time, heterogeneity, and energy minded scheduling for fair multi-tenant use in FPGAs. IEEE Transactions on Computers, 74(7):2515–2528, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2023:WCA

[KMH^{+}23] Sungkeun Kim, Farabi Mahmud, Jiayi Huang, Pritam Majumder, Chia-Che Tsai, Abdullah Muzahid, and Eun Jung Kim. WHISTLE: CPU abstractions for hardware and software memory safety invariants. IEEE Transactions on Computers, 72(3): 811–825, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kirkland:2022:USI

[KMVD22] Paul Kirkland, Davide Manna, Alex Vicente, and Gaetano Di Caterina. Unsupervised spiking instance segmentation on event data using STDP features. IEEE Transactions on Computers, 71(11):2728– 2739, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kundu:2020:PDS

[KNP^{+}20] S. Kundu, M. Nazemi, M. Pedram, K. M. Chugg, and P. A. Beerel. Pre-defined sparsity for low-complexity convolutional neural networks. IEEE Transactions on Computers, 69(7):1045–1058, July 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Koc:2020:AIM

[Koç20]  Çetin Kaya Koç.  Algo-

rithms for inversion mod p^{k}. IEEE Transactions on Computers, 69(6):907–913, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Khan:2023:DTS

[KOH^{+}23] Asif Ali Khan, Sebastien Ollivier, Fazal Hameed, Jeronimo Castrillon, and Alex K. Jones. DownShift: Tuning shift reduction with reliability for racetrack memories. IEEE Transactions on Computers, 72(9):2585– 2599, September 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Koo:2023:FEI

[KOT^{+}23] Gunjae Koo, Yunho Oh, Hung-Wei Tseng, Won Woo Ro, and Murali Annavaram. FLIXR: Embedding index into flash translation layer in SSDs. IEEE Transactions on Computers, 72(1):250–263, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2023:ALS

[KPD^{+}23] Ji-Hoon Kim, Yeo-Reum Park, Jaeyoung Do, Soo-Young Ji, and Joo-Young Kim. Accelerating large-scale graph-based nearest neighbor search on a computational storage platform. IEEE Transactions on Computers, 72(1): 278–290, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2025:SRF

[KPH^{+}25] Kevin Kim, Katherine Parry, David Harris, Cedar Turek, Alessandro Maiuolo, Rose Thompson, and James Stine. Shared recurrence floatingpoint divide/sqrt and integer divide/remainder with early termination. IEEE Transactions on Computers, 74(2): 740–748, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kim:2022:CCP

[KPL^{+}22] Namhyung Kim, Hanmin Park, Dongwoo Lee, Sungbum Kang, Jinho Lee, and Kiyoung Choi. ComPreEND: Computation pruning through predictive early negative detection for ReLU in a deep neural network accelerator. IEEE Transactions on Computers, 71(7): 1537–1550, July 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2024:TAG

[KPS^{+}24] Myeonggu Kang, Junyoung Park, Hyein Shin, Jaekang Shin, and Lee-Sup Kim. ToEx: Accelerating generation stage of transformer-based language models via token-adaptive early exit. IEEE Transactions on Computers, 73(9): 2248–2261, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kurth:2022:OSP

[KRB^{+}22] Andreas Kurth, Wolfgang Rönninger, Thomas Benz, Matheus Cavalcante, Fabian Schuiki, Florian Zaruba, and Luca Benini. An open-source platform for high-performance non-coherent on-chip communication. IEEE Transactions on Computers, 71(8):1794– 1809, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Koltuk:2024:UIH

[KS24] Furkan Koltuk and Ece Güran Schmidt. Uniformity and independence of H3 hash functions for Bloom filters. IEEE Transactions on Computers, 73(8): 1913–1923, August 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Khataei:2024:SSS

[KSB24] Alireza Khataei, Gaurav Singh, and Kia Bazargan. SimBU: Self-similarity-based hybrid binary-unary computing for nonlinear functions. IEEE Transactions on Computers, 73(9):2192– 2205, September 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2023:MFE

[KSKK23] Myeonggu Kang, Hyein Shin, Junkyum Kim, and Lee-Sup Kim. MGen: a framework for energy-efficient In-ReRAM acceleration of multitask BERT. IEEE Transactions on Computers, 72(11): 3140–3152, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kazemi:2022:FMB

[KSL^{+}22] Arman Kazemi, Mohammad Mehdi Sharifi, Ann Franchesca Laguna, Franz Müller, Xunzhao Yin, Thomas Kämpfe, Michael Niemier, and X. Sharon Hu. FeFET multi-bit contentaddressable memories for inmemory nearest neighbor search. IEEE Transactions on Computers, 71(10):2565– 2576, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kalali:2022:NPP

[KvL22] Ercan Kalali and Rene van Leuken. Near-precise parameter approximation for multiple multiplications on a single DSP block. IEEE Transactions on Computers, 71(9): 2036–2047, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Koliogeorgi:2022:PGD

[KXGS22] Konstantina Koliogeorgi, Sotirios Xydis, Georgi Gaydadjiev, and Dimitrios Soudris. GANDAFL: Dataflow acceleration for short read alignment on NGS data. IEEE Transactions on Computers, 71(11):3018– 3031, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Krishnan:2022:EMS

[KYS^{+}22] Gokul Krishnan, Li Yang, Jingbo Sun, Jubin Hazra, Xiaocong Du, Maximilian Liehr, Zheng Li, Karsten Beckmann, Rajiv V. Joshi, Nathaniel C. Cady, Deliang Fan, and Yu Cao. Exploring model stability of deep neural networks for reliable RRAM-based in-memory acceleration. IEEE Transactions on Computers, 71(11):2740– 2752, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kokkinis:2025:EPM

[KZS^{+}25] Argyris Kokkinis, Georgios Zervakis, Kostas Siozios, Mehdi Baradaran Tahoori, and Jörg Henkel. Enabling printed multilayer perceptrons realization via area-aware neural minimization. IEEE Transactions on Computers, 74(4):1461–1469, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kang:2025:RHC

[KZXR25] Jaeyoung Kang, Minxuan Zhou, Weihong Xu, and Tajana Rosing. RelHDx: Hyperdimensional computing for learning on graphs with Fe-FET acceleration. IEEE Transactions on Computers, 74(5):1730–1742, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2020:ESI

[LAKS20] Z. Liu, R. Azarderakhsh, H. Kim, and H. Seo. Efficient software implementation of ring-LWE encryption on IoT processors. IEEE Transactions on Computers, 69(10): 1424–1433, October 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lumpp:2021:TMS

[LAPB21] Francesco Lumpp, Stefano Aldegheri, Hiren D. Patel, and Nicola Bombieri. Task mapping and scheduling for OpenVX applications on heterogeneous multi/many-core architectures. IEEE Transactions on Computers, 70(8): 1148–1159, August 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2022:BMM

[LB22] Doowon Lee and Valeria Bertacco. Bypassing multicore memory bugs with coarse-grained reconfigurable logic. IEEE Transactions on Computers, 71(9):2191– 2204, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:RRF

[LBC24] YongGang Li, Yu Bao, and Yeh-Ching Chung. Randomize the running function when it is disclosed. IEEE Transactions on Computers, 73(6): 1516–1530, June 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lv:2024:SDL

[LCC^{+}24] Zhihan Lv, Dongliang Chen, Bin Cao, Houbing Song, and Haibin Lv. Secure deep learning in defense in deeplearning-as-a-service computing systems in digital twins. IEEE Transactions on Computers, 73(3):656–668, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2025:TEL

[LCGH25] Chuan Luo, Shenghua Cao, Shanyu Guo, and Chunming Hu. Towards effective local search for qubit mapping. IEEE Transactions on Computers, 74(6):1897–1910, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lien:2022:EER

[LCH22] Yi-Han Lien, Yi-Hua Chen, and Po-Chun Huang. Enabling efficient random data insertion/deletion on blockbased file systems. IEEE Transactions on Computers, 71(6):1479–1494, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2022:PGE

[LCHK22] Hyoeun Lee, Youngjoon Choi, Taeho Han, and Kanghee Kim. Probabilistically guaranteeing end-to-end latencies in autonomous vehicle computing systems. IEEE Transactions on Computers, 71(12): 3361–3374, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2021:VWF

[LCHL21] Yong-Gang Li, Yeh-Ching Chung, Kai Hwang, and Yue-Jin Li. Virtual Wall: Filtering rootkit attacks to protect Linux kernel functions. IEEE Transactions on Computers, 70(10):1640–1653, October 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:DPD

[LCJ^{+}24] Fuliang Li, Songlin Chen, Xingxin Jia, Chengxi Gao, Pengfei Wang, Xingwei Wang, and Jiannong Cao. Distributed program deployment for resource-aware programmable switches. IEEE Transactions on Computers, 73(5):1357–1370, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lim:2025:NEC

[LCJ^{+}25] Jaeil Lim, Jaewon Chung, Donghun Jeong, Daegeun Jee, and Euicheol Lim. A new ECC configuration method for DRAM system considering metadata. IEEE Transactions on Computers, 74(4): 1293–1305, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2020:SBC

[LCL^{+}20] R. Liu, X. Chen, D. Liu, Y. Ling, W. Wang, Y. Tan, C. Xiao, C. Yang, R. Zhang, and L. Liang. Separable binary convolutional neural network on embedded systems. IEEE Transactions on Computers, 69(10):1474–1486, October 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:SAA

[LCM25] Huize Li, Dan Chen, and Tulika Mitra. SADIMM: Accelerating sparse attention using DIMM-based near-memory processing. IEEE Transactions on Computers, 74(2): 542–554, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:PCG

[LCS^{+}25] Fuliang Li, Qin Chen, Jiaxing Shen, Xingwei Wang, and Jiannong Cao. Performance characteristics and guidelines of offloading middleboxes onto BlueField-2 DPU. IEEE Transactions on Computers, 74(2):609–622, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2021:ZDF

[LCX21] Guoxi Li, Wenzhi Chen, and Yang Xiang. Zweilous: A decoupled and flexible memory management framework. IEEE Transactions on Computers, 70(9):1350– 1362, September 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:DCS

[LCY^{+}25] Zerun Li, Xiaoming Chen, Yuxin Yang, Feng Min, Xiaoyu Zhang, and Yinhe Han. A data-centric softwarehardware co-designed architecture for large-scale graph processing. IEEE Transactions on Computers, 74(4): 1109–1122, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:ECB

[LCZ22] Yin Li, Xinyuan Cui, and Yu Zhang. An efficient CRT-based bit-parallel multiplier for special pentanomials. IEEE Transactions on Computers, 71(3):736–742, March 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:ECP

[LD22] Borui Li and Wei Dong. Edgecentric programming for IoT applications with automatic code partitioning. IEEE Transactions on Computers, 71(10):2408–2422, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2024:HOK

[LDF^{+}24] Junyi Liu, Aleksandar Dragojevi , Shane Fleming, Antonios Katsarakis, Dario Korolija, Igor Zablotchi, Ho-Cheung Ng, Anuj Kalia, and Miguel Castro. Honeycomb: Ordered key–value store acceleration on an FPGA-based SmartNIC. IEEE Transactions on Computers, 73(3): 857–871, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:HRM

[LDF^{+}25] Xuanli Liu, Zhenjiang Dong, Weibei Fan, Mengjie Lv, Xueli Sun, Jin Qi, and Sun-Yuan Hsieh. A highly reliable multiplexing scheme in hypercubestructured hierarchical networks. IEEE Transactions on Computers, 74(10):3462– 3475, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:ELS

[LDG^{+}22] Yuxuan Li, Xiaohui Duan, Lin Gan, Wubing Wan, Yuhu Chen, Kai Xu, Jinzhe Yang, Weiguo Liu, Wei Xue, Haohuan Fu, and Guangwen Yang. Enabling large-scale simulation of CAM on the Sunway TaihuLight supercomputer. IEEE Transactions on Computers, 71(4):824–837, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:HMB

[LDLK22] Cunlu Li, Dezun Dong, Xiangke Liao, and John Kim. Hybrid memory buffer microarchitecture for high-radix routers. IEEE Transactions on Computers, 71(11):2888– 2902, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2025:BGU

[LDT^{+}25] Yunzhen Luo, Yan Ding, Zhuo Tang, Keqin Li, Kenli Li, and Chubo Liu. BEAST-GNN: a united bit sparsity-aware accelerator for graph neural networks. IEEE Transactions on Computers, 74(7):2402–2416, July 2025. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2025:BAF

[LDW^{+}25] Ping Luo, Xiaoge Deng, Ziqing Wen, Tao Sun, and Dongsheng Li. BHerd: Accelerating federated learning by selecting beneficial herd of local gradients. IEEE Transactions on Computers, 74(9): 2977–2990, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2023:TTA

[LDZ^{+}23] Jie Li, Yuhui Deng, Yi Zhou, Zhen Zhang, Geyong Min, and Xiao Qin. Towards thermalaware workload distribution in cloud data centers based on failure models. IEEE Transactions on Computers, 72(2): 586–599, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:WTF

[LFGD25] Borui Li, Hongchang Fan, Yi Gao, and Wei Dong. Wa-WoT: Towards flexible and efficient Web of Things services via WebAssembly on resourceconstrained IoT devices. IEEE Transactions on Computers, 74(3):1094–1108, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lurbe:2022:DDL

[LFP^{+}22] Manel Lurbe, Josué Feliu, Salvador Petit, Maria E. Gómez, and Julio Sahuquillo. DeepP: Deep learning multi-program prefetch configuration for the IBM POWER 8. IEEE Transactions on Computers, 71(10): 2646–2658, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liao:2021:APM

[LFW21] Zhiheng Liao, Jingyan Fu, and Jinhui Wang. Ameliorate performance of memristor-based ANNs in edge computing. IEEE Transactions on Computers, 70(8):1299–1310, August 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2021:EDN

[LFX^{+}21] J. Lu, C. Fang, M. Xu, J. Lin, and Z. Wang. Evaluations on deep neural networks training using posit number system. IEEE Transactions on Computers, 70(2):174–187, February 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:DSA

[LG22] Congmiao Li and Jean-Luc Gaudiot. Detecting spectre attacks using hardware performance counters. IEEE Transactions on Computers, 71(6): 1320–1331, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:GRD

[LGC^{+}23] Liang Liu, Yanan Guo, Yueqiang Cheng, Youtao Zhang, and Jun Yang. Generating robust DNN with resistance to bit-flip based adversarial weight attack. IEEE Transactions on Computers, 72(2): 401–413, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:MAU

[LGL^{+}24] Jing Li, Song Guo, Weifa Liang, Jianping Wang, Quan Chen, Wenchao Xu, Kang Wei, and Xiaohua Jia. Mobilityaware utility maximization in digital twin-enabled serverless edge computing. IEEE Transactions on Computers, 73(7): 1837–1851, July 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lou:2022:OHT

[LGW^{+}22] Wenqi Lou, Lei Gong, Chao Wang, Zidong Du, and Xuehai Zhou. OctCNN: a high throughput FPGA accelerator for CNNs using Octave convolution algorithm. IEEE Transactions on Computers, 71(8): 1847–1859, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2022:ECT

[LGX^{+}22] Chunchi Liu, Hechuan Guo, Minghui Xu, Shengling Wang, Dongxiao Yu, Jiguo Yu, and Xiuzhen Cheng. Extending on-chain trust to off-chain trustworthy blockchain data collection using trusted execution environment (TEE). IEEE Transactions on Computers, 71(12):3268–3280, December 2022. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:CNA

[LGZ^{+}25] Yi Liu, Song Guo, Jie Zhang, Zicong Hong, Yufeng Zhan, and Qihua Zhou. Collaborative neural architecture search for personalized federated learning. IEEE Transactions on Computers, 74(1): 250–262, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:DCP

[LHA^{+}25] Taotao Li, Huawei Huang, Parhat Abla, Zhihong Deng, Qinglin Yang, Anke Xie, Debiao He, and Zibin Zheng. DataFly: a confidentialitypreserving data migration across heterogeneous blockchains. IEEE Transactions on Computers, 74(6):1814–1828, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:PPD

[LHFW24] Luyi Li, Jiayi Huang, Lang Feng, and Zhongfeng Wang. Prefender: a prefetching defender against cache side channel attacks as a pretender. IEEE Transactions on Computers, 73(6):1457–1471, June 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2022:FSM

[LHK^{+}22] Sunjung Lee, Seunghwan Hwang, Michael Jaemin Kim, Jaewan Choi, and Jung Ho Ahn. Future scaling of memory hierarchy for tensor cores and eliminating redundant shared memory traffic using inter-warp multicasting. IEEE Transactions on Computers, 71(12):3115–3126, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2021:CSC

[LHL^{+}21] Dongsheng Li, Zhiyao Hu, Zhiquan Lai, Yiming Zhang, and Kai Lu. Coordinative scheduling of computation and communication in dataparallel systems. IEEE Transactions on Computers, 70(12): 2182–2197, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:EER

[LHL^{+}23] Qiang Liu, Yuhui Hao, Weizhuang Liu, Bo Yu, Yiming Gan, Jie Tang, Shao-Shan Liu, and Yuhao Zhu. An energy efficient and runtime reconfigurable accelerator for robotic localization. IEEE Transactions on Computers, 72(7):1943–1957, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2022:BCT

[LHN^{+}22] Dongxiao Liu, Cheng Huang, Jianbing Ni, Xiaodong Lin, and Xuemin Sherman Shen. Blockchain-cloud transparent data marketing: Consortium management and fairness. IEEE Transactions on Computers, 71(12):3322– 3335, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:ESS

[LHR^{+}22] Jingwei Li, Suyu Huang, Yanjing Ren, Zuoru Yang, Patrick P. C. Lee, Xiaosong Zhang, and Yao Hao. Enabling secure and space-efficient metadata management in encrypted deduplication. IEEE Transactions on Computers, 71(4): 959–970, April 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:BBD

[LHR^{+}23] Yizhi Liu, Xiaohan Hao, Wei Ren, Ruoting Xiong, Tianqing Zhu, Kim-Kwang Raymond Choo, and Geyong Min. A blockchain-based decentralized, fair and authenticated information sharing scheme in zero trust Internet-of-Things. IEEE Transactions on Computers, 72(2):501–512, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:FPF

[LHS^{+}25] Junchao Li, Runsheng Hou, Guangyong Shang, Huanle Zhang, Xiuzhen Cheng, and Runyu Pan. FVM: Practical feather-weight virtualization on commodity microcontrollers. IEEE Transactions on Computers, 74(7): 2389–2401, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2021:NMN

[LHW^{+}21] Limei Lin, Yanze Huang, Dajin Wang, Sun-Yuan Hsieh, and Li Xu. A novel measurement for network reliability. IEEE Transactions on Computers, 70(10):1719–1731, October 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:UAM

[LHW^{+}25] Ji Li, Qiang He, Xingwei Wang, Ammar Hawbani, Keping Yu, Yuanguo Bi, and Liang Zhao. UAV-assisted microservice mobile edge computing architecture: Addressing post-disaster emergency medical rescue. IEEE Transactions on Computers, 74(8): 2635–2648, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:HPM

[LHX^{+}25] Chao Li, Xuchu Huang, Zhicheng Xu, Bo Wen, Ruibin Mao, Min Zhou, Thomas Kämpfe, Kai Ni, Can Li, Xunzhao Yin, and Cheng Zhuo. High-performance in-memory Bayesian inference with multibit ferroelectric FET. IEEE Transactions on Computers, 74(9):2923–2935, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2022:BAM

[LHXH22] Limei Lin, Yanze Huang, Li Xu, and Sun-Yuan Hsieh. Better adaptive malicious users detection algorithm in human contact networks. IEEE Transactions on Computers, 71(11):2968–2981, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:CCD

[LHXH24] Wenjie Li, Aokun Hu, Ningyi Xu, and Guanghui He. CoDA: a co-design framework for versatile and efficient attention accelerators. IEEE Transactions on Computers, 73(8):

1924–1938, August 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2021:LFD

[LHY^{+}21] Jianghua Liu, Jingyu Hou, Wenjie Yang, Yang Xiang, Wanlei Zhou, Wei Wu, and Xinyi Huang. Leakagefree dissemination of authenticated tree-structured data with multi-party control. IEEE Transactions on Computers, 70(7):1120–1131, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2024:HPT

[LHZ^{+}24] Xiao-Yang Liu, Hao Hong, Zeliang Zhang, Weiqin Tong, Jean Kossaifi, Xiaodong Wang, and Anwar Walid. High-performance tensor-train primitives using GPU tensor cores. IEEE Transactions on Computers, 73(11):2634– 2648, November 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:PNA

[LHZR25] Yuchen Liu, Ligang He, Zhigao Zhang, and Shenyuan Ren. PFed-NS: an adaptive personalized federated learning scheme through neural network segmentation. IEEE Transactions on Computers, 74(6):1936–1948, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:FFI

[LJH^{+}25] Jingjin Li, Weixiong Jiang, Yuting He, Qingyu Yang, Anqi Gao, Yajun Ha, Ender Özcan, Ruibin Bai, Tianxiang Cui, and Heng Yu. FiDRL: Flexible invocation-based deep reinforcement learning for DVFS scheduling in embedded systems. IEEE Transactions on Computers, 74(1):71–85, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2021:RTS

[LJY21] D. Lee, H. Jung, and H. Yang.

Real-time schedulability analysis and enhancement of transiently powered processors with NVMs. IEEE Transactions on Computers, 70(3): 372–383, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2024:CBL

[LJY^{+}24] Haikun Liu, Xiaozhong Jin, Chencheng Ye, Xiaofei Liao, Hai Jin, and Yu Zhang. I/O causality based in-line data deduplication for non-volatile memory enabled storage systems. IEEE Transactions on Computers, 73(5):1327–1340, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:RHV

[LJZ^{+}25] Jiguo Li, Licheng Ji, Yicheng Zhang, Yang Lu, and Jianting Ning. Response-hiding and volume-hiding verifiable searchable encryption with conjunctive keyword search. IEEE Transactions on Computers, 74(2):455–467, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2021:IBP

[LKK^{+}21] H. Lee, H. Kim, C. Kim, H. Han, and E. Seo. Idempotencebased preemptive GPU kernel scheduling for embedded systems. IEEE Transactions on Computers, 70(3): 332–346, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Longofono:2021:CTR

[LKMJ21] Stephen Longofono, Donald Kline, Rami Melhem, and Alex K. Jones. A CAS-TLE with TOWERs for reliable, secure phase-change memory. IEEE Transactions on Computers, 70(9):1311– 1324, September 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2021:ACG

[LL21] Xianfeng Li and Gengchao Li. An adaptive CPU–GPU governing framework for mobile games on big.LITTLE architectures. IEEE Transactions on Computers, 70(9): 1472–1483, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2022:PMF

[LL22] Jaewoo Lee and Jinkyu Lee.

MC-FLEX: Flexible mixedcriticality real-time scheduling by task-level mode switch. IEEE Transactions on Computers, 71(8):1889–1902, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2023:TPB

[LL23] Jin-Yi Lin and Shu-Yen Lin.

Temperature-prediction based rate-adjusted time and space mapping algorithm for 3D CNN accelerator systems. IEEE Transactions on Computers, 72(10):2767–2780, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2024:CCD

[LLC^{+}24] Zhiyong Lin, Hai Liu, Xiaowen Chu, Yiu-Wing Leung, and Ivan Stojmenovic. Constructing connected-dominating-set with maximum lifetime in cognitive radio networks. IEEE Transactions on Computers, 73(4):1165–1179, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2023:CMC

[LLCJ23] Wanling Lin, Xiao-Yan Li, Jou-Ming Chang, and Xiaohua Jia. Constructing multiple CISTs on BCube-based data center networks in the occurrence of switch failures. IEEE Transactions on Computers, 72(7):1971–1984, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:FJS

[LLD^{+}25] Yulong Li, Wenxin Li, Yuxuan Du, Yinan Yao, Song Zhang, Linxuan Zhong, and Keqiu Li. Flexible job scheduling with spatial-temporal compatibility for in-network aggregation. IEEE Transactions on Computers, 74(4):1322–1333, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2023:DSE

[LLFT23] He Li, Jiawei Liang, Hongxiang Fan, and Yongming Tang. Design space exploration for efficient quantum most-significant digit-first arithmetic. IEEE Transactions on Computers, 72(6): 1822–1829, June 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2023:CCO

[LLJ^{+}23] Xin Li, Zhi Li, Yaqi Ju, Xiaofei Zhang, Rongyao Wang, and Wei Zhou. COP: a combinational optimization power budgeting method for manycore systems in dark silicon. IEEE Transactions on Computers, 72(5):1356–1370, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2023:SCS

[LLK^{+}23] Xiangzhong Luo, Di Liu, Hao Kong, Shuo Huai, Hui Chen, and Weichen Liu. SurgeNAS: a comprehensive surgery on hardware-aware differentiable neural architecture search. IEEE Transactions on Computers, 72(4):1081–1094, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2020:OLM

[LLL^{+}20] H. Liu, R. Liu, X. Liao, H. Jin, B. He, and Y. Zhang. Objectlevel memory allocation and migration in hybrid memory systems. IEEE Transactions on Computers, 69(9): 1401–1413, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2023:LIC

[LLL^{+}23] Yida Li, Huizhang Luo, Fenfang Li, Junqi Wang, and Kenli Li. LAMP: Improving compression ratio for AMR applications via level associated mapping-based preconditioning. IEEE Transactions on Computers, 72(12):3370– 3382, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:AHH

[LLL^{+}25a] Weijie Liu, Kai Lu, Zhiquan Lai, Shengwei Li, Keshi Ge, Dongsheng Li, and Xicheng Lu. AutoPipe-H: a heterogeneity-aware dataparalleled pipeline approach on commodity GPU servers. IEEE Transactions on Computers, 74(4):1196–1209, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lyu:2025:EMB

[LLL25b] Fei Lyu, Yuanyong Luo, and Weiqiang Liu. An efficient methodology for binary logarithmic computations of floating-point numbers with normalized output within one ulp of accuracy. IEEE Transactions on Computers, 74(5): 1800–1813, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:ALR

[LLM^{+}25] Tong Li, Wei Liu, Xinyu Ma, Shuaipeng Zhu, Jingkun Cao, Duling Xu, Zhaoqi Yang, Senzhen Liu, Taotao Zhang, Yinfeng Zhu, Bo Wu, Kezhi Wang, and Ke Xu. Accelerating loss recovery for content delivery network. IEEE Transactions on Computers, 74(7):2223–2237, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2025:ALS

[LLR25] Hao Lu, Jian Liu, and Kui Ren. Aurora: Leaderless statemachine replication with high throughput. IEEE Transactions on Computers, 74(5): 1690–1701, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2022:BEP

[LLS^{+}22] Jizhao Liu, Jing Lian, Julien Clinton Sprott, Qidong Liu, and Yide Ma. The butterfly effect in primary visual cortex. IEEE Transactions on Computers, 71(11):2803– 2815, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2023:MDB

[LLS^{+}23] Hyokeun Lee, Seungyong Lee, Byeongki Song, Moonsoo Kim, Seokbo Shim, Hyuk-Jae Lee, and Hyun Kim. An in-module disturbance barrier for mitigating write disturbance in phase-change memory. IEEE Transactions on Computers, 72(4):1150–1162, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:QSV

[LLS^{+}24] Jiaye Li, Yangding Li, Jiagang Song, Jian Zhang, and Shichao Zhang. Quantum support vector machine for classifying noisy data. IEEE Transactions on Computers, 73(9): 2233–2247, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lee:2025:NAC

[LLS^{+}25] Seungyong Lee, Sanghyun Lee, Minseok Seo, Chunmyung Park, Woojae Shin, Hyuk-Jae Lee, and Hyun Kim. NPC: A non-conflicting processing-in-memory controller in DDR memory systems. IEEE Transactions on Computers, 74(3):1025– 1039, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2023:BEB

[LLT^{+}23] Shipeng Li, Jingwei Li, Yuxing Tang, Xiapu Luo, Zheyuan He, Zihao Li, Xi Cheng, Yang Bai, Ting Chen, Yuzhe Tang, Zhe Liu, and Xiaosong Zhang. BlockExplorer: Exploring blockchain big data via parallel processing. IEEE Transactions on Computers, 72(8): 2377–2389, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:OCS

[LLW^{+}24] Zhenzheng Li, Jiong Lou, Jianfei Wu, Jianxiong Guo, Zhiqing Tang, Ping Shen, Weijia Jia, and Wei Zhao. Online container scheduling with fast function startup and low memory cost in edge computing. IEEE Transactions on Computers, 73(12):2747–2760, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2023:FPK

[LLWZ23] Kaizhan Lin, Jianming Lin, Weize Wang, and Chang-An Zhao. Faster public-key compression of SIDH with less memory. IEEE Transactions on Computers, 72(9):2668– 2676, September 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2024:QPA

[LLX^{+}24a] Wei Liang, Yuhui Li, Jianlong Xu, Zheng Qin, Dafang Zhang, and Kuan-Ching Li. QoS prediction and adversarial attack protection for distributed services under DLaaS. IEEE Transactions on Computers, 73(3): 669–682, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2024:RCH

[LLX^{+}24b] Yu Liu, Qi Luo, Mengbai Xiao, Dongxiao Yu, Huashan Chen, and Xiuzhen Cheng. Reordering and compression for hypergraph processing. IEEE Transactions on Computers, 73(6):1486–1499, June 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2022:FBV

[LLY22] Yandong Luo, Yuan-Chun Luo, and Shimeng Yu. A ferroelectric-based Volatile/Non-Volatile dual-mode buffer memory for deep neural network accelerators. IEEE Transactions on Computers, 71(9):2088–2101, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2025:CCA

[LLZ^{+}25] Weiwei Lin, Jinhui Lin, Haotong Zhang, Wentai Wu, Weizheng Wu, Zhetao Li, and Keqin Li. Cacomp: a cloud-assisted collaborative deep learning compiler framework for DNN tasks on edge. IEEE Transactions on Computers, 74(8):2663–2674, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lyu:2021:SCT

[LM21] Yangdi Lyu and Prabhat Mishra. Scalable concolic testing of RTL models. IEEE Transactions on Computers, 70(7):979–991, July 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2021:DSI

[LMDC21] He Li, Ian McInerney, James J.

Davis, and George A. Constantinides. Digit stability inference for iterative methods using redundant number representation. IEEE Transactions on Computers, 70(7): 1074–1080, July 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:GGP

[LMH^{+}25] Bin Liu, Yongyao Ma, Zijian Hu, Zeyu Ji, Zhenli He, and Keqin Li. GroPipe: a grouped pipeline hybrid parallel method for accelerating DCNNs training. IEEE Transactions on Computers, 74(7): 2487–2500, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:DDV

[LMM^{+}22] Yingying Li, Jianfeng Ma, Yinbin Miao, Huizhong Li, Qiang Yan, Yue Wang, Ximeng Liu, and Kim-Kwang Raymond Choo. DVREI: Dynamic verifiable retrieval over encrypted images. IEEE Transactions on Computers, 71(8):1755–1769, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2023:PPB

[LMM^{+}23] Yanrong Liang, Jianfeng Ma, Yinbin Miao, Da Kuang, Xiangdong Meng, and Robert H. Deng. Privacy-preserving Bloom filter-based keyword search over large encrypted cloud data. IEEE Transactions on Computers, 72(11): 3086–3098, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:HHA

[LMW^{+}24] Hongyi Li, Songchen Ma, Taoyi Wang, Weihao Zhang, Guanrui Wang, Chenhang Song, Huanyu Qu, Junfeng Lin, Cheng Ma, Jing Pei, and Rong Zhao. HASP: Hierarchical asynchronous parallelism for multi-NN tasks. IEEE Transactions on Computers, 73(2):366–379, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:SHP

[LMZ^{+}25] Song Liu, Jie Ma, Zengyuan Zhang, Xinhe Wan, Bo Zhao, and Weiguo Wu. Scalpel: High performance contention-aware task co-scheduling for shared cache hierarchy. IEEE Transactions on Computers, 74(2): 678–690, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Long:2025:HLP

[LOM^{+}25] Jiangshan Long, Changhai Ou, Yajun Ma, Yifan Fan, Hua Chen, and Shihui Zheng. How to launch a powerful side-channel collision attack? IEEE Transactions on Computers, 74(3):835–847, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Louri:2020:SJ

[Lou20] A. Louri. State of the journal.

IEEE Transactions on Computers, 69(4):466–467, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lima:2021:PHD

[LPC^{+}21] F. D. S. Lima, F. L. F. Pereira, I. C. Chaves, J. C. Machado, and J. P. P. Gomes. Predicting the health degree of hard disk drives with asymmetric and ordinal deep neural models. IEEE Transactions on Computers, 70(2):188–198, February 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2021:RAE

[LPD^{+}21] Y. Liang, R. Pan, Y. Du, C. Fu, L. Shi, T.-W. Kuo, and C. J. Xue. Readahead efficiency on mobile devices: Observation, characterization, and optimization. IEEE Transactions on Computers, 70(1):99–110, January 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2020:XSI

[LPW20] Z. Li, B. Peng, and C. Weng. XeFlow: Streamlining interprocessor pipeline execution for the discrete CPU–GPU platform. IEEE Transactions on Computers, 69(6):819–831, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

LeCompte:2022:PSM

[LPYT22] Travis LeCompte, Lu Peng, Xu Yuan, and Nian-Feng Tzeng. Protecting synchronization mechanisms of parallel big data kernels via logging. IEEE Transactions on Computers, 71(9):2156– 2162, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2022:DSA

[LQC^{+}22] Liu Liu, Zheng Qu, Zhaodong Chen, Fengbin Tu, Yufei Ding, and Yuan Xie. Dynamic sparse attention for scalable transformer acceleration. IEEE Transactions on Computers, 71(12):3165–3178, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:DAF

[LQL^{+}24] Chunpei Li, Wangjie Qiu, Xianxian Li, Chen Liu, and Zhiming Zheng. A dynamic adaptive framework for practical Byzantine fault tolerance consensus protocol in the Internet of Things. IEEE Transactions on Computers, 73(7): 1669–1682, July 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2024:SCH

[LQM^{+}24] Junfeng Lin, Huanyu Qu, Songchen Ma, Xinglong Ji, Hongyi Li, Xiaochuan Li, Chenhang Song, and Weihao Zhang. SongC: A compiler for hybrid near-memory and in-memory many-core architecture. IEEE Transactions on Computers, 73(10):2420– 2433, October 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2021:PSS

[LQN^{+}21] Jinwen Liang, Zheng Qin, Jianbing Ni, Xiaodong Lin, and Xuemin Shen. Practical and secure SVM classification for cloud-based remote clinical decision services. IEEE Transactions on Computers, 70(10): 1612–1625, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2020:BBA

[LQY^{+}20] Q. Liu, S. Qin, B. Yu, J. Tang, and S. Liu. π-BA: Bundle adjustment hardware accelerator based on distribution of 3D-point observations. IEEE Transactions on Computers, 69(7):1083–1095, July 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2022:IBC

[LR22] Jian Liu and Kui Ren. Improving blockchains with client-assistance. IEEE Transactions on Computers, 71(5): 1230–1236, May 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Libano:2023:EED

[LRB23] Fabiano Libano, Paolo Rech, and John Brunhaver. Efficient error detection for matrix multiplication with systolic arrays on FPGAs. IEEE Transactions on Computers, 72(8): 2390–2403, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2022:SNR

[LRL22] Shanshan Liu, Pedro Reviriego, and Fabrizio Lombardi. Selective neuron re-computation (SNRC) for error-tolerant neural networks. IEEE Transactions on Computers, 71(3):684– 695, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lopez-Randulfe:2022:TCS

[LRRK^{+}22] Javier López-Randulfe, Nico Reeb, Negin Karimi, Chen Liu, Hector A. Gonzalez, Robin Dietrich, Bernhard Vogginger, Christian Mayr, and Alois Knoll. Timecoded spiking Fourier transform in neuromorphic hardware. IEEE Transactions on Computers, 71(11):2792– 2802, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2020:EAE

[LSCX20] Q. Li, L. Shi, Y. Cui, and C. J.

Xue. Exploiting asymmetric errors for LDPC decoding optimization on 3D NAND flash memory. IEEE Transactions on Computers, 69(4):475–488, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2025:RRS

[LSH25] Jianwen Luo, Yuhao Shu, and Yajun Ha. RSQC: Recursive sparse QUBO construction for quantum annealing machines. IEEE Transactions on Computers, 74(6):2114–2128, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:FOM

[LSL^{+}25] Wei Li, Zicheng Shen, Xiulong Liu, Chuntao Ding, and Jiaxing Shen. Fed-OGD: Mitigating straggler effects in federated learning via orthogonal gradient descent. IEEE Transactions on Computers, 74(9): 3018–3031, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:EQS

[LSS25] Wenjie Liu, Bingmei Su, and Feiyang Sun. Efficient quantum secure vector dominance and its applications in computational geometry. IEEE Transactions on Computers, 74(6):2129–2143, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2023:TAF

[LSU^{+}23] Ching-Chi Lin, Junjie Shi, Niklas Ueter, Mario Günzel, Jan Reineke, and Jian-Jia Chen. Type-aware federated scheduling for typed DAG tasks on heterogeneous multicore platforms. IEEE Transactions on Computers, 72(5): 1286–1300, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:CHR

[LSW^{+}23] Zixuan Liu, Xiaoyu Song, Zhuowei Wang, Yan Wang, and Jian-Tao Zhou. Constructing high radix quotient digit selection tables for SRT division and square root. IEEE Transactions on Computers, 72(7):2111–2119, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:RBD

[LSW25] Cong Li, Qingni Shen, and Zhonghai Wu. Redactable blockchain from decentralized chameleon hash functions, revisited. IEEE Transactions on Computers, 74(6):1911–1920, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lv:2021:ERA

[LSXZ21] M. Lv, H. Sun, J. Xin, and N. Zheng. Efficient repair analysis algorithm exploration for memory with redundancy and in-memory ECC. IEEE Transactions on Computers, 70(5):775–788, May 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:GSB

[LT25] Chengqing Li and Kai Tan.

The graph structure of baker’s maps implemented on a computer. IEEE Transactions on Computers, 74(5):1524–1537, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:GSG

[LTFL22] Chengqing Li, Kai Tan, Bingbing Feng, and Jinhu Lü. The graph structure of the generalized discrete Arnold’s cat map.

IEEE Transactions on Computers, 71(2):364–377, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lloret-Talavera:2022:EHE

[LTJS^{+}22] Guillermo Lloret-Talavera, Marc Jorda, Harald Servat, Fabian Boemer, Chetan Chauhan, Shigeki Tomishima, Nilesh N. Shah, and Antonio J. Peña. Enabling homomorphically encrypted inference for large DNN models. IEEE Transactions on Computers, 71(5):1145–1155, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:DOO

[LTL^{+}25] Chuang Li, Changyao Tan, Gang Liu, Yanhua Wen, Yan Wang, and Kenli Li. DC-ORAM: an ORAM scheme based on dynamic compression of data blocks and position map. IEEE Transactions on Computers, 74(5):1495–1509, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lu:2021:GEI

[Lu21] Zhonghai Lu. Guest editorial: IEEE TC Special Issue On Communications for Many-core Processors and Accelerators. IEEE Transactions on Computers, 70(6):817–818, June 2021. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:TLS

[LV23] Qingyue Liu and Peter Varman. Telepathy: a lightweight silent data access protocol for NVRAM+RDMA enabled distributed storage. IEEE Transactions on Computers, 72(3):839–852, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2022:RHD

[LW22] Dave Y.-W. Lin and Charles H.-P. Wen. Rad-hard designs by automated latchingdelay assignment and timeborrowable D-flip-flop. IEEE Transactions on Computers, 71(5):1008–1020, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2022:MOM

[LWC^{+}22] Xin Liang, Ben Whitney, Jieyang Chen, Lipeng Wan, Qing Liu, Dingwen Tao, James Kress, David Pugmire, Matthew Wolf, Norbert Podhorszki, and Scott Klasky. MGARD+: Optimizing multilevel methods for error-bounded scientific data reduction. IEEE Transactions on Computers, 71(7): 1522–1536, July 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lunglmayr:2020:DAE

[LWH20] M. Lunglmayr, D. Wiesinger, and W. Haselmayr. Design and analysis of efficient maximum/minimum circuits for stochastic computing. IEEE Transactions on Computers, 69(3):402–409, March 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:NND

[LWH^{+}24] Shiyu Li, Yitu Wang, Edward Hanson, Andrew Chang, Yang Seok Ki, Hai Li, and Yiran Chen. NDRec: a near-data processing system for training large-scale recommendation models. IEEE Transactions on Computers, 73(5):1248–1261, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2021:EHT

[LWL^{+}21] Shengwen Liang, Ying Wang, Cheng Liu, Lei He, Huawei Li, Dawen Xu, and Xiaowei Li. EnGN: A high-throughput and energy-efficient accelerator for large graph neural networks. IEEE Transactions on Computers, 70(9): 1511–1525, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2022:RLB

[LWL^{+}22] Yusen Li, Xiwei Wang, Haoyuan Liu, Lingjun Pu, Shanjiang Tang, Gang Wang, and Xiaoguang Liu. Reinforcement learning-based resource partitioning for improving responsiveness in cloud gaming. IEEE Transactions on Computers, 71(5):1049–1062, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2023:LFP

[LWL^{+}23] Wen Li, Ying Wang, Cheng Liu, Yintao He, Lian Liu, Huawei Li, and Xiaowei Li. On-line fault protection for ReRAM-based neural networks. IEEE Transactions on Computers, 72(2):423– 437, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liao:2024:DDU

[LWL^{+}24] Yunkun Liao, Jingya Wu, Wenyan Lu, Xiaowei Li, and Guihai Yan. DPU-Direct: Unleashing remote accelerators via enhanced RDMA for disaggregated datacenters. IEEE Transactions on Computers, 73(8):2081–2095, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:EHP

[LWL^{+}25] Hongmin Li, Si Wu, Zhipeng Li, Qianli Wang, Yongkun Li, and Yinlong Xu. Enabling high performance and resource utilization in clustered cache via hotness identification, data copying, and instance merging. IEEE Transactions on Computers, 74(2): 371–385, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lao:2022:BCS

[LWNC22] Bin Lao, Yi Wu, Ge Nong, and Wai Hong Chan. Building and checking suffix array simultaneously by induced sorting method. IEEE Transactions on Computers, 71(4):756–765, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lo:2024:IAC

[LWW^{+}24] Yun-Chen Lo, Jun-Shen Wu, Chia-Chun Wang, Yu-Chih Tsai, Chih-Chen Yeh, Wen-Chien Ting, and Ren-Shuo Liu. ISSA: Architecting CNN accelerators using input-skippable, set-associative computing-in-memory. IEEE Transactions on Computers, 73(9):2136– 2149, September 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:HIC

[LWW^{+}25] Qianhui Liu, Jiadong Wang, Yang Wang, Xin Yang, Gang Pan, and Haizhou Li. Humaninspired computing for robust and efficient audio-visual speech recognition. IEEE Transactions on Computers, 74(9):2950–2961, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:LHS

[LWX^{+}25] Zijun Li, Chenyang Wu, Chuhao Xu, Quan Chen, Shuo Quan, Bin Zha, Qiang Wang, Weidong Han, Jie Wu, and Minyi Guo. Lightweight and holistic-scalable serverless secure container runtime for high-density deployment and high-concurrency startup. IEEE Transactions on Computers, 74(8):2621–2634, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lv:2023:FDD

[LWYJ23] Xianwei Lv, Qianqian Wang, Chen Yu, and Hai Jin. A feedback-driven DNN inference acceleration system for edge-assisted video analytics. IEEE Transactions on Computers, 72(10):2902–2912, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2025:OMA

[LWZ^{+}25] Na Lin, Zhijiang Wang, Liang Zhao, Ammar Hawbani, Zhi Liu, and Mohsen Guizani. Optimizing Multi-AAV cooperative tracking for real-time applications in network-challenged environments. IEEE Transactions on Computers, 74(7):2461–2472, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2025:DDP

[LWZZ25] Chenhong Luo, Yong Wang, Yanjun Zhang, and Leo Yu Zhang. Distributed differentially private matrix factorization for implicit data via secure aggregation. IEEE Transactions on Computers, 74(2):705–716, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:DGP

[LXL^{+}25] Zhetao Li, Yong Xiao, Haolin Liu, Xiaofei Liao, Ye Yuan, and Junzhao Du. Dynamic graph publication with differential privacy guarantees for decentralized applications. IEEE Transactions on Computers, 74(5):1771–1785, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:AAA

[LXW^{+}23] Juan Liu, Guoqing Xiao, Fan Wu, Xiangke Liao, and Kenli Li. AAPP: an accelerative and adaptive path planner for robots on GPU. IEEE Transactions on Computers, 72(8): 2336–2349, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2020:ADN

[LY20] Y. Luo and S. Yu. Accelerating deep neural network in-situ training with nonvolatile and volatile memory based hybrid precision synapses. IEEE Transactions on Computers, 69(8): 1113–1127, August 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2021:AAC

[LY21] Yandong Luo and Shimeng Yu. AILC: Accelerate onchip incremental learning with compute-in-memory technology. IEEE Transactions on Computers, 70(8):1225– 1238, August 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liang:2022:PMM

[LYC22] Yuhong Liang, Ming-Chang Yang, and Shuo-Han Chen. MAGIC: Making IMR-based HDD perform like CMRbased HDD. IEEE Transactions on Computers, 71(3): 643–657, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2023:ETM

[LYC^{+}23] Qi Luo, Dongxiao Yu, Xiuzhen Cheng, Hao Sheng, and Weifeng Lyu. Exploring truss maintenance in fully dynamic graphs: a mixed structurebased approach. IEEE Transactions on Computers, 72(3): 707–718, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2022:RTT

[LYF^{+}22] Shengzhong Liu, Shuochao Yao, Xinzhe Fu, Huajie Shao, Rohan Tabish, Simon Yu, Ayoosh Bansal, Heechul Yun, Lui Sha, and Tarek Abdelzaher. Real-time task scheduling for machine perception in intelligent cyber-physical systems. IEEE Transactions on Computers, 71(8):1770–1783, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lin:2024:DTO

[LYGC24] Hongcai Lin, Lei Yang, Hao Guo, and Jiannong Cao. Decentralized task offloading in edge computing: An offlineto-online reinforcement learning approach. IEEE Transactions on Computers, 73(6): 1603–1615, June 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2024:HHP

[LYH^{+}24] Bihui Liu, Zhenyu Ye, Qiao Hu, Yupeng Hu, Yuchong Hu, Yang Xu, and Keqin Li. HPDK: a hybrid PM-DRAM key–value store for high I/O throughput. IEEE Transactions on Computers, 73(6): 1575–1587, June 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:LAS

[LYW^{+}23] Jianghua Liu, Jian Yang, Wei Wu, Xinyi Huang, and Yang Xiang. Lightweight authentication scheme for data dissemination in cloud-assisted healthcare IoT. IEEE Transactions on Computers, 72(5): 1384–1395, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:DGD

[LYW^{+}25] Lian Liu, Jinxin Yu, Mengdi Wang, Xiaowei Li, Yinhe Han, and Ying Wang. DNA: a general dynamic neural network accelerator. IEEE Transactions on Computers, 74(9): 3210–3222, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2021:PLS

[LZC^{+}21] X. Li, M. Zhang, K. Chen, Y. Wu, X. Qian, and W. Zheng. 3-D partitioning for large-scale graph processing. IEEE Transactions on Computers, 70(1):111– 127, January 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Luo:2024:CLR

[LZC^{+}24] Chuanwen Luo, Jian Zhang, XiaoLu Cheng, Yi Hong, Zhibo Chen, and Xiaoshuang Xing. Computation offloading in resource-constrained edge computing systems based on deep reinforcement learning. IEEE Transactions on Computers, 73(1):109– 122, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2021:QMB

[LZF21] Sanjiang Li, Xiangzhen Zhou, and Yuan Feng. Qubit mapping based on subgraph isomorphism and filtered depthlimited search. IEEE Transactions on Computers, 70(11): 1777–1788, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2024:EED

[LZG^{+}24] Jiesong Liu, Feng Zhang, Jiawei Guan, Hsin-Hsuan Sung, Xiaoguang Guo, Saiqin Long, Xiaoyong Du, and Xipeng Shen. Enabling efficient deep learning on MCU with transient redundancy elimination. IEEE Transactions on Computers, 73(12):2649–2663, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2024:BBP

[LZS^{+}24] Shiyu Li, Yuan Zhang, Yaqing Song, Nan Cheng, Kan Yang, and Hongwei Li. Blockchainbased portable authenticated data transmission for mobile edge computing: a universally composable secure solution. IEEE Transactions on Computers, 73(4):1114–1125, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2021:LBM

[LZW^{+}21] L. Li, J. Zhou, T. Wei, M. Chen, and X. S. Hu. Learning-based modeling and optimization for real-time system availability. IEEE Transactions on Computers, 70(4): 581–594, April 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:DDH

[LZW23a] Song Liu, Zengyuan Zhang, and Weiguo Wu. DHTS: a dynamic hybrid tiling strategy for optimizing stencil computation on GPUs. IEEE Transactions on Computers, 72(10): 2795–2807, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2023:HPT

[LZW^{+}23b] Xiao-Yang Liu, Zeliang Zhang, Zhiyuan Wang, Han Lu, Xiaodong Wang, and Anwar Walid. High-performance tensor learning primitives using GPU tensor cores. IEEE Transactions on Computers, 72(6):1733–1746, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2024:EBB

[LZW^{+}24a] Fangxin Liu, Wenbo Zhao, Zongwu Wang, Yongbiao Chen, Xiaoyao Liang, and Li Jiang. ERA-BS: Boosting the efficiency of ReRAMbased PIM accelerator with fine-grained bit-level sparsity. IEEE Transactions on Computers, 73(9):2320– 2334, September 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2024:TCE

[LZW^{+}24b] Xiulong Liu, Zhiyuan Zheng, Wenbin Wang, Hao Xu, Fengjun Xiao, and Keqiu Li. Towards cost-effective and robust packaging in multileader BFT blockchain systems. IEEE Transactions on Computers, 73(11):2590– 2604, November 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lv:2025:SNP

[LZW25] Hao Lv, Lei Zhang, and Ying Wang. In-situ NAS: a plugand-search neural architecture search framework across hardware platforms. IEEE Transactions on Computers, 74(9): 2856–2869, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liu:2025:ECS

[LZX^{+}25] Xiulong Liu, Zhiyuan Zheng, Hao Xu, Zhelin Liang, Gaowei Shi, Chenyu Zhang, and Keqiu Li. Enabling consistent sensing data sharing among IoT edge servers via lightweight consensus. IEEE Transactions on Computers, 74(6): 2045–2057, June 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2025:ADF

[LZY^{+}25] Qinglun Li, Miao Zhang, Nan Yin, Quanjun Yin, Li Shen, and Xiaochun Cao. Asymmetrically decentralized federated learning. IEEE Transactions on Computers, 74(8): 2745–2756, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Liao:2022:SAS

[LZZ^{+}22] Xiaofei Liao, Jin Zhao, Yu Zhang, Bingsheng He, Ligang He, Hai Jin, and Lin Gu. A structure-aware storage optimization for out-of-core concurrent graph processing. IEEE Transactions on Computers, 71(7):1612–1625, July 2022. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-

9956 (electronic).

Morais:2024:EHB

[MÁJG^{+}24] Lucas Morais, Carlos Álvarez, Daniel Jiménez-González, Juan Miguel de Haro, Guido Araujo, Michael Frank, Alfredo Goldman, and Xavier Martorell. Enabling HW-based task scheduling in large multicore architectures. IEEE Transactions on Computers, 73(1): 138–151, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Manocha:2023:GSF

[MAM23] Aninda Manocha, Juan L. Aragón, and Margaret Martonosi. Graphfire: Synergizing fetch, insertion, and replacement policies for graph analytics. IEEE Transactions on Computers, 72(1):291–304, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

McCrabb:2021:OVP

[MB21] Andrew McCrabb and Valeria Bertacco. Optimizing vertex pressure dynamic graph partitioning in many-core systems. IEEE Transactions on Computers, 70(6):936–949, June 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Medina:2021:GMC

[MBP21] R. Medina, E. Borde, and L. Pautet. Generalized mixedcriticality static scheduling for periodic directed acyclic graphs on multi-core processors. IEEE Transactions on Computers, 70(3): 457–470, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mohamed:2023:UPC

[MC23] Nadya A. Mohamed and Joseph R. Cavallaro. A unified parallel CORDIC-based hardware architecture for LSTM network acceleration. IEEE Transactions on Computers, 72(10):2752–2766, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Magnani:2025:SMO

[MCD^{+}25] Gabriele Magnani, Daniele Cattaneo, Lev Denisov, Giuseppe Tagliavini, Giovanni Agosta, and Stefano Cherubin. Synergistic memory optimisations: Precision tuning in heterogeneous memory hierarchies. IEEE Transactions on Computers, 74(9):3168– 3180, September 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Minutoli:2022:PSH

[MCS^{+}22] Marco Minutoli, Vito Giovanni Castellana, Nicola Saporetti, Stefano Devecchi, Marco Lattuada, Pietro Fezzardi, Antonino Tumeo, and Fabrizio Ferrandi. Svelto: High-level synthesis of multithreaded accelerators for graph analytics. IEEE Transactions on Computers, 71(3): 520–533, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Makovenko:2022:ROC

[MCT22] Mykyta Makovenko, Min Cheng, and Chao Tian. Revisiting the optimization of Cauchy Reed–Solomon coding matrix for fault-tolerant data storage. IEEE Transactions on Computers, 71(8): 1839–1846, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mehrabi:2020:ECC

[MDJ20] M. A. Mehrabi, C. Doche, and A. Jolfaei. Elliptic curve cryptography point multiplication core for hardware security module. IEEE Transactions on Computers, 69(11): 1707–1718, November 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mukherjee:2022:TPB

[MDM22] Anandarup Mukherjee, Pallav Kumar Deb, and Sudip Misra. Tremors: Privacy-breaching inference of computing tasks using vibration-based condition monitors. IEEE Transactions on Computers, 71(10): 2620–2631, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mallasen:2024:BPE

[MDPM24] David Mallasén, Alberto A.

Del Barrio, and Manuel Prieto-Matias. Big-PERCIVAL: Exploring the native use of 64bit posit arithmetic in scientific computing. IEEE Transactions on Computers, 73(6): 1472–1485, June 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mondal:2024:IFA

[MDR^{+}24] Sandip Kumar Mondal, Prakash Dey, Himadry Sekhar Roy, Avishek Adhikari, and Subhamoy Maitra. Improved fault analysis on Subterranean 2.0. IEEE Transactions on Computers, 73(6):1631–1639, June 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Moran:2020:EEP

[MFRR20] A. Morán, C. F. Frasser, M. Roca, and J. L. Rosselló. Energy-efficient pattern recognition hardware with elementary cellular automata. IEEE Transactions on Computers, 69(3):392–401, March 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Merlin:2024:DDI

[MGFY24] Itay Merlin, Esteban Garzón, Alexander Fish, and Leonid Yavits. DIPER: Detection and identification of pathogens using edit distance-tolerant resistive CAM. IEEE Transactions on Computers, 73(10): 2463–2473, October 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mohammadi:2020:EED

[MHA^{+}20] M. Mohammadi, S. Han, E. Atoofian, A. Baniasadi, T. M. Aamodt, and W. J. Dally. Energy efficient ondemand dynamic branch prediction models. IEEE Transactions on Computers, 69(3): 453–465, March 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mahmud:2022:QDR

[MHDMEA22] Naveed Mahmud, Bennett Haase-Divine, Andrew MacGillivray, and Esam El-Araby. Quantum dimension reduction for pattern recognition in high-resolution spatio-spectral data. IEEE Transactions on Computers, 71(1):1–12, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mei:2021:ZEJ

[MHJ^{+}21] Linyan Mei, Pouya Houshmand, Vikram Jain, Sebastian Giraldo, and Marian Verhelst. ZigZag: Enlarging joint architecture-mapping design space exploration for DNN accelerators. IEEE Transactions on Computers, 70(8): 1160–1174, August 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Multanen:2022:EEI

[MHK^{+}22] Joonas Multanen, Kari Hepola, Asif Ali Khan, Jeronimo Castrillon, and Pekka Jääskeläinen. Energy-efficient instruction delivery in embedded systems with domain wall memory. IEEE Transactions on Computers, 71(9): 2010–2021, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mills:2023:AFL

[MHM^{+}23] Jed Mills, Jia Hu, Geyong Min, Rui Jin, Siwei Zheng, and Jin Wang. Accelerating federated learning with a global biased optimiser. IEEE Transactions on Computers, 72(6):1804–1814, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Meng:2020:BSC

[MHS^{+}20] D. Meng, R. Hou, G. Shi, B. Tu, A. Yu, Z. Zhu, X. Jia, Y. Wen, and Y. Yang. Builtin security computer: Deploying security-first architecture using active security processor. IEEE Transactions on Computers, 69(11):1571– 1583, November 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Moussa:2025:CTP

[MHT25] Dina A. Moussa, Michael Hefenbrock, and Mehdi Tahoori. Compressed test pattern generation for deep neural networks. IEEE Transactions on Computers, 74(1):307– 315, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mikaitis:2024:MMT

[Mik24] Mantas Mikaitis. Monotonicity of multi-term floatingpoint adders. IEEE Transactions on Computers, 73(6): 1531–1543, June 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Marotta:2022:NNB

[MIPQ22] Romolo Marotta, Mauro Ianni, Alessandro Pellegrini, and Francesco Quaglia. NBBS: A non-blocking buddy system for multi-core machines. IEEE Transactions on Computers, 71(3):599–612, March 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Miwa:2020:FBD

[MIY^{+}20] S. Miwa, M. Ishihara, H. Yamaki, H. Honda, and M. Schulz. Footprint-based DIMM hotplug. IEEE Transactions on Computers, 69(2):172– 184, February 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Majumder:2021:RCS

[MKH^{+}21] Pritam Majumder, Sungkeun Kim, Jiayi Huang, Ki Hwan Yum, and Eun Jung Kim. Remote control: A simple deadlock avoidance scheme for modular systems-on-chip. IEEE Transactions on Computers, 70(11):1928–1941, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mert:2022:ESF

[MKÖ^{+}22] Ahmet Can Mert, Emre Karabulut, Erdinç Öztürk, Erkay Savaş, and Aydin Aysu. An extensive study of flexible design methods for the number theoretic transform. IEEE Transactions on Computers, 71(11):2829–2843, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2024:SDG

[MKY^{+}24] Lianbo Ma, Haidong Kang, Guo Yu, Qing Li, and Qiang He. Single-domain generalized predictor for neural architecture search system. IEEE Transactions on Computers, 73(5):1400–1413, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Myung:2021:ESE

[MKYP21] Kihyeon Myung, Sunggon Kim, Heon Young Yeom, and Jiwoong Park. Efficient and scalable external sort framework for NVMe SSD. IEEE Transactions on Computers, 70(12):2211–2217, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2024:GGF

[MLL^{+}24] Liyuan Ma, Xiulong Liu, Yuhan Li, Chenyu Zhang, Gaowei Shi, and Keqiu Li. GFBE: a generalized and finegrained blockchain evaluation framework. IEEE Transactions on Computers, 73(3): 942–955, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2025:HTD

[MLL^{+}25] Peijun Ma, Jie Li, Hongjin Liu, Jiangyi Shi, Shaolin Zhang, Weitao Pan, and Yue Hao. Hardware Trojan detection methods for gate-level netlists based on graph neural networks. IEEE Transactions on Computers, 74(5):

1470–1481, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mu:2023:HHA

[MLW^{+}23] Pengyu Mu, Yi Liu, Rui Wang, Guoxiang Liu, Zhonghao Sun, Hailong Yang, Zhongzhi Luan, and Depei Qian. HAOTuner: a hardware adaptive operator autotuner for dynamic shape tensor compilers. IEEE Transactions on Computers, 72(11): 3178–3190, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mu:2025:DLO

[MLW^{+}25] Pengyu Mu, Yi Liu, Rui Wang, Guoxiang Liu, Hangcheng An, Qianhe Zhao, Hailong Yang, Chenhao Xie, Zhongzhi Luan, Chunye Gong, and Depei Qian. Deep learning operators performance tuning for changeable sized input data on tensor accelerate hardware. IEEE Transactions on Computers, 74(6):2101–2113, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Majumder:2020:PRT

[MNB20] S. Majumder, J. F. D. Nielsen, and T. Bak. PaRTAA: A realtime multiprocessor for mixedcriticality airborne systems. IEEE Transactions on Computers, 69(8):1221–1232, August 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mert:2022:LLA

[MÖS22] Ahmet Can Mert, Erdinç Öztürk, and Erkay Savaş. Low-latency ASIC algorithms of modular squaring of large integers for VDF evaluation. IEEE Transactions on Computers, 71(1):107–120, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Min:2025:MCA

[MPYJ25] Cheolgi Min, Jiwoong Park, Heon Young Yeom, and Hyungsoo Jung. MDC+: a cooperative approach to memory-efficient fork-based checkpointing for in-memory database systems. IEEE Transactions on Computers, 74(9):3059–3071, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Minaeva:2021:CPO

[MRA^{+}21] Anna Minaeva, Debayan Roy, Benny Akesson, Zdenêk Hanzálek, and Samarjit Chakraborty. Control performance optimization for application integration on automotive architectures. IEEE Transactions on Computers, 70(7): 1059–1073, July 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Motetti:2024:JPC

[MRB^{+}24] Beatrice Alessandra Motetti, Matteo Risso, Alessio Burrello, Enrico Macii, Massimo Poncino, and Daniele Jahier Pagliari. Joint pruning and channel-wise mixed-precision quantization for efficient deep neural networks. IEEE Transactions on Computers, 73(11): 2619–2633, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mishty:2025:CGO

[MS25] Kaniz Mishty and Mehdi Sadi. Chiplet-Gym: Optimizing chiplet-based AI accelerator design with reinforcement learning. IEEE Transactions on Computers, 74(1): 43–56, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mao:2024:JVN

[MSLY24] Yingling Mao, Xiaojun Shang, Yu Liu, and Yuanyuan Yang. Joint virtual network function placement and flow routing in edge-cloud continuum. IEEE Transactions on Computers, 73(3):872–886, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Menshchikov:2021:RTD

[MSP^{+}21] Alexander Menshchikov, Dmitrii Shadrin, Viktor Prutyanov, Daniil Lopatkin, Sergey Sosnin, Evgeny Tsykunov, Evgeny Iakovlev, and Andrey Somov. Real-time detection of hogweed: UAV platform empowered by deep learning. IEEE Transactions on Computers, 70(8):1175–1188, August 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mao:2021:LCE

[MSSL21] Haiyu Mao, Jiwu Shu, Mingcong Song, and Tao Li. Lr-GAN: A compact and energy efficient PIM-based architecture for GAN training. IEEE Transactions on Computers, 70(9):1427–1442, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2021:TAR

[MSW^{+}21] C. Ma, Z. Shen, J. Wang, Y. Wang, R. Chen, Y. Guan, and Z. Shao. Tiler: An autonomous region-based scheme for SMR storage. IEEE Transactions on Computers, 70(2): 291–304, February 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mottaghi:2022:FFI

[MSZ22] Mohammad Hadi Mottaghi, Mehdi Sedighi, and Morteza Saheb Zamani. FIFA: a fully invertible FPGA architecture to reduce BTI-induced aging effects. IEEE Transactions on Computers, 71(9):2277– 2286, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mirzaei-Teshnizi:2025:PMM

[MTK25] Shahab Mirzaei-Teshnizi and Parviz Keshavarzi. Parallel modular multiplication using variable length algorithms. IEEE Transactions on Computers, 74(1):143–154, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mukhanov:2021:RDO

[MTV^{+}21] Lev Mukhanov, Konstantinos Tovletoglou, Hans Vandierendonck, Dimitrios S. Nikolopoulos, and Georgios Karakonstantis. Revealing DRAM operating GuardBands through workload-aware error predictive modeling. IEEE Transactions on Computers, 70(11): 1976–1987, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mo:2024:LEF

[MWJ^{+}24] Hai Mo, Yong Wu, Honglan Jiang, Zining Ma, Fabrizio Lombardi, Jie Han, and Leibo Liu. Learning the error features of approximate multipliers for neural network applications. IEEE Transactions on Computers, 73(3): 842–856, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2023:MEP

[MXY^{+}23] Tengchao Ma, Changqiao Xu, Shujie Yang, Yiting Huang, Qingzhao An, Xiaohui Kuang, and Luigi Alfredo Grieco. A mutation-enabled proactive defense against serviceoriented man-in-the-middle attack in Kubernetes. IEEE Transactions on Computers, 72(7):1843–1856, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mandal:2020:AEI

[MYGA20] K. Mandal, B. Yang, G. Gong, and M. Aagaard. Analysis and efficient implementations of a class of composited de Bruijn sequences. IEEE Transactions on Computers, 69(12):1835– 1848, December 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mizutani:2021:OLF

[MYUK21] Kenji Mizutani, Hiroshi Yamaguchi, Yutaka Urino, and Michihiro Koibuchi. OPTWEB: A lightweight fully connected inter-FPGA network for efficient collectives. IEEE Transactions on Computers, 70(6): 849–862, June 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ma:2025:DMM

[MZM^{+}25] Lianbo Ma, Yuee Zhou, Ye Ma, Guo Yu, Qing Li, Qiang He, and Yan Pei. Defying multi-model forgetting in one-shot neural architecture search using orthogonal gradient learning. IEEE Transactions on Computers, 74(5): 1678–1689, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Mi:2022:GFI

[MZZC22] Zeyu Mi, Haoqi Zhuang, Binyu Zang, and Haibo Chen. General and fast inter-process communication via bypassing privileged software. IEEE Transactions on Computers, 71(10):2435–2448, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nabeel:2020:RTS

[NAP^{+}20] M. Nabeel, M. Ashraf, S. Patnaik, V. Soteriou, O. Sinanoğlu, and J. Knechtel. 2.5D root of trust: Secure systemlevel integration of untrusted chiplets. IEEE Transactions on Computers, 69(11):1611– 1625, November 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nannipieri:2025:HDA

[NCD^{+}25] Pietro Nannipieri, Luca Crocetti, Stefano Di Matteo, Luca Fanucci, and Sergio Saponara. Hardware design of an advanced-feature cryptographic tile within the European Processor Initiative. IEEE Transactions on Computers, 74(3):762–775, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Niu:2024:GBA

[NHW^{+}24] Xin Niu, Yajing Huang, Zhiwei Wang, Chen Yu, and Hai Jin. Game-based adaptive FLOPs and partition point decision mechanism with latency and energy-efficient tradeoff for edge intelligence. IEEE Transactions on Computers, 73(4):1099–1113, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nguyen:2024:HHR

[NICY24] Thai-Hoang Nguyen, Muhammad Imran, Jaehyuk Choi, and Joon-Sung Yang. HY-DRA: a hybrid resistance drift resilient architecture for phase change memory-based neural network accelerators. IEEE Transactions on Computers, 73(9):2123–2135, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nath:2022:SCW

[NK22] Arijit Nath and Hemangee K. Kapoor. SWEL-COFAE: Wear leveling and adaptive encoding assisted compression of frequent words in non-volatile main memories. IEEE Transactions on Computers, 71(9): 2263–2276, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Novkin:2024:AQA

[NKA24] Rodion Novkin, Florian Klemme, and Hussam Amrouch. Approximationand quantization-aware training for graph neural networks. IEEE Transactions on Computers, 73(2):599–612, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ni:2023:HHP

[NKeSK^{+}23] Ziying Ni, Ayesha Khalid, Dur e Shahwar Kundi, Máire O Neill, and Weiqiang Liu. HPKA: a high-performance CRYSTALS-Kyber accelerator exploring efficient pipelining. IEEE Transactions on Computers, 72(12):3340– 3353, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Noh:2023:FFD

[NKL^{+}23] Seock-Hwan Noh, Jahyun Koo, Seunghyun Lee, Jongse Park, and Jaeha Kung. FlexBlock: a flexible DNN training accelerator with multi-mode block floating point support. IEEE Transactions on Computers, 72(9):

2522–2535, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nwobodo:2025:HAF

[NKN^{+}25] Onyeka Josephine Nwobodo, Godlove Suila Kuaban, Valery Nkemeni, Kamil Wereszczynski, and Krzysztof Adam Cyran. A hybrid adaptive filter for head tracking in augmented reality (AR)-based flight simulators. IEEE Transactions on Computers, 74(8): 2581–2592, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Niu:2025:CTS

[NLC^{+}25] Xin Niu, Xianwei Lv, Wang Chen, Chen Yu, and Hai Jin. Computing tasks saving schemes through early exit in edge intelligence-assisted systems. IEEE Transactions on Computers, 74(5):1565–1576, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nguyen:2022:EPL

[NM22] Thien Nguyen and Alexander McCaskey. Enabling pulselevel programming, compilation, and execution in XACC. IEEE Transactions on Computers, 71(3):547–558, March 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Niaz:2025:ARE

[NML25] Hafiz Adnan Niaz, Ravi Reddy Manumachu, and Alexey Lastovetsky. Accurate and reliable energy measurement and modelling of data transfer between CPU and GPU in parallel applications on heterogeneous hybrid platforms. IEEE Transactions on Computers, 74(3):1011–1024, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nguyen:2025:SHS

[NNH^{+}25] Nang Hung Nguyen, Truong Thao Nguyen, Trong Nghia Hoang, Hieu H. Pham, Thanh Hung Nguyen, and Phi Le Nguyen. SAFA: Handling sparse and scarce data in federated learning with accumulative learning. IEEE Transactions on Computers, 74(6):1844–1856, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Newell:2020:IBB

[NP20] A. Newell and S. Pupyrev.

Improved basic block reordering. IEEE Transactions on Computers, 69(12):1784– 1794, December 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nath:2022:EWV

[NS22] Kaushik Nath and Palash Sarkar. Efficient 4-way vectorizations of the Montgomery ladder. IEEE Transactions on Computers, 71(3):712– 723, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nguyen:2023:DSV

[NT23] Truc Nguyen and My T. Thai. Denial-of-service vulnerability of hash-based transaction sharding: Attack and countermeasure. IEEE Transactions on Computers, 72(3): 641–652, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ngoc:2025:VNB

[NTDH25] Tu Dinh Ngoc, Boris Teabe, Georges Da Costa, and Daniel Hagimont. Virtual NVMebased storage function framework with fast I/O request state management. IEEE Transactions on Computers, 74(7):2253–2266, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Nie:2024:VSP

[NTL^{+}24] Chen Nie, Chenyu Tang, Jie Lin, Huan Hu, Chenyang Lv, Ting Cao, Weifeng Zhang, Li Jiang, Xiaoyao Liang, Weikang Qian, Yanan Sun, and Zhezhi He. VSPIM: SRAM processing-in-memory DNN acceleration via vector– scalar operations. IEEE Transactions on Computers, 73(10):2378–2390, October 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Navarro-Torres:2025:CEL

[NTPAB^{+}25] Agustín Navarro-Torres, Biswabandan Panda, Jesús Alastruey-Benedé, Pablo Ibáñez, Víctor Viñals-Yúfera, and Alberto Ros. A complexity-effective local delta prefetcher. IEEE Transactions on Computers, 74(5):1482–1494, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Neves:2021:CAD

[NTR21] N. Neves, P. Tomás, and N. Roma. Compiler-assisted data streaming for regular code structures. IEEE Transactions on Computers, 70(3): 483–494, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Naji:2025:NTA

[NWH^{+}25] Abdulbary Naji, Xingfu Wang, Ammar Hawbani, Aiman Ghannami, Liang Zhao, XiaoHua Xu, and Wei Zhao. NetMod: Toward accelerating cloud RAN distributed unit modulation within programmable switches. IEEE Transactions on Computers, 74(2):665–677, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Naji:2025:NNN

[NWL^{+}25] Abdulbary Naji, Xingfu Wang, Ping Liu, Ammar Hawbani, Liang Zhao, Xiaohua Xu, and Fuyou Miao. NetCRC-NR: Innetwork 5G NR CRC accelerator. IEEE Transactions on Computers, 74(4):1418–1430, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ortega:2023:APS

[OAB^{+}23] Cristobal Ortega, Lluc Alvarez, Alper Buyuktosunoglu, Ramon Bertran, Todd Rosedahl, Pradip Bose, and Miquel Moreto. Adaptive power shifting for power-constrained heterogeneous systems. IEEE Transactions on Computers, 72(3):627–640, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ortega:2021:IAH

[OAC^{+}21] C. Ortega, L. Alvarez, M. Casas, R. Bertran, A. Buyuktosunoglu, A. E. Eichenberger, P. Bose, and M. Moretó. Intelligent adaptation of hardware knobs for improving performance and power consumption. IEEE Transactions on Computers, 70(1):1–16, January 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Oh:2023:RDP

[OAK^{+}23] Byoungchan Oh, Nilmini Abeyratne, Nam Sung Kim, Jeongseob Ahn, Ronald G. Dreslinski, and Trevor Mudge. Rethinking DRAM’s page mode with STT-MRAM. IEEE Transactions on Computers, 72(5):1503–1517, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ojha:2023:PCS

[OD23] Divya Ojha and Sandhya Dwarkadas. Preventing coherence state side channel leaks using TimeCache. IEEE Transactions on Computers, 72(2):374–385, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Omar:2020:OSC

[ODK20] H. Omar, B. D’Agostino, and O. Khan. OPTIMUS: A security-centric dynamic hardware partitioning scheme for processors that prevent microarchitecture state attacks. IEEE Transactions on Computers, 69(11):1558– 1570, November 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Owahid:2023:IPB

[OJ23] Abdullah A. Owahid and Eugene B. John. Instruction profiling based predictive throttling for power and performance. IEEE Transactions on Computers, 72(12):3532– 3545, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Oh:2020:POR

[OKC^{+}20] M. Oh, K. Kim, D. Choi, H. Lee, and E. Chung. Peroperation reusability based allocation and migration policy for hybrid cache. IEEE Transactions on Computers, 69(2): 158–171, February 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Orosa:2024:EEC

[OKU^{+}24] Lois Orosa, Skanda Koppula, Yaman Umuroglu, Konstantinos Kanellopoulos, Juan Gómez-Luna, Michaela Blott, Kees Vissers, and Onur Mutlu. EcoFlow: Efficient convolutional dataflows on lowpower neural network accelerators. IEEE Transactions on Computers, 73(9):2275– 2289, September 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Olivier:2022:SLB

[OLC^{+}22] Pierre Olivier, Hugo Lefeuvre, Daniel Chiba, Stefan Lankes, Changwoo Min, and Binoy Ravindran. A syscall-level binary-compatible unikernel. IEEE Transactions on Computers, 71(9):2116–2127, September 2022. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Ollivier:2023:TCS

[OLD^{+}23] Sebastien Ollivier, Stephen Longofono, Prayash Dutta, Jingtong Hu, Sanjukta Bhanja, and Alex K. Jones. Toward comprehensive shifting fault tolerance for domain-wall memories with PIETT. IEEE Transactions on Computers, 72(4):1095–1109, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ou:2020:LDA

[OLZ^{+}20] C. Ou, S.-K. Lam, C. Zhou, G. Jiang, and F. Zhang. A lightweight detection algorithm for collision-optimized divide-and-conquer attacks. IEEE Transactions on Computers, 69(11):1694–1706, November 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Oku:2022:HRB

[OTTT22] Daisuke Oku, Masashi Tawada, Shu Tanaka, and Nozomu Togawa. How to reduce the bitwidth of an Ising model by adding auxiliary spins. IEEE Transactions on Computers, 71(1):223–234, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Paim:2022:FCT

[PAR^{+}22] Guilherme Paim, Hussam Amrouch, Leandro M. G. Rocha, Brunno Abreu, Eduardo Antônio César da Costa, Sergio Bampi, and Jörg Henkel. A framework for crossing temperature-induced timing errors underlying hardware accelerators to the algorithm and application layers. IEEE Transactions on Computers, 71(2):349–363, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Paul:2023:PCC

[PB23a] Shubhra Deb Paul and Swarup Bhunia. CurIAs: Currentbased IC authentication by exploiting supply current variations. IEEE Transactions on Computers, 72(2):466– 479, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Prasad:2023:MDP

[PB23b] Ananth Krishna Prasad and Mahdi Nazm Bojnordi. Monarch: a durable polymorphic memory for data intensive applications. IEEE Transactions on Computers, 72(2): 535–547, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Paguada:2025:BPP

[PBBA25] Servio Paguada, Lejla Batina, Ileana Buhan, and Igor Armendariz. Being patient and persistent: Optimizing an early stopping strategy for deep learning in profiled attacks. IEEE Transactions on Computers, 74(3): 875–886, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Papaphilippou:2024:EDA

[PC24] Philippos Papaphilippou and Thiem Van Chu. Efficient deadlock avoidance for 2-D mesh NoCs that use OQ or VOQ routers. IEEE Transactions on Computers, 73(5): 1414–1426, May 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pratihar:2023:BSF

[PCA^{+}23] Kuheli Pratihar, Urbi Chatterjee, Manaar Alam, Rajat Subhra Chakraborty, and Debdeep Mukhopadhyay. Birds of the same feather flock together: a dual-mode circuit candidate for strong PUF-TRNG functionalities. IEEE Transactions on Computers, 72(6):1636–1651, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Perotti:2024:AES

[PCA^{+}24] Matteo Perotti, Matheus Cavalcante, Renzo Andri, Lukas Cavigelli, and Luca Benini. Ara2: Exploring single- and multi-core vector processing with an efficient RVV 1.0 compliant open-source processor. IEEE Transactions on Computers, 73(7):1822–1836, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pazzaglia:2023:OIC

[PCBD23] Paolo Pazzaglia, Daniel Casini, Alessandro Biondi, and Marco Di Natale. Optimizing intercore communications under the LET paradigm using DMA engines. IEEE Transactions on Computers, 72(1): 127–139, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2022:RCR

[PCCK22] Kangkyu Park, Seungkyu Choi, Yeongjae Choi, and Lee-Sup Kim. Rare computing: Removing redundant multiplications from sparse and repetitive data in deep neural networks. IEEE Transactions on Computers, 71(4):795–808, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pagliari:2021:CID

[PCMP21] Daniele Jahier Pagliari, Roberta Chiaro, Enrico Macii, and Massimo Poncino. CRIME: Input-dependent collaborative inference for recurrent neural networks. IEEE Transactions on Computers, 70(10): 1626–1639, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Patel:2021:SLK

[PD21] Chintan Patel and Nishant Doshi. Secure lightweight key exchange using ECC for user-gateway paradigm. IEEE Transactions on Computers, 70(11):1789–1803, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2022:WAC

[PE22] Jonggyu Park and Young Ik Eom. Weight-aware cache for application-level proportional I/O sharing. IEEE Transactions on Computers, 71(10): 2395–2407, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Picornell:2021:EPM

[PFHD21] T. Picornell, J. Flich, C. Hernández, and J. Duato. Enforcing predictability of many-cores with DCFNoC. IEEE Transactions on Computers, 70(2): 270–283, February 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Papadimitriou:2023:SDC

[PG23] George Papadimitriou and Dimitris Gizopoulos. Silent data corruptions: Microarchitectural perspectives. IEEE Transactions on Computers, 72(11):3072–3085, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2024:BBD

[PHC24] Mingeon Park, Seokjin Hwang, and Hyungmin Cho. BiRD: Bi-directional input reuse dataflow for enhancing depthwise convolution performance on systolic arrays. IEEE Transactions on Computers, 73(12):2708–2721, 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pan:2025:AMB

[PHL^{+}25] Zhe Pan, Shuibing He, Xu Li, Xuechen Zhang, Rui Wang, Yanlong Yin, and Gang Chen. Advanced maximal biclique enumeration on GPUs using bitmaps. IEEE Transactions on Computers, 74(8): 2552–2566, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Piliposyan:2023:PHT

[PK23] Gor Piliposyan and Saqib Khursheed. PCB hardware Trojan run-time detection through machine learning.

IEEE Transactions on Computers, 72(7):1958–1970, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pradhan:2023:MBV

[PKPR23] Ankit Pradhan, Jonathan King, Srinivas Pinisetty, and Partha S. Roop. Model based verification of spiking neural networks in cyber physical systems. IEEE Transactions on Computers, 72(9): 2426–2439, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2021:RCR

[PL21] M. C. Park and D. H. Lee.

Random CFI (RCFI): Efficient fine-grained control-flow integrity through random verification. IEEE Transactions on Computers, 70(5):733–745, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Papaphilippou:2022:FFL

[PLB22] Philippos Papaphilippou, Wayne Luk, and Chris Brooks. FLiMS: a fast lightweight 2way merger for sorting. IEEE Transactions on Computers, 71(12):3215–3226, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pan:2024:AAM

[PLH^{+}24] Zhe Pan, Xu Li, Shuibing He, Xuechen Zhang, Rui Wang, Yunjun Gao, Gang Chen, and Xian-He Sun. AMBEA: Aggressive maximal biclique enumeration in large bipartite graph computing. IEEE Transactions on Computers, 73(12):2664–2677, 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Phung:2020:LPM

[PLZ20] J. Phung, Y. C. Lee, and A. Y. Zomaya. Lightweight power monitoring framework for virtualized computing environments. IEEE Transactions on Computers, 69(1):14–25, January 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pan:2023:AIN

[PLZ^{+}23] Yuqian Pan, Zhaojun Lu, Haichun Zhang, Haoming Zhang, Md Tanvir Arafin, Zhenglin Liu, and Gang Qu. ADLPT: Improving 3D NAND flash memory reliability by adaptive lifetime prediction techniques. IEEE Transactions on Computers, 72(6): 1525–1538, June 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2020:ISR

[PM20] H. Park and J. Moon. Improving SSD read latency via coding. IEEE Transactions on Computers, 69(12):1809– 1822, December 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pan:2025:ATA

[PM25] Zhixin Pan and Prabhat Mishra. AI trojan attack for evading machine learningbased detection of hardware trojans. IEEE Transactions on Computers, 74(3): 860–874, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Podlubne:2024:MBG

[PMA^{+}24] Ariel Podlubne, Johannes Mey, Andreas Andreou, Sergio Pertuz, Uwe Aßmann, and Diana Göhringer. Modelbased generation of hardware/ software architectures with hybrid schedulers for robotics systems. IEEE Transactions on Computers, 73(7): 1640–1654, July 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pogue:2024:FIP

[PN24] Trevor E. Pogue and Nicola Nicolici. Fast inner-product algorithms and architectures for deep neural network accelerators. IEEE Transactions on Computers, 73(2): 495–509, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pogue:2025:KMM

[PN25] Trevor E. Pogue and Nicola Nicolici. Karatsuba matrix multiplication and its efficient custom hardware implementations. IEEE Transactions on Computers, 74(4):1377–1391, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Paul:2023:TBM

[PNK^{+}23] Bikram Paul, Angana Nath, Srinivasan Krishnaswamy, Jan Pidanic, Zdenek Nemec, and Gaurav Trivedi. Tensor based multivariate polynomial modulo multiplier for cryptographic applications. IEEE Transactions on Computers, 72(6):1581–1594, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ponzina:2021:EEC

[PPQBA21] Flavio Ponzina, Miguel Peón-Quirós, Andreas Burg, and David Atienza. E2CNNs: Ensembles of convolutional neural networks to improve robustness against memory errors in edge-computing devices. IEEE Transactions on Computers, 70(8):1199– 1212, August 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pedrero:2022:SBT

[PQG^{+}22] Manuel Pedrero, Ricardo Quislant, Eladio Gutierrez, Emilio L. Zapata, and Oscar Plata. Speculative barriers with transactional memory. IEEE Transactions on Computers, 71(1):197–208, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pietrykowski:2022:PEP

[PS22] Michael Pietrykowski and Carol Smidts. Predictive execution of parallel simulations in hard real-time systems. IEEE Transactions on Computers, 71(12):3227– 3241, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Perina:2021:FRT

[PSBB21] André B. Perina, Arthur Silitonga, Jürgen Becker, and Vanderlei Bonato. Fast resource and timing aware design optimisation for highlevel synthesis. IEEE Transactions on Computers, 70(12): 2070–2082, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pan:2022:HAM

[PSM22] Zhixin Pan, Jennifer Sheldon, and Prabhat Mishra. Hardware-assisted malware detection and localization using explainable machine learning. IEEE Transactions on Computers, 71(12):3308– 3321, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Perez:2021:SLL

[PVB21] Iván Pérez, Enrique Vallejo, and Ramón Beivide. S-SMART++: A low-latency NoC leveraging speculative bypass requests. IEEE Transactions on Computers, 70(6): 819–832, June 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Peng:2022:MNM

[PYDG22] Bo Peng, Jianguo Yao, Yaozu Dong, and Haibing Guan. MDev-NVMe: Mediated passthrough NVMe virtualization solution with adaptive polling. IEEE Transactions on Computers, 71(2):251–265, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Park:2020:PRB

[PYS20] J. Park, H. Yeom, and Y. Son. Page reusability-based cache partitioning for multi-core systems. IEEE Transactions on Computers, 69(6):812–818, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Paul:2023:RES

[PYS^{+}23] Bikram Paul, Tarun Kumar Yadav, Balbir Singh, Srinivasan Krishnaswamy, and Gaurav Trivedi. A resource efficient software-hardware codesign of lattice-based homomorphic encryption scheme on the FPGA. IEEE Transactions on Computers, 72(5): 1247–1260, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Peng:2022:PLL

[PYW^{+}22] Hao Peng, Renyu Yang, Zheng Wang, Jianxin Li, Lifang He, Philip S. Yu, Albert Y. Zomaya, and Rajiv Ranjan. Lime: Low-cost and incremental learning for dynamic heterogeneous information networks. IEEE Transactions on Computers, 71(3): 628–642, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Peng:2025:HEP

[PYW^{+}25] Xiaosong Peng, Laurence T.

Yang, Xiaokang Wang, Debin Liu, and Jie Li. A high-efficiency parallel mechanism for canonical polyadic decomposition on heterogeneous computing platform. IEEE Transactions on Computers, 74(10):3377–3389, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Peng:2021:TON

[PYYG21] Bo Peng, Ming Yang, Jianguo Yao, and Haibing Guan. A throughput-oriented NVMe storage virtualization with workload-aware management. IEEE Transactions on Computers, 70(12):2112–2124, December 2021. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Pan:2023:LPF

[PZY^{+}23] Yuqian Pan, Haichun Zhang, Runze Yu, Zhaojun Lu, Haoming Zhang, and Zhenglin Liu. LightWarner: Predicting failure of 3D NAND flash memory using reinforcement learning. IEEE Transactions on Computers, 72(3): 853–867, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qi:2023:SHB

[QCX^{+}23] Huayi Qi, Ye Cheng, Minghui Xu, Dongxiao Yu, Haipeng Wang, and Weifeng Lyu. Split: a hash-based memory optimization method for zero-knowledge succinct noninteractive argument of knowledge (zk-SNARK). IEEE Transactions on Computers, 72(7):1857–1870, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qiu:2024:GPA

[QHT^{+}24] Yudi Qiu, Tao Huang, Yuxin Tang, Yanwei Liu, Yang Kong, Xulin Yu, Xiaoyang Zeng, and Yibo Fan. Gem5Tune: a parameter auto-tuning framework for gem5 simulator to reduce errors. IEEE Transactions on Computers, 73(3): 902–914, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qureshi:2021:GSA

[QHZ^{+}21] Yasir Mahmood Qureshi, Jose Manuel Herruzo, Marina Zapater, Katzalin Olcoz, Sonia Gonzalez-Navarro, Oscar Plata, and David Atienza. Genome sequence alignment — design space exploration for optimal performance and energy architectures. IEEE Transactions on Computers, 70(12):2218–2233, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qu:2025:FAF

[QJY^{+}25] Zhihao Qu, Ninghui Jia, Baoliu Ye, Shihong Hu, and Song Guo. FedQClip: Accelerating federated learning via quantized clipped SGD. IEEE Transactions on Computers, 74(2):717–730, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qi:2024:LEC

[QLH^{+}24] Jiaxing Qi, Zhongzhi Luan, Shaohan Huang, Carol Fung, and Hailong Yang. LogSay: an efficient comprehension system for log numerical reasoning. IEEE Transactions on Computers, 73(7):1809–1821, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qiu:2025:OLN

[QSC^{+}25] Tie Qiu, Jingchen Sun, Ning Chen, Songwei Zhang, Weisheng Si, and Xingwei Wang. Olive-like networking: a uniformity driven robust topology generation scheme for IoT system. IEEE Transactions on Computers, 74(1): 86–100, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qian:2020:GEI

[QWK20] X. Qian, Y. Wang, and A. Karanth. Guest editors introduction to the special issue on machine learning architectures and accelerators. IEEE Transactions on Computers, 69(7):929–930, July 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qian:2023:OAE

[QWT^{+}23] Bin Qian, Zhenyu Wen, Junqi Tang, Ye Yuan, Albert Y. Zomaya, and Rajiv Ranjan.

OsmoticGate: Adaptive edgebased real-time video analytics for the Internet of Things. IEEE Transactions on Computers, 72(4):1178–1193, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Qiu:2024:EPB

[QZZ^{+}24] Han Qiu, Yi Zeng, Qinkai Zheng, Shangwei Guo, Tianwei Zhang, and Hewu Li. An efficient preprocessingbased approach to mitigate advanced adversarial attacks. IEEE Transactions on Computers, 73(3):645–655, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rispo:2024:RTA

[RACB24] Veronica Rispo, Federico Aromolo, Daniel Casini, and Alessandro Biondi. Responsetime analysis of bundled gang tasks under partitioned FP scheduling. IEEE Transactions on Computers, 73(11): 2534–2547, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ramezanpour:2020:SPS

[RAD20] K. Ramezanpour, P. Ampadu, and W. Diehl. SCAUL: Power side-channel analysis with unsupervised learning. IEEE Transactions on Computers, 69(11):1626–1638, November 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Risso:2023:LNA

[RBC^{+}23] Matteo Risso, Alessio Burrello, Francesco Conti, Lorenzo Lamberti, Yukai Chen, Luca Benini, Enrico Macii, Massimo Poncino, and Daniele Jahier Pagliari. Lightweight neural architecture search for temporal convolutional networks at the edge. IEEE Transactions on Computers, 72(3): 744–758, March 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roy:2021:DFA

[RBM21] Dibyendu Roy, Bhagwan Bathe, and Subhamoy Maitra. Differential fault attack on Kreyvium & FLIP. IEEE Transactions on Computers, 70(12):2161–2167, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Richter-Brockmann:2022:FBS

[RBMG22] Jan Richter-Brockmann, Johannes Mono, and Tim Güneysu. Folding BIKE: Scalable hardware implementation for reconfigurable devices. IEEE Transactions on Computers, 71(5):1204–1215, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Richter-Brockmann:2023:RFA

[RBSG23] Jan Richter-Brockmann, Pascal Sasdrich, and Tim Güneysu. Revisiting fault adversary models hardware faults in theory and practice. IEEE Transactions on Computers, 72(2): 572–585, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Riedel:2023:MSM

[RCAB23] Samuel Riedel, Matheus Cavalcante, Renzo Andri, and Luca Benini. MemPool: a scalable manycore architecture with a low-latency shared L1 memory. IEEE Transactions on Computers, 72(12): 3561–3575, December 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Resch:2025:CSM

[RCC^{+}25] Salonik Resch, Hüsrev Cılasun, Zamshed I. Chowdhury, Masoud Zabihi, Yang Lv, Jian-Ping Wang, Sachin S. Sapatnekar, Ismail Akturk, and Ulya R. Karpuzcu. The case for secure miniservers beyond the edge. IEEE Transactions on Computers, 74(10):3448– 3461, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rathore:2021:LFL

[RCS^{+}21] Vijeta Rathore, Vivek Chaturvedi, Amit K. Singh, Thambipillai Srikanthan, and Muhammad Shafique. Longevity framework: Leveraging online integrated aging-aware hierarchical mapping and VF-selection for lifetime reliability optimization in manycore processors. IEEE Transactions on Computers, 70(7):1106–1119, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roy:2023:SSM

[RDS23] Sanjit Kumar Roy, Rajesh Devaraj, and Arnab Sarkar. SAFLA: Scheduling multiple real-time periodic task graphs on heterogeneous systems. IEEE Transactions on Computers, 72(4):1067–1080, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reviriego:2024:SQF

[RGD^{+}24] Pedro Reviriego, Miguel González, Niv Dayan, Gabriel Huecas, Shanshan Liu, and Fabrizio Lombardi. On the security of quotient filters: Attacks and potential countermeasures. IEEE Transactions on Computers, 73(9):2165– 2177, September 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ray:2022:MEH

[RGS22] Sanchita Saha Ray, Surajeet Ghosh, and Bhaskar Sardar. Memory efficient hash-based longest prefix matching architecture with zero false +ve and nearly zero false ve rate for IP processing. IEEE Transactions on Computers, 71(6): 1261–1275, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ravipati:2024:CCA

[RGvS^{+}24] Divya Praneetha Ravipati, Ramanuj Goel, Victor M. van Santen, Hussam Amrouch, and Preeti Ranjan Panda. CAPE: Criticality-aware performance and energy optimization policy for NCFETbased caches. IEEE Transactions on Computers, 73 (12):2830–2843, 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rasolroveicy:2024:BBP

[RHF24] Mohammadreza Rasolroveicy, Wejdene Haouari, and Marios Fokaefs. BlockCompass: a benchmarking platform for blockchain performance. IEEE Transactions on Computers, 73(8):2111– 2122, August 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ros:2024:WPA

[RJ24] Alberto Ros and Alexandra Jimborean. Wrong-path-aware entangling instruction prefetcher. IEEE Transactions on Computers, 73(2): 548–559, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Radheshwar:2023:DFA

[RKMR23] R Radheshwar, Meenakshi Kansal, Pierrick Méaux, and Dibyendu Roy. Differential fault attack on Rasta and FiLIPdsm. IEEE Transactions on Computers, 72(8): 2418–2425, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Restuccia:2023:FDV

[RMKO23] Francesco Restuccia, Andres Meza, Ryan Kastner, and Jason Oberg. A framework for design, verification, and management of SoC access control systems. IEEE Transactions on Computers, 72(2): 386–400, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reviriego:2021:SET

[RMO21] P. Reviriego, J. Martínez, and M. Ottavi. Soft error tolerant count min sketches. IEEE Transactions on Computers, 70(2):284–290, February 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ro:2022:HFS

[RMR22] Jin Woo Ro, Avinash Malik, and Partha Roop. High fidelity simulation of hybrid systems using higher order hybrid automata. IEEE Transactions on Computers, 71(7): 1668–1680, July 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reyhani-Masoleh:2020:NLA

[RMTA20] A. Reyhani-Masoleh, M. Taha, and D. Ashmawy. New lowarea designs for the AES forward, inverse and combined S-boxes. IEEE Transactions on Computers, 69(12):1757– 1773, December 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rodriguez:2022:EHB

[ROPdlT22] Alfonso Rodríguez, Andrés Otero, Marco Platzner, and Eduardo de la Torre. Exploiting hardware-based dataparallel and multithreading models for smart edge computing in reconfigurable FP-GAs. IEEE Transactions on Computers, 71(11):2903– 2914, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roth:2024:EDS

[Rot24] Ron M. Roth. Error-detection schemes for analog contentaddressable memories. IEEE Transactions on Computers, 73(7):1795–1808, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Restuccia:2023:BMA

[RPB^{+}23] Francesco Restuccia, Marco Pagani, Alessandro Biondi, Mauro Marinoni, and Giorgio Buttazzo. Bounding memory access times in multiaccelerator architectures on FPGA SoCs. IEEE Transactions on Computers, 72(1): 154–167, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rapp:2021:NNB

[RPMH21] Martin Rapp, Anuj Pathania, Tulika Mitra, and Jörg Henkel. Neural network-based performance prediction for task migration on S-NUCA manycores. IEEE Transactions on Computers, 70(10):1691– 1704, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rahimi:2021:DSH

[RPS^{+}21] Hamed Rahimi, Yvan Picaud, Kamal Deep Singh, Giyyarpuram Madhusudan, Salvatore Costanzo, and Olivier Boissier. Design and simulation of a hybrid architecture for edge computing in 5G and beyond. IEEE Transactions on Computers, 70(8):

1213–1224, August 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rezaalipour:2020:AMA

[RRDB20] M. Rezaalipour, M. Rezaalipour, M. Dehyadegari, and M. N. Bojnordi. AxMAP: Making approximate adders aware of input patterns. IEEE Transactions on Computers, 69(6):868–882, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ruospo:2025:EIS

[RRMS25] Annachiara Ruospo, Matteo Sonza Reorda, Riccardo Mariani, and Ernesto Sanchez. An effective iterative statistical fault injection methodology for deep neural networks. IEEE Transactions on Computers, 74(7):2431–2444, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Roohi:2020:AAG

[RSA^{+}20] A. Roohi, S. Sheikhfaal, S. Angizi, D. Fan, and R. F. De-Mara. ApGAN: Approximate GAN for robust low energy learning from imprecise components. IEEE Transactions on Computers, 69(3): 349–360, March 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reviriego:2023:APA

[RSMMG^{+}23] Pedro Reviriego, Alfonso Sánchez-Macián, Elena Merino-Gómez, Ori Rottenstreich, Shanshan Liu, and Fabrizio Lombardi. Attacking the privacy of approximate membership check filters by positive concentration. IEEE Transactions on Computers, 72(5): 1409–1419, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rapp:2020:PCA

[RSP^{+}20] M. Rapp, M. Sagi, A. Pathania, A. Herkersdorf, and J. Henkel. Power- and cacheaware task mapping with dynamic power budgeting for many-cores. IEEE Transactions on Computers, 69(1): 1–13, January 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rosenfeld:2022:SGA

[RSR22] Bleema Rosenfeld, Osvaldo Simeone, and Bipin Rajendran. Spiking generative adversarial networks with a neural network discriminator: Local training, Bayesian models, and continual metalearning. IEEE Transactions on Computers, 71(11):2778– 2791, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Reddy:2023:DAR

[RSZ23] Sathi Sarveswara Reddy, Sharad Sinha, and Wei Zhang. Design and analysis of RSA and Paillier homomorphic cryptosystems using PSObased evolutionary computation. IEEE Transactions on Computers, 72(7):1886–1900, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ravipati:2025:CCC

[RvSP^{+}25] Divya Praneetha Ravipati, Victor M. van Santen, Shivendra Singh Parihar, Yogesh Singh Chauhan, Preeti Ranjan Panda, and Hussam Amrouch. Cryo-CACTI: Cryogenic-aware CACTI for cache modeling down to 10K in advanced 7nm FinFETs. IEEE Transactions on Computers, 74(8):2567–2580, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rodrigues:2023:SRM

[RWCC23] Walber M. Rodrigues, Felipe N. Walmsley, George D. C. Cavalcanti, and Rafael M. O. Cruz. Security relevant methods of android’s API classification: a machine learning empirical evaluation. IEEE Transactions on Computers, 72(11):3273–3285, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Segura:2022:EES

[SAG22] Albert Segura, Jose-Maria Arnau, and Antonio González. Energy-efficient stream compaction through filtering and coalescing accesses in GPGPU memory partitions. IEEE Transactions on Computers, 71(7):1711–1723, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Seo:2021:SIK

[SAJA21] Hwajeong Seo, Mila Anastasova, Amir Jalali, and Reza Azarderakhsh. Supersingular Isogeny Key Encapsulation (SIKE) Round 2 on ARM Cortex-M4. IEEE Transactions on Computers, 70(10): 1705–1718, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Singh:2020:CAE

[SBP^{+}20] A. K. Singh, K. R. Basireddy, A. Prakash, G. V. Merrett, and B. M. Al-Hashimi. Collaborative adaptation for energyefficient heterogeneous mobile SoCs. IEEE Transactions on Computers, 69(2): 185–197, February 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Saha:2025:PAR

[SCA^{+}25] Sangeet Saha, Shounak Chakraborty, Sukarn Agarwal, Magnus Själander, and Klaus McDonald-Maier. PRECIOUS: Approximate real-time computing in MLC-MRAM based heterogeneous CMPs. IEEE Transactions on Computers, 74(10): 3476–3489, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sadok:2021:SDC

[SCC21] Hugo Sadok, Miguel Elias M.

Campista, and Luís Henrique M. K. Costa. Stateful DRF: Considering the past in a multi-resource allocation. IEEE Transactions on Computers, 70(7):1094–1105, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Saez:2022:LFL

[SCFPM22] Juan Carlos Saez, Fernando Castro, Graziano Fanizzi, and Manuel Prieto-Matias. LFOC+: a fair OS-level cacheclustering policy for commodity multicore systems. IEEE Transactions on Computers, 71(8):1952–1967, August 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shi:2024:ATC

[SCL^{+}24] Jianqi Shi, Yinghao Chen, Qin Li, Yanhong Huang, Yang Yang, and Mengyan Zhao. Automated test cases generator for IEC 61131-3 structured text based dynamic symbolic execution. IEEE Transactions on Computers, 73(4): 1048–1059, April 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sheng:2021:EBU

[SCY21] F. Sheng, Q. Cao, and J. Yao. Exploiting buffered updates for fast streaming graph analysis. IEEE Transactions on Computers, 70(2):255– 269, February 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2023:ERK

[SCY^{+}23] Yongqian Sun, Daguo Cheng, Tiankai Yang, Yuhe Ji, Shenglin Zhang, Man Zhu, Xiao Xiong, Qiliang Fan, Minghan Liang, Dan Pei, Tianchi Ma, and Yu Chen. Efficient and robust KPI outlier detection for large-scale datacenters. IEEE Transactions on Computers, 72(10): 2858–2871, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Siracusa:2022:CMO

[SDR^{+}22] Marco Siracusa, Emanuele Del Sozzo, Marco Rabozzi, Lorenzo Di Tucci, Samuel Williams, Donatella Sciuto, and Marco Domenico Santambrogio. A comprehensive methodology to optimize FPGA designs via the roofline model. IEEE Transactions on Computers, 71(8):1903– 1915, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shukla:2023:CCI

[SEM23] Aditya Shukla, Mikhail Erementchouk, and Pinaki Mazumder. Custom CMOS Ising machine based on relaxed Burer– Monteiro–Zhang heuristic. IEEE Transactions on Computers, 72(10):2835–2846, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2020:RTS

[SGL^{+}20] J. Sun, N. Guan, F. Li, H. Gao, C. Shi, and W. Yi. Real-time scheduling and analysis of OpenMP DAG tasks supporting nested parallelism. IEEE Transactions on Computers, 69(9):1335– 1348, September 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2021:ACW

[SGS^{+}21] J. Sun, N. Guan, J. Sun, X. Zhang, Y. Chi, and F. Li. Algorithms for computing the WCRT bound of OpenMP task systems with conditional branches. IEEE Transactions on Computers, 70(1): 57–71, January 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2024:LLS

[SHZ^{+}24] Mingyang Song, Zhongyun Hua, Yifeng Zheng, Hejiao Huang, and Xiaohua Jia. LSDedup: Layered secure deduplication for cloud storage. IEEE Transactions on Computers, 73(2):422–435, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2025:EVS

[SHZ^{+}25] Mingyang Song, Zhongyun Hua, Yifeng Zheng, Qing Liao, and Xiaohua Jia. Enabling verifiable search and integrity auditing in encrypted decentralized storage using one proof. IEEE Transactions on Computers, 74(8): 2731–2744, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salamat:2020:AHC

[SIR20] S. Salamat, M. Imani, and T. Rosing. Accelerating hyperdimensional computing on FPGAs by exploiting computational reuse. IEEE Transactions on Computers, 69(8): 1159–1171, August 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2025:RGE

[SJYQ25] Hui Sun, Xiangxiang Jiang, Yinliang Yue, and Xiao Qin. RGKV: a GPGPU-empowered compaction framework for LSM-tree-based KV stores with optimized data transfer and parallel processing. IEEE Transactions on Computers, 74(5):1605–1619, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Surabhi:2022:TDE

[SKA^{+}22] Virinchi Roy Surabhi, Prashanth Krishnamurthy, Hussam Amrouch, Jörg Henkel, Ramesh Karri, and Farshad Khorrami. Trojan detection in embedded systems with FinFET technology. IEEE Transactions on Computers, 71(11):3061– 3071, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shin:2025:AAT

[SKH^{+}25] Jaekang Shin, Myeonggu Kang, Yunki Han, Junyoung Park, and Lee-Sup Kim. AToM: Adaptive token merging for efficient acceleration of vision transformer. IEEE Transactions on Computers, 74(5):1620–1633, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Seol:2021:ADP

[SKK^{+}21] H. Seol, M. Kim, T. Kim, Y. Kim, and L.-S. Kim. Amnesiac DRAM: a proactive defense mechanism against cold boot attacks. IEEE Transactions on Computers, 70(4): 539–551, April 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shin:2023:FFF

[SKK23] Hyein Shin, Myeonggu Kang, and Lee-Sup Kim. Fault-free: a framework for analysis and mitigation of stuck-at-fault on realistic ReRAM-based DNN accelerators. IEEE Transactions on Computers, 72(7): 2011–2024, July 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shamsa:2022:CAB

[SKLR22] Elham Shamsa, Anil Kanduri, Pasi Liljeberg, and Amir M. Rahmani. Concurrent application bias scheduling for energy efficiency of heterogeneous multi-core platforms. IEEE Transactions on Computers, 71(4):743–755, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Symvoulidis:2023:UMB

[SKM^{+}23] Chrysostomos Symvoulidis, Athanasios Kiourtis, George Marinos, Jean-Didier Totow Tom-Ata, George Manias, Argyro Mavrogiorgou, and Dimosthenis Kyriazis. A user mobility-based data placement strategy in a hybrid cloud/edge environment using a causal-aware deep learning network. IEEE Transactions on Computers, 72(12):3603– 3616, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sakalis:2020:USD

[SKR^{+}20] C. Sakalis, S. Kaxiras, A. Ros, A. Jimborean, and M. Själander. Understanding selective delay as a method for efficient secure speculative execution. IEEE Transactions on Computers, 69(11):1584–1595, November 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shi:2023:QSP

[SL23] Run-Hua Shi and Yi-Fei Li.

Quantum secret permutating protocol. IEEE Transactions on Computers, 72(5): 1223–1235, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sui:2025:PTP

[SLDZ25] Xiao Sui, Qichang Liu, Sisi Duan, and Haibin Zhang. Pike: Two-phase BFT with linearity and flexible view change. IEEE Transactions on Computers, 74(8):2772– 2784, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2025:PSF

[SLLS25] Yunpeng Song, Yujiong Liang, Jialin Liu, and Liang Shi. Prophet: SSD failure analysis and prediction guided by flash reliability characteristics in data centers. IEEE Transactions on Computers, 74(8): 2529–2541, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2023:TPS

[SLOM^{+}23] Bosheng Song, Kenli Li, David Orellana-Martín, Xiangxiang Zeng, and Mario J. Pérez-Jiménez. Tissue P systems with states in cells. IEEE Transactions on Computers, 72(9):2561–2570, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2021:CAS

[SLS^{+}21] Zhirong Shen, Shiyao Lin, Jiwu Shu, Chengxin Xie, Zhijie Huang, and Yingxun Fu. Cluster-aware scattered repair in erasure-coded storage: Design and analysis. IEEE Transactions on Computers, 70(11):1861–1874, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shang:2022:OSF

[SLY22a] Xiaojun Shang, Zhenhua Liu, and Yuanyuan Yang. Online service function chain placement for cost-effectiveness and network congestion control. IEEE Transactions on Computers, 71(1):27–39, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2022:IAS

[SLY^{+}22b] Qingxiao Sun, Yi Liu, Hailong Yang, Ming Dun, Zhongzhi Luan, Lin Gan, Guangwen Yang, and Depei Qian. Inputaware sparse tensor storage format selection for optimizing MTTKRP. IEEE Transactions on Computers, 71(8): 1968–1981, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Spliet:2022:PSD

[SM22] Roy Spliet and Robert D.

Mullins. Sim-D: a SIMD accelerator for hard real-time systems. IEEE Transactions on Computers, 71(4):851–865, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sugiura:2024:IFA

[SM24] Keisuke Sugiura and Hiroki Matsutani. An integrated FPGA accelerator for deep learning-based 2D/3D path planning. IEEE Transactions on Computers, 73(6):

1442–1456, June 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Symons:2025:SDS

[SMC^{+}25] Arne Symons, Linyan Mei, Steven Colleman, Pouya Houshmand, Sebastian Karl, and Marian Verhelst. Stream: Design space exploration of layer-fused DNNs on heterogeneous dataflow accelerators. IEEE Transactions on Computers, 74(1):237–249, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salahvarzi:2021:NNO

[SMFS21] A. Salahvarzi, A. M. H. Monazzah, M. Fazeli, and K. Skadron. NOSTalgy: Near-optimum run-time STT-MRAM quality-energy knob management for approximate computing applications. IEEE Transactions on Computers, 70(3):414–427, March 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sa:2022:FLR

[SMP22] Bruno Sá, José Martins, and Sandro Pinto. A first look at RISC-V virtualization from an embedded systems perspective. IEEE Transactions on Computers, 71(9):2177– 2190, September 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shi:2022:AME

[SMY22] Xinming Shi, Leandro L.

Minku, and Xin Yao. Adaptive memory-enhanced time delay reservoir and its memristive implementation. IEEE Transactions on Computers, 71(11):2766–2777, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Servadei:2020:ACE

[SMZ^{+}20] L. Servadei, E. Mosca, E. Zennaro, K. Devarajegowda, M. Werner, W. Ecker, and R. Wille. Accurate cost estimation of memory systems utilizing machine learning and solutions from computer vision for design automation. IEEE Transactions on Computers, 69(6):856–867, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sehatbakhsh:2020:RRE

[SNA^{+}20] N. Sehatbakhsh, A. Nazari, M. Alam, F. Werner, Y. Zhu, A. Zajic, and M. Prvulovic. REMOTE: Robust external malware detection framework by using electromagnetic signals. IEEE Transactions on Computers, 69(3): 312–326, March 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salami:2021:FAE

[SNN21] B. Salami, H. Noori, and M. Naghibzadeh. Fairnessaware energy efficient scheduling on heterogeneous multicore processors. IEEE Transactions on Computers, 70(1): 72–82, January 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

SLPSK:2023:SCR

[SNRB23] Patanjali SLPSK, Abhishek Anil Nair, Chester Rebeiro, and Swarup Bhunia. SIGNED: a challenge-response scheme for electronic hardware watermarking. IEEE Transactions on Computers, 72(6): 1763–1777, June 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Schober:2022:HAM

[SNT22] Peter Schober, M. Hassan Najafi, and Nima TaheriNejad. High-accuracy multiplyaccumulate (MAC) technique for unary stochastic computing. IEEE Transactions on Computers, 71(6):1425–1439, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sutradhar:2023:PPC

[SO23] Kartick Sutradhar and Hari Om. A privacy-preserving comparison protocol. IEEE Transactions on Computers, 72(6):1815–1821, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Seyoum:2021:STO

[SPB^{+}21] Biruk Seyoum, Marco Pagani, Alessandro Biondi, Sara Balleri, and Giorgio Buttazzo. Spatio-temporal optimization of deep neural networks for reconfigurable FPGA SoCs. IEEE Transactions on Computers, 70(11): 1988–2000, November 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Silvestri:2022:ERM

[SPDQ22] Emiliano Silvestri, Alessandro Pellegrini, Pierangelo Di Sanzo, and Francesco Quaglia. Effective runtime management of tasks and priorities in GNU OpenMP applications. IEEE Transactions on Computers, 71(10):2632–2645, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shi:2023:WGM

[SPH^{+}23] Xuanhua Shi, Xuan Peng, Ligang He, Yunfei Zhao, and Hai Jin. Waterwave: a GPU memory flow engine for concurrent DNN training. IEEE Transactions on Computers, 72(10): 2938–2950, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sousa:2020:TIR

[SPMP20] L. Sousa, R. Paludo, P. Martins, and H. Pettenghi. Towards the integration of reverse converters into the RNS channels. IEEE Transactions on Computers, 69(3): 342–348, March 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Simon:2020:BCC

[SQR^{+}20] W. A. Simon, Y. M. Qureshi, M. Rios, A. Levisse, M. Zapater, and D. Atienza. BLADE: An in-cache computing architecture for edge devices. IEEE Transactions on Computers, 69(9):1349–1363, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

SLPSK:2023:TSA

[SRB23] Patanjali SLPSK, Sandip Ray, and Swarup Bhunia. TREEHOUSE: a secure asset management infrastructure for protecting 3DIC designs. IEEE Transactions on Computers, 72(8):2306– 2320, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salamin:2021:PEH

[SRP^{+}21] Sami Salamin, Martin Rapp, Anuj Pathania, Arka Maity, Jörg Henkel, Tulika Mitra, and Hussam Amrouch. Powerefficient heterogeneous manycore design with NCFET technology. IEEE Transactions on Computers, 70(9):1484– 1497, September 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2025:IUS

[SSCK25] Wenhao Sun, Wendi Sun, Song Chen, and Yi Kang. IOPS: a unified SpMM accelerator based on inner-outer-hybrid product. IEEE Transactions on Computers, 74(7): 2210–2222, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sharma:2021:MCM

[SSJ21] Mayank Sharma, Sumit Soman, and Jayadeva. Minimal complexity machines under weight quantization. IEEE Transactions on Computers, 70(8):1189–1198, August 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Senapati:2022:PPA

[SSK22] Debabrata Senapati, Arnab Sarkar, and Chandan Karfa. PRESTO: a penalty-aware real-time scheduler for task graphs on heterogeneous platforms. IEEE Transactions on Computers, 71(2):421– 435, February 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shepherd:2023:IBB

[SSM23] Carlton Shepherd, Benjamin Semal, and Konstantinos Markantonakis. Investigating black-box function recognition using hardware performance counters. IEEE Transactions on Computers, 72(7): 2065–2079, July 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sadhukhan:2024:VEF

[SSP^{+}24] Rajat Sadhukhan, Sayandeep Saha, Sudipta Paria, Swarup Bhunia, and Debdeep Mukhopadhyay. VALIANT: an EDA flow for side-channel leakage evaluation and tailored protection. IEEE Transactions on Computers, 73(2): 436–450, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2024:MLE

[SST^{+}24] Hui Sun, Chen Sun, Haoqiang Tong, Yinliang Yue, and Xiao Qin. A machine learning-empowered cache management scheme for highperformance SSDs. IEEE Transactions on Computers, 73(8):2066–2080, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2024:BBD

[SSW^{+}24] Jiahao Shen, Hao Sheng, Shuai Wang, Ruixuan Cong, Da Yang, and Yang Zhang. Blockchain-based distributed multiagent reinforcement learning for collaborative multiobject tracking framework. IEEE Transactions on Computers, 73(3):778–788, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shao:2021:TDM

[SSY^{+}21] Huajie Shao, Dachun Sun, Shuochao Yao, Lu Su, Zhibo Wang, Dongxin Liu, Shengzhong Liu, Lance Kaplan, and Tarek Abdelzaher. Truth discovery with multi-modal data in social sensing. IEEE Transactions on Computers, 70(9): 1325–1337, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2020:APM

[SSZ^{+}20] J. Sun, G. Sun, S. Zhan, J. Zhang, and Y. Chen. Automated performance modeling of HPC applications using machine learning. IEEE Transactions on Computers, 69(5):749–763, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.
org/document/8956059.

Shirai:2023:MSF

[ST23a] Tatsuhiko Shirai and Nozomu Togawa. Multi-spin-flip engineering in an Ising machine. IEEE Transactions on Computers, 72(3):759–771, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shirai:2023:SVR

[ST23b] Tatsuhiko Shirai and Nozomu Togawa. Spin-variable reduction method for handling linear equality constraints in Ising machines. IEEE Transactions on Computers, 72(8): 2151–2164, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shirazi:2023:ERR

[STK23] Mahmoud Shirazi, Lothar Thiele, and Mehdi Kargahi. Energy-resilient realtime scheduling. IEEE Transactions on Computers, 72(1): 69–81, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2024:EGS

[STQ^{+}24] Binqi Sun, Mirco Theile, Ziyuan Qin, Daniele Bernardini, Debayan Roy, Andrea Bastoni, and Marco Caccamo. Edge generation scheduling for DAG tasks using deep reinforcement learning. IEEE Transactions on Computers, 73(4):1034–1047, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shao:2021:NOH

[STW^{+}21] En Shao, Guangming Tan, Zhan Wang, Guojun Yuan, Zheng Cao, and Ninghui Sun. A new optoelectronic hybrid network based on scheduling optimization of optical links. IEEE Transactions on Computers, 70(6):863–876, June 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2024:LWI

[STYQ24] Hui Sun, Haoqiang Tong, Yinliang Yue, and Xiao Qin. LAC: a workload intensityaware caching scheme for highperformance SSDs. IEEE Transactions on Computers, 73(7):1738–1752, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2024:AAM

[STZ^{+}24] Zhaoyan Shen, Qingxiang Tang, Tianren Zhou, Yuhao Zhang, Zhiping Jia, Dongxiao Yu, Zhiyong Zhang, and Bingzhe Li. ASHL: an adaptive multi-stage distributed deep learning training scheme for heterogeneous environments. IEEE Transactions on Computers, 73(1):30–43, January 2024. CODEN ITCOB4.

ISSN 0018-9340 (print), 1557-9956 (electronic).

Slimani:2023:ARF

[SWR^{+}23] Camélia Slimani, Chun-Feng Wu, Stéphane Rubini, Yuan-Hao Chang, and Jalil Boukhobza. Accelerating random forest on memory-constrained devices through data storage optimization. IEEE Transactions on Computers, 72(6): 1595–1609, June 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sirri:2025:ESS

[SWR^{+}25] Scott Sirri, Zhe Wang, Netanel Raviv, Jeremy Fineman, and Kunal Agrawal. Efficient static schedules for fault-tolerant transmissions on shared media. IEEE Transactions on Computers, 74(9): 2882–2895, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Su:2024:LIE

[SXC^{+}24] Chao Su, Xiaoshuang Xing, Xiaolu Cheng, Rui Guo, and Chuanwen Luo. LPAH: Illustrating efficient live patching with alignment holes in kernel data. IEEE Transactions on Computers, 73(10): 2434–2448, October 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2024:RSA

[SXH^{+}24] Wei Song, Zihan Xue, Jinchi Han, Zhenzhen Li, and Peng Liu. Randomizing set-associative caches against conflict-based cache sidechannel attacks. IEEE Transactions on Computers, 73(4): 1019–1033, April 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Song:2024:PTC

[SXXL24] Wei Song, Da Xie, Zihan Xue, and Peng Liu. A parallel tag cache for hardware managed tagged memory in multicore processors. IEEE Transactions on Computers, 73(11): 2488–2503, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shi:2024:RCP

[SXZJ24] Zhanhui Shi, Jie Xiao, Weidong Zhu, and Jianhui Jiang. A reliability-critical path identifying method with local and global adjacency probability matrix in combinational circuits. IEEE Transactions on Computers, 73(1): 123–137, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2024:LGH

[SYD^{+}24] Shiyu Shen, Hao Yang, Wangchen Dai, Lu Zhou, Zhe Liu, and Yunlei Zhao. Leveraging GPU in homomorphic encryption: Framework design and analysis of BFV variants. IEEE Transactions on Computers, 73(12):2817–2829, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shen:2023:CCA

[SYL^{+}23] Shiyu Shen, Hao Yang, Yu Liu, Zhe Liu, and Yunlei Zhao. CARM: CUDA-Accelerated RNS Multiplication in word-wise homomorphic encryption schemes for Internet of Things. IEEE Transactions on Computers, 72(7):1999–2010, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2022:MNA

[SYW^{+}22] Gongjian Sun, Mingyu Yan, Duo Wang, Han Li, Wenming Li, Xiaochun Ye, Dongrui Fan, and Yuan Xie. Multinode acceleration for largescale GCNs. IEEE Transactions on Computers, 71(12): 3140–3152, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Stauffer:2022:PSS

[SZ22] Jake Stauffer and Qingxue Zhang. SpikeBASE: Spiking neural learning algorithm with backward adaptation of synaptic efflux. IEEE Transactions on Computers, 71(11): 2707–2716, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sadiqbatcha:2022:RTF

[SZAT22] Sheriff Sadiqbatcha, Jinwei Zhang, Hussam Amrouch, and Sheldon X.-D. Tan. Real-time full-chip thermal tracking: a post-silicon, machine learning perspective. IEEE Transactions on Computers, 71(6): 1411–1424, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Schuiki:2021:SSR

[SZHB21] F. Schuiki, F. Zaruba, T. Hoefler, and L. Benini. Stream semantic registers: A lightweight RISC-V ISA extension achieving full compute utilization in single-issue cores. IEEE Transactions on Computers, 70(2):212–227, February 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Salamin:2022:INT

[SZK^{+}22] Sami Salamin, Georgios Zervakis, Florian Klemme, Hammam Kattan, Yogesh Chauhan, Jörg Henkel, and Hussam Amrouch. Impact of NCFET technology on eliminating the cooling cost and boosting the efficiency of Google TPU. IEEE Transactions on Computers, 71(4):906–918, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sha:2022:AAT

[SZL^{+}22] Sai Sha, Yi Zhang, Yingwei Luo, Xiaolin Wang, and Zhenlin Wang. Accelerating address translation for virtualization by leveraging hardware mode. IEEE Transactions on Computers, 71(11):3047– 3060, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sun:2024:BBM

[SZL^{+}24] Wenhao Sun, Zhiwei Zou, Deng Liu, Wendi Sun, Song Chen, and Yi Kang. Bitbalance: Model-hardware codesign for accelerating NNs by exploiting bit-level sparsity. IEEE Transactions on Computers, 73(1):152–163, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Su:2022:DOM

[SZS^{+}22] Ya Su, Youjian Zhao, Ming Sun, Shenglin Zhang, Xidao Wen, Yongsu Zhang, Xian Liu, Xiaozhou Liu, Junliang Tang, Wenfei Wu, and Dan Pei. Detecting outlier machine instances through Gaussian mixture variational autoencoder with one dimensional CNN. IEEE Transactions on Computers, 71(4): 892–905, April 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Titopoulos:2025:OSS

[TAP^{+}25] Vasileios Titopoulos, Kosmas Alexandridis, Christodoulos Peltekis, Chrysostomos Nicopoulos, and Giorgos Dimitrakopoulos. Optimizing structured-sparse matrix multiplication in RISC-V vector processors. IEEE Transactions on Computers, 74(4): 1446–1460, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tahir:2023:LBC

[TARK23] Anam Tahir, Bastian Alt, Amr Rizk, and Heinz Koeppl. Load balancing in compute clusters with delayed feedback. IEEE Transactions on Computers, 72(6):1610–1622, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tamimi:2025:CUC

[TBS^{+}25] Sajjad Tamimi, Arthur Bernhardt, Florian Stock, Ilia Petrov, and Andreas Koch. CINDA: Using cache-coherent interconnects for accelerating databases by enabling neardata processing of update transactions. IEEE Transactions on Computers, 74(7): 2238–2252, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tan:2021:OSE

[TC21] Bochen Tan and Jason Cong. Optimality study of existing quantum computing layout synthesis tools. IEEE Transactions on Computers, 70(9): 1363–1373, September 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tuli:2023:SCR

[TCJ23] Shreshth Tuli, Giuliano Casale, and Nicholas R. Jennings. SciNet: Codesign of resource management in cloud computing environments. IEEE Transactions on Computers, 72(12):3590–3602, December 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tao:2023:BRF

[TCX^{+}23] Youming Tao, Sijia Cui, Wenlu Xu, Haofei Yin, Dongxiao Yu, Weifa Liang, and Xiuzhen Cheng. Byzantineresilient federated learning at edge. IEEE Transactions on Computers, 72(9):2600– 2614, September 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tran:2023:TFP

[TDH^{+}23] Thai-Ha Tran, Ba-Anh Dao, Trong-Thuc Hoang, Van-Phuc Hoang, and Cong-Kha Pham. Transition factors of power consumption models for CPA attacks on cryptographic RISC-V SoC. IEEE Transactions on Computers, 72(9): 2689–2700, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tapwal:2023:SBV

[TDMP23] Riya Tapwal, Pallav Kumar Deb, Sudip Misra, and Surjya Kanta Pal. Shadows: Blockchain virtualization for interoperable computations in IIoT environments. IEEE Transactions on Computers, 72(3):868–879, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2022:AAE

[TDZ^{+}22] Zhuo Tang, Lifan Du, Xuedong Zhang, Li Yang, and Kenli Li. AEML: an acceleration engine for multi-GPU load-balancing in distributed heterogeneous environment. IEEE Transactions on Computers, 71(6): 1344–1357, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2025:LRA

[TDZ^{+}25] Dan Tang, Rui Dai, Chenguang Zuo, Jingwen Chen, Keqin Li, and Zheng Qin. A lowrate DoS attack mitigation scheme based on port and traffic state in SDN. IEEE Transactions on Computers, 74(5): 1758–1770, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Thomann:2023:HSC

[TGA23] Simon Thomann, Paul R. Genssler, and Hussam Amrouch. HW/SW co-design for reliable TCAM-based inmemory brain-inspired hyperdimensional computing. IEEE Transactions on Computers, 72(8):2404–2417, August 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2022:WOM

[TGS^{+}22] Yufeng Tang, Zheng Gong, Tao Sun, Jinhai Chen, and Zhe Liu. WBMatrix: an optimized matrix library for white-box block cipher implementations. IEEE Transactions on Computers, 71(12): 3375–3388, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2023:CCP

[TJG^{+}23] Yue Tang, Xu Jiang, Nan Guan, Dong Ji, Xiantong Luo, and Wang Yi. Comparing communication paradigms in cause-effect chains. IEEE Transactions on Computers, 72(1):82–96, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsukada:2020:NNB

[TKM20] M. Tsukada, M. Kondo, and H. Matsutani. A neural network-based on-device learning anomaly detector for edge devices. IEEE Transactions on Computers, 69(7): 1027–1044, July 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Taheri:2023:AAC

[TKN23] Ebadollah Taheri, Ryan Gary Kim, and Mahdi Nikdast. AdEle+: an adaptive congestionand-energy-aware elevator selection for partially connected 3D networks-on-chip. IEEE Transactions on Computers, 72(8):2278–2292, August 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tan:2024:RPS

[TLC^{+}24] Wende Tan, Chenyang Li, Yangyu Chen, Yuan Li, Chao Zhang, and Jianping Wu. ROLoad-PMP: Securing sensitive operations for kernels and bare-metal firmware. IEEE Transactions on Computers, 73(12):2722–2733, 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tian:2025:CTA

[TLLL25] Minghao Tian, Yue Liang, Bowen Liu, and Dajiang Liu. CoSpMV: Towards agile software and hardware co-design for SpMV computation. IEEE Transactions on Computers, 74(6):1921–1935, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tosoni:2024:MFM

[TOF^{+}24] Francesco Tosoni, Nicola Dall Ora, Enrico Fraccaroli, Sara Vinco, and Franco Fummi. Multidomain fault models covering the analog side of a smart or cyber physical system. IEEE Transactions on Computers, 73(3): 829–841, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Towhidy:2023:DIA

[TOM23] Ahmad Towhidy, Reza Omidi, and Karim Mohammadi. On the design of iterative approximate floating-point multipliers. IEEE Transactions on Computers, 72(6):1623–1635, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tan:2023:SAQ

[TPWY23] Jingweijia Tan, Liqi Ping, Qixiang Wang, and Kaige Yan. Saca-AVF: a quantitative approach to analyze the architectural vulnerability factors of CNN accelerators. IEEE Transactions on Computers, 72(11):3042–3056, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2022:VRE

[TQL^{+}22] Yutao Tang, Zhengrui Qin, Zhiqiang Lin, Yue Li, Shanhe Yi, Fengyuan Xu, and Qun Li. vTrust: Remotely executing mobile apps transparently with local untrusted OS. IEEE Transactions on Computers, 71(12):3349–3360, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tarapore:2022:OIL

[TRBM22] Dharmesh Tarapore, Shahin Roozkhosh, Steven Brzozowski, and Renato Mancuso. Observing the invisible: Live cache inspection for high-performance embedded systems. IEEE Transactions on Computers, 71(3): 559–572, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Takeshita:2024:AFF

[TRG^{+}24] Jonathan Takeshita, Dayane Reis, Ting Gong, Michael T. Niemier, Xiaobo Sharon Hu, and Taeho Jung. Accelerating finite-field and torus fully homomorphic encryption via compute-enabled (s)RAM. IEEE Transactions on Computers, 73(10):2449–2462, October 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Turan:2020:HAH

[TRV20] F. Turan, S. S. Roy, and I. Verbauwhede. HEAWS: An accelerator for homomorphic encryption on the Amazon AWS FPGA. IEEE Transactions on Computers, 69(8): 1185–1196, August 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Talebi:2021:RRH

[TSM^{+}21] Mahdi Talebi, Arash Salahvarzi, Amir Mahdi Hosseini Monazzah, Kevin Skadron, and Mahdi Fazeli. ROCKY: A robust hybrid on-chip memory kit for the processors with STT-MRAM cache technology. IEEE Transactions on Computers, 70(12):2198– 2210, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tsiokanos:2023:AAE

[TTG^{+}23] Ioannis Tsiokanos, Styliani Tompazi, Giorgis Georgakoudis, Lev Mukhanov, and Georgios Karakonstantis. ARETE: Accurate error assessment via machine learning-guided dynamic-timing analysis. IEEE Transactions on Computers, 72(4):1026–1040, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tian:2023:VVR

[TWaKo^{+}23] Guohua Tian, Jianghong Wei, Miros aw Kuty owski, Willy Susilo, Xinyi Huang, and Xiaofeng Chen. VRBC: a verifiable redactable blockchain with efficient query and integrity auditing. IEEE Transactions on Computers, 72(7): 1928–1942, July 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tu:2022:HSD

[TWJ^{+}22] Kuan-Hua Tu, Hung-En Wang, Jie-Hong R. Jiang, Natalia Kushik, and Nina Yevtushenko. Homing sequence derivation with quantified Boolean satisfiability. IEEE Transactions on Computers, 71(3):696–711, March 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tian:2022:ESI

[TWL^{+}22] Jing Tian, Piaoyang Wang, Zhe Liu, Jun Lin, Zhongfeng Wang, and Johann Großschädl. Efficient software implementation of the SIKE protocol using a new data representation. IEEE Transactions on Computers, 71(3):670–683, March 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tan:2025:EGI

[TWY^{+}25] Jingweijia Tan, Jiashuo Wang, Kaige Yan, Xiaohui Wei, and Xin Fu. Evaluating GPU’s instruction-level error characteristics under low supply voltages. IEEE Transactions on Computers, 74(2): 555–568, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tan:2023:HSV

[TWZ^{+}23] Weihang Tan, Antian Wang, Xinmiao Zhang, Yingjie Lao, and Keshab K. Parhi. Highspeed VLSI architectures for modular polynomial multiplication via fast filtering and applications to lattice-based cryptography. IEEE Transactions on Computers, 72(9): 2454–2466, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2022:FDR

[TZ22] Yok Jye Tang and Xinmiao Zhang. Fast en/decoding of Reed–Solomon codes for failure recovery. IEEE Transactions on Computers, 71(3): 724–735, March 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tan:2024:PAA

[TZY^{+}24] Kai Tan, Dongyang Zhan, Lin Ye, Hongli Zhang, and Binxing Fang. A practical adversarial attack against sequencebased deep learning malware classifiers. IEEE Transactions on Computers, 73(3): 708–721, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Tang:2021:FLE

[TZZ^{+}21] Xiongchao Tang, Chen Zhang, Jidong Zhai, Xuehai Qian, Wenguang Chen, and Yong Jiang. A fast lock for explicit message passing architectures. IEEE Transactions on Computers, 70(10):1555–1568, October 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ueter:2023:PPP

[UGvdBC23] Niklas Ueter, Mario Günzel, Georg von der Brüggen, and Jian-Jia Chen. Parallel path progression DAG scheduling. IEEE Transactions on Computers, 72(10):3002–3016, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ueno:2020:HTG

[UMM^{+}20] R. Ueno, S. Morioka, N. Miura, K. Matsuda, M. Nagata, S. Bhasin, Y. Mathieu, T. Graba, J. Danger, and N. Homma. High throughput/ gate AES hardware architectures based on datapath compression. IEEE Transactions on Computers, 69(4):534–548, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ullah:2021:AOA

[USS^{+}21] S. Ullah, H. Schmidl, S. S. Sahoo, S. Rehman, and A. Kumar. Area-optimized accurate and approximate softcore signed multiplier architectures. IEEE Transactions on Computers, 70(3): 384–392, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ugurlu:2022:PPB

[UYZP22] Elvan Mert Ugurlu, Baki Berkay Yilmaz, Alenka Zajić, and Milos Prvulovic. PITEM: Permutations-based instruction tracking via electromagnetic side-channel signal analysis. IEEE Transactions on Computers, 71(5):1156–1169, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vasan:2020:MCA

[VAV^{+}20] D. Vasan, M. Alazab, S. Venkatraman, J. Akram, and Z. Qin. MTHAEL: Cross-architecture IoT malware detection based on neural network advanced ensemble learning. IEEE Transactions on Computers, 69(11):1654–1667, November 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Varsamopoulos:2020:CNN

[VBA20] S. Varsamopoulos, K. Bertels, and C. G. Almudever. Comparing neural network based decoders for the surface code. IEEE Transactions on Computers, 69(2):300–311, February 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vandebon:2021:EHL

[VCLN21] Jessica Vandebon, Jose G. F.

Coutinho, Wayne Luk, and Eriko Nurvitadhi. Enhancing high-level synthesis using a meta-programming approach. IEEE Transactions on Computers, 70(12):2043– 2055, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Volkova:2020:AAR

[VHL20] A. Volkova, T. Hilaire, and C. Lauter. Arithmetic approaches for rigorous design of reliable fixed-point LTI filters. IEEE Transactions on Computers, 69(4):489–504, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vyas:2021:KMM

[VJWZ^{+}21] V. Vyas, L. Jiang-Wei, P. Zhou, X. Hu, and J. S. Friedman. Karnaugh map method for memristive and spintronic asymmetric basis logic functions. IEEE Transactions on Computers, 70(1):

128–138, January 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vairam:2022:FBF

[VKRK22] Prasanna Karthik Vairam, Pratyush Kumar, Chester Rebeiro, and V. Kamakoti. FadingBF: a Bloom filter with consistent guarantees for online applications. IEEE Transactions on Computers, 71(1): 40–52, January 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Valero:2025:DFT

[VLPS25] Alejandro Valero, Vicente Lorente, Salvador Petit, and Julio Sahuquillo. Dual fasttrack cache: Organizing ringshaped racetracks to work as L1 caches. IEEE Transactions on Computers, 74(8): 2812–2826, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Valente:2024:TTO

[VRR^{+}24] Luca Valente, Francesco Restuccia, Davide Rossi, Ryan Kastner, and Luca Benini. TOP: Towards open & predictable heterogeneous SoCs. IEEE Transactions on Computers, 73(12):2678–2692, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

vanSanten:2023:MPC

[vSDHA23] Victor M. van Santen, Fu Lam Florian Diep, Jörg Henkel, and Hussam Amrouch. Massively parallel circuit setup in GPU-SPICE. IEEE Transactions on Computers, 72(8):2127– 2138, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vaidhun:2023:PMC

[VSG^{+}23] Sudharsan Vaidhun, Tianning She, Qijun Gu, Sajal K. Das, Kecheng Yang, and Zhishan Guo. Precise mixed-criticality scheduling on varying-speed multiprocessors. IEEE Transactions on Computers, 72(1): 43–54, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Vasselle:2020:LIF

[VTM^{+}20] Aurélien Vasselle, Hugues Thiebeauld, Quentin Maouhoub, Adèle Morisset, and Sébastien Ermeneux. Laser-induced fault injection on smartphone bypassing the secure bootextended version. IEEE Transactions on Computers, 69(10):1449–1459, October 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2025:QTL

[WAN^{+}25] Kaijie Wei, Hideharu Amano, Ryohei Niwase, Yoshiki Yamaguchi, and Takefumi Miyoshi. Qu-Trefoil: Large-scale quantum circuit simulator working on FPGA with SATA storages. IEEE Transactions on Computers, 74(4):1306–1321, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:CTB

[WBJC22] Chundong Wang, Gunavaran Brihadiswarn, Xingbin Jiang, and Sudipta Chattopadhyay. Circ-Tree: a B+-tree variant with circular design for persistent memory. IEEE Transactions on Computers, 71(2): 296–308, February 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:ACB

[WCB23] Xiaogang Wang, Jian Cao, and Rajkumar Buyya. Adaptive cloud bundle provisioning and multi-workflow scheduling via coalition reinforcement learning. IEEE Transactions on Computers, 72(4): 1041–1054, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2025:CCP

[WCD25] Yixun Wei, Zhichao Cao, and David H. C. Du. CPI: a collaborative partial indexing design for large-scale deduplication systems. IEEE Transactions on Computers, 74(2):

483–494, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:ZSG

[WCL^{+}23] Jinzhen Wang, Qi Chen, Tong Liu, Qing Liu, and Xubin He. zPerf: a statistical gray-box approach to performance modeling and extrapolation for scientific lossy compression. IEEE Transactions on Computers, 72(9): 2641–2655, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:CCI

[WCQW22] Yifeng Wang, Baolei Cheng, Yu Qian, and Dajin Wang. Constructing completely independent spanning trees in a family of line-graph-based data center networks. IEEE Transactions on Computers, 71(5):1194–1203, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2020:JMC

[WCYK20] C. Wu, Y. Chang, M. Yang, and T. Kuo. Joint management of CPU and NVDIMM for breaking down the Great Memory Wall. IEEE Transactions on Computers, 69(5): 722–733, May 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://

                                        ieeexplore.ieee.org/document/
                                        8950228.

Wang:2024:TSR

[WCZ^{+}24] Yanling Wang, Xiaolin Chang, Haoran Zhu, Jianhua Wang, Yanwei Gong, and Lin Li. Towards secure runtime customizable trusted execution environment on FPGA-SoC. IEEE Transactions on Computers, 73(4):1138–1151, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:CEG

[WDC^{+}25] Ziheng Wang, Xiaoshe Dong, Heng Chen, Yan Kang, and Qiang Wang. CUSPX: Efficient GPU implementations of post-quantum signature SPHINCS+. IEEE Transactions on Computers, 74(1): 15–28, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2020:LLF

[WDCC20] E. Wang, J. J. Davis, P. Y. K. Cheung, and G. A. Constantinides. LUTNet: Learning FPGA configurations for highly efficient neural network inference. IEEE Transactions on Computers, 69(12):1795– 1808, December 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:CAH

[WDL^{+}25] Haotian Wang, Yan Ding, Yumeng Liu, Weichen Liu, Chubo Liu, Wangdong Yang, and Kenli Li. A contextawareness and hardwarefriendly sparse matrix multiplication kernel for CNN inference acceleration. IEEE Transactions on Computers, 74(4):1182–1195, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:EPS

[WDQ^{+}22] Bangyan Wang, Lei Deng, Zheng Qu, Shuangchen Li, Zheng Zhang, and Yuan Xie. Efficient processing of sparse tensor decomposition via unified abstraction and PEinteractive architecture. IEEE Transactions on Computers, 71(2):266–281, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2023:POC

[WDW^{+}23] Jiashu Wu, Hao Dai, Yang Wang, Yong Zhang, Dong Huang, and Chengzhong Xu. PackCache: an online costdriven data caching algorithm in the cloud. IEEE Transactions on Computers, 72(4): 1208–1214, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2022:DEC

[WDZ^{+}22] Suzhen Wu, Chunfeng Du, Weiwei Zhang, Bo Mao, and Hong Jiang. DedupHR: Exploiting content locality to alleviate read/write interference in deduplication-based flash storage. IEEE Transactions on Computers, 71(6): 1332–1343, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2023:EEA

[WDZ^{+}23] Suzhen Wu, Chunfeng Du, Weidong Zhu, Jindong Zhou, Hong Jiang, Bo Mao, and Lingfang Zeng. EaD: ECC-Assisted deduplication with high performance and low memory overhead for ultralow latency flash storage. IEEE Transactions on Computers, 72(1):208–221, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2024:FBS

[WDZ^{+}24] Zhaorui Wu, Yuhui Deng, Yi Zhou, Jie Li, Shujie Pang, and Xiao Qin. FaaSBatch: Boosting serverless efficiency with in-container parallelism and resource multiplexing. IEEE Transactions on Computers, 73(4):1071–1085, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:DHS

[WFH^{+}24] Yueyao Wang, Samuel Furman, Nicolas Hardy, Margaret Ellis, Godmar Back, Yili Hong, and Kirk Cameron. A detailed historical and statistical analysis of the influence of hardware artifacts on SPEC integer benchmark performance. IEEE Transactions on Computers, 73(5): 1262–1274, May 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:DDE

[WFH25] Zhican Wang, Hongxiang Fan, and Guanghui He. DESA: Dataflow efficient systolic array for acceleration of transformers. IEEE Transactions on Computers, 74(6): 2058–2072, June 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wen:2025:TER

[WFL^{+}25] Xiaoqing Wen, Quanbi Feng, Hanzheng Lyu, Jianyu Niu, Yinqian Zhang, and Chen Feng. TeeRollup: Efficient rollup design using heterogeneous TEE. IEEE Transactions on Computers, 74(10): 3546–3558, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2021:IWP

[WFT^{+}21] C. Wang, D. Feng, W. Tong, J. Liu, B. Wu, W. Zhao, Y. Zhang, and Y. Chen. Improving write performance on cross-point RRAM arrays by leveraging multidimensional non-uniformity of cell effective voltage. IEEE Transactions on Computers, 70(4):566–580, April 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2020:EEE

[WFW^{+}20] J. Wang, X. Fu, X. Wang, S. Liu, L. Gao, and W. Zhang. Enabling energy-efficient and reliable neural network via neuron-level voltage scaling. IEEE Transactions on Computers, 69(10):1460–1473, October 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:WWL

[WFZ^{+}25] Chong Wang, Wanyi Fu, Jiangwei Zhang, Shiyao Li, Rui Hou, Jian Yang, and Yu Wang. WOLF: Weightlevel OutLier and fault integration for reliable LLM deployment. IEEE Transactions on Computers, 74(10):3390– 3403, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wen:2022:EOS

[WGD^{+}22] Yuanbo Wen, Qi Guo, Zidong Du, Jianxing Xu, Zhenxing Zhang, Xing Hu, Wei Li, Rui Zhang, Chao Wang, Xuehai Zhou, and Tianshi Chen. Enabling one-size-fits-all compilation optimization for inference across machine learning computers. IEEE Transactions on Computers, 71(9): 2313–2326, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2021:FBA

[WGJZ21] Chao Wang, Lei Gong, Fahui Jia, and Xuehai Zhou. An FPGA based accelerator for clustering algorithms with custom instructions. IEEE Transactions on Computers, 70(5):725–732, May 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2020:FSA

[WGL^{+}20] T. Wang, T. Geng, A. Li, X. Jin, and M. Herbordt. FPDeep: Scalable acceleration of CNN training on deeply-pipelined FPGA clusters. IEEE Transactions on Computers, 69(8):1143– 1158, August 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2020:WUA

[WGM^{+}20] C. Wang, L. Gong, X. Ma, X. Li, and X. Zhou. WooKong: A ubiquitous accelerator for recommendation algorithms with custom instruction sets on FPGA. IEEE Transactions on Computers, 69(7): 1071–1082, July 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:HAG

[WGT^{+}22] Haozhao Wang, Song Guo, Bin Tang, Ruixuan Li, Yutong Yang, Zhihao Qu, and Yi Wang. Heterogeneity-aware gradient coding for tolerating and leveraging stragglers. IEEE Transactions on Computers, 71(4):779–794, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2020:SEC

[WHC20] X. Wang, F. Huang, and H. Chen. Secure and efficient control data isolation with register-based data cloaking. IEEE Transactions on Computers, 69(2):226–238, February 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:DTC

[WHC^{+}23] Xiaohang Wang, Hengli Huang, Ruolin Chen, Yingtao Jiang, Amit Kumar Singh, Mei Yang, and Letian Huang. Detection of thermal covert channel attacks based on classification of components of the thermal signal features. IEEE Transactions on Computers, 72(4):

971–983, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2024:CCD

[WHK24] Wan-Ling Wu, Jen-Wei Hsieh, and Hao-Yu Ku. CDS: Coupled data storage to enhance read performance of 3D TLC NAND flash memory. IEEE Transactions on Computers, 73(3):694–707, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2021:SSA

[WHL^{+}21] W. Wu, L. He, W. Lin, R. Mao, C. Maple, and S. Jarvis. SAFA: a semiasynchronous protocol for fast federated learning with low overhead. IEEE Transactions on Computers, 70(5):655–668, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:RBL

[WHL^{+}23] Wei-Chen Wang, Chien-Chung Ho, Yung-Chun Li, Liang-Chi Chen, and Yu-Ming Chang. Reaping both latency and reliability benefits with elaborate sanitization design for 3D TLC NAND flash. IEEE Transactions on Computers, 72(11):3029– 3041, November 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:DTO

[WHM^{+}22] Jin Wang, Jia Hu, Geyong Min, Wenhan Zhan, Albert Y. Zomaya, and Nektarios Georgalas. Dependent task offloading for edge computing based on deep reinforcement learning. IEEE Transactions on Computers, 71(10):2449– 2461, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:MAA

[WHQ^{+}24] Juzhen Wang, Yiqi Hu, Yiren Qi, Ziwen Peng, and Changjia Zhou. Mitigating adversarial attacks based on denoising & reconstruction with finance authentication system case study. IEEE Transactions on Computers, 73(2): 314–326, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:HHD

[WHX^{+}25] Jinquan Wang, Zhisheng Huo, Limin Xiao, Jinqian Yang, Jiantong Huo, and Minyi Guo. Hierarchical hashing: a dynamic hashing method with low write amplification and high performance for nonvolatile memory. IEEE Transactions on Computers, 74(4): 1138–1151, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2022:EBF

[WHY^{+}22] Yuhan Wu, Jintao He, Shen Yan, Jianyu Wu, Tong Yang, Olivier Ruas, Gong Zhang, and Bin Cui. Elastic Bloom filter: Deletable and expandable filter using elastic fingerprints. IEEE Transactions on Computers, 71(4):984–991, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2020:MTM

[WJL^{+}20] C. Wu, C. Ji, Q. Li, C. Gao, R. Pan, C. Fu, L. Shi, and C. J. Xue. Maximizing I/O throughput and minimizing performance variation via reinforcement learning based I/O merging for SSDs. IEEE Transactions on Computers, 69(1):72–86, January 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:GFE

[WJLC24] Ting Wang, Xin Jiang, Qin Li, and Haibin Cai. GreedW: a flexible and efficient decentralized framework for distributed machine learning. IEEE Transactions on Computers, 73(3):801–814, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2020:AVC

[WL20] B. Wang and Z. Lu. Advance virtual channel reservation. IEEE Transactions on Computers, 69(9):1320– 1334, September 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:AAP

[WL24] Yun Wang and Qiang Liu. AQA: an adaptive posttraining quantization method for activations of CNNs. IEEE Transactions on Computers, 73(8):2025–2035, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2024:TSF

[WLC^{+}24] Mingyu Wu, Zhe Li, Haibo Chen, Binyu Zang, Shaojun Wang, Lei Yu, Sanhong Li, and Haitao Song. Toward an SGX-friendly Java runtime. IEEE Transactions on Computers, 73(1):44–57, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:DAA

[WLD^{+}22] Yang Wang, Min Li, Hao Dai, Kenneth B. Kent, Kejiang Ye, and Chengzhong Xu. Deadlock avoidance algorithms for recursion-tree modeled requests in parallel executions. IEEE Transactions on Computers, 71(9): 2073–2087, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:EDA

[WLD^{+}25] Liangyuan Wang, Xudong Liu, Haonan Ding, Yi Hu, Kai Peng, and Menglan Hu. Energy-delay-aware joint microservice deployment and request routing with DVFS in edge: a reinforcement learning approach. IEEE Transactions on Computers, 74(5): 1589–1604, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:OSP

[WLF^{+}25a] Pengwei Wang, Yi Li, Chao Fang, Yichen Zhong, and Zhijun Ding. Optimizing serverless performance through game theory and efficient resource scheduling. IEEE Transactions on Computers, 74(6):1990–2002, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2025:KKP

[WLF^{+}25b] Jingya Wu, Wenyan Lu, Haishuang Fan, Hao Kong, Xiaowei Li, and Guihai Yan. KPU: Kernel processing unit for in-memory analytical query processing. IEEE Transactions on Computers, 74(8):2702–2716, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2025:RIS

[WLHW25] Qianqian Wu, Qiang Liu, Ying He, and Zefan Wu. Reconfigurable intelligent surface assisted UAV-MCS based on transformer enhanced deep reinforcement learning. IEEE Transactions on Computers, 74(9):3143–3155, September 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2021:RPO

[WLQ^{+}21] Hai Wang, Wei Li, Wenjie Qi, Diya Tang, Letian Huang, and He Tang. Runtime performance optimization of 3-D microprocessors in dark silicon. IEEE Transactions on Computers, 70(10):1539–1554, October 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2020:PAH

[WLR20] S. Wang, X. Li, and R. Ruiz. Performance analysis for heterogeneous cloud servers using queueing theory. IEEE Transactions on Computers, 69(4): 563–576, April 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2021:EEI

[WLW^{+}21] Tian Wang, Yucheng Lu, Jianhuang Wang, Hong-Ning Dai, Xi Zheng, and Weijia Jia. EIHDP: Edge-intelligent hierarchical dynamic pricing based on cloud-edge-client collaboration for IoT systems. IEEE Transactions on Computers, 70(8):1285–1298, August 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:AAA

[WLW^{+}22a] Jihe Wang, Jun Liu, Danghui Wang, Jianfeng An, and Xiaoya Fan. An automaticaddressing architecture with fully serialized access in racetrack memory for energyefficient CNNs. IEEE Transactions on Computers, 71(1): 235–250, January 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:MRR

[WLW^{+}22b] Jihe Wang, Jun Liu, Danghui Wang, Shengbing Zhang, and Xiaoya Fan. MemUnison: a Racetrack-ReRAM-Combined pipeline architecture for energy-efficient inmemory CNNs. IEEE Transactions on Computers, 71(12): 3281–3294, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:CTM

[WLW^{+}22c] Yuze Wang, Peng Liu, Weidong Wang, Xiaohang Wang, and Yingtao Jiang. On a consistency testing model and strategy for revealing RISC processor’s dark instructions and vulnerabilities. IEEE Transactions on Computers, 71(7):1586–1597, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:NHP

[WLW^{+}25] Zhonghua Wang, Kai Lu, Jiguang Wan, Hong Jiang, Zeyang Zhao, Peng Xu, Biliang Lai, Guokuan Li, and Changsheng Xie. NStore: a high-performance NUMAaware key–value store for hybrid memory. IEEE Transactions on Computers, 74(3): 929–943, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:MBE

[WLY^{+}23] Yi Wang, Jing Liao, Jing Yang, Zhengda Li, Chenlin Ma, and Rui Mao. Meta-Block: Exploiting cross-layer and direct storage access for decentralized blockchain storage systems. IEEE Transactions on Computers, 72(7): 2052–2064, July 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2021:TDM

[WLZ^{+}21] F. Wu, B. Li, B. Zhang, Z. Cao, J. Diehl, H. Wen, and D. H. C. Du. Track-Lace: Data management for interlaced magnetic recording. IEEE Transactions on Computers, 70(3):347–358, March 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:EML

[WLZ^{+}23] Hengrui Wang, Huiping Lin, Zheng Zhong, Tong Yang, and Muhammad Shahzad. Enhanced machine learning sketches for network measurements. IEEE Transactions on Computers, 72(4):957–970, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:SAS

[WLZ^{+}25] Yan Wang, Bo Lv, Quan Zhou, Junfei Li, and Tan Tan. Schedulability analysis for self-suspending tasks under EDF-Like scheduling. IEEE Transactions on Computers, 74(7):2364–2375, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:BAM

[WMG^{+}25] Xun Wang, Xiangyu Meng, Zhuoqiang Guo, Mingzhen Li, Lijun Liu, Mingfan Li, Qian Xiao, Tong Zhao, Ninghui Sun, Guangming Tan, and Weile Jia. 29-billion atoms molecular dynamics simulation with ab initio accuracy on 35 million cores of new Sunway supercomputer. IEEE Transactions on Computers, 74(5): 1634–1648, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:CMB

[WMH^{+}24] Ruixuan Wang, Sabrina Hassan Moon, Xiaobo Sharon Hu, Xun Jiao, and Dayane Reis. A computing-in-memory-based one-class hyperdimensional computing model for outlier detection. IEEE Transactions on Computers, 73(6): 1559–1574, June 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:TSN

[WNL^{+}23] Ziheng Wang, Farzad Niknia, Shanshan Liu, Pedro Reviriego, Paolo Montuschi, and Fabrizio Lombardi. Tolerance of Siamese Networks (SNs) to memory errors: Analysis and design. IEEE Transactions on Computers, 72(4):1136–1149, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2022:PTP

[WNP^{+}22] Hao Wu, Krishnendra Nathella, Matthew Pabst, Dam Sunwoo, Akanksha Jain, and Calvin Lin. Practical temporal prefetching with compressed on-chip metadata. IEEE Transactions on Computers, 71(11):2858–2871, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:ASE

[WPL^{+}23] Yuchao Wang, Yanguo Peng, Ximeng Liu, Zuobin Ying, Jiangtao Cui, Dongyao Niu, and Xiaofang Xia. aChain: a SQL-empowered analytical blockchain as a database. IEEE Transactions on Computers, 72(11):3099–3112, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2022:CMT

[WRT^{+}22] Lizhou Wu, Siddharth Rao, Mottaqiallah Taouil, Erik Jan Marinissen, Gouri Sankar Kar, and Said Hamdioui. Characterization, modeling, and test of intermediate state defects in STT-MRAMs. IEEE Transactions on Computers, 71(9): 2219–2233, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:PDP

[WRW^{+}23] Dan Wang, Ju Ren, Zhibo Wang, Yichuan Wang, and Yaoxue Zhang. PrivAim: a dual-privacy preserving and quality-aware incentive mechanism for federated learning. IEEE Transactions on Computers, 72(7):1913–1927, July 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wills:2020:GEI

[WS20] L. W. Wills and K. Swaminathan. Guest editorial: IEEE TC special issue on domain-specific architectures for emerging applications. IEEE Transactions on Computers, 69(8):1096–1098, August 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2024:RWD

[WSC^{+}24] Ronglong Wu, Zhirong Shen, Jianqiang Chen, Chengshuo Zheng, Zhiwei Yang, and Jiwu Shu. Relieving write disturbance for phase change memory with RESET-aware data encoding. IEEE Transactions on Computers, 73(8): 1939–1952, August 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wen:2025:ALE

[WSC^{+}25] Yu Wen, Aamir Bader Shah, Ruizhi Cao, Chen Zhang, Jiefu Chen, Xuqing Wu, Chenhao Xie, and Xin Fu. AR-Light: Enabling fast and lightweight multi-user augmented reality via semantic segmentation and collaborative view synchronization. IEEE Transactions on Computers, 74(6): 2073–2086, June 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wistoff:2023:SPC

[WSG^{+}23] Nils Wistoff, Moritz Schneider, Frank K. Gürkaynak, Gernot Heiser, and Luca Benini. Systematic prevention of on-core timing channels by full temporal partitioning. IEEE Transactions on Computers, 72(5):1420–1430, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2025:LSL

[WSG^{+}25] Minghui Wu, Dawei Sun, Shang Gao, Keqin Li, and Rajkumar Buyya. Ls-Stream: Lightening stragglers in join operators for skewed data stream processing. IEEE Transactions on Computers, 74(8):2841–2855, August 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2023:TIE

[WSHJ23] Wenchao Wu, Xuanhua Shi, Ligang He, and Hai Jin. TurboGNN: Improving the end-to-end performance for sampling-based GNN training on GPUs. IEEE Transactions on Computers, 72(9): 2571–2584, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:TDR

[WSLX24] Fuyu Wang, Minghua Shen, Yutong Lu, and Nong Xiao. TensorMap: a deep RLbased tensor mapping framework for spatial accelerators. IEEE Transactions on Computers, 73(8):1899–1912, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Willis:2024:PCE

[WSM^{+}24] Benjamin R. Willis, Aviral Shrivastava, Joshua Mack, Shail Dave, Chaitali Chakrabarti, and John Brunhaver. Cyclebite: Extracting task graphs from unstructured compute-programs. IEEE Transactions on Computers, 73(1):221–234, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:RRT

[WSQ^{+}25] Xuhang Wang, Zhuoran Song, Chunyu Qi, Fangxin Liu, Naifeng Jing, Li Jiang, and Xiaoyao Liang. RTSA: a run-through sparse attention framework for video transformer. IEEE Transactions on Computers, 74(6):1949–1962, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2020:CSH

[WSS^{+}20] Y. Wang, Y. Shen, C. Su, J. Ma, L. Liu, and X. Dong. CryptSQLite: SQLite with high data security. IEEE Transactions on Computers, 69(5):666–678, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

                                        https://ieeexplore.ieee.
                                        org/document/8946540.

Wang:2025:DFA

[WT25] Weizhe Wang and Deng Tang.

Differential fault attack on HE-friendly stream ciphers: Masta, Pasta, and Elisabeth. IEEE Transactions on Computers, 74(7):2267–2277, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2021:HHA

[WTL^{+}21] Xi Wang, Antonino Tumeo, John D. Leidel, Jie Li, and Yong Chen. HAM: Hotspotaware manager for improving communications with 3Dstacked memory. IEEE Transactions on Computers, 70(6): 833–848, June 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:HCC

[WTL^{+}24] Hao Wang, Bo Tang, Chi Harold Liu, Shangqin Mao, Jiahong Zhou, Zipeng Dai, Yaqi Sun, Qianlong Xie, Xingxing Wang, and Dong Wang. HiBid: a cross-channel constrained bidding system with budget allocation by hierarchical offline deep reinforcement learning. IEEE Transactions on Computers, 73(3):815–828, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2021:MMD

[WWC21] T. Wei, C. Wang, and C. W.

Chen. Modularized morphing of deep convolutional neural networks: a graph approach. IEEE Transactions on Computers, 70(2):305–315, February 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:MAT

[WWJ^{+}23] Shengjie Wang, Xiaohang Wang, Yingtao Jiang, Amit Kumar Singh, Mei Yang, and Letian Huang. Modeling and analysis of thermal covert channel attacks in many-core systems. IEEE Transactions on Computers, 72(2): 494–500, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:TMM

[WWJ^{+}25] Xiaohang Wang, Yifan Wang, Yingtao Jiang, Amit Kumar Singh, and Mei Yang. On task mapping in multi-chiplet based many-core systems to optimize inter- and intrachiplet communications. IEEE Transactions on Computers, 74(2):510–525, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wong:2023:KNS

[WWL^{+}23] Zheng-Yan Wong, Denis C.- K. Wong, Wai-Kong Lee, Kai-Ming Mok, Wun-She Yap, and Ayesha Khalid. KaratSaber: New speed records for Saber polynomial multiplication using efficient Karatsuba FPGA architecture. IEEE Transactions on Computers, 72(7): 1830–1842, July 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2025:AOA

[WWL^{+}25] Xiaoqian Wu, Peng Wang, Shaoquan Li, Huaxiao Liu, and Lei Liu. An area optimization approach for largescale RM-TB dual logic circuits based on a multitasking optimization algorithm. IEEE Transactions on Computers, 74(7):2348–2363, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:EHE

[WWM^{+}23] Lening Wang, Qiyu Wan, Peixun Ma, Jing Wang, Minsong Chen, Shuaiwen Leon Song, and Xin Fu. Enabling high-efficient ReRAM-based CNN training via exploiting crossbar-level insignificant writing elimination. IEEE Transactions on Computers, 72(11):3218–3230, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wen:2022:POM

[WWS^{+}22] Shengyan Wen, Xiaohang Wang, Amit Kumar Singh, Yingtao Jiang, and Mei Yang. Performance optimization of many-core systems by exploiting task migration and dark core allocation. IEEE Transactions on Computers, 71(1): 92–106, January 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:OCC

[WWX^{+}24] Shihang Wang, Xingbo Wang, Zhiyuan Xu, Bingzhen Chen, Chenxi Feng, Qi Wang, and Terry Tao Ye. Optimizing CNN computation using RISC-V custom instruction sets for edge platforms. IEEE Transactions on Computers, 73(5):1371–1384, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:OGB

[WXL^{+}23] Pengyu Wang, Cheng Xu, Chao Li, Jing Wang, Taolei Wang, Lu Zhang, Xiaofeng Hou, and Minyi Guo. Optimizing GPU-based graph sampling and random walk for efficiency and scalability. IEEE Transactions on Computers, 72(9):2508–2521, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2025:NTR

[WY25] Jin-Tao Wang and Tian-Yu Ye. A novel two-round twoparty quantum private comparison protocol based on quantum walks. IEEE Transactions on Computers, 74(8):

2542–2551, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2024:FAR

[WYB^{+}24] Chunshu Wu, Chen Yang, Sahan Bandara, Tong Geng, Anqi Guo, Pouya Haghi, Ang Li, and Martin Herbordt. FPGA-accelerated range-limited molecular dynamics. IEEE Transactions on Computers, 73(6):1544–1558, June 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:PIR

[WYSL22] Hao Wang, Xiangyu Yang, Yuanming Shi, and Jun Lin. A proximal iteratively reweighted approach for efficient network sparsification. IEEE Transactions on Computers, 71(1):185–196, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:RFM

[WYX^{+}24] Hulin Wang, Donglin Yang, Yaqi Xia, Zheng Zhang, Qigang Wang, Jianping Fan, Xiaobo Zhou, and Dazhao Cheng. Raptor-T: a fused and memory-efficient sparse transformer for long and variablelength sequences. IEEE Transactions on Computers, 73(7):1852–1865, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:TCA

[WYZ^{+}22] Xueyan Wang, Jianlei Yang, Yinglin Zhao, Xiaotao Jia, Rong Yin, Xuhang Chen, Gang Qu, and Weisheng Zhao. Triangle counting accelerations: From algorithm to in-memory computing architecture. IEEE Transactions on Computers, 71(10):2462– 2472, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:HCE

[WZCM23] Guang Wang, Ziyuan Zhu, Xu Cheng, and Dan Meng. A high-coverage and efficient instruction-level testing approach for x86 processors. IEEE Transactions on Computers, 72(11):3203– 3217, November 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

West:2020:TUS

[WZD^{+}20] B. L. West, J. Zhou, R. G. Dreslinksi, O. D. Kripfgans, J. B. Fowlkes, C. Chakrabarti, and T. F. Wenisch. Tetris: Using software/hardware codesign to enable handheld, physics-limited 3D plane-wave ultrasound imaging. IEEE Transactions on Computers, 69(8):1209–1220, August 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:ENN

[WZF^{+}24] Jing Wang, Jinbin Zhu, Xin Fu, Di Zang, Keyao Li, and Weigong Zhang. Enhancing neural network reliability: Insights from hardware/ software collaboration with neuron vulnerability quantization. IEEE Transactions on Computers, 73(8):1953– 1966, August 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:ODP

[WZG^{+}23] Haojie Wang, Jidong Zhai, Mingyu Gao, Feng Zhang, Tuowei Wang, Zixuan Ma, Shizhi Tang, Liyan Zheng, Wen Wang, Kaiyuan Rong, Yuanyong Chen, and Zhihao Jia. Optimizing DNNs with partially equivalent transformations and automated corrections. IEEE Transactions on Computers, 72(12):3546– 3560, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2024:UFA

[WZG^{+}24] Qianmei Wu, Fan Zhang, Shize Guo, Kun Yang, and Haoting Shen. A unified and fully automated framework for wavelet-based attacks on random delay. IEEE Transactions on Computers, 73(9):

2206–2219, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wu:2022:IIA

[WZGT22] Yulong Wu, Weizhe Zhang, Nan Guan, and Yue Tang. Improving interference analysis for real-time DAG tasks under partitioned scheduling. IEEE Transactions on Computers, 71(7):1495–1506, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:OSD

[WZH^{+}23] Ne Wang, Ruiting Zhou, Ling Han, Hao Chen, and Zongpeng Li. Online scheduling of distributed machine learning jobs for incentivizing sharing in multi-tenant systems. IEEE Transactions on Computers, 72(3):653–667, March 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wei:2024:RBA

[WZJ^{+}24] Zheng Wei, Xingjun Zhang, Zeyu Ji, Jingbo Li, and Jia Wei. Revisit and benchmarking of automated quantization toward fair comparison. IEEE Transactions on Computers, 73(1):18–29, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:CPC

[WZSL22] Yongzhi Wang, Yu Zou, Yulong Shen, and Yao Liu. CFHider: Protecting control flow confidentiality with Intel SGX. IEEE Transactions on Computers, 71(9): 2128–2141, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2023:HOS

[WZW^{+}23] Yu Wang, You Zhou, Fei Wu, Yu Zhong, Jian Zhou, Zhonghai Lu, Shu Li, Zhengyong Wang, and Changsheng Xie. Holistic and opportunistic scheduling of background I/Os in flash-based SSDs. IEEE Transactions on Computers, 72(11):3127– 3139, November 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2024:LOR

[WZW^{+}24] Yingjia Wang, You Zhou, Fei Wu, Jie Zhang, and Ming-Chang Yang. Land of Oz: Resolving orderless writes in zoned namespace SSDs. IEEE Transactions on Computers, 73(11):2520–2533, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:2022:PVE

[WZX^{+}22] Tianchen Wang, Jiawei Zhang, Jinjun Xiong, Song Bian, Zheyu Yan, Meiping Huang, Jian Zhuang, Takashi Sato, Xiaowei Xu, and Yiyu Shi. VisualNet: An end-to-end human visual system inspired framework to reduce inference latency of deep neural networks. IEEE Transactions on Computers, 71(11):2717– 2727, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2020:SAG

[XAP20] J. Xiao, S. Altmeyer, and A. D. Pimentel. Schedulability analysis of global scheduling for multicore systems with shared caches. IEEE Transactions on Computers, 69(10): 1487–1499, October 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2025:LBF

[XCG^{+}25] Shiyuan Xu, Xue Chen, Yu Guo, Yuer Yang, Shengling Wang, Siu-Ming Yiu, and Xiuzhen Cheng. Lattice-based forward secure multi-user authenticated searchable encryption for cloud storage systems. IEEE Transactions on Computers, 74(5):1663–1677, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xin:2022:HTH

[XCZ^{+}22] Yao Xin, Donglong Chen, Chongyang Zeng, Weichen Zhang, Yi Wang, and Ray C. C. Cheung. High throughput hardware/software heterogeneous system for RRPNbased scene text detection. IEEE Transactions on Computers, 71(7):1507–1521, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xie:2025:APH

[XGMJ25] Hongcheng Xie, Yu Guo, Yinbin Miao, and Xiaohua Jia. Access-pattern hiding search over encrypted databases by using distributed point functions. IEEE Transactions on Computers, 74(3):1066– 1078, March 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xia:2024:SSC

[XGZ^{+}24] Zhihua Xia, Qi Gu, Wenhao Zhou, Lizhi Xiong, Jian Weng, and Naixue Xiong. STR: Secure computation on additive shares using the share-transform-reveal strategy. IEEE Transactions on Computers, 73(2):340– 352, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xia:2022:CPT

[XHY^{+}22] Yubin Xia, Zhichao Hua, Yang Yu, Jinyu Gu, Haibo Chen, Binyu Zang, and Haibing Guan. Colony: a privileged trusted execution environment with extensibility. IEEE Transactions on Computers, 71(2):479–492, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hu:2024:MSI

[xHzLH^{+}24] He xuan Hu, Zhen zhou Lin, Qiang Hu, Ye Zhang, Wei Wei, and Wei Wang. Multi-source information fusion based DLaaS for traffic flow prediction. IEEE Transactions on Computers, 73(4): 994–1003, April 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xin:2025:HAA

[XJL^{+}25] Yao Xin, Chengjun Jia, Wenjun Li, Ori Rottenstreich, Yang Xu, Gaogang Xie, Zhihong Tian, and Jun Li. A heterogeneous and adaptive architecture for decision-tree-based ACL engine on FPGA. IEEE Transactions on Computers, 74(1):263–277, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xue:2024:AGO

[XJY^{+}24] Yongqi Xue, Jinlun Ji, Xinming Yu, Shize Zhou, Siyue Li, Xinyi Li, Tong Cheng, Shiping Li, Kai Chen, Zhonghai Lu, Li Li, and Yuxiang Fu. Automatic generation and optimization framework of NoCbased neural network accelerator through reinforcement learning. IEEE Transactions on Computers, 73(12):2882– 2896, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiong:2021:LIT

[XKS21] W. Xiong, S. Katzenbeisser, and J. Szefer. Leaking information through cache LRU states in commercial processors and secure caches. IEEE Transactions on Computers, 70(4):511–523, April 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xie:2025:ELB

[XL25] Yujun Xie and Yuan Liu. An efficient LUT6-based Montgomery modular multiplication using radix-16 Booth method. IEEE Transactions on Computers, 74(9):3223– 3237, September 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2022:DLF

[XLL^{+}22] Liangliang Xu, Min Lyu, Zhipeng Li, Cheng Li, and Yinlong Xu. A data layout and fast failure recovery scheme for distributed storage systems with mixed erasure codes. IEEE Transactions on Computers, 71(8): 1740–1754, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xie:2024:GNN

[XLS^{+}24] Guorui Xie, Qing Li, Zhenning Shi, Hanbin Fang, Shengpeng Ji, Yong Jiang, Zhenhui Yuan, Lianbo Ma, and Mingwei Xu. Generating neural networks for diverse networking classification tasks via hardware-aware neural architecture search. IEEE Transactions on Computers, 73(2): 481–494, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiang:2025:AAD

[XLT^{+}25] Debin Xiang, Liqiang Lu, Siwei Tan, Xinghui Jia, Zhe Zhou, Guangyu Sun, Mingshuai Chen, and Jianwei Yin. AdaptDQC: Adaptive distributed quantum computing with quantitative performance analysis. IEEE Transactions on Computers, 74(10):3277– 3290, October 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2020:AGN

[XLW^{+}20] D. Xu, C. Liu, Y. Wang, K. Tu, B. He, and L. Zhang. Accelerating generative neural networks on unmodified deep learning processors: a software approach. IEEE Transactions on Computers, 69(8): 1172–1184, August 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xi:2023:PSS

[XLWO23] Xiaodan Xi, Ge Li, Ye Wang, and Michael Orshansky. A provably secure strong PUF based on LWE: Construction and implementation. IEEE Transactions on Computers, 72(2):346–359, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2022:CCB

[XLY^{+}22] Minghui Xu, Shuo Liu, Dongxiao Yu, Xiuzhen Cheng, Shaoyong Guo, and Jiguo Yu. CloudChain: a cloud blockchain using shared memory consensus and RDMA. IEEE Transactions on Computers, 71(12):3242–3253, December 2022. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xing:2025:HRM

[XLY^{+}25] Yile Xing, Guangyan Li, Zewen Ye, Ryan W. L. Luk, Donglong Chen, Hong Yan, and Ray C. C. Cheung. High-radix/mixed-radix NTT multiplication algorithm/architecture co-design over Fermat modulus. IEEE Transactions on Computers, 74(10):3519–3533, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xie:2025:FFL

[XLZ^{+}25] Yujun Xie, Yuan Liu, Xin Zheng, Bohan Lan, Dengyun Lei, Dehao Xiang, Shuting Cai, and Xiaoming Xiong. FLALM: a flexible low arealatency Montgomery modular multiplication on FPGA. IEEE Transactions on Computers, 74(1):29–42, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2021:PCD

[XNB21] Yao Xiao, Shahin Nazarian, and Paul Bogdan. Plasticityon-chip design: Exploiting self-similarity for data communications. IEEE Transactions on Computers, 70(6):950–962, June 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2025:RTS

[XNH^{+}25] Yuankai Xu, Yinchen Ni, Tiancheng He, Ruiqi Sun, Yier Jin, and An Zou. Realtime scheduling and analysis of fixed-priority tasks on a basic heterogeneous architecture with multiple CPUs and many PEs. IEEE Transactions on Computers, 74(8): 2785–2798, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xue:2023:BBF

[XNL^{+}23] Liang Xue, Jianbing Ni, Dongxiao Liu, Xiaodong Lin, and Xuemin Shen. Blockchain-based fair and finegrained data trading with privacy preservation. IEEE Transactions on Computers, 72(9):2440–2453, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xie:2020:SSS

[XNLX20] J. Y. Xie, G. Nong, B. Lao, and W. Xu. Scalable suffix sorting on a multicore machine. IEEE Transactions on Computers, 69(9): 1364–1375, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2022:MSC

[XPR^{+}22] Zhuang Xu, Owen Pemberton, Sujoy Sinha Roy, David Oswald, Wang Yao, and Zhiming Zheng. Magnifying sidechannel leakage of latticebased cryptosystems with chosen ciphertexts: the case study of Kyber. IEEE Transactions on Computers, 71(9): 2163–2176, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2022:DAP

[XQC^{+}22] Fei Xu, Yiling Qin, Li Chen, Zhi Zhou, and Fangming Liu.

DNN: Achieving predictable distributed DNN training with serverless architectures. IEEE Transactions on Computers, 71(2):450–463, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiang:2020:BNC

[XST20] X. Xiang, P. Sigdel, and N. Tzeng. Bufferless networkon-chips with bridged multiple subnetworks for deflection reduction and energy savings. IEEE Transactions on Computers, 69(4):577–590, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2022:BRB

[XSYL22] Jie Xiao, Zhanhui Shi, Xuhua Yang, and Jungang Lou. BM-RCGL: Benchmarking approach for localization of reliability-critical gates in combinational logic blocks. IEEE Transactions on Computers, 71(5):1063–1076, May 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2023:REE

[XTWG23] Ke Xu, Ming Tang, Han Wang, and Sylvain Guilley. Reverse-engineering and exploiting the frontend bus of Intel processor. IEEE Transactions on Computers, 72(2): 360–373, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2025:MAM

[XTWW25] Ke Xu, Ming Tang, Quancheng Wang, and Han Wang. Microarchitectural attacks and mitigations on retire resources in modern processors. IEEE Transactions on Computers, 74(4):1253–1266, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2024:RAG

[Xu24] Jianfeng Xu. Research and application of general information measures based on a unified model. IEEE Transactions on Computers, 73(3): 915–927, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2024:AAD

[XWL^{+}24] Zichuan Xu, Lin Wang, Weifa Liang, Qiufen Xia, Wenzheng Xu, Pan Zhou, and Omer F. Rana. Age-aware data selection and aggregator placement for timely federated continual learning in mobile edge computing. IEEE Transactions on Computers, 73(2): 466–480, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2021:POQ

[XWP^{+}21] Siyuan Xiao, Xiaohang Wang, Maurizio Palesi, Amit Kumar Singh, Liang Wang, and Terrence Mak. On performance optimization and quality control for approximatecommunication-enabled networkson-chip. IEEE Transactions on Computers, 70(11):1817– 1830, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2025:DDA

[XXC^{+}25] Guoqing Xiao, Li Xia, Yuedan Chen, Hongyang Chen, and Wangdong Yang. DCGG: a dynamically adaptive and hardware-software coordinated runtime system for GNN acceleration on GPUs. IEEE Transactions on Computers, 74(7):2293–2305, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2024:OSB

[XXJ^{+}24] Guangquan Xu, Guohua Xin, Litao Jiao, Jian Liu, Shaoying Liu, Meiqi Feng, and Xi Zheng. OFEI: a semi-black-box Android adversarial sample attack framework against DLaaS. IEEE Transactions on Computers, 73(4):956–969, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2023:EAE

[XXL^{+}23] Hao Xu, Bin Xiao, Xiulong Liu, Li Wang, Shan Jiang, Weilian Xue, Jianrong Wang, and Keqiu Li. Empowering authenticated and efficient queries for STK transaction-based blockchains. IEEE Transactions on Computers, 72(8):2209–2223, August 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xiao:2025:ECE

[XXZ^{+}25] Danlei Xiao, Shaobo Xu, Chuan Zhang, Licheng Wang, Xiulong Liu, and Liehuang Zhu. EC2P: Cost-effective cross-chain payments via hubs resisting the abort attack. IEEE Transactions on Computers, 74(10):3504–3518, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xue:2023:SES

[XYM23] Zihui Xue, Yuedong Yang, and Radu Marculescu. SUGAR: Efficient subgraph-level training via resource-aware graph partitioning. IEEE Transactions on Computers, 72(11): 3167–3177, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2023:SBE

[XZC^{+}23] Minghui Xu, Zongrui Zou, Ye Cheng, Qin Hu, Dongxiao Yu, and Xiuzhen Cheng. SPDL: a blockchain-enabled secure and privacy-preserving decentralized learning system. IEEE Transactions on Computers, 72(2):548–558, February 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2021:AAV

[XZL^{+}21] Zichuan Xu, Zhiheng Zhang, John C. S. Lui, Weifa Liang, Qiufen Xia, Pan Zhou, Wenzheng Xu, and Guowei Wu. Affinity-aware VNF placement in mobile edge clouds via leveraging GPUs. IEEE Transactions on Computers, 70(12):2234–2248, December 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2023:SSA

[XZL^{+}23] Zichuan Xu, Lizhen Zhou, Weifa Liang, Qiufen Xia, Wenzheng Xu, Wenhao Ren, Haozhe Ren, and Pan Zhou. Stateful serverless application placement in MEC with function and state dependencies. IEEE Transactions on Computers, 72(9):2701– 2716, September 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Xu:2025:TEE

[XZL^{+}25] Hao Xu, Jiaqi Zhang, Xiulong Liu, Zhimin Yu, Tingyu Fan, Baochao Chen, and Keqiu Li. Tangram: Enabling efficient and balanced dynamic storage extension on sharding blockchain systems. IEEE Transactions on Computers, 74(6):2031–2044, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yazdani:2020:LLA

[YAG20] R. Yazdani, J. Arnau, and A. González. LAWS: Locality-AWare Scheme for automatic speech recognition. IEEE Transactions on Computers, 69(8):1197–1208, August 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yuan:2025:TOP

[YBF^{+}25] Zimeng Yuan, Yuanguo Bi, Yanbo Fan, Yuheng Liu, Lianbo Ma, Liang Zhao, and Qiang He. Trajectory optimization and power allocation for multi-UAV wireless networks: a communicationbased multi-agent deep reinforcement learning approach. IEEE Transactions on Computers, 74(10):3404–3418, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yayla:2022:FBB

[YBG^{+}22] Mikail Yayla, Sebastian Buschjäger, Aniket Gupta, Jian-Jia Chen, Jörg Henkel, Katharina Morik, Kuan-Hsun Chen, and Hussam Amrouch. FeFET-based binarized neural networks under temperature-dependent bit errors. IEEE Transactions on Computers, 71(7): 1681–1695, July 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2021:MMP

[YBW21] Chao Yu, Yuebin Bai, and Rui Wang. MIPSGPU: Minimizing pipeline stalls for GPUs with non-blocking execution. IEEE Transactions on Computers, 70(11):1804– 1816, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2020:RFC

[YCKW20] M. Yang, Y. Chang, T. Kuo, and C. Wu. Request flow coordination for growing-scale solid-state drives. IEEE Transactions on Computers, 69(6):832–843, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2024:GER

[YCL^{+}24] Dongxiao Yu, Ruopeng Chen, Xin Li, Mengbai Xiao, Guanghui Zhang, and Yao Liu. A GPUenabled real-time framework for compressing and rendering volumetric videos. IEEE Transactions on Computers, 73(3):789–800, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2024:ONA

[YCS^{+}24] Da Yang, Zhenglong Cui, Hao Sheng, Rongshan Chen, Ruixuan Cong, Shuai Wang, and Zhang Xiong. An occlusion and noise-aware stereo framework based on light field imaging for robust disparity estimation. IEEE Transactions on Computers, 73(3): 764–777, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yuan:2024:DLL

[YCY^{+}24] Yuan Yuan, Shuzhen Chen, Dongxiao Yu, Zengrui Zhao, Yifei Zou, Lizhen Cui, and Xiuzhen Cheng. Distributed learning for large-scale models at edge with privacy protection. IEEE Transactions on Computers, 73(4):1060–1070, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ye:2024:MIA

[YDG^{+}24] Kui Ye, Shengxin Dai, Bing Guo, Yan Shen, Chuanjie Liu, Kejun Bi, Fei Chen, Yuchuan Hu, and Mingjie Zhao. A mutual-influence-aware heuristic method for quantum circuit mapping. IEEE Transactions on Computers, 73(12):2855–2867, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yuan:2025:AIF

[YDL^{+}25] Shijing Yuan, Beiyu Dong, Jie Li, Song Guo, Hongyang Chen, Chentao Wu, Jie Wu, and Wei Zhao. Adaptive incentivize for federated learning with cloud-edge collaboration under multi-level information sharing. IEEE Transactions on Computers, 74(7): 2445–2460, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2025:HHC

[YDW^{+}25] Lingxiao Yang, Xuewen Dong, Zhiguo Wan, Di Lu, Yushu Zhang, and Yulong Shen. HiCoCS: High concurrency crosssharding on permissioned blockchains. IEEE Transactions on Computers, 74(7): 2168–2182, July 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2022:SEN

[YFC^{+}22] Jianlei Yang, Wenzhi Fu, Xingzhou Cheng, Xucheng Ye, Pengcheng Dai, and Weisheng Zhao. S2 Engine: a novel systolic architecture for sparse convolutional neural networks. IEEE Transactions on Computers, 71(6):1440–1452, June 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ye:2023:CIC

[YGW^{+}23] Yuejin Ye, Heng Guo, Bingzhuo Wang, Pengxiao Wang, Dexun Chen, and Fang Li. Coupled incomplete Cholesky and Jacobi preconditioned conjugate gradient on the new generation of Sunway many-core architecture. IEEE Transactions on Computers, 72(11):3326– 3339, November 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2020:MSM

[YH20] D. K. Yu and J. Hsieh. A management scheme of multilevel retention-time queues for improving the endurance of flash-memory storage devices. IEEE Transactions on Computers, 69(4):549–562, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2024:CRS

[YH24] Leilei Yu and Yunghsiang S.

Han. Construction of Reed– Solomon erasure codes with four parities based on systematic Vandermonde matrices. IEEE Transactions on Computers, 73(7):1875–1882, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yan:2020:IAT

[YHC^{+}20] A. Yan, Y. Hu, J. Cui, Z. Chen, Z. Huang, T. Ni, P. Girard, and X. Wen. Information assurance through redundant design: A novel TNU error-resilient latch for harsh radiation environment. IEEE Transactions on Computers, 69(6):789–799, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ye:2025:PAN

[YHH^{+}25] Zewen Ye, Junhao Huang, Tianshun Huang, Yudan Bai, Jinze Li, Hao Zhang, Guangyan Li, Donglong Chen, Ray C. C. Cheung, and Kejie Huang. PQNTRU: Acceleration of NTRU-based schemes via customized post-quantum processor. IEEE Transactions on Computers, 74(5): 1649–1662, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2021:DBQ

[YHV^{+}21] H. Yu, Y. Ha, B. Veeravalli, F. Chen, and H. El-Sayed. DVFS-based quality maximization for adaptive applications with diminishing return. IEEE Transactions on Computers, 70(5):803–816, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2025:KBA

[YHW^{+}25] Kaiqi Yang, Qiang He, Xingwei Wang, Zhi Liu, Yufei Liu, Min Huang, and Liang Zhao.

KDN-based adaptive computation offloading and resource allocation strategy optimization: Maximizing user satisfaction. IEEE Transactions on Computers, 74(5):1743–1757, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yao:2020:PEP

[YL20] Y. Yao and Z. Lu. Pursuing extreme power efficiency with PPCC guided NoC DVFS. IEEE Transactions on Computers, 69(3):410–426, March 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2021:SDD

[YLC^{+}21] Y. Yu, Y. Li, S. Che, N. K.

Jha, and W. Zhang. Softwaredefined design space exploration for an efficient DNN accelerator architecture. IEEE Transactions on Computers, 70(1):45–56, January 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yao:2023:THP

[YLG^{+}23] Lulu Yao, Yongkun Li, Fan Guo, Si Wu, Yinlong Xu, and John C. S. Lui. Towards high performance and efficient memory deduplication via mixed pages. IEEE Transactions on Computers, 72(4): 926–940, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2023:RSC

[YLHL23] Leilei Yu, Sian-Jheng Lin, Hanxu Hou, and Zhengrui Li. Reed–Solomon coding algorithms based on Reed– Muller transform for any number of parities. IEEE Transactions on Computers, 72(9): 2677–2688, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2020:FEA

[YLL^{+}20] L. Yu, Z. Lin, S. Lin, Y. S. Han, and N. Yu. Fast encoding algorithms for Reed– Solomon codes with between four and seven parity symbols. IEEE Transactions on Computers, 69(5):699–705, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.
org/document/8949691.

Yao:2024:ABS

[YLL^{+}24] Lulu Yao, Yongkun Li, Patrick P. C. Lee, Xiaoyang Wang, and Yinlong Xu. AdaptMD: Balancing space and performance in NUMA architectures with adaptive memory deduplication. IEEE Transactions on Computers, 73(6): 1588–1602, June 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yao:2023:EOG

[YLT^{+}23] Jianguo Yao, Qiumin Lu, Run Tian, Keqin Li, and Haibing Guan. An economy-oriented GPU virtualization with dynamic and adaptive oversubscription. IEEE Transactions on Computers, 72(5): 1371–1383, May 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2024:FFB

[YLZ^{+}24] Yongkui Yang, Zhenyan Lu, Jingwei Zeng, Xingguo Liu, Xuehai Qian, and Zhibin Yu. Falic: An FPGA-based multi-scalar multiplication accelerator for zero-knowledge proof. IEEE Transactions on Computers, 73(12):2791–2804, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2021:PRA

[YNJS21] L. Yang, B. Nie, A. Jog, and E. Smirni. Practical resilience analysis of GPGPU applications in the presence of single- and multi-bit faults. IEEE Transactions on Computers, 70(1):30–44, January 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yun:2024:GEN

[YNP^{+}24] Sungmin Yun, Hwayong Nam, Jaehyun Park, Byeongho Kim, Jung Ho Ahn, and Eojin Lee. GraNDe: Efficient neardata processing architecture for graph neural networks. IEEE Transactions on Computers, 73(10):2391–2404, October 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yao:2024:PFT

[YPD^{+}24] Dezhong Yao, Wanning Pan, Yutong Dai, Yao Wan, Xiaofeng Ding, Chen Yu, Hai Jin, Zheng Xu, and Lichao Sun. FedGKD: Toward heterogeneous federated learning via global knowledge distillation. IEEE Transactions on Computers, 73(1):3–17, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ye:2025:ACB

[YPL^{+}25] Jin Ye, Yajun Peng, Yijun Li, Zhaoyi Li, and Jiawei Huang. Asynchronous control based aggregation transport protocol for distributed deep learning. IEEE Transactions on Computers, 74(4):1362–1376, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yi:2025:GEG

[YQX^{+}25] Shiyan Yi, Yudi Qiu, Guohao Xu, Lingfei Lu, Xiaoyang Zeng, and Yibo Fan. GATe: Efficient graph attention network acceleration with near-memory processing. IEEE Transactions on Computers, 74(10):3419–3432, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yin:2021:MMT

[YTD^{+}21] J. Yin, Y. Tang, S. Deng, B. Zheng, and A. Y. Zomaya. MUSE: a multi-tierd and SLAdriven deduplication framework for cloud storage systems. IEEE Transactions on Computers, 70(5):759–774, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2021:MID

[YWC^{+}21] M.-C. Yang, C.-F. Wu, S.- H. Chen, Y.-L. Lin, C.-W. Chang, and Y.-H. Chang. On minimizing internal data migrations of flash devices via lifetime-retention harmonization. IEEE Transactions on Computers, 70(3): 428–439, March 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2024:FLH

[YWC^{+}24a] Tianyang Yu, Bi Wu, Ke Chen, Gong Zhang, and Weiqiang Liu. Fully learnable hyperdimensional computing framework with ultratiny accelerator for edge-side applications. IEEE Transactions on Computers, 73(2):574–585, February 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2024:MBA

[YWC^{+}24b] Tianyang Yu, Bi Wu, Ke Chen, Gong Zhang, and Weiqiang Liu. Memristor-based approximate query architecture for in-memory hyperdimensional computing. IEEE Transactions on Computers, 73(11): 2605–2618, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2025:NHP

[YWF^{+}25] Weiling Yang, Pengyu Wang, Jianbin Fang, Dezun Dong, Zhengbin Pang, Runxi He, Peng Zhang, Tao Tang, Chun Huang, Yonggang Che, and Jie Ren. nDirect2: a highperformance library for direct convolutions on multicore CPUs. IEEE Transactions on Computers, 74(6):1829–1843, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yilmaz:2023:MMC

[YWP^{+}23] Baki Berkay Yilmaz, Frank Werner, Sunjae Y. Park, Elvan Mert Ugurlu, Erik Jorgensen, Milos Prvulovic, and Alenka Zaji . MarCNNet: a Markovian convolutional neural network for malware detection and monitoring multicore systems. IEEE Transactions on Computers, 72(4): 1122–1135, April 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2025:RSU

[YWQ^{+}25] Xu Yang, Qiuhao Wang, Saiyu Qi, Ke Li, and Yong Qi. RO(SE)^{2}: Search-efficient robust searchable encryption with forward and backward security. IEEE Transactions on Computers, 74(9):3114– 3128, September 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2023:HDA

[YWX^{+}23] Zichao Yang, Heng Wu, Yuanjia Xu, Yuewen Wu, Hua Zhong, and Wenbo Zhang. Hydra: Deadline-aware and efficiency-oriented scheduling for deep learning jobs on heterogeneous GPUs. IEEE Transactions on Computers, 72(8):2224–2236, August 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2024:CCD

[YYCR24] Jin Yang, Zhenkun Yang, Jeremy Casas, and Sandip Ray. Correct-by-construction design of custom accelerator microarchitectures. IEEE Transactions on Computers, 73(1):278–291, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2024:ESD

[YYQ^{+}24] Jiguo Yu, Biwei Yan, Huayi Qi, Shengling Wang, and Wei Cheng. An efficient and secure data sharing scheme for edgeenabled IoT. IEEE Transactions on Computers, 73(1): 178–191, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

You:2024:ESF

[YYW^{+}24] Xin You, Hailong Yang, Siqi Wang, Tao Peng, Chen Ding, Xinyuan Li, Bangduo Chen, Zhongzhi Luan, Tongxuan Liu, Yong Li, and Depei Qian. Exploiting structured feature and runtime isolation for highperformant recommendation serving. IEEE Transactions on Computers, 73(11):2474– 2487, November 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2025:PBA

[YYZ^{+}25] Dongxiao Yu, Yuan Yuan, Yifei Zou, Xiao Zhang, Yu Liu, Lizhen Cui, and Xiuzhen Cheng. Pruning-based adaptive federated learning at the edge. IEEE Transactions on Computers, 74(5):1538–1548, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2025:HPE

[YZG^{+}25] Guangkuo Yang, Meng Zhang, Peng Guo, Xuepeng Zhan, Shaoqi Yang, Xiaohuan Zhao, Xinyi Guo, Pengpeng Sang, Jixuan Wu, Fei Wu, and Jiezhi Chen. High-precision error bit prediction for 3D QLC NAND flash memory: Observations, analysis, and modeling. IEEE Transactions on Computers, 74(4):1392–1404, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yan:2023:LSS

[YZJ23] Baoyue Yan, Jinbin Zhu, and Bo Jiang. Limon: a scalable and stable key–value engine for fast NVMe devices. IEEE Transactions on Computers, 72(10):3017–3028, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yang:2025:SEC

[YZM^{+}25] Li Yang, Wei Zhang, Yinbin Miao, Yanrong Liang, Xinghua Li, Kim-Kwang Raymond Choo, and Robert H. Deng. Secure and efficient cross-modal retrieval over encrypted multimodal data. IEEE Transactions on Computers, 74(4):1405–1417, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yan:2024:FPB

[YZX^{+}24] Biwei Yan, Hongliang Zhang, Minghui Xu, Dongxiao Yu, and Xiuzhen Cheng. FedRFQ: Prototype-based federated learning with reduced redundancy, minimal failure, and enhanced quality. IEEE Transactions on Computers, 73(4):1086–1098, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Yu:2025:ESS

[YZY^{+}25] Haiyang Yu, Hui Zhang, Zhen Yang, Yuwen Chen, and Huan Liu. Efficient and secure storage verification in cloudassisted industrial IoT networks. IEEE Transactions on Computers, 74(5):1702–1716, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zagalo:2023:RTS

[ZABHCG23] Kevin Zagalo, Yasmina Abdeddaïm, Avner Bar-Hen, and Liliana Cucu-Grosjean. Response time stochastic analysis for fixed-priority stable real-time systems. IEEE Transactions on Computers, 72(1): 3–14, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zervakis:2022:TAD

[ZAS^{+}22] Georgios Zervakis, Iraklis Anagnostopoulos, Sami Salamin, Ourania Spantidi, Isai Roman-Ballesteros, Jörg Henkel, and Hussam Amrouch. Thermalaware design for approximate DNN accelerators. IEEE Transactions on Computers, 71(10):2687–2697, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zijlstra:2022:LBC

[ZBT22] Timo Zijlstra, Karim Bigou, and Arnaud Tisserand. Latticebased cryptosystems on FPGA: Parallelization and comparison using HLS. IEEE Transactions on Computers, 71(8): 1916–1927, August 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2024:VIL

[ZC24] Ying Zhao and Jinjun Chen.

Vector-indistinguishability: Location dependency based privacy protection for successive location data. IEEE Transactions on Computers, 73(4): 970–979, April 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zini:2023:AAM

[ZCB23] Matteo Zini, Daniel Casini, and Alessandro Biondi. Analyzing ARM’s MPAM from the perspective of time predictability. IEEE Transactions on Computers, 72(1): 168–182, January 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2023:ICU

[ZCC^{+}23] Han Zhao, Weihao Cui, Quan Chen, Jingwen Leng, Deze Zeng, and Minyi Guo. Improving cluster utilization through adaptive resource management for deep neural network and CPU jobs colocation. IEEE Transactions on Computers, 72(12):3458– 3472, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2023:IEI

[ZCCG23] Han Zhao, Weihao Cui, Quan Chen, and Minyi Guo. ISPA: Exploiting intra-SM parallelism in GPUs via finegrained resource management. IEEE Transactions on Computers, 72(5):1473–1487, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:RBM

[ZCD^{+}22] Jiarui Zhang, Yukun Cheng, Xiaotie Deng, Bo Wang, Jan Xie, Yuanyuan Yang, and Mengqian Zhang. A reputation-based mechanism for transaction processing in blockchain systems. IEEE Transactions on Computers, 71(10):2423–2434, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zoni:2020:ADC

[ZCF20] D. Zoni, L. Cremona, and W. Fornaciari. All-digital control-theoretic scheme to optimize energy budget and allocation in multi-cores. IEEE Transactions on Computers, 69(5):706–721, May 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL

https://ieeexplore.ieee.
org/document/8949548.

Zhang:2024:EEA

[ZCH^{+}24] Jianting Zhang, Wuhui Chen, Zicong Hong, Gang Xiao, Linlin Du, and Zibin Zheng. Efficient execution of arbitrarily complex cross-shard contracts for blockchain sharding. IEEE Transactions on Computers, 73(5):1190–1205, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2020:FFM

[ZCJ^{+}20] C. Zhang, Q. Cao, H. Jiang, W. Zhang, J. Li, and J. Yao. A fast filtering mechanism to improve efficiency of large-scale video analytics. IEEE Transactions on Computers, 69(6): 914–928, June 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2020:NFM

[ZCK20] H. Zhang, D. Chen, and S. Ko. New flexible multipleprecision multiply-accumulate unit for deep neural network training and inference. IEEE Transactions on Computers, 69(1):26–38, January 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:HRD

[ZCP22] Bo Zhang, Zeming Cheng, and Massoud Pedram. Highradix design of a scalable Montgomery modular multiplier with low latency. IEEE Transactions on Computers, 71(2):436–449, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:IMM

[ZCP23] Bo Zhang, Zeming Cheng, and Massoud Pedram. An iterative Montgomery modular multiplication algorithm with low area–time product. IEEE Transactions on Computers, 72(1):236–249, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:GCO

[ZCR22] Jiaqi Zhang, Xiangru Chen, and Sandip Ray. GCONV Chain: Optimizing the wholelife cost in end-to-end CNN acceleration. IEEE Transactions on Computers, 71(9): 2300–2312, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:AAI

[ZCR23] Jiaqi Zhang, Xiangru Chen, and Sandip Ray. AINNS: All-inclusive neural network scheduling via accelerator formalization. IEEE Transactions on Computers, 72(2): 559–571, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zou:2025:RVT

[ZCS^{+}25] Xiaofeng Zou, Cen Chen, Hongen Shao, Qinyu Wang, Xiaobin Zhuang, Yangfan Li, and Keqin Li. ReViT: Vision transformer accelerator with reconfigurable semantic-aware Differential Attention. IEEE Transactions on Computers, 74(3):1079–1093, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zokaee:2023:PSS

[ZCSJ23] Farzaneh Zokaee, Fan Chen, Guangyu Sun, and Lei Jiang. Sky-Sorter: a processing-in-memory architecture for largescale sorting. IEEE Transactions on Computers, 72(2): 480–493, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2021:PAP

[ZCW^{+}21] Shuai Zhao, Wanli Chang, Ran Wei, Weichen Liu, Nan Guan, Alan Burns, and Andy Wellings. Priority assignment on partitioned multiprocessor systems with shared resources. IEEE Transactions on Computers, 70(7):1006–1018, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:AEM

[ZCWC23] Sunrui Zhang, Xiaole Cui, Feng Wei, and Xiaoxin Cui. An area-efficient in-memory implementation method of arbitrary Boolean function based on SRAM array. IEEE Transactions on Computers, 72(12):3416–3430, December 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2020:PPD

[ZCX^{+}20] J. Zhou, Y. Chen, W. Xie, D. Dai, S. He, and W. Wang. PRS: A pattern-directed replication scheme for heterogeneous object-based storage. IEEE Transactions on Computers, 69(4):591–605, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhong:2023:MBI

[ZCX^{+}23] Fangtian Zhong, Zekai Chen, Minghui Xu, Guoming Zhang, Dongxiao Yu, and Xiuzhen Cheng. Malware-on-the-brain: Illuminating malware byte codes with images for malware classification. IEEE Transactions on Computers, 72(2): 438–451, February 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhong:2024:MCA

[ZCY^{+}24] Fangtian Zhong, Xiuzhen Cheng, Dongxiao Yu, Bei Gong, Shuaiwen Song, and Jiguo Yu. MalFox: Camouflaged adversarial malware example generation based on Conv-GANs against black-box detectors. IEEE Transactions on Computers, 73(4):980–993, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:DBD

[ZCY^{+}25] Yu Zhang, Renhai Chen, Hangyu Yan, Hongyue Wu, and Zhiyong Feng. DCAS-BMT: Dynamic construction and adjustment of skewed bonsai Merkle tree for performance enhancement in secure non-volatile memory. IEEE Transactions on Computers, 74(7):2183–2194, July 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:TQA

[ZCZ^{+}22] Wei Zhang, Quan Chen, Ningxin Zheng, Weihao Cui, Kaihua Fu, and Minyi Guo. Toward QoS-awareness and improved utilization of spatial multitasking GPUs. IEEE Transactions on Computers, 71(4):866–879, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2023:DDH

[ZCZW23] Jiang Zhou, Yong Chen, Mai Zheng, and Weiping Wang. Data distribution for heterogeneous storage systems. IEEE Transactions on Computers, 72(6):1747–1762, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:AKF

[ZDC^{+}25] Han Zhao, Junxiao Deng, Weihao Cui, Quan Chen, Youtao Zhang, Deze Zeng, and Minyi Guo. Adaptive kernel fusion for improving the GPU utilization while ensuring QoS. IEEE Transactions on Computers, 74(2):386–400, February 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2022:LML

[ZDV^{+}22] Jiawei Zhao, Steve Dai, Rangharajan Venkatesan, Brian Zimmer, Mustafa Ali, Ming-Yu Liu, Brucek Khailany, William J. Dally, and Anima Anandkumar. LNS-Madam: Low-precision training in logarithmic number system using multiplicative weight update. IEEE Transactions on Computers, 71(12):3179– 3190, December 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2023:VAQ

[ZDW^{+}23] Pengcheng Zhu, Weiping Ding, Lihua Wei, Xueyun Cheng, Zhijin Guan, and Shiguang Feng. A variationaware quantum circuit mapping approach based on multiagent cooperation. IEEE Transactions on Computers, 72(8):2237–2249, August 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2023:PPT

[ZDY^{+}23] Xinghui Zhu, Zijie Di, Qingsong Yao, Xuewen Dong, Jiandong Wang, and Yulong Shen. Performance-power tradeoff in heterogeneous SaaS clouds with trustworthiness guarantee. IEEE Transactions on Computers, 72(6):1554–1567, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2023:SMD

[ZDZ^{+}23] Shulin Zeng, Guohao Dai, Niansong Zhang, Xinhao Yang, Haoyu Zhang, Zhenhua Zhu, Huazhong Yang, and Yu Wang. Serving Multi-DNN workloads on FPGAs: a coordinated architecture, scheduling, and mapping perspective. IEEE Transactions on Computers, 72(5):1314–1328, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2020:MLC

[ZFD^{+}20] Y. Zhao, Z. Fan, Z. Du, T. Zhi, L. Li, Q. Guo, S. Liu, Z. Xu, T. Chen, and Y. Chen. Machine learning computers with fractal von Neumann architecture. IEEE Transactions on Computers, 69(7): 998–1014, July 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2023:MMB

[ZFH23] Hai Zhou, Dan Feng, and Yuchong Hu. MDTUpdate: a multi-block double tree update technique in heterogeneous erasure-coded clusters. IEEE Transactions on Computers, 72(10):2808–2821, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2025:FGC

[ZFH^{+}25] Hai Zhou, Dan Feng, Yuchong Hu, Wei Wang, and Huadong Huang. Fast garbage collection in erasure-coded storage clusters. IEEE Transactions on Computers, 74(8): 2827–2840, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:TNE

[ZFL^{+}22] Lu Zhang, Weiqi Feng, Chao Li, Xiaofeng Hou, Pengyu Wang, Jing Wang, and Minyi Guo. Tapping into NFV environment for opportunistic serverless edge function deployment. IEEE Transactions on Computers, 71(10):2698– 2704, October 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:TPM

[ZFQ^{+}23] Long Zhang, Gang Feng, Shuang Qin, Xiaoqian Li, Yao Sun, and Bin Cao. Trust-preserving mechanism for blockchain assisted mobile crowdsensing. IEEE Transactions on Computers, 72(11): 3113–3126, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2021:EHE

[ZFZ^{+}21] X. Zhang, X. Fu, D. Zhuang, C. Xie, and S. L. Song. Enabling highly efficient capsule networks processing through software-hardware co-design. IEEE Transactions on Computers, 70(4):495–510, April 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhong:2023:CTP

[ZG23] Yadi Zhong and Ujjwal Guin.

A comprehensive test pattern generation approach exploiting the SAT attack for logic locking. IEEE Transactions on Computers, 72(8): 2293–2305, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2021:DSV

[ZGB^{+}21] X. Zeng, S. Garg, M. Barika, S. Bista, D. Puthal, A. Y. Zomaya, and R. Ranjan. Detection of SLA violation for big data analytics applications in cloud. IEEE Transactions on Computers, 70(5):746–758, May 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:PRB

[ZGD23] Baoquan Zhang, Haoyu Gong, and David H. C. Du. PMDB: a range-based key-value store on hybrid NVM-storage systems. IEEE Transactions on Computers, 72(5):1274–1285, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:TDI

[ZGG^{+}23] Jie Zhang, Song Guo, Jingcai Guo, Deze Zeng, Jingren Zhou, and Albert Y. Zomaya. Towards dataindependent knowledge transfer in model-heterogeneous federated learning. IEEE Transactions on Computers, 72(10):2888–2901, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zoni:2025:FBO

[ZGG25] Davide Zoni, Andrea Galimberti, and Davide Galli.

An FPGA-based open-source hardware–software framework for side-channel security research. IEEE Transactions on Computers, 74(6):2087–2100, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zyarah:2020:NSS

[ZGK20] A. M. Zyarah, K. Gomez, and D. Kudithipudi. Neuromorphic system for spatial and temporal information processing. IEEE Transactions on Computers, 69(8): 1099–1112, August 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2021:FAS

[ZGL^{+}21] Q. Zhou, S. Guo, H. Lu, L. Li, M. Guo, Y. Sun, and K. Wang. Falcon: Addressing stragglers in heterogeneous parameter server via multiple parallelism. IEEE Transactions on Computers, 70(1): 139–155, January 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhan:2020:DRL

[ZGLZ20] Y. Zhan, S. Guo, P. Li, and J. Zhang. A deep reinforcement learning based offloading game in edge computing. IEEE Transactions on Computers, 69(6):883–893, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:AFL

[ZGQ^{+}22] Jie Zhang, Song Guo, Zhihao Qu, Deze Zeng, Yufeng Zhan, Qifeng Liu, and Rajendra Akerkar. Adaptive federated learning on non-IID data with resource constraint. IEEE Transactions on Computers, 71(7):1655–1667, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2025:DMV

[ZGR^{+}25] Jinhua Zhu, Zhen Gao, Pedro Reviriego, Shanshan Liu, and Fabrizio Lombardi. Dependability of the K minimum values sketch: Protection and comparative analysis. IEEE Transactions on Computers, 74(1):210–221, January 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:NCA

[ZGWY22] Bowen Zhang, Huaxi Gu, Kun Wang, and Yintang Yang. A novel CONV acceleration strategy based on logical PE set segmentation for row stationary dataflow. IEEE Transactions on Computers, 71(6): 1466–1478, June 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:EDN

[ZHL^{+}24] Qinglong Zhang, Rui Han, Chi Harold Liu, Guoren Wang, and Lydia Y. Chen. ElasticDNN: On-device neural network remodeling for adapting evolving vision domains at edge. IEEE Transactions on Computers, 73(6):1616–1630, June 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:TMT

[ZHLR22] Jipeng Zhang, Junhao Huang, Zhe Liu, and Sujoy Sinha Roy. Time-memory tradeoffs for Saber+ on memoryconstrained RISC-V platform. IEEE Transactions on Computers, 71(11):2996– 3007, November 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2020:HBA

[ZHM20] L. Zhou, Y. Hu, and Y. Makris.

A hardware-based architectureneutral framework for realtime IoT workload forensics. IEEE Transactions on Computers, 69(11):1668– 1680, November 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:SSR

[ZHY^{+}25] Zhixin Zhao, Yitao Hu, Guotao Yang, Ziqi Gong, Chen Shen, Laiping Zhao, Wenxin Li, Xiulong Liu, and Wenyu Qu. SLOpt: Serving real-time inference pipeline with strict latency constraint. IEEE Transactions on Computers, 74(4):1431–1445, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zou:2021:ADR

[ZHYJ21] J. Zou, T. Hao, C. Yu, and H. Jin. A3C-DO: a regional resource scheduling framework based on deep reinforcement learning in edge scenario. IEEE Transactions on Computers, 70(2): 228–239, February 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:EHC

[ZJJ25] Sizhe Zhang, Kyle Juretus, and Xun Jiao. Exploring hyperdimensional computing robustness against hardware errors. IEEE Transactions on Computers, 74(6):1963–1977, June 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:ADC

[ZJW^{+}24] Yucheng Zhang, Hong Jiang, Chunzhi Wang, Wei Huang, Meng Chen, Yongxuan Zhang, and Le Zhang. Applying delta compression to packed datasets for efficient data reduction. IEEE Transactions on Computers, 73(1):

73–85, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:EEL

[ZLC^{+}22] Runyu Zhang, Duo Liu, Xianzhang Chen, Xiongxiong She, Chaoshu Yang, Yujuan Tan, Zhaoyan Shen, Zili Shao, and Lei Qiao. ELOFS: an extensible low-overhead flash file system for resource-scarce embedded devices. IEEE Transactions on Computers, 71(9): 2327–2340, September 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2023:NCT

[ZLC^{+}23a] Liehuang Zhu, Qi Liu, Zhuo Chen, Can Zhang, Feng Gao, and Zhongliang Yang. A novel covert timing channel based on bitcoin messages. IEEE Transactions on Computers, 72(10):2913–2924, October 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhuang:2023:EHP

[ZLC^{+}23b] Hongbin Zhuang, Xiao-Yan Li, Jou-Ming Chang, Cheng-Kuan Lin, and Ximeng Liu. Embedding Hamiltonian paths in k-ary ncubes with exponentiallymany faulty edges. IEEE Transactions on Computers, 72(11):3245–3258, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2021:ESD

[ZLH^{+}21] Lutan Zhao, Peinan Li, Rui Hou, Michael C. Huang, Peng Liu, Lixin Zhang, and Dan Meng. Exploiting security dependence for conditional speculation against Spectre attacks. IEEE Transactions on Computers, 70(7):963–978, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2022:RRA

[ZLL^{+}22a] Jianwei Zheng, Yu Liu, Xuejiao Liu, Luhong Liang, Deming Chen, and Kwang-Ting Cheng. ReAAP: a reconfigurable and algorithm-oriented array processor with compilerarchitecture co-design. IEEE Transactions on Computers, 71(12):3088–3100, December 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2022:FDF

[ZLL^{+}22b] Chunyang Zhou, Guohui Li, Jianjun Li, Quan Zhou, and Bing Guo. FAS-DQN: Freshness-aware scheduling via reinforcement learning for latency-sensitive applications. IEEE Transactions on Computers, 71(10):2381–2394, October 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:AGT

[ZLL^{+}23] Lizhi Zhang, Kai Lu, Zhiquan Lai, Yongquan Fu, Yu Tang, and Dongsheng Li. Accelerating GNN training by adapting large graphs to distributed heterogeneous architectures. IEEE Transactions on Computers, 72(12):3473– 3488, December 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2024:BFE

[ZLL^{+}24a] Honghong Zeng, Jie Li, Jiong Lou, Shijing Yuan, Chentao Wu, Wei Zhao, Sijin Wu, and Zhiwen Wang. BSR-FL: an efficient byzantinerobust privacy-preserving federated learning framework. IEEE Transactions on Computers, 73(8):2096–2110, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:GGP

[ZLL^{+}24b] Xiaoyu Zhang, Zerun Li, Rui Liu, Xiaoming Chen, and Yinhe Han. GAS: General-purpose in-memory-computing accelerator for sparse matrix multiplication. IEEE Transactions on Computers, 73(6):1427–1441, June 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2024:NLM

[ZLM^{+}24] Liang Zhao, Tianyu Li, Guiying Meng, Ammar Hawbani, Geyong Min, Ahmed Y. Al-Dubai, and Albert Y. Zomaya. Novel Lagrange multipliersdriven adaptive offloading for vehicular edge computing. IEEE Transactions on Computers, 73(12):2868–2881, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:LCL

[ZLS^{+}24] Chong Zhang, Songfan Li, Yihang Song, Qianhe Meng, Li Lu, Hongzi Zhu, and Xin Wang. A lightweight and chiplevel reconfigurable architecture for next-generation IoT end devices. IEEE Transactions on Computers, 73(3): 747–763, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:HPS

[ZLT^{+}24] Haiyan Zhang, Xinghua Li, Jiawei Tang, Chunlei Peng, Yunwei Wang, Ning Zhang, Yingbin Miao, Ximeng Liu, and Kim-Kwang Raymond Choo. Hiding in plain sight: Adversarial attack via style transfer on image borders. IEEE Transactions on Computers, 73(10):2405–2419, October 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:MUC

[ZLT^{+}25] Liang Zhao, Shuo Li, Zhiyuan Tan, Ammar Hawbani, Stelios Timotheou, and Keping Yu. A multi-UAV cooperative task scheduling in dynamic environments: Throughput maximization. IEEE Transactions on Computers, 74(2): 442–454, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:RTR

[ZLW^{+}24] Kai Zhang, Xuejia Lai, Lei Wang, Jie Guan, Bin Hu, Senpeng Wang, and Tairong Shi. Real-time related-key attack on full-round shadow designed for IoT nodes. IEEE Transactions on Computers, 73(2): 613–620, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhan:2022:PLE

[ZLWG22] Yufeng Zhan, Peng Li, Leijie Wu, and Song Guo. L4L: Experience-driven computational resource control in federated learning. IEEE Transactions on Computers, 71(4): 971–983, April 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:RHP

[ZLWJ23] Zhendong Zhang, Peng Liu, Weidong Wang, and Yingtao Jiang. RUPA: a high performance, energy efficient accelerator for rule-based password generation in heterogeneous password recovery system. IEEE Transactions on Computers, 72(4):900–913, April 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2023:HAO

[ZLZ^{+}23] Jingsen Zhu, Mengming Li, Xingjian Zhang, Kai Bu, Miao Zhang, and Tianqi Song. Hitchhiker: Accelerating ORAM with dynamic scheduling. IEEE Transactions on Computers, 72(8): 2321–2335, August 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:TFA

[ZLZ24a] Jie-Fang Zhang, Cheng-Hsun Lu, and Zhengya Zhang. TetriX: Flexible architecture and optimal mapping for tensorized neural network processing. IEEE Transactions on Computers, 73(5):1219–1232, May 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2024:LAS

[ZLZ24b] Tianming Zhao, Wei Li, and Albert Y. Zomaya. Learningaugmented scheduling. IEEE Transactions on Computers, 73(11):2548–2562, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:BPA

[ZMH^{+}25] Benteng Zhang, Yingchi Mao, Xiaoming He, Huawei Huang, and Jie Wu. Balancing privacy and accuracy using significant gradient protection in federated learning. IEEE Transactions on Computers, 74(1): 278–292, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:ASC

[ZML^{+}24] Tianze Zhang, Xuhong Miao, Yibin Li, Lei Jia, and Yinghao Zhuang. AUV surfacing control with adversarial attack against DLaaS framework. IEEE Transactions on Computers, 73(2): 327–339, February 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:IIL

[ZML^{+}25] Yunping Zhao, Sheng Ma, Tiejun Li, Jianmin Zhang, and Yuhua Tang. Intra- and inter-layer scheduling exploration and optimization for ReRAM-based DNN accelerators. IEEE Transactions on Computers, 74(9):3003– 3017, September 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2023:ECC

[ZMS^{+}23] Zhipeng Zhao, Joseph Melber, Siddharth Sahay, Shashank Obla, Eriko Nurvitadhi, and James C. Hoe. Exploiting the common case when accelerating input-dependent stream processing by FPGA. IEEE Transactions on Computers, 72(5):1343–1355, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:HEC

[ZNW^{+}24] Jiangjiang Zhang, Zhenhu Ning, Muhammad Waqas, Hisham Alasmary, Shanshan Tu, and Sheng Chen. Hybrid edge-cloud collaborator resource scheduling approach based on deep reinforcement learning and multiobjective optimization. IEEE Transactions on Computers, 73(1): 192–205, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:DSM

[ZOH^{+}25] Liang Zhang, Zhanrong Ou, Changhui Hu, Haibin Kan, and Jiheng Zhang. Data sharing in the metaverse with key abuse resistance based on decentralized CP-ABE. IEEE Transactions on Computers, 74(3):901–914, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:BBP

[ZQC^{+}25] Qingyang Zhang, Shuai Qian, Jie Cui, Hong Zhong, Fengqun Wang, and Debiao He. Blockchain-based privacypreserving deduplication and integrity auditing in cloud storage. IEEE Transactions on Computers, 74(5): 1717–1729, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2024:PSD

[ZQG^{+}24] Yue Zeng, Zhihao Qu, Song Guo, Baoliu Ye, Jie Zhang, Jing Li, and Bin Tang. SafeDRL: Dynamic microservice provisioning with reliability and latency guarantees in edge environments. IEEE Transactions on Computers, 73(1):235–248, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2020:NAE

[ZQY^{+}20] N. Zhang, Q. Qin, H. Yuan, C. Zhou, S. Yin, S. Wei, and L. Liu. NTTU: An area-efficient low-power NTTuncoupled architecture for NTT-based multiplication. IEEE Transactions on Computers, 69(4):520–533, April 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:EIA

[ZSC^{+}23] Qingyang Zhang, Dongfang Sui, Jie Cui, Chengjie Gu, and Hong Zhong. Efficient integrity auditing mechanism with secure deduplication for blockchain storage. IEEE Transactions on Computers, 72(8):2365–2376, August 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:GEF

[ZSC^{+}25] Mingyuan Zhao, Hao Sheng, Rongshan Chen, Ruixuan Cong, Tun Wang, Zhenglong Cui, Da Yang, Shuai Wang, and Wei Ke. A GPU-enabled framework for light field efficient compression and realtime rendering. IEEE Transactions on Computers, 74(4): 1168–1181, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zaruba:2021:STP

[ZSHB21] Florian Zaruba, Fabian Schuiki, Torsten Hoefler, and Luca Benini. Snitch: A tiny pseudo dual-issue processor for area and energy efficient execution of floating-point intensive workloads. IEEE Transactions on Computers, 70(11):1845– 1860, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2020:FFC

[ZSS20] R. K. Zhao, R. Steinfeld, and A. Sakzad. FACCT: FAst, Compact, and Constant-Time discrete Gaussian sampler over integers. IEEE Transactions on Computers, 69(1): 126–137, January 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:VSB

[ZSS^{+}22] Jiliang Zhang, Chaoqun Shen, Haihan Su, Md Tanvir Arafin, and Gang Qu. Voltage overscaling-based lightweight authentication for IoT security. IEEE Transactions on Computers, 71(2):323–336, February 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:OQC

[ZSWS24] Mengyuan Zhang, Tairong Shi, Wenling Wu, and Han Sui. Optimized quantum circuit of AES with interlacinguncompute structure. IEEE Transactions on Computers, 73(11):2563–2575, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zou:2024:VIC

[ZSX^{+}24] Yifei Zou, Shikun Shen, Mengbai Xiao, Peng Li, Dongxiao Yu, and Xiuzhen Cheng. Value of information: a comprehensive metric for client selection in federated edge learning. IEEE Transactions on Computers, 73(4):1152–1164, April 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2023:LLH

[ZTLW23] Danyang Zhu, Jing Tian, Minghao Li, and Zhongfeng Wang. Low-latency hardware architecture for VDF evaluation in class groups. IEEE Transactions on Computers, 72(6):1706–1717, June 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zaman:2022:PPL

[ZTT22] Mashiyat Zaman, Kotaro Tanahashi, and Shu Tanaka. PyQUBO: Python library for mapping combinatorial optimization problems to QUBO form. IEEE Transactions on Computers, 71(4):838–850, April 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhan:2023:ARB

[ZTY^{+}23] Dongyang Zhan, Kai Tan, Lin Ye, Xiangzhan Yu, Hongli Zhang, and Zheng He. An adversarial robust behavior sequence anomaly detection approach based on critical behavior unit learning. IEEE Transactions on Computers, 72(11):3286–3299, November 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2025:ECH

[ZTY^{+}25] Yingkun Zhou, Zhengshuyuan Tian, Wenhao Yang, Tingting Zhang, Jinpeng Ye, Chenji Han, Tianyi Liu, and Fuxin Zhang. ETBench: Characterizing hybrid vision transformer workloads across edge devices. IEEE Transactions on Computers, 74(6): 1857–1871, June 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:LLO

[ZTZ^{+}25] Sanle Zhao, Yujuan Tan, Zhaoyang Zeng, Jing Yu, Zhuoxin Bai, Ao Ren, Xianzhang Chen, and Duo Liu. LAShards: Low-overhead and self-adaptive MRC construction for non-stack algorithms. IEEE Transactions on Computers, 74(10):3490–3503, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2022:DMB

[ZWB^{+}22] Zhigao Zheng, Tao Wang, Ali Kashif Bashir, Mamoun Alazab, Shahid Mumtaz, and Xiaoyan Wang. A decentralized mechanism based on differential privacy for privacypreserving computation in smart grid. IEEE Transactions on Computers, 71(11): 2915–2926, November 2022.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zou:2022:CCA

[ZWC^{+}22] Kaiwei Zou, Ying Wang, Long Cheng, Songyun Qu, Huawei Li, and Xiaowei Li. CAP: Communication-aware automated parallelization for deep learning inference on CMP architectures. IEEE Transactions on Computers, 71(7): 1626–1639, July 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2023:FTM

[ZWC^{+}23] Ziming Zhao, Mingyu Wu, Xujie Cao, Haibo Chen, and Binyu Zang. Flock: Towards multitasking virtual machines for function-as-a-service. IEEE Transactions on Computers, 72(11):3153– 3166, November 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2025:LFT

[ZWJ^{+}25] Zhigao Zheng, Guojia Wan, Jiawei Jiang, Chuang Hu, Hao Liu, Shahid Mumtaz, and Bo Du. Lock-free triangle counting on GPU. IEEE Transactions on Computers, 74(3):1040–1052, March 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:SEO

[ZWL^{+}25] Qianhe Zhao, Rui Wang, Yi Liu, Hailong Yang, Zhongzhi Luan, and Depei Qian. Sifter: an efficient operator autotuner with speculative design space exploration for deep learning compiler. IEEE Transactions on Computers, 74(10):3251–3262, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2020:FEN

[ZWM20] X. Zhou, L. Wang, and A. Mishchenko. Fast exact NPN classification by co-designing canonical form and its computation algorithm. IEEE Transactions on Computers, 69(9):1293– 1307, September 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:SRS

[ZWO^{+}25] Jianshun Zhang, Fang Wang, Jiaxin Ou, Yi Wang, Ming Zhao, Sheng Qiu, Junxun Huang, Baoquan Li, Peng Fang, and Dan Feng. Scavenger+: Revisiting space-time tradeoffs in key–value separated LSM-Trees. IEEE Transactions on Computers, 74(10):3332–3346, October 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2024:EDR

[ZWSF24] Yang Zhou, Fang Wang, Zhan Shi, and Dan Feng. An efficient deep reinforcement learning-based automatic cache replacement policy in cloud block storage systems. IEEE Transactions on Computers, 73(1): 164–177, January 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2022:LTI

[ZWWY22] Pengzhan Zhou, Xin Wei, Cong Wang, and Yuanyuan Yang. k-level truthful incentivizing mechanism and generalized k-MAB problem. IEEE Transactions on Computers, 71(7):1724–1739, July 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:DDS

[ZWX^{+}25] Chen Zhang, Yang Wang, Zhiqiang Xie, Cong Guo, Yunxin Liu, Jingwen Leng, Zhigang Ji, Yuan Xie, and Ru Huang. DSTC: Dual-side sparse tensor core for DNNs acceleration on modern GPU architectures. IEEE Transactions on Computers, 74(2): 341–355, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2023:OFC

[ZWY^{+}23] Fang Zhou, Song Wu, Jianhui Yue, Hai Jin, and Jiangqiu Shen. Object fingerprint cache for heterogeneous memory system. IEEE Transactions on Computers, 72(9): 2496–2507, September 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:SKA

[ZWZZ24] Yuhan Zhang, Wenling Wu, Lei Zhang, and Yafei Zheng. Single-key attack on fullround shadow designed for IoT nodes. IEEE Transactions on Computers, 73(12):2776–2790, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2024:BPD

[ZXD^{+}24] Mengying Zhao, Shuo Xu, Lihao Dong, Chun Jason Xue, Dongxiao Yu, Xiaojun Cai, and Zhiping Jia. Branch predictor design for energy harvesting powered nonvolatile processors. IEEE Transactions on Computers, 73(3): 722–734, March 2024. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2024:EML

[ZXG^{+}24] Xiaotian Zhao, Ruge Xu, Yimin Gao, Vaibhav Verma, Mircea R. Stan, and Xinfei Guo. Edge-MPQ: Layer-wise mixed-precision quantization with tightly integrated versatile inference units for edge computing. IEEE Transactions on Computers, 73(11): 2504–2519, November 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2023:BFA

[ZXL^{+}23] Shihui Zheng, Ruihao Xing, Junlong Lai, Junkai Liu, Haofeng Wang, and Changhai Ou. Breaking fault attack countermeasures with sidechannel information. IEEE Transactions on Computers, 72(5):1396–1408, May 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:IAB

[ZXL^{+}24] Zongpu Zhang, Chenbo Xia, Cunming Liang, Jian Li, Chen Yu, Tiwei Bie, Roberts Martin, Daly Dan, Xiao Wang, Yong Liu, and Haibing Guan. Un-IOV: Achieving baremetal level I/O virtualization performance for cloud usage with migratability, scalability and transparency. IEEE Transactions on Computers, 73(7):1655–1668, July 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:SBA

[ZXMX25] Pan Zhang, Lei Xu, Lin Mei, and Chungen Xu. Sketchbased adaptive communication optimization in federated learning. IEEE Transactions on Computers, 74(1): 170–184, January 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:OCE

[ZXW^{+}24] Chen Zhang, Qingyuan Xie, Mingyue Wang, Yu Guo, and Xiaohua Jia. Optimal compression for encrypted key– value store in cloud systems. IEEE Transactions on Computers, 73(3):928–941, March 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:AFT

[ZXX^{+}25] Sirong Zhao, Guoqi Xie, Chenglai Xiong, Kenli Li, Xuejun Yu, Bo Wan, and Yiwen Jiang. AVL function table for LeafHooks insertion with obfuscated control flow integrity. IEEE Transactions on Computers, 74(4):1334–1347, April 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:DTA

[ZXY^{+}24] Ruirui Zhang, Zhenzhen Xie, Dongxiao Yu, Weifa Liang, and Xiuzhen Cheng. Digital twin-assisted federated learning service provisioning over mobile edge networks. IEEE Transactions on Computers, 73(2):586–598, February 2024.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zong:2021:PPT

[ZXZ^{+}21] Pengchen Zong, Tian Xia, Haoran Zhao, Jianming Tong, Zehua Li, Wenzhe Zhao, Nanning Zheng, and Pengju Ren. PIT: Processing-in-transmission with fine-grained data manipulation networks. IEEE Transactions on Computers, 70(6):877–891, June 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:REB

[ZXZ^{+}24] Qingyang Zhang, Chang Xu, Hong Zhong, Chengjie Gu, and Jie Cui. Revocable and efficient blockchain-based finegrained access control against EDoS attacks in cloud storage. IEEE Transactions on Computers, 73(8):2012–2024, August 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2020:SCA

[ZYD^{+}20] F. Zhang, B. Yang, X. Dong, S. Guilley, Z. Liu, W. He, F. Zhang, and K. Ren. Sidechannel analysis and countermeasure design on ARM-Based quantum-resistant SIKE. IEEE Transactions on Computers, 69(11):1681–1693, November 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2022:PPB

[ZYL^{+}22] Yiming Zhang, Lujia Yin, Dongsheng Li, Yuxing Peng, and Kai Lu. ParaX: Bandwidthefficient instance assignment for DL on Multi-NUMA many-core CPUs. IEEE Transactions on Computers, 71(11):3032–3046, November 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhou:2024:HHA

[ZYQ^{+}24] Ao Zhou, Jianlei Yang, Yingjie Qi, Tong Qiao, Yumeng Shi, Cenlin Duan, Weisheng Zhao, and Chunming Hu. HGNAS: Hardwareaware graph neural architecture search for edge devices. IEEE Transactions on Computers, 73(12):2693–2707, 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2020:IWC

[ZYXD20] B. Zhang, M. Yang, X. Xie, and D. H. C. Du. Idler: I/O workload controlling for better responsiveness on host-aware shingled magnetic recording drives. IEEE Transactions on Computers, 69(6):777–788, June 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2023:JJP

[ZYZ^{+}23] Yangming Zhao, Cheng Yang, Gongming Zhao, Yunfei Hou, Ting Wang, and Chunming Qiao. JointPS: Joint parameter server placement and flow scheduling for machine learning clusters. IEEE Transactions on Computers, 72(12): 3503–3518, December 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:TAA

[ZYZ^{+}25a] Jinkai Zhang, Yinghao Yang, Zhe Zhou, Zhicheng Hu, Xin Zhao, Liang Chang, Hang Lu, and Xiaowei Li. Trident: The acceleration architecture for high-performance private set intersection. IEEE Transactions on Computers, 74(4): 1152–1167, April 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2025:DPA

[ZYZ^{+}25b] Yilan Zhu, Honghui You, Wei Zhang, Jiming Xu, Qian Lou, Shoumeng Yan, and Lei Ju. DAHE: Parameter-adaptive and memory efficient FPGA acceleration of homomorphic encryption. IEEE Transactions on Computers, 74(8): 2687–2701, August 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2025:STM

[ZZ25] Yi-Wen Zhang and Hui Zheng. Slack time management for imprecise mixed-criticality systems with reliability constraints. IEEE Transactions on Computers, 74(5): 1577–1588, May 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zou:2023:CBC

[ZZC^{+}23] Qiang Zou, Yifeng Zhu, Jianxi Chen, Yuhui Deng, and Xiao Qin. Characterization of I/O behaviors in cloud storage workloads. IEEE Transactions on Computers, 72(10): 2726–2739, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2024:NTO

[ZZF^{+}24] Liang Zhang, Hongzi Zhu, Wen Fei, Yunzhe Li, Mingjin Zhang, Jiannong Cao, and Minyi Guo. Novas: Tackling online dynamic video analytics with service adaptation at mobile edge servers. IEEE Transactions on Computers, 73(9): 2220–2232, September 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2020:SFG

[ZZG20] C. Zhang, Y. Zeng, and X. Guo. Scrabble: A finegrained cache with adaptive merged block. IEEE Transactions on Computers, 69(1): 112–125, January 2020. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2023:EES

[ZZG^{+}23] Deze Zeng, Andong Zhu, Lin Gu, Peng Li, Quan Chen, and Minyi Guo. Enabling efficient spatio-temporal GPU sharing for network function virtualization. IEEE Transactions on Computers, 72(10):2963– 2977, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhao:2025:DCD

[ZZH^{+}25] Liang Zhao, Zijia Zhao, Ammar Hawbani, Zhi Liu, Zhiyuan Tan, and Keping Yu. Dynamic caching dependencyaware task offloading in mobile Edge Computing. IEEE Transactions on Computers, 74(5):1510–1523, May 2025. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2021:EPE

[ZZL21] Wei Zhang, Hang Zhang, and John Lach. Extending performance-energy tradeoffs via dynamic core scaling. IEEE Transactions on Computers, 70(11):1875– 1886, November 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zheng:2022:MRV

[ZZM^{+}22] Jiaqi Zheng, Zixuan Zhang, Qiufang Ma, Xiaofeng Gao, Chen Tian, and Guihai Chen. Multi-resource VNF deployment in a heterogeneous cloud. IEEE Transactions on Computers, 71(1):81–91, January 2022. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhu:2025:BSA

[ZZW^{+}25] Zixuan Zhu, Xiaolong Zhou, Chundong Wang, Li Tian, Zunkai Huang, and Yongxin Zhu. Bit-sparsity aware acceleration with compact CSD code on generic matrix multiplication. IEEE Transactions on Computers, 74(2): 414–426, February 2025. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zeng:2020:AIS

[ZZZ^{+}20] X. Zeng, T. Zhi, X. Zhou, Z. Du, Q. Guo, S. Liu, B. Wang, Y. Wen, C. Wang, X. Zhou, L. Li, T. Chen, N. Sun, and Y. Chen. Addressing irregularity in sparse neural networks through a cooperative Software/Hardware approach. IEEE Transactions on Computers, 69(7):968–985, July 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zhang:2023:MPA

[ZZZ^{+}23] Xianglong Zhang, Huanle Zhang, Guoming Zhang, Hong Li, Dongxiao Yu, Xiuzhen Cheng, and Pengfei Hu. Model poisoning attack on neural network without reference data. IEEE Transactions on Computers, 72(10):2978– 2989, October 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).